A Complete Bibliography of Publications in IMA Journal of Applied Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

06 October 2025 Version 1.122

Title word cross-reference

(2 + 1) [1026]. (2, 2) [460]. 1 [1304, 1000, 1225, 1155, 832, 1289, 1006, 1437, 932, 880, 1745, 1037]. 1: 1 [851]. 2 [1304, 1741, 1003, 1391, 1475, 1289, 858, 994, 1411, 1570]. 2 × 2 [857, 1557]. 3 [1528, 1225, 1522, 1034, 1645, 1457, 542, 1074, 1289, 1042, 1285, 1388, 1036, 1440, 1022, 505]. v [1813]. Ai(x) [166]. C^{∞} [935]. EC′ [493]. G_{x_{0}}^{ϵ} [832]. I_{1} [1336]. I_{4} [1336]. I_{6} [1336]. J_{v}(t) [165]. J_{v}(t), I_{v}(t) [165]. Ji_{v}(t) [165]. l^{1} [1826]. L^{2} [954, 1818, 1361, 1745]. λ [1543, 477]. m [1611]. m > −1 [292]. R^{N} [1134, 1138]. KL_{∗} [1542]. m → +∞ [228]. N [882, 1493, 545, 1451, 790, 1075, 1332]. ∨_{0} [763]. ω [477]. p [1843, 1286]. P − N [48]. pnpn [327]. q [64, 89, 145]. R^{2} [359]. T [1549]. θ [1069]. u_{t} = (|u|^{m−1}u_{x})_{x} [292]. u_{t} = ∇ · (u^{m}∇u) [228]. W [1418]. x^{3/2} [166]. z(x, y) [881].

-analogue [64]. -boundedness [763]. -convexity [1437]. -critical [1745]. -D [1304, 542]. -dimensional [1075, 1026]. -gradient [1818]. -irreducibility [1843]. -Junctions [48]. -Laplacian [1286]. -methods [1069]. -mode [545]. -order [832]. -part [790]. -peakon [1493]. -polynomial [89, 145]. -resonant [851]. -stability [954, 1542, 763]. -Sylvester [1549]. -Symmetries [1543, 935]. -torus [1451]. -weighted [1826].

1 [1312]. 107’ [1133]. 1D [1837].

2 [1077]. 27 [37]. 2D [1012, 1769, 1767]. 2nd [1324].

35 [219]. 37 [278]. 3D [1770, 1767].

4 [1209]. 46 [437]. 49 [450].

502 [1783]. 55 [569].

60th [1469]. 67 [760].

70th [1329].

82j [37]. 87c [219].

90a [278]. 92b [437]. 93j [450]. 97b [569].

A. [522]. abdominal [1345]. above [1765]. abrasion [1125]. absence [1403]. absolute [1697]. absorbers [1768]. Absorbing [773, 377, 1404, 399]. absorption [256, 281, 417, 117, 544, 1312]. absorptive [1744]. absorptivity [653]. abundance [41]. Accelerated [1739, 1863]. accelerating [375]. Acceleration [1779, 1482, 1677, 896]. accidents [1284]. according [900]. account [1479, 1132]. accretion [201, 296, 270]. accurate [1685, 348, 92]. Acoustic [413, 1556, 1656, 1867, 586, 1176, 1218, 1703, 1774, 106, 302, 428, 663, 422, 1074, 1007, 31, 336, 691, 971, 1039, 638, 1652, 1626, 56, 1470, 412, 113, 1803, 508]. acoustic-gravity [1470]. acoustically [930]. acoustics [95]. actin [1325]. Active [921, 796, 1728, 1778, 1308, 1351]. active-dissipative [1778]. actuated [1649]. Adapting [1479, 1470]. Adaptive [1305]. Addendum [219, 278, 137]. addition [1626]. adhesion [1681, 28, 27]. adhesive [1491]. adjacent [216]. adjoint [695, 785, 1622]. adjointness [1389]. adsorption [293]. Advances [1469]. Advection [352, 591, 1516, 555, 526, 1611, 1018, 1316, 709]. Advection-diffusion [352, 591, 709]. advection-dispersion [1316]. Adversarial [1828]. Aerodynamic [135]. aerodynamics [1246]. aerofoil [1508]. aerofoils [1114, 1008]. aerosol [93, 137, 225, 217]. affine [1071]. age [1118, 657, 1458, 1539, 1600, 1525, 1794]. age-dependent [657]. age-distribution [1118]. age-structured [1600, 1525, 1794]. agent [846]. aggregation [981]. Air [1271, 1707, 1180, 1327, 1742, 722]. Air-cushioning [1271]. air-gap [1707]. aircraft [722, 1183]. airfoil [316]. Airy [166]. algal [1588]. Algebra [23]. Algebraic [562, 378, 683]. algebraically [24]. algebras [728]. Algebro [1501]. Algebro-geometric [1501]. Algorithm [1498, 101, 1538, 1361, 1753, 1548]. algorithms [1739, 1779, 90, 1052]. Allee [1836]. allelopathy [1251]. Allen [1506, 1373]. Allowed [518]. alloy [924, 1759, 279, 185, 806, 391]. alloys [74, 794, 186, 1132]. almost [1429]. along [1380, 733, 1799, 24, 944]. along-track [1380]. alpha [1076]. alternating [1739, 1027, 765, 1764]. alternative [1229, 626, 1576]. ambient [305, 1356, 1289]. Ampère [1247]. amplifier [1240, 1511]. Amplitude [69, 1269, 390, 1609, 307, 477, 977, 1267]. Amplitude-dependent [69]. analogous [1641]. analogue [430, 64]. analogues [1337]. analogy [113]. analyses [1208, 1020]. Analysing [1249, 1268]. Analysis [1627, 1850, 1024, 1801, 84, 1511, 1670, 1647, 666, 1352, 1006, 651, 1662, 48, 1333, 1337, 1179, 1150, 1069, 1600, 1663, 737, 861, 57, 1810, 1498, 924, 891, 1400, 255, 1404, 331, 609, 646, 1303, 1500, 1770, 934, 954, 33, 623, 621, 1754, 1775, 811, 863, 1096, 85, 1871, 928, 822, 775, 1211, 1383, 1832, 583, 231, 242, 546, 564, 907, 1282, 1462, 1192, 1447, 1497, 220, 161, 1698, 943, 1120, 1256, 1672, 826, 1559, 1872, 253, 1345, 894, 162, 1363, 1526, 757, 722, 358, 327, 634, 784, 1173, 1624, 1101, 1548, 386, 391, 730, 5, 1393, 1057, 1137]. analysis [1434]. Analytic [1061, 406, 645, 1508, 1747, 1789, 1052, 1085, 595]. Analytical [354, 483, 745, 1394, 1304, 57, 1442, 699, 742, 972, 550, 1129]. Analyticity [1092, 1107, 1391, 1778]. Analyzing [1831]. anchoring [1634]. Andy [978]. aneurysms [1345]. Anger [165]. angiogenesis [1224]. angle [912, 595, 1596, 1817]. angled [1855, 1839]. Angles [152]. angular [1556, 1656, 270]. Animal [1521]. Animals [1, 37]. anisotropic [1478, 255, 955, 994, 1708, 1751, 694, 1345, 914]. anisotropically [175]. Annular [52, 820, 429, 444, 904, 1052, 1119, 1555]. anomalies [602]. ansatz [581]. ansatz-based [581]. antennas [421]. anti [1495, 836]. anti-phase [1495]. anti-plane [836]. Antiplane [729, 819]. antiretroviral [1600]. antisymmetric [1188]. any [732]. anymore [1635]. aortic [1345]. aperiodic [132]. apertures [538, 569]. apparatus [374]. Appl [219, 278, 450, 569, 437, 37]. Applicable [23, 1343]. Application [1365, 240, 1853, 704, 19, 819, 1538, 68, 423, 450, 863, 1437, 1216, 290, 1345, 875, 1346, 751]. applications [1863, 1490, 139, 1685, 1468, 974, 642, 1501, 1044, 1743, 790, 1314, 1541, 971, 515, 1841, 526, 1631, 1413]. Applied [1461, 1821, 1133, 791, 1648, 503, 204, 1643, 1142, 1469, 1028, 1377, 1147, 846, 1008, 310]. applying [1289, 374]. Approach [115, 1863, 992, 1432, 484, 1861, 1263, 810, 891, 708, 212, 1532, 565, 1072, 1699, 340, 378, 1023, 1743, 1007, 1042, 1191, 1696, 691, 1497, 468, 1622, 1595, 44, 174, 1310, 1288, 575, 1181, 1277, 1575, 1268, 1679, 655, 683, 1825]. approaches [1563, 811, 1140]. approximants [704]. Approximate [862, 371, 270, 163, 710, 196, 685, 530, 1798, 1389]. approximated [22]. Approximating [60, 949]. Approximation [1450, 1663, 1785, 1858, 1538, 785, 1370, 323, 832, 240, 1052, 344, 1054, 1671, 1316]. approximations [1781, 515, 348, 1660, 147, 1862]. arbitrarily [342]. arbitrary [779, 1263, 1519, 1572, 415, 689, 501, 595, 1596, 446, 826, 713, 1046, 359]. arc [1764, 143]. arc-length [143]. area [1712]. areas [1139]. argument [166, 324, 575]. arising [771, 1188, 208, 318, 1549, 319, 507, 96, 1530, 1381, 1254, 266, 1179, 389, 358, 76, 634, 237, 100, 558, 1382]. array [1078, 868, 62, 474, 636, 543, 1234, 550, 950, 587, 688]. Arrays [16, 927, 1816]. Arrhenius [1750, 653]. arsenic [242]. arterial [1336]. arteries [1103]. article [1107]. Articular [1365, 1346]. articulated [105]. artificial [1785, 1006]. aspect [304, 1150]. aspects [1746, 184]. assembly [1359]. Assessment [412]. assigned [1038]. assignment [14]. assist [912]. associated [1169, 664, 944, 651, 1841, 1766, 222, 330, 382]. asymmetric [1678, 369]. Asymptotic [692, 714, 20, 1404, 1051, 739, 1065, 733, 1079, 1770, 623, 98, 272, 546, 564, 1765, 749, 427, 750, 1497, 515, 239, 434, 1120, 1672, 311, 491, 894, 654, 1233, 722, 327, 1236, 1624, 1101, 402, 783, 799, 833, 257, 102, 891, 1400, 333, 1802, 609, 646, 385, 1707, 1500, 33, 811, 1370, 497, 323, 378, 495, 231, 242, 1421, 841, 171, 220, 203, 1800, 181, 442, 568, 330, 433, 1657, 1845, 162]. Asymptotically [1844, 1807, 1786]. Asymptotics [1741, 1640, 1855, 486, 1384, 1563, 678, 1775, 767, 1665, 1314, 1839, 1264, 1656, 1833]. asynchronous [1688]. Atanacković [63]. atmosphere [21, 697]. atmospheres [1180]. Atmospheric [303]. atomic [1020]. attacks [1828, 1824]. Attenuation [215]. attitude [1377]. attraction [1452, 1868]. attraction-repulsion [1452]. attractor [1475]. attractors [1476, 1804, 845]. audio [1240]. autocatalator [506]. autocatalysis [858]. autocatalytic [521, 441, 532, 579, 614, 846, 471, 496, 547, 896]. autocatalytic-type [896]. autoignition [578]. automatic [282]. autonomous [1381, 1841, 838, 1434]. autoregressive [1631]. Autosoliton [669].

Averaged [1474, 491]. averaging [551, 933, 977]. avoid [466]. avoiding [1553]. Avrami [1531]. awareness [1848]. axial [820, 187, 1086, 1846]. axially [1732, 879, 436]. Axisymmetric [429, 563, 276, 658, 814, 1040, 1078, 489, 1509, 211, 1840, 1327, 1811, 1559, 28]. azimuthal [1338].

B [278, 1094]. background [1708, 1751]. backing [234]. Bäcklund [1156, 583, 697]. backtracking [1662]. backward [1828, 1446, 1088]. Bailey [64]. Balance [190, 545]. ball [1850, 1801]. balls [1134]. BAM [1288]. bar [1837, 1002, 937]. barrier [1287, 321, 382]. barriers [868, 198, 158]. bars [726, 955]. base [1029]. based [668, 891, 945, 811, 1042, 581, 1842, 1052, 1140, 685, 1230, 404, 1125, 1548, 1616]. basic [203]. Basis [36, 1014, 842, 751]. batch [1381]. Bayesian [945]. beaches [86]. beam [681, 444, 774, 1635, 651, 1141, 1830, 1029, 751, 1130, 1773, 1033]. beam-columns [1130]. beam-like [444]. beams [62, 1155, 162, 842]. bean [1591]. bearing [1321, 1420, 1628, 944, 543, 1137]. beat [1827]. bed [318, 206]. Bedrosian [1085]. behavior [949, 1845]. Behaviour [959, 1745, 799, 1459, 1619, 692, 833, 257, 714, 1051, 453, 553, 733, 1079, 1601, 567, 951, 1174, 1336, 354, 1210, 1676, 171, 427, 434, 1812, 1714, 902, 1233, 181, 1082, 1134]. behind [1407]. Bell [1191]. bells [1644]. below [211]. Beltrami [1747, 1722]. Bénard [476]. Bending [1870, 155, 716, 1338, 1375, 822, 215, 1102, 1337, 1492, 1618]. Benjamin [364]. bent [342, 289]. Bernoulli [1281, 651, 842]. Berry [430]. Bessel [165, 971, 324]. bets [926]. between [1227, 230, 1763, 309, 211, 1819, 531, 1705, 859, 77, 1386, 474, 549, 1319, 1427, 674, 1166, 535, 392, 92, 1363, 1331, 602]. beyond [1731, 1146]. bi [879, 1621, 1331, 1723]. bi-axially [879]. bi-Laplacian [1723]. bi-material [1331]. bi-stochastic [1621]. BIAcore [974]. bianisotropic [1299]. biased [400]. biaxial [1471, 1517]. bidirectional [1683]. BIE [986]. BIE-FE [986]. bifurcated [510]. bifurcating [1536]. Bifurcation [1475, 1752, 1174, 775, 558, 1057, 1137, 1731, 924, 38, 1674, 945, 1495, 1847, 438, 267, 1245, 1727, 1124, 1101, 1041]. Bifurcations [770, 1763, 1476, 1472, 1102, 1736]. big [1463]. biharmonic [1062, 1597, 76]. bilayer [1767]. bilevel [1831, 1825]. bilinear [121, 64, 104]. bimaterials [914]. Binary [1191, 1205, 1684, 186, 1759, 279, 185, 806, 391]. Bingham [643]. bio [1337]. bio-inspired [1337]. biofilm [1792]. biological [1243, 1350, 800]. biology [1144, 469]. bioluminescence [1360, 928]. biomass [829]. biomechanical [738]. biomechanics [1793]. biorthogonal [1320, 330]. birefringence [342]. Birth [1522]. birthday [1329, 1469]. bispherical [225, 217]. Bivariational [263]. black [1774, 1535]. Blake [1578, 1661]. blebbing [1326]. Block [541]. Block-size [541]. blocks [742]. Blow [1744, 620, 1528, 1247, 831, 567, 612, 675, 767, 1291, 295, 1311, 1214, 1440]. blow-up [1528, 1247, 831, 567, 612, 675, 767, 1291, 295, 1311, 1214, 1440]. blowflies [1304]. Blowing [1348]. Blowing-up [1348]. Blowup [801]. bodies [1490, 1275, 150, 1414, 1072, 501, 1272, 841, 1714]. body [1091, 1653, 1252, 332, 754, 596, 776, 434, 1209, 1128, 1399, 654, 713, 1246, 1667]. Bogdanov [1731]. boiler [1087]. boiling [1181]. Boltzmann [22, 1163, 1164, 1167, 1166, 246, 1168, 1159, 1165]. bond [687]. bonded [1544]. bone [1793]. Book [23]. books [258]. boost [1474]. border [1475, 1618]. Born [1237]. borne [1795]. Bose [1609]. both [1441]. bottom [989, 1411]. bounce [1850, 1801]. bound [269, 12, 956, 1742]. boundaries [249, 1516, 1638, 944, 1652, 1611, 297, 1738, 399]. Boundary [84, 1569, 1769, 465, 157, 835, 446, 748, 291, 540, 701, 833, 1744, 1606, 1815, 178, 1420, 216, 196, 759, 1861, 1121, 500, 69, 1213, 73, 1028, 599, 639, 787, 1487, 14, 1685, 485, 318, 458, 856, 55, 739, 385, 918, 206, 79, 665, 1011, 1226, 1789, 917, 740, 169, 502, 1344, 916, 1749, 514, 497, 863, 1031, 58, 116, 1088, 761, 1164, 144, 306, 1842, 384, 948, 474, 968, 1694, 1623, 354, 451, 1425, 685, 938, 24, 97, 1208, 1220, 536, 1004, 598, 698, 1690, 1556, 1656, 396, 154, 975, 1835, 988, 1449, 387, 266, 635, 1661, 768, 207]. boundary [1806, 1502, 462, 641, 586, 1374, 39, 556, 1233, 1013, 915, 237, 865, 773, 399, 374, 1486, 81, 163, 382, 343, 1870, 1566, 1632, 1444, 275, 147, 751, 1709, 1756, 359, 682, 1494, 383, 459, 1045, 1455]. boundary-element [384]. boundary-finite [1344]. Boundary-Integral [84]. boundary-layer [216, 500, 69, 385, 206, 948, 641, 1374, 1486]. Boundary-value [748, 1028, 599, 639, 938, 536]. bounded [1749, 1035, 1393, 682]. boundedness [208, 1820, 763]. bounding [1435]. Bounds [958, 265, 248, 240, 121, 512, 566]. Boussinesq [1805, 620, 1030, 906, 1533, 767]. Boussinesq-type [906]. Bragg [1497]. brain [1796]. branch [304, 312, 350, 319, 582]. branches [38]. Break [1185, 1561, 1065]. Break-up [1185, 1561, 1065]. breakdown [1711]. breaking [514]. breakup [1273, 837, 867]. breast [1853]. Bresse [1117]. bridge [1435, 398, 662, 1447, 758]. bridges [1599]. brief [1161]. Brinkman [994, 1641, 404]. broad [1486]. broadband [1551]. Brownian [1583, 501]. Broyden [1382]. Bubble [1579, 367, 1015, 1596, 1813, 1001]. bubbles [1579, 1220, 1661, 372, 1302, 1710]. bubbly [1037]. buckled [1444]. Buckling [774, 11, 877, 1097, 826, 827, 1129, 1101, 1130]. Buffer [1752]. bulging [1105]. bulging/necking [1105]. Bulk [6, 15, 1227, 1495, 525]. Bulk-Reacting [6, 15]. Bulkley [560, 465]. bundle [1581]. Bunsen [946]. buoyancy [1291, 582]. buoyancy-driven [1291]. burden [1604]. Burgers [337, 1197, 1111, 518, 933]. buried [1409, 1399]. Burton [1034]. business [611, 1323]. BV [1295]. BVP [469]. BZ [1059, 1254].

C [522, 724, 728]. C-symmetries [728].

cables [201, 1599]. Cahn

[703, 333, 243, 1729, 1506, 1373]. Cahn-like [1506]. cake [564]. calculate [92]. calculating [78]. calculation [240, 443]. calculations [148]. calculus [426, 1841]. Calderón [1032]. calibration [1364, 44]. call [1535]. calorimetry [727, 880]. Camassa [1214, 1493]. Can [1827]. cancer [1683, 1853]. canned [673]. Canonical [1310]. cap [1674]. capacitance [159]. capacities [1415]. capacitor [1300]. capacity [46, 323, 1800, 39, 283]. Capillaries [120]. Capillary [1398, 867, 807, 182, 849, 1681, 1226, 592, 439, 972, 1119, 1833, 525]. capillary-gravity [439]. capital [1323]. caps [748]. car [1677, 730]. car-following [1677, 730]. carbon [317, 1281, 1629, 509, 539, 1392]. carbon-climate [1392]. carbon-coated [539]. cardiac [1431]. Cardiovascular [1225]. carrier [1402]. carrying [1415]. cartilage [1346, 1365]. cascade [1514]. case [1483, 528, 1634, 778, 994, 745, 636, 455, 1848]. cases [777, 1063, 1635, 1579]. Casting [112, 1670, 1707, 311]. catalysed [982]. catalyst [941]. catalytic [830]. catalyzed [1173]. cathode [170]. Cauchy [1739, 73, 1280, 281, 702, 884, 1027, 530, 835, 1064, 1397]. Cauchy-Stefan [73]. caused [1209, 376]. cavitating [1037]. cavities [196, 1318, 1727]. cavity [1261, 408, 214, 1020, 1756]. celebrating [1329]. cell [1118, 883, 1262, 1325, 1477, 941, 1600]. cells [656, 1314, 1208, 372]. cellular [981]. cellulosic [435, 455]. center [1522]. central [836]. Centre [1639, 124]. centre-modes [124]. centred [718]. centres [1536]. ceramic [1261, 539, 601, 679]. ceramics [707]. Certain [89, 145, 791, 901, 695, 1285, 744, 1279, 247, 310]. CH [1397]. chain [1689, 1550]. change [286, 1143, 806, 391]. changes [989, 236]. changing [701, 1791, 528, 1455]. channel [989, 1617, 1776, 1664, 1665, 1844, 911, 1219]. Channell [1592]. channels [1019, 427, 1587, 1813, 1593]. chaos [1303, 258, 855]. Chaotic [1068, 668, 227, 1174, 104]. characteristic [712]. characteristics [946, 1377, 143]. Characterization [1105, 1500, 1723]. Characterizing [1222, 1436]. charge [407, 1507, 268, 379, 419, 341, 923, 264]. Charged [303]. chatter [1128]. Chebyshev [589]. chemical [288, 676, 1591, 858, 441, 579, 686, 471, 1812, 1255, 432, 772, 1293]. Chemically [381, 1693, 1165]. chemostat [1137]. chemotactic [1795]. chemotaxis [1483, 831, 942, 981, 1584, 1452, 1820, 1868]. chemotaxis-fluid [1820]. chimney [909]. China [745]. chiral [585, 690, 793]. chirp [1024]. Chlamydia [1186]. choking [1307]. chordae [1601]. ciliary [111]. circle [777, 1192]. Circular [84, 401, 196, 927, 1443, 1513, 1380, 138, 724, 276, 680, 1597, 1109, 1315, 1587, 1234, 510, 1374, 607, 1615, 1085, 718, 1723]. circulatory [1262, 535]. circumferential [613, 494]. clamped [1870]. Clarkson [689]. class [527, 487, 197, 1557, 65, 1438, 14, 1240, 1511, 1446, 821, 869, 695, 1362, 790, 1285, 282, 931, 1542, 1301, 324, 1714, 1652, 141, 76, 960, 70, 238, 1046, 326, 247, 1434, 292, 533]. class-D [1240, 1511]. classes [1414, 518]. Classical [1146, 984, 780, 1677, 83, 890, 1233, 1279, 499, 1416]. classically [1666]. Classification [900, 881, 1114]. clearances [1872]. Climate [1124, 1143, 1698, 1672, 1392]. climbing [957]. cloaking [1541]. close [1747, 490, 1702, 1246]. Closed [1513, 1196, 836, 1140, 418]. Closed-form [1513, 836, 1140, 418]. closely [159, 950]. closing [1766]. closure [491, 933]. Clouds [303, 1001]. clustering [1651, 1659]. clutch [1004]. CMath [1578]. coal [1180]. coalescence [912, 1221]. coated [759, 1404, 539]. coating [144, 1009]. coaxial [1725]. cocoa [1810]. cod [658]. cod-ends [658]. Coefficient [188, 1791, 331, 528, 1002, 1489, 1191, 501, 1529, 604, 651, 475, 518, 842]. coefficients [746, 1446, 1410, 1056, 171, 1315, 1234, 1610, 1803, 1455]. Coexistence [1112, 800, 286]. coffee [1591]. Cohen [1035]. Coherent [1193, 1203, 1694]. cold [788, 1090]. collagen [1343]. collapse [1532]. collision [1475, 1872, 1125]. collision-based [1125]. collisionless [1592, 788]. Column [16, 11, 699]. columns [1130]. combinations [288]. Combined [820, 546, 986]. combining [1576]. combustible [540, 468, 511, 578]. combustion [318, 1182, 319, 507, 1750, 395, 266, 1179, 237, 558, 655, 383, 459]. comment [837]. Comments [126]. common [808]. communication [1230]. commutative [1755]. compact [1793]. compactly [1320]. Comparative [811, 1178]. compared [1395]. Comparison [271, 1819, 726, 1013]. compartment [1495]. compartment-bulk [1495]. compatible [1319]. competing [1868]. competition [1227, 1516, 674, 1058, 1259, 1407, 1137, 94, 456, 1455, 533]. competition-diffusion [533]. competitive [1483, 346, 1056, 1112, 1179, 1610, 1149]. complement [1545]. complementarity [1474, 1408, 875]. Complementary [218, 642]. Complete [1362, 295, 176, 331, 1219]. completed [462]. completeness [695, 71]. Complex [1752, 1619, 1528, 148, 139, 1643, 819, 560, 1377, 1743, 1265, 324, 1602, 1413]. complexing [846]. complexity [26]. compliant [1850, 1801, 144, 968]. component [1164, 1214, 1609, 1493, 1709]. components [1585, 806]. componentwise [1828]. composite [19, 1252, 49, 937, 512, 566]. composites [1053, 679, 265, 1279, 326]. Composition [1083]. compositional [1232]. compound [1185]. compressed [1732, 1129]. compressibility [1676, 911]. Compressible [1701, 487, 500, 458, 818, 895, 502, 597, 908, 267, 449, 822, 1210, 454, 1579, 1388, 418, 635, 1363, 556, 889, 1486, 1632, 1692]. compression [775, 840, 1101]. compressive [1111]. Computation [294, 143, 1095, 356, 710, 772, 1026]. computational [1340, 883, 1468, 1147, 26, 1717]. computationally [598]. computations [1572]. compute [1382]. computed [1830]. Computer [622, 129, 1528]. computers [1688]. Computing [1785, 1854]. concentrated [39]. concentrating [965]. Concentration [280, 242]. concentrations [288, 443]. concentric [230, 1873]. concept [139]. conceptual [1672]. condensates [1609]. condensed [322, 1178, 993]. condensed-phase [1178, 993]. condition [833, 625, 710, 1823, 249, 1364, 739, 385, 1011, 917, 415, 1164, 474, 410, 1425, 95, 295, 826, 1374, 607, 412, 1566, 508, 1756]. Conditional [1156, 660]. Conditions [1307, 193, 864, 1420, 759, 1861, 419, 1569, 1519, 1685, 347, 79, 105, 1155, 916, 1749, 514, 761, 1314, 285, 451, 1208, 1004, 1556, 1656, 975, 1449, 644, 1502, 1013, 1626, 925, 284, 1608, 773, 374, 571, 655, 359, 682, 383, 459]. Conducting [1012, 692, 759, 726, 323, 1378, 765, 654, 315, 1110, 1394, 1711]. Conduction [84, 886, 1356, 498, 282, 1318, 392, 636, 905, 225, 217]. conductive [1399, 1756]. conductivity [32, 1716, 435, 156]. conductor [1115, 296, 920]. conductors [379, 1691, 296]. conduits [1113]. cone [809, 1577, 1364, 980, 615, 1109, 1315, 1830, 1486]. cone-beam [1830]. coned [1321]. Conference [1193]. confidence [1749]. configuration [1634]. configurations [1430]. confinement [1840]. conflict [460]. confluent [431]. confocal [1564]. Conformal [573, 859, 96, 1800, 2, 1008]. Congruent [264]. conical [996]. conically [693]. conjecture [430]. conjugate [702, 271]. conjugated [1379]. connected [787, 940, 1443, 301, 1597, 1447, 1614, 781, 359]. connections [839]. conservation [857, 813, 1577, 1642, 893, 801, 1464, 1050, 1295, 606, 1389]. conserved [333, 951]. conserving [463]. consistent [1690, 1818, 265]. consolidation [1506]. constant [732, 451, 956, 1374, 270, 557]. constants [288, 665, 1723]. constitutive [1414, 1864, 192, 524, 326]. constrained [1863, 726, 1622, 1029]. constraint [1131, 524]. constraints [182, 1237]. Construction [1071, 1014, 1085, 1320, 290]. constructions [1501]. Constructive [1460, 397, 1829]. consumption [180, 219, 1250, 383, 459]. Contact [1453, 861, 1850, 870, 1354, 762, 1801, 979, 732, 1467, 912, 934, 1042, 898, 1771, 1510, 584, 1518, 1473, 1331, 1295, 28, 27, 1616]. containers [1151, 1742]. Containing [303, 1762, 1541]. contaminated [668]. context [494]. continuation [1737, 143, 1268]. continued [769]. Continuous [112, 663, 517, 1471, 1517, 886, 1670, 1707, 1446, 645, 311, 902, 1343, 1854]. continuous-casting [311]. continuously [1215, 1358]. continuum [1571, 1507, 1164, 61]. contract [1287]. contracting [622]. contrast [1855, 1839, 1671]. contrasting [841]. contribution [881]. Control [1846, 673, 325, 1474, 1553, 36, 1549, 1504, 1044, 183, 1431, 1381, 1848, 1649, 1351]. Controllability [1322]. controllable [821, 869]. controlled [992, 1511, 589, 495, 409]. controlling [1202]. controls [282]. convected [929]. Convection [740, 376, 673, 1790, 1215, 476, 1784, 590, 211, 856, 665, 1520, 281, 961, 948, 1075, 51, 1318, 909, 988, 976, 404, 641, 1374, 1371, 391]. convection-diffusion [665]. Convective [1697, 1629, 98, 185, 361, 298, 358]. convective-diffusive [358]. conventional [947]. convergence [1503, 1871, 1296, 1230, 195, 1277, 1060]. convergences [311]. convergent [1807, 1786]. converter [1474]. convex [1863, 1061, 809, 642, 1370, 1842, 501, 900]. Convexity [1838, 542, 1437]. Convolution [875, 1149]. cooled [35, 403, 502]. cooling [133]. cooperation [1058]. cooperative [1149]. coordinate [3, 1626, 677]. coordinates [1423, 1136]. copper [745]. Coriolis [1219]. corner [1580, 1081, 299, 1416]. corners [339]. Correction [1790, 1815, 1807, 1850, 1855, 1856, 1869]. corrector [1230]. correlation [1303]. Corrigendum [450, 569, 1751, 437, 1133, 37, 225]. corrugated [1157]. Cosserat [616]. cost [1604, 1437]. costs [1400]. Couette [492, 929, 1203, 370]. Coulomb [664, 1448]. Coulomb-like [1448]. counterions [1834]. couple [1625]. Coupled [441, 436, 1459, 1154, 874, 1474, 1281, 565, 1567, 1729, 1495, 1533, 1377, 506, 1167, 439, 1068, 532, 471, 1622, 1496, 1547, 561, 568, 1679, 238, 1064, 1046, 1515, 1441]. coupled-inductor [1474]. couples [1454, 1582]. coupling [1232, 883, 521, 1441]. couplings [1679]. covered [1860]. crack [734, 647, 934, 1409, 796, 388, 836, 1695]. cracks [1042, 907, 1559, 1334]. creep [975]. criteria [1138, 1440]. criterion [180, 219, 926, 776]. Critical [1721, 1868, 193, 304, 69, 305, 1194, 1745, 7]. Criticality [156, 180, 219]. Cross [954, 1219, 201, 196, 1712, 839, 973, 1253, 1593]. Cross-diffusion [954, 839, 973, 1253]. Cross-equatorial [1219]. cross-section [196]. cross-sectional [1712]. crowd [1638]. crust [1856, 1851]. crystal [1663, 729, 187, 80, 1809, 1728, 183, 1440, 1081, 207, 189, 995]. crystalline [231, 242]. crystallization [1505]. crystals [958, 252, 1471, 819, 1650, 1634, 592, 301, 724, 872, 634, 1517]. CTL [1312]. cube [889]. cubic [506, 390, 579]. cuboid [1025]. current [1066, 923, 629, 260, 1764, 143, 224, 400, 327]. current-voltage [143, 327]. currents [153]. Curvature [1735, 249, 517, 1562, 1494]. curvature-dependent [1494]. curve [1730, 90, 239, 327]. curve- [90]. curved [1607, 1760, 849, 206, 1596, 1652, 181, 1008, 1150, 1544]. curved-ground [1150]. Curves [1752, 857, 60]. Curvilinear [3, 1494, 573, 907]. cushioning [1271, 1327]. cusp [1334]. cut [723]. cut-off [723]. cutting [853]. cycle [1763, 1359]. cycles [1536, 1522, 390, 1342, 1323]. cyclic [1737, 1339]. cyclically [772]. cycling [829]. cylinder [630, 708, 133, 138, 613, 457, 802, 1374, 607, 1135, 939, 1207]. cylinders [1335, 1668, 1816, 707, 602, 587]. Cylindrical [449, 6, 15, 734, 820, 733, 643, 923, 1732, 822, 1614, 636, 826, 1234, 214, 608, 743]. cylindrically [1008]. Czochralski [187, 80, 207, 188, 189].

D [1528, 1000, 1034, 932, 1304, 1741, 1225, 1522, 1003, 1240, 1511, 1645, 1391, 1155, 1457, 1475, 542, 1074, 1042, 1006, 858, 1285, 994, 880, 1388, 1036, 1440, 1745, 1411, 1037, 1022, 505, 1570]. d’Alembert [1780, 1201]. dam [344]. damage [861]. damaged [1129]. damp [468, 435, 455]. Damped [1648, 1567, 1723]. Damping [538, 569, 1438, 1210, 842, 903, 386, 1045]. Darboux [1779]. Darcy [365]. Data [1858, 1688, 461, 945, 1463, 332, 774, 1689, 1070, 1111, 1436, 1025, 1295, 443, 1686]. Davey [817, 1063, 1311, 925]. DC [1270]. Dead [1347]. Dead-time [1347]. Dean [241]. debond [57]. debt [611]. decagonal [1279]. Decay [1155, 628, 440, 1117, 824, 858, 153, 897, 24, 1070, 1043, 1141, 973, 562, 1704]. deck [149]. decomposition [1423, 1386, 555, 1545]. decompositions [561]. decontamination [1812]. deconvolution [1405, 1076]. decreasing [1456]. deep [1828, 1822, 1668, 78, 1267]. defects [1471, 1517, 592, 434, 1738, 640]. defined [515]. defining [90]. definition [1789]. defocusing [1063]. deformable [1328, 1818]. Deformation [1262, 1471, 1517, 599, 716, 1338, 1098, 1840, 594, 720, 1202, 454, 1343, 1334]. Deformations [351, 487, 345, 821, 869, 335, 452, 488, 580, 667, 847, 162]. deformed [1106, 513, 107, 608, 1341]. Deforming [1512]. Degenerate [1419, 1658, 949, 575]. degenerating [1790, 1784]. degrading [1333]. degree [1872]. delamination [1129]. delay [866, 1781, 860, 1352, 33, 1094, 551, 1537, 843, 1635, 873, 1056, 1112, 1381, 1865, 829, 1527, 783, 1323, 1069, 1175, 1441]. delay-dependent [1069]. delay-integro-differential [1069]. Delayed [1752, 747, 1810, 1674, 1003, 14, 33, 1369, 1417, 1804, 1677, 1035, 1212, 632, 1288, 829, 1312, 1251, 1294, 1406, 1057]. delays [1058, 1358, 1421, 1514, 1422]. delta [685, 1309]. dendritic [301]. denoising [1648]. dense [1809]. denseness [233]. densities [452, 488, 1361]. density [1318, 205, 1836, 1715, 391, 520, 1165]. density-suppressed [1836, 1715]. dental [1830]. dependence [886, 4]. dependencies [481]. dependent [540, 1478, 416, 1176, 1703, 1354, 69, 657, 787, 1699, 1278, 1298, 1646, 967, 1859, 1256, 1724, 1343, 1069, 1494]. depletion [355]. deployment [852]. deposit [1579]. deposits [745]. depth [363, 1572, 906, 348, 807, 557]. Derivation [1573, 938, 606]. derivative [139, 1276, 1538, 1719, 1560, 793, 1566]. derivatives [165, 1036]. derive [1818]. described [1771]. describes [755]. Describing [119, 1477, 418]. description [1218]. Design [170, 199, 290, 1439]. desingularization [1748]. destabilization [1200]. destabilizing [105]. destructive [1553]. Detachment [1681]. Detection [1265, 1853, 1700, 388]. Determination [305, 1356, 1002, 474, 1389, 288, 545, 1023, 475, 541, 645]. determining [332, 1630, 655]. developed [312]. developing [312]. development [73, 1617, 1111, 1529, 990, 1209]. developmental [469]. device [143, 203, 253, 327]. Devices [16, 220, 400]. Diagnostics [1616]. diagonalizable [1233]. diamond [1512]. Dielectric [1711, 759, 713, 1342, 963, 566, 1555]. difference [797, 964, 695, 1259, 244, 173, 1535]. differences [526]. different [1218, 1665, 1610]. differentiability [663]. differential [559, 1169, 257, 274, 1781, 1802, 1087, 1446, 785, 1062, 764, 1549, 99, 463, 301, 378, 1147, 581, 439, 1348, 727, 880, 711, 1496, 991, 1237, 1100, 290, 755, 123, 1429, 1153, 865, 70, 1394, 100, 238, 1633, 1527, 783, 1069, 1434]. differential-algebraic [378]. differential-delay [1527]. differentiation [602]. differing [761, 1624]. Diffraction [1481, 1839, 964, 194, 377, 1218, 708, 215, 1378, 31, 1315, 321, 1671, 1867, 999, 158, 334, 382, 687, 1855]. diffuse [1687, 1387]. diffuse-interface [1387]. diffuser [1488]. Diffusion [1252, 495, 409, 231, 280, 909, 381, 637, 992, 1733, 1227, 1459, 1619, 1790, 473, 1157, 278, 328, 1784, 1655, 1643, 1516, 893, 676, 593, 739, 1298, 528, 665, 228, 715, 1721, 954, 1094, 1495, 221, 256, 67, 98, 82, 272, 281, 1479, 1743, 8, 417, 1006, 689, 824, 839, 306, 1369, 1417, 1804, 555, 1075, 1156, 1258, 1296, 352, 591, 242, 546, 1734, 97, 295, 51, 604, 544, 684, 890, 1245, 1212, 1192, 552, 313, 338, 441, 532, 471, 496, 547, 896, 1630, 973, 526, 660, 1333, 1611, 800, 414, 561, 477, 562, 1013, 1082, 141, 40, 298, 568]. diffusion [915, 273, 432, 773, 677, 1236, 578, 1679, 1292, 1253, 998, 709, 1294, 1515, 1406, 383, 459, 1523, 456, 1433, 1862, 533, 670]. Diffusion-controlled [495, 409, 992]. diffusion-encoding [1479]. diffusion-reaction [272]. diffusions [1393]. diffusiophoresis [1834]. diffusive [1304, 1366, 1226, 863, 1537, 183, 1058, 1239, 293, 1256, 632, 1426, 358, 1016, 1610, 1149, 1455, 1480]. diffusive-kinetic [183]. diffusivities [893, 1624]. diffusivity [435]. Digital [1287, 1148]. dilations [777]. dilute [74, 185, 1710]. dimension [778, 253, 683, 1524, 1455]. dimensional [196, 304, 393, 798, 176, 73, 1802, 66, 1412, 458, 347, 1030, 151, 700, 1572, 675, 1808, 394, 498, 1075, 282, 596, 650, 316, 83, 454, 604, 653, 742, 765, 129, 220, 1641, 1081, 694, 370, 1090, 723, 462, 1095, 845, 1714, 719, 556, 1114, 2, 814, 27, 1266, 1267, 709, 1026]. dimensionality [1822, 356]. dimensions [672, 50, 79, 1332, 1241, 136]. dip [1442]. dipole [615, 1110]. Dirac [777, 1799]. Direct [1425, 1173, 1273, 529, 663, 1401, 689, 1687, 1614, 644]. director [1634]. Dirichlet [1061, 1739, 65, 1404, 1285, 698, 1865, 1419, 1013, 752, 1385, 383]. Dirichlet-to-Neumann [1385]. disappearance [156]. disassembly [1359]. disc [968, 375]. Discharge [381, 260, 355]. Discontinuities [6, 15, 916, 1295]. discontinuity [1519, 43]. Discontinuous [744, 1536, 1642, 537, 1111, 1025, 865].

Discrete

[1595, 703, 777, 1121, 860, 1721, 332, 351, 1174, 1554, 808, 999, 300, 1854, 1723]. discrete-map [1721]. discrete-time [1174, 808]. discretization [91]. discs [405]. disease [1344, 1795, 1175]. disintegration [1519]. disks [1741, 602]. disordered [1222]. dispersal [1866]. Dispersion [303, 1357, 1377, 534, 1396, 1859, 674, 1018, 1316, 13]. dispersive [1727, 1308, 1669]. dispersive-dissipative [1308]. displacement [332, 1086, 1492, 247]. displacements [214]. dissipation [1654, 1623, 1076, 1141, 356]. dissipative [833, 1805, 1778, 1235, 1308]. dissociative [546]. dissolution [1484, 1624, 995]. distance [323]. distances [1395]. distancing [1604]. distinguished [1652]. distributed [866, 1598, 1781, 1352, 1591, 843, 1358, 1660, 455, 1393]. distributed-order [1393]. Distribution [188, 1118, 1442, 495, 1514, 1110, 1343, 1499, 1862]. distributional [167]. distributions [379, 210, 1355, 1531, 750, 232, 1524]. distributive [998]. disturbances [1411, 609, 646]. divergent [934]. Dixit [1376]. DNLS [1202]. domain [1061, 38, 663, 1789, 1721, 1614, 1545, 793, 682]. domains [1229, 1638, 940, 1443, 1858, 1721, 1842, 1052, 1597, 1540, 743, 1393]. dominance [152, 1737]. dominated [400]. Dopant [280, 231]. Dopants [188]. doped [1202]. dose [1830]. dots [1083]. double [1863, 1300, 1574, 1495, 239, 1502, 462, 1611]. double-layered [1574]. double-stator [1300]. doubly [1446, 1816, 1477, 1614, 781]. doubly-periodic [1816, 1477]. down [834, 1207]. downstream [443]. DP [1397]. drainage [1790, 1784, 1166]. draining [735, 834]. Drift [400, 555, 1192]. drift-diffusion [1192]. Drift-dominated [400]. drilling [1846]. driven [979, 810, 980, 1582, 1291, 1477, 1767, 1834, 1342, 1116, 81, 1129, 1692, 1873]. Drop [1561, 1184]. droplet [1673, 1185]. droplets [1847, 1693, 1165]. drops [1012, 912]. drug [1250, 1624]. dry [1180]. drying [1834]. dryland [1758]. dryout [1087]. Dual [948, 70, 796, 907, 1716]. Duality [642, 498, 1310]. duct [643, 276, 1725]. Ducts [6, 15, 944]. due [135, 597, 77, 549, 1623, 1734, 1264, 543, 1333, 1608, 1486]. Duffing [227, 307]. Dupuit [344]. duration [754]. During [103, 188, 1564, 1083, 185, 727, 880, 1339, 1682]. Dynamic [818, 1344, 1164, 1554, 1872, 861, 1604, 912, 1650, 1770, 1601, 394, 1425, 1439, 94, 1773]. dynamic-type [1425]. Dynamical [258, 1746, 1282, 1305, 1322, 681, 589, 1243, 1542, 610, 1698, 12, 1310, 1253, 1548]. Dynamics [1420, 1257, 1328, 1004, 795, 1794, 1321, 946, 1588, 333, 1643, 1474, 657, 1187, 1250, 860, 1749, 1620, 1866, 839, 932, 1220, 1430, 1662, 756, 1050, 370, 1812, 1077, 1133, 1127, 1424, 1540, 1186, 1458, 1632, 1717, 1795, 1175, 1406, 1525, 1787].

each [94]. Eagle [278]. early [1812]. early-time [1812]. earth [1399]. Ebola [1848]. eclipse [754]. economic [1244]. eddy [153]. eddy-currents [153]. Edge [183, 182, 1799, 369]. edges [1799]. edited [23]. Editorial [1217, 1726, 1158, 1821, 1260, 952, 815, 1223]. EEG [1303]. Effect [1628, 572, 1119, 187, 73, 804, 528, 1617, 1749, 1198, 1537, 352, 1325, 1676, 1220, 686, 582, 600, 13, 1836, 578, 1150, 1293, 1868]. Effective [937, 188, 759, 1811, 1793, 512]. Effects [1226, 1508, 1507, 910, 1513, 1819, 974, 588, 923, 1735, 472, 521, 824, 183, 1576, 1764, 614, 846, 1347, 1428, 556, 911, 405, 661, 1860, 1266, 1550, 191, 209, 1539, 391]. Efficient [545, 598, 1361, 772, 1382]. Eigen [841]. Eigen-oscillations [841]. Eigenanalysis [1281]. eigenfunction [944, 299]. eigenfunctions [668, 100]. eigenmodes [1640]. Eigensolutions [216].

Eigenvalue [14, 1275, 1404, 1541, 402, 1041]. Eigenvalues [136, 841, 1413, 1524]. eigenvectors [695]. eight [390]. eighth [1362]. eighth-order [1362]. Einstein [1609]. elasmoid [1337]. Elastic [1157, 234, 1490, 736, 155, 487, 820, 1275, 150, 1262, 958, 732, 1414, 1837, 1864, 1569, 639, 1097, 1072, 609, 646, 1513, 1699, 1574, 603, 816, 1278, 345, 1053, 821, 869, 733, 1079, 818, 1840, 332, 1086, 594, 1375, 351, 467, 267, 563, 335, 513, 774, 822, 955, 538, 569, 825, 753, 913, 879, 1762, 454, 595, 922, 1612, 841, 750, 1357, 87, 434, 943, 616, 600, 635, 748, 937, 723, 205, 786, 1333, 1714, 1652, 505, 1102, 802, 580, 1331, 1313, 1640, 1492, 608, 1055, 1341, 1231, 687, 605, 1870, 903, 1625, 1444, 1657, 1695, 1761]. elastic [1675]. elastic-plastic [1762]. elastic-walled [1444]. elastica [1581]. elasticity [1532, 1030, 1770, 741, 96, 449, 1336, 698, 835, 962, 418, 768, 828, 719, 1526, 76, 1279]. elasto [594, 621, 1419]. elasto-plastic [594, 621, 1419]. elastodynamic [1769, 343]. elastodynamics [157, 996]. elastohydrodynamical [1582]. elastomeric [613]. elastoplastic [1002]. elastoplasticity [1330]. elastostatic [938]. elastostatics [25]. electric [1760, 1012, 629, 1270, 614, 686, 846, 972, 1555]. Electrical [883, 1691]. electrically [765]. electrified [531, 1207]. electro [260]. electro-discharge [260]. electroactive [1106]. Electrochemical [199, 260, 499]. electrochemistry [715, 254]. electrode [317]. electrodeposition [910]. electrodes [629]. Electrodiffusion [656]. Electrodischarge [103]. Electroelastic [1106]. electroencephalography [1564]. Electroforming [10]. electrohydraulic [1044]. electrohydrodynamics [461, 504, 1270]. Electrohydrostatic [419]. Electrolysis [479]. electrolyte [941]. electromachining [170]. Electromagnetic [585, 690, 809, 1541, 1315, 607, 692, 833, 797, 708, 1299, 1378, 277, 524, 1395, 1109, 226, 388, 1614, 638, 1652, 1686, 1047]. electromagnetism [696]. electropaint [130]. electrophoretic [453]. electrophysiology [1431]. electroplating [470]. electrorheological [1004]. electrostatic [204, 323, 1430, 872, 654, 264, 1113]. electrostatics [537]. element [1432, 1017, 1344, 1360, 1827, 384, 835, 1661, 1345]. element-Landweber [835]. elementary [1503, 64]. elements [1823, 235]. elevated [21]. elimination [105]. ellipsoid [652, 1399]. ellipsoids [1747, 1587]. Elliptic [249, 831, 1063, 514, 1002, 1701, 697, 931, 1658, 291, 1593, 1385]. elliptic-hyperbolic [1701]. elliptical [630]. elongated [1586]. embedded [259, 31, 802, 713, 1492]. emerge [1424, 1540]. enclosure [181]. encoding [1479]. encountered [1218]. end [77, 824, 76, 1870]. end-strip [76]. endothermic [1179, 789, 803]. endpoint [1563]. ends [658, 1444]. energies [1115, 1228, 1491]. Energy [1076, 1043, 1141, 523, 16, 525, 1244, 681, 1435, 1072, 463, 1579, 931, 750, 1081, 205, 452, 488, 1342, 920, 1046]. energy-conserving [463]. energy-deposit [1579]. energy-economic [1244]. energy-harvesting [1342]. Energy-minimizing [525]. energy-preserving [920]. engineering [1177]. Enhanced [1586]. enhancement [839]. Enhancing [1182, 1021]. ensemble [1777, 1717]. ensembles [491]. entanglement [1782]. entering [133, 169, 840]. enthalpy [1047]. Entropy [1367, 703, 969]. Entry [284]. envelopes [308]. environment [690, 1516, 1016]. enzyme [1173]. enzyme-catalyzed [1173]. epidemic [1503, 1604, 1402, 1456, 1765, 1662, 1539, 1525, 1480]. Epidemics [1418]. epidemiological [1458]. epitaxial [1387]. epithelial [1683, 389]. epithelial-mesenchymal [1683]. Epstein [749]. equal [1635]. Equation [1779, 1752, 799, 1304, 364, 1744, 1815, 797, 22, 480, 473, 3, 178, 1024, 965, 448, 1636, 1730, 1739, 1797, 503, 268, 227, 204, 708, 208, 1051, 1802, 406, 882, 1280, 1519, 1565, 331, 940, 1443, 1087, 620, 1196, 1769, 337, 918, 1030, 108, 243, 1062, 917, 1197, 36, 782, 33, 567, 1775, 1549, 1504, 319, 507, 67, 98, 272, 281, 612, 1871, 301, 702, 1489, 1743, 1289, 1164, 689, 1720, 936, 1804, 843, 662, 1269, 1777, 788, 884, 1314, 1396, 959, 1425, 697, 1248, 994, 691, 1202, 97, 295, 1111, 1290, 1529, 1152, 1004, 1214, 1560]. equation [475, 427, 1597, 1594, 396, 554, 1573, 71, 229, 1745, 1373, 1841, 949, 329, 1806, 75, 246, 1502, 462, 371, 1611, 251, 250, 1615, 1368, 1018, 1316, 562, 1286, 755, 40, 568, 933, 1575, 123, 634, 283, 814, 399, 1055, 175, 1236, 1613, 558, 1371, 1527, 386, 783, 519, 1857, 885, 356, 359, 1026, 1523, 1845, 1535, 292]. equations [1863, 703, 559, 624, 1169, 416, 257, 1606, 964, 1503, 102, 197, 484, 196, 888, 1121, 274, 1528, 1273, 1154, 1000, 278, 1247, 328, 489, 537, 1748, 1781, 901, 1438, 1412, 787, 461, 1567, 385, 1446, 79, 1770, 320, 1138, 68, 1391, 764, 1646, 114, 1729, 1457, 1549, 221, 906, 1533, 514, 99, 675, 463, 908, 1170, 378, 1088, 1506, 240, 1147, 8, 360, 1701, 1191, 581, 1799, 790, 1696, 1075, 1156, 439, 277, 1348, 1199, 354, 618, 95, 1210, 1282, 1750, 51, 1690, 986, 244, 1755, 203, 134, 975, 313, 1573, 711, 1496, 1658, 1630, 991, 1778, 146]. equations [976, 518, 660, 1333, 1547, 253, 263, 307, 290, 414, 921, 436, 421, 575, 1013, 141, 661, 933, 977, 1575, 1153, 693, 300, 865, 773, 677, 70, 314, 100, 920, 238, 939, 1064, 1046, 1292, 173, 1397, 1069, 1709, 1692, 1434]. equatorial [1219, 1669]. equidiffusional [483]. equilibria [1592, 1044, 1808, 1678, 368, 1090, 424]. Equilibrium [411, 38, 419, 1759, 583, 395, 1081, 1376]. equipartition [59]. equivalence [900]. equivalent [1490]. ergodic [1419]. Errata [128]. Erratum [1659, 1517, 1365, 1667]. error [1828, 1360, 694]. errors [1564, 1380, 1690]. Essential [1487]. estimate [663, 1360, 832]. estimates [681, 1391, 841, 1778, 265, 752, 1704]. Estimating [665, 1800, 190]. estimation [657, 945, 1243, 1080, 694, 1548]. estuaries [466]. Euler [1779, 1281, 1701, 1210, 651, 1811, 1448, 842]. European [1400, 1535]. eutectic [1809]. eutectoid [509]. Evaluation [1140, 30, 1249]. evanescent [804]. evaporating [1328]. Evaporation [120]. evasion [1820]. even [1434]. even-order [1434]. eversion [594]. evidence [1009]. Evolution [1646, 115, 1856, 1571, 1796, 1121, 503, 1247, 1131, 764, 848, 114, 502, 1375, 67, 58, 1083, 360, 1531, 439, 1282, 1070, 985, 1841, 1778, 436, 1343, 1851]. evolving [1463]. Ewald [1249]. Exact [268, 1477, 1835, 1301, 660, 1873, 36, 576, 594, 1664, 335, 417, 513, 1396, 583, 697, 1844, 1277, 141, 805, 960, 939]. exactly [444]. Examination [1380]. exchange [883, 643, 374]. exchanges [189]. excitation [1024, 1188, 529, 62, 369, 1029, 950, 1773]. excited [1289, 161, 1419]. exhibiting [438, 1291]. Existence [416, 762, 965, 325, 252, 341, 1650, 1758, 1239, 1496, 991, 469, 75, 290, 237, 963, 1089, 499, 1603, 1459, 278, 350, 831, 882, 942, 1280, 886, 1086, 408, 221, 1170, 394, 83, 1820, 1493, 786, 925, 1836, 693, 1376, 1046, 275, 533]. exit [1355, 1398, 284]. exocytotic [1314]. exothermic [705, 803]. exothermic-endothermic [803]. expanding [622]. Expansion [1241, 176, 323, 788, 239, 608]. expansions [1802, 944, 535, 749, 431, 299, 442, 330, 17, 89, 145]. expectations [1410]. Experimental [1009, 307, 945, 161, 855, 250]. experiments [302, 1102]. Explicit [1320, 306, 913, 1722, 550, 933, 1631, 1862, 197, 746, 895, 778, 1841, 223]. exploited [1251]. exploration [200]. explosion [747, 362, 156]. explosions [180, 219]. explosive [1550]. explosives [1178, 993]. exponent [678]. Exponential [1567, 1410, 1264, 584, 845, 1287, 1348, 1070, 1358, 873, 935, 1833, 842]. exponentially [72, 497, 224]. exponentially-stratified [72]. expression [895]. Extended [1383, 622, 460, 554, 1047]. extensible [1599, 161]. Extension [1279, 820, 1320, 818, 1789, 1201, 136]. extensional [958, 723]. Extensions [777, 1780]. extensive [969]. exterior [1061, 639, 1721]. external [1390, 860, 1035, 826, 713]. extinction [1410, 1056, 1112]. extraction [1216, 311]. extreme [1195, 1108, 1150]. Extremely [280]. extremum [642]. extrusion [1645]. exudation [1530].

F [278]. fabrication [1757]. factorizable [127]. Factorization [1557, 118, 527, 704, 734, 1460, 1566, 1718, 1761, 1756]. failure [1754, 1170]. Faithful [854]. falling [1697, 1189, 1205]. families [513]. family [497]. Far [1587, 736, 1855, 1218, 930, 233, 1839, 1039, 1656, 1686]. Far-field [1587, 736, 1855, 930, 1839, 1039, 1656, 1686]. Faraday [1860]. Fast [710, 1655, 1685, 577, 1080]. FE [986, 1208]. feasibility [1824]. features [813]. fed [844, 1381]. fed-batch [1381]. feedback [14, 1504, 947, 1677, 698, 1640, 1392, 1649, 751]. ferrite [1377]. ferrite-coupled [1377]. ferrofluid [1394]. ferromagnetic [1190]. fiber [1170]. fibers [1760]. fibre [289, 1202, 1497, 9, 1343]. fibre-optic [289]. fibre-reinforced [9]. Fibres [1365, 486, 1338, 733, 342, 539, 601, 436, 1492, 1346]. Fickett [395]. Fickett-Majda [395]. Fickian [1379]. field [1663, 1760, 736, 1855, 1869, 710, 1503, 731, 1852, 1390, 1188, 419, 255, 333, 176, 187, 1522, 1283, 332, 951, 423, 450, 1728, 930, 1005, 233, 1678, 1839, 1039, 765, 1656, 134, 1587, 226, 614, 686, 872, 543, 1485, 654, 1277, 1492, 478, 1351, 1686, 1773]. fields [1218, 1655, 204, 1012, 1299, 597, 1743, 1395, 1270, 823, 846, 1818, 1766, 972, 264, 1394, 1555]. filament [1757, 1581, 1582]. filament-bundle [1581]. filaments [1170]. Filippov [1476]. Filippov-type [1476]. filled [603, 1086, 1442, 1742, 222]. filling [315]. film [1606, 979, 924, 1372, 629, 1166, 1693, 375, 709, 1207]. film/substrate [966]. films [1379, 1083, 1205, 1834, 1268, 1544, 447]. filter [1676, 1717]. filtered [947, 1419]. filtering [1548]. filters [839, 564]. filtration [567, 1371]. FIMA [1578]. final [132, 1023, 1386]. finding [576]. fine [909]. finger [814]. fingering [565, 570]. fingers [372]. Finite [177, 59, 588, 370, 9, 494, 1345, 162, 1334, 1535, 54, 363, 1432, 1017, 927, 1098, 695, 848, 1780, 594, 1344, 1360, 1827, 1479, 335, 449, 241, 1291, 276, 598, 392, 962, 235, 418, 1216, 1102, 1337, 667, 1526, 977, 807, 1268, 1267, 1709, 1499]. finite-amplitude [977]. Finite-dimensional [370]. finite-element [1017]. Finite-length [588]. finite-product [1216]. finite-strain [1344, 962]. finite-time [1291, 1268]. finitely [1106, 513]. fire [1257]. first [624, 257, 576, 502]. first-order [624, 257, 576]. fish [1337]. Fisher [1070, 985, 1611, 575, 100]. Fisher-KPP [575]. FitzHugh [1619]. fixed [1599, 1475, 680, 855, 1798, 1686, 1382]. flame [946, 810, 483, 982, 789, 1717]. flames [810, 803]. flap [123]. flat [641, 1626, 35, 807]. flatness [44]. flattening [1674]. flaw [226]. flexible [1331]. flexural [1572, 1267]. flexural-gravity [1572, 1267]. flight [784]. floe [1040]. floes [1078]. flotation [1790, 1642, 1784]. Flow [6, 15, 718, 381, 482, 1232, 216, 1648, 1688, 304, 630, 128, 110, 230, 979, 1263, 1507, 1091, 1015, 1262, 69, 1390, 622, 492, 590, 80, 1048, 980, 347, 206, 1011, 929, 1840, 1242, 1442, 712, 1617, 1776, 597, 1664, 767, 1423, 77, 1743, 1189, 1488, 144, 259, 538, 569, 306, 948, 1291, 1623, 1713, 1140, 1325, 1372, 1816, 577, 316, 454, 1036, 792, 1596, 1264, 1203, 1009, 387, 1284, 370, 855, 232, 457, 1639, 446, 1037, 299, 462, 18, 1387, 641, 400, 1428, 1349, 1374, 1363, 911, 1181, 1159, 1772, 722, 284, 358, 443, 915, 1297, 76]. flow [1219, 124, 373, 1608, 374, 142, 1116, 1486, 1394, 1008, 1134, 587, 939, 464, 1593, 1307, 1207, 520, 570, 1873]. flow-exact [1664]. Flows [1114, 735, 1227, 1606, 312, 989, 500, 1586, 560, 1012, 1668, 1580, 385, 1534, 984, 248, 643, 1501, 1705, 1476, 1585, 357, 1665, 254, 1871, 117, 1666, 1167, 149, 1200, 1160, 1832, 352, 591, 770, 395, 1440, 425, 427, 1641, 235, 990, 1818, 122, 1168, 1722, 1844, 1570, 45, 876, 7, 1161, 399, 35, 81, 1692, 591]. fluctuation [1053]. fluctuations [1222]. fluid [1606, 1321, 1628, 849, 1273, 1390, 1276, 603, 72, 62, 929, 133, 1086, 1442, 617, 597, 563, 1006, 215, 306, 1269, 1200, 1771, 24, 595, 535, 745, 1820, 795, 1477, 387, 1209, 972, 1333, 641, 1428, 586, 911, 222, 1317, 315, 977, 297, 373, 1221, 903, 1717, 640, 834, 1177]. fluid-filled [1086, 222]. fluid-loaded [62]. fluid-lubricated [1321, 1628]. fluid-solid [586, 640]. fluidized [318]. fluids [363, 1213, 1684, 531, 1242, 643, 1713, 1140, 511, 235, 446, 1204, 1466, 92, 578, 1555]. Flux [112, 1577, 313, 338, 857, 1642, 1513, 1374]. Flux-ratio [313]. fluxes [606, 1868]. focal [900]. Focusing [1063, 479, 1745, 1502, 290]. Fokas [1775, 1720]. Fokas-Lenells [1775]. Fokker [432, 399, 1692]. fold [1670]. fold-type [1670]. folding [1131]. foliated [1734]. following [1377, 1677, 730]. food [1550]. foods [673]. force [128, 110, 860, 1219]. Forced [72, 108, 438, 490, 1698, 543, 1722, 1723, 1382]. forces [1281, 246, 1544]. forcing [1628, 1078, 227]. forest [1253]. Foreword [1329, 850, 1324, 953, 978]. form [20, 1513, 1196, 1475, 1140, 1108, 418, 496, 969, 836, 175]. formation [1663, 1561, 924, 1674, 1151, 1670, 1707, 987, 186, 572, 506, 521, 858, 1554, 601, 707, 1185, 1767, 896, 1562, 828, 784, 1184, 1715, 223, 1020, 570, 1433]. forming [1019, 745]. forms [683]. formula [1538, 1201, 1120]. formulae [197, 1780, 1241]. formulas [1386, 1844]. formulation [178, 1712, 721, 1330, 1510, 1622, 505, 432]. formulations [915]. Forward [1821, 1788, 1802, 400]. forward-biased [400]. foundation [1770, 774]. Four [1251, 1520]. Fourier [147, 626, 262, 1595]. fourth [448, 1774, 503, 1755, 12]. fourth-order [1774, 503, 1755, 12]. fractal [1432, 1627, 1071, 1332, 430]. fraction [1596]. Fractional [1663, 1815, 1322, 1276, 626, 1719, 256, 1348, 1036, 262, 1332, 1595, 1806, 1018, 1316, 1429, 1393, 1523, 1535, 1060]. fractions [769]. fracture [729, 819, 1494]. fractured [745, 541]. framework [1733, 1829, 1849, 1630]. Franklin [1644]. Fréchet [663]. freckles [186]. Fredholm [1229]. Free [1115, 1263, 1015, 206, 1694, 988, 574, 641, 1455, 1232, 1215, 312, 989, 73, 1028, 1516, 485, 609, 1534, 594, 306, 294, 97, 296, 1264, 1835, 1209, 266, 631, 18, 1611, 39, 867, 693, 297, 237, 374, 81, 1137]. free-boundary [306, 237, 81]. Free-stream [1694]. Free-surface [1263, 1015, 206, 574, 312, 989, 294]. free-vortex [693]. Freedericksz [724]. freedom [1872]. freezing [74, 1206, 391]. frequencies [625, 717, 649, 723]. frequency [486, 1188, 517, 1699, 1858, 1457, 1370, 1038, 1584, 1700, 1399, 1652, 1686, 781]. fresh [309]. Friction [1128, 666, 959, 1448]. Friction-induced [1128]. frictional [1850, 1801, 1510, 823]. fringes [152]. front [812, 1464, 756, 686, 1558, 902, 1253]. fronts [866, 565, 1379, 985, 669, 1610, 783, 1480]. froth [1790, 1784]. FRS [109]. fuel [883, 941, 1183, 383, 459]. full [202]. full-wave [202]. fully [312, 565, 420]. function [165, 1468, 664, 166, 1014, 689, 994, 1265, 146, 1602, 969, 1551, 64, 914, 655]. functional [777, 257, 797, 790, 290, 763, 865, 1236, 1089]. functional-difference [797]. functional-differential [257, 290]. functionally [955, 1559]. functionals [240, 121]. Functions [451, 791, 1807, 165, 164, 1249, 857, 131, 1071, 139, 212, 940, 36, 1014, 971, 515, 808, 119, 1332, 1786, 324, 431, 1460, 1408, 900, 17, 89, 145]. Fundamental [781, 277, 95, 996, 1018, 125]. fungal [593]. furnace [1182, 1764]. Further [68]. fused [1757]. fusion [1314, 687]. future [1145].

G [278]. Galerkin [240, 34, 633]. gallery [1741]. Game [1604]. games [1437]. gamma [164, 1781]. Gap [1204, 1707, 276]. gaps [215]. Gardner [1026]. gas [1697, 200, 230, 1481, 474, 549, 1160, 1220, 633, 1161, 93, 137, 1162, 1165]. gas-kinetic [1162]. gas-solid [633]. gaseous [909]. gases [154]. gating [1813]. gauge [1807, 1786]. General [1366, 819, 1857, 1706, 699, 858, 498, 555, 1816, 684, 1295, 1640, 1679, 1125, 998, 1709, 1794, 1433, 1603, 533]. generalization [1152, 1373]. generalizations [689]. Generalized [677, 1032, 791, 1478, 1730, 131, 1099, 1071, 1276, 1519, 331, 801, 620, 337, 817, 1063, 1504, 1248, 1282, 1311, 749, 313, 769, 658, 1766, 1722, 660, 847, 1368, 983, 925, 401, 892, 440, 1862]. generalized-affine [1071]. generated [706, 671, 645]. generating [988, 64]. Generation [1411, 1500, 68, 840, 1367, 1013]. generator [1787]. generators [559, 1342]. Gennes [1808]. gently [86]. geochemical [565]. Geometric [910, 1637, 565, 1501, 1605].

Geometrical [1472, 194, 813, 413, 1345]. geometries [54, 705, 560, 1767]. Geometry [857, 1508, 1555]. Gierer [1678]. Giesekus [1561, 1771]. Gilbert [1565]. Ginzburg [1598, 1752, 371, 356]. Gipps [730]. given [1250, 28, 27]. glacial [1763, 357]. Glass [1590]. glasses [169]. Global [1402, 831, 1280, 871, 612, 542, 961, 1358, 1421, 1820, 1493, 47, 1295, 1836, 1312, 1679, 1795, 1175, 1045, 1744, 194, 1571, 1797, 882, 553, 1014, 1170, 1108, 758, 1233, 925, 1046, 783, 1089]. glycolysis [1433]. Gordon [1196, 1152, 1046]. Görtler [458, 58, 380]. gossip [1779]. governed [637, 385]. governing [1333]. Gower [1057]. Grad [583]. Grad-Shafranov-type [583]. grade [1276, 1140]. graded [955, 1559]. Gradient [1673, 702, 549, 1818, 1548]. Gradient-induced [1673]. gradients [645]. Graham [278]. Graham-Eagle [278]. grain [1531, 1738]. grain-size [1531]. granular [1187, 1654, 1554, 823]. Graph [1651, 1659]. grating [1873]. gratings [1378]. Graves [415]. gravitating [1090]. Gravity [979, 182, 1534, 1572, 439, 171, 972, 478, 1470, 81, 1267]. gravity-capillary [182]. Gravity-driven [979, 81]. Gray [278]. Grazing [1763, 1476, 659]. grazing-sliding [1476]. Great [829]. Green [1837, 940, 938, 994, 1120, 914]. Green’s [1249]. Greitzer [775]. grid [878, 887]. grooving [503]. Gross [882]. Grossberg [1035]. ground [1051, 1150]. groundwater [309]. group [559, 822, 805, 386]. groups [1591]. growing [252, 1103]. grown [187]. Growth [188, 1118, 924, 1792, 647, 1366, 1104, 1759, 1809, 1719, 301, 1083, 564, 183, 1477, 1835, 207, 1387, 829, 1600]. Guided [627, 486, 515]. guides [627]. gust [1508].

habitats [1866]. Half [649, 377, 135, 732, 1819, 1574, 918, 1283, 1453, 1780, 1140, 938, 1152, 1560, 635, 802, 315, 1116, 605, 1657]. half-line [1152, 1560]. half-plane [377, 135, 918, 1283, 1116]. half-space [1453, 635, 802, 315]. half-spaces [732, 1819, 938, 605]. Ham [1140]. Hamiltonian [606]. Hammerstein [263]. Handbook [23]. hangers [1599]. Hankel [146]. hard [1850, 1801, 529, 412]. hardening [594, 394]. Harmonic [726, 710, 1653, 302, 545, 1299, 438, 490, 1378, 277, 264, 1238, 1334, 1625, 508, 1675, 1773, 1667]. harmonically [1289]. harmonics [1594, 1551]. harvesting [1112, 1342, 1251]. Hashin [512]. Hashin-Shtrikman [512]. Havelock [1470]. having [1668, 923, 1624]. health [1344]. hearing [682]. heart [1766]. Heat [84, 392, 39, 190, 359, 799, 1115, 1815, 1252, 886, 80, 1513, 1298, 726, 1182, 36, 408, 236, 1356, 445, 884, 967, 1314, 282, 1425, 83, 1690, 988, 636, 905, 1806, 1374, 1367, 830, 189, 283, 1608, 1613, 93, 137]. Heat-Conduction [84]. heat-generating [988]. heat-source [282]. heated [1276, 211, 617, 601, 707, 502]. Heated/cooled [502]. heater [1590]. heating [1261, 572, 539, 653, 742, 756, 679, 481, 435, 455, 743]. hedgehogs [1471, 1517]. hedging [1309]. Hele [333, 1705, 1031, 254, 1477, 372, 1134, 570]. helicity [1076]. Helmholtz [3, 1739, 1034, 918, 1743, 1027, 71, 229, 1615]. Helmholtz-type [1027]. Hénon [1051]. hepatitis [1094]. Hermite [20, 863]. Hermitian [1775]. Herschel [560, 465]. heteroclinic [1737]. heterogeneity [1569]. heterogeneous [1402, 1516, 1819, 248, 825, 1693, 552, 1016, 1165]. heuristic [827]. hexagonal [1647, 996]. Hidden [559, 463, 1436]. hierarchies [983]. hierarchy [1501, 1793]. High [486, 1457, 360, 242, 280, 1671, 273, 625, 1321, 1420, 759, 476, 717, 1188, 1370, 706, 614, 1652, 311, 45, 300, 435]. High-contrast [1671]. High-frequency [486, 1370, 1652]. High-low [1457]. High-order [360, 273, 759]. high-speed [1321, 1420, 706]. high-wavenumber [476]. high/low [300]. high/low-wave-number [300]. higher [545, 1030, 1391, 378, 936, 246, 1860, 571, 136, 1455, 1041]. higher-index [378]. higher-order [545, 1030, 936, 1860, 571]. highest [78]. highlights [1461]. highly [732, 461, 906, 823, 1399]. highway [730]. HII [115]. Hilbert [1427, 1054, 154, 1301]. Hilliard [703, 333, 243, 1729]. hindrance [974]. Hirota [1845]. History [1354, 1286]. History-dependent [1354]. HIV [1312, 1600, 1794]. HIV-1 [1312]. hodograph [268]. Hogan [1469]. Hohenberg [1730]. Hölder [1515]. holding [466]. hole [366, 437]. hole-pressure [366, 437]. Holes [84, 1774, 1598, 636]. Holling [1245, 1016]. hollow [1109]. Holm [1214, 1493]. holonomic [1202]. Homoclinic [1726, 1735]. homogeneous [1176, 690, 958, 1819, 332, 1409, 1027, 451, 841, 1841, 802]. Homogenization [1227, 19, 1000, 1484, 1660, 1787, 1692, 1053, 778, 1363, 291, 1231]. homotopy [1268]. honour [1469, 1578]. hood [840]. Hookean [847, 401]. Hopf [127, 234, 527, 704, 1519, 1752, 1495, 438, 790, 118, 1427, 1245, 1216]. horizontal [1574, 1629, 1776, 1374]. host [1186]. hostile [1721]. Hot [601]. hotspot [572]. Hpm [1140]. HSV [1133, 1077]. HSV-2 [1133, 1077]. Hubbell [749]. Hugoniot [857]. human [1344, 1793]. Huxley [1197]. Hybrid [1863, 1370, 1721, 997, 1542]. hydraulic [990]. Hydrodynamic [912, 117]. hydrodynamical [520]. hydrodynamics [1164, 125, 1001]. Hydromagnetic [929]. hydromechanics [397]. hydrostatic [1095]. Hyperasymptotic [431]. hyperbolic [771, 857, 1073, 942, 14, 801, 59, 576, 236, 1023, 1701, 897, 1070, 1750, 744, 1622, 985, 1206, 921, 1233, 1295, 1603]. hyperbolic-parabolic [1206]. hyperbolicity [1450]. hyperelastic [720, 1101]. hypergeometric [431]. hypergradients [1831]. hypersingular [691]. hypocoercivity [1692]. hypoexponential [1781]. hypotheses [694]. hypothesis [1635, 229]. hysteresis [1636, 325, 1511]. hysteresis-controlled [1511].

I-502 [1783]. I/mode [1494]. Ice

[201, 1856, 1040, 1078, 874, 296, 270, 1851]. ideal [387, 235, 297]. idealized [993]. identical [733]. identification [1577, 997, 604, 651, 1080, 1399]. Identifying [1796, 967]. identity [1085]. ignition [540, 705, 305, 32, 271]. II [339, 646, 869, 528, 96, 1296, 1656, 547, 455, 145, 1494, 459, 1862]. III [819, 1070, 1092, 1107, 1043]. ILDM [891]. ill [1830]. ill-posed [1830]. illicit [1250]. illustration [732]. IMA [219, 278, 1324, 450, 569, 437, 1133, 37, 1680, 1193, 1461]. Image [1626, 1110, 1648, 630, 1274]. imaging [1687, 1612, 644, 1825]. IMBq [1064]. imbricated [1337]. immersed [1344, 53]. immiscible [1273, 1428]. immune [1312]. immunity [1456]. immunoassay [1552]. Impact [1171, 1677, 871, 1840, 1472, 353, 1271, 1559, 875, 1848]. Impact-induced [1171, 871]. impacts [1327, 1128]. Impedance [725, 809, 895, 1743, 1315, 1556, 1656, 359]. impenetrable [1708, 1751, 793, 1695]. imperfect [914, 687]. imperfectly [1544]. impermeable [374]. Implementation [1365, 1811, 1346]. implication [373]. implicit [1322, 689, 1060]. Importance [1585]. improperly [384, 530]. Improved [906, 1138, 1007, 410, 1594, 1162]. improving [1182, 512]. Impulse [128, 110]. impulsive [843, 1358, 873, 1056, 1112, 1429, 1153, 1633]. impulsively [1623]. in-host [1186]. in-phase [1495]. in/PWM [1240]. incidence [659, 1614, 1175, 1525]. incident [294, 930, 478]. Incipient [1428]. inclination [1617]. inclined [72, 520, 1773]. including [1508, 1442, 1006, 577, 1573, 694, 911]. inclusion [1703, 1157, 605, 1625]. inclusions [1691, 1687]. incompressible [1340, 1338, 609, 646, 816, 345, 821, 869, 513, 1006, 613, 753, 879, 1612, 792, 387, 600, 723, 1570, 1313, 977, 1341, 1101]. incorporating [1750, 1349, 1113, 1445]. incremental [819]. indefinite [842]. indentation [1675]. indentations [1615]. Index [760, 1741, 378]. indices [1803]. indirect [1820]. induce [1555]. induced [1673, 1194, 871, 1171, 962, 1128, 828, 1172, 1317, 225, 217, 482]. induction [1436, 405]. inductor [1474]. industrial [1651, 1659]. inequalities [254, 218, 61]. inequality [956, 1288]. inertial [1269]. inertialess [837]. inevitability [1824]. inexact [1831]. Inextensible [107, 523]. infected [1662]. infection [1604, 1094, 1312, 1458, 1600, 1794]. infectious [1480]. Infeld [1237]. infiltration [213]. Infinite [66, 234, 735, 30, 681, 62, 495, 1804, 409, 1816, 276, 680, 1421, 1556, 1656, 246, 1615, 222, 557, 688]. Infinite-dimensional [66]. infinite-range [246]. infinitely [1668]. infinitesimal [523]. inflated [1105]. Inflation [155, 267, 449, 418]. Influence [989, 1390, 1591, 954, 1814, 298]. information [1553, 413]. informed [1827]. infused [1844]. Ingham [1215]. Inhibition [789, 1767, 902]. inhibitor [1137]. inhibitory [1550]. Inhomogeneities [877, 1097]. inhomogeneity [779, 1513, 821, 869, 508]. Inhomogeneous [444, 1041, 1768, 302, 1699, 567, 741, 1708, 1751, 1541, 1357, 694, 75, 436, 580, 667, 960, 179, 334, 175, 1371, 1718, 1803]. Initial [888, 1121, 97, 396, 1806, 701, 461, 1789, 780, 82, 863, 295, 1070, 1111, 1290, 1529, 547, 1209, 943, 1233, 1295, 1046, 655, 1709, 1815, 1111].

Initial-boundary

[396, 1806, 1789, 863, 1233, 1709, 1815].

Initial-boundary-value [1121]. Initial-layer [888]. initial-value [1070, 1111, 1290, 1529, 943]. initially [609, 1380, 1313]. initiation [993]. Initiative [1783]. injecting [574]. Injection [297, 947, 948, 1477, 1134, 482]. injection-driven [1477]. ink [1828]. inner [597]. innovation [1548]. inpainting [1274]. input [1386, 774, 475]. input-output [1386]. inputs [1035]. insight [1472]. insights [1868]. insoluble [65]. inspired [1337, 1546]. instabilities [213, 565, 185, 361, 570]. Instability [472, 551, 1558, 1697, 1760, 304, 1335, 987, 1599, 1734, 1245, 1119, 1172, 1714, 632, 380, 1295, 1608, 1544, 184, 391, 1704, 1616]. Instantaneous [87, 746, 1633]. insulated [403]. insulation [578]. insulators [1785]. Integrability [1755, 331, 1543]. Integrable [1720, 503, 335, 1152, 1026]. Integral [165, 204, 84, 320, 277, 1055, 190, 791, 964, 30, 102, 178, 484, 196, 810, 891, 489, 537, 708, 901, 787, 1487, 1509, 347, 1769, 811, 240, 1348, 354, 95, 691, 749, 239, 146, 462, 505, 263, 421, 175, 920, 310]. Integral-equation [204]. integrals [30, 165, 164, 934, 157, 442]. integrated [627]. integrating [581]. integro [1087, 1259, 123, 1394, 1069]. integro-difference [1259]. integro-differential [1087, 123, 1394]. integrodifferential [251, 314]. intensity [1859]. intensity-dependent [1859]. Interacting [1520, 15, 439]. Interaction [1495, 1733, 1869, 1508, 1852, 1154, 936, 501, 185, 535, 1820, 1491, 554, 1241, 776, 1209, 371, 1135, 1308, 783, 1293, 1706]. interactions [500, 346, 1205, 361, 966, 300, 1669]. Interactive [1130]. interatomic [755]. intercellular [1230]. interface [1861, 187, 309, 1513, 816, 919, 1383, 1430, 1387, 92, 1331, 914, 640, 1494]. interfaces [567, 1705, 1742, 1649, 1371]. Interfacial [1228, 1809, 570, 1832, 1860]. Interference [151, 1426, 1550]. interior [599, 1299, 1541]. intermediate [183, 1318]. internal [363, 1198, 1582, 1594, 1672, 903, 1137, 1873]. internally [1732]. interpolation [878, 1071, 332, 863, 1285, 887]. Interpretable [1825]. interval [735, 1815, 132, 848, 1780, 1804, 1806, 1709]. intervals [1646]. intervisibility [645]. introduced [91]. Introduction [1122, 1589, 1680, 1098, 1158]. invariance [90]. Invariant [798, 1450, 747, 1156]. invariants [1336]. invasion [671, 1415]. invasions [1407]. Inventory [101]. Inverse [1792, 1768, 1592, 731, 1263, 1802, 106, 302, 428, 529, 1699, 1858, 25, 1360, 1409, 542, 467, 741, 1386, 1074, 445, 1052, 336, 1425, 651, 1612, 971, 1716, 388, 644, 638, 1602, 2, 56, 358, 291, 1631, 1753, 343, 1566, 1515, 1695, 1761, 1686, 1756]. inverted [1423]. inverter [1066, 1347]. Investigation [946, 1550, 250, 805]. Investigations [303]. Inviscid [124, 1221, 603, 1242, 767, 294, 770, 232, 7]. involved [1258]. involving [178, 249, 1802, 573, 465, 497, 1296, 536]. ionic [614, 686]. ionized [154]. ions [656]. Irradiation [615]. irreducibility [1843]. irregular [150, 461, 1652]. irrotational [1585]. isoelectric [479]. isogeometric [1383]. isolated [1599]. isolation [588]. isoperimetric [254]. isospectral [429, 859, 904, 1191]. isostatic [1439]. isothermal [908, 806, 697, 441, 532, 471, 1367]. isotropic [821, 869, 1453, 332, 1330, 616, 205, 894, 889, 1341, 605]. isotropy [1453]. Issue [1329, 1142, 1324, 1821, 1122, 1469, 1680, 1098, 1578, 1177]. issues [1457]. Iterating [780]. iteration [1837]. iterative [148, 1364, 884, 1679].

J [219, 278, 450, 569, 437, 1133, 37]. Jacobi [64]. James [109]. jet [1065, 1189, 294, 353, 235, 1209, 574, 123, 478]. jet-flap [123]. jets [1697, 1561, 1607, 1509, 1185, 837, 52, 1184]. John [1578, 1469]. Johnson [1531, 1248]. joined [605]. Journal [1461]. jump [975]. jumps [1309]. Junctions [48]. Justification [694].

Kadomtsev [331]. Kärger [1479].

Karlotoxin [1814]. Kawasaki [1571]. KdV [1489, 1396]. Keller [932]. Kelly [926]. Kelvin [1048, 596, 650, 125]. keratin [1359]. Kermack [1418]. kernel [1797, 146]. kernels [527, 320, 1621]. Kerr [1727]. key [1279]. keyhole [738]. kick [546]. kick-out [546]. kidney [1682]. kind [1096, 1171]. kinematical [813]. kinematically [1319]. kinematics [631]. kinetic [183, 154, 1162]. kinetics [85, 425, 493]. King [978]. Kirchhoff [1438, 694]. Klein [1468, 1046]. Knops [953]. Knowles [1099]. known [288]. Kolmogorov [1531, 1417, 1089, 456, 1862]. Kolmogorov-type [1417, 1089]. Kono [1654]. Korteweg [1111, 1290, 1529, 1388, 396, 554, 933]. KP [1191, 1573]. KPP [575]. Krawiec [1323]. Krugman [1376]. Kruskal [689]. KS [1489]. KS-KdV [1489]. Kubo [1120]. Kuramoto [1504, 1777, 1368, 440]. Kuwabara [1654].

Ladyzhenskaya [956]. lagging [511]. Lagrangian [373]. lake [1748]. Lamb [879, 1762]. Lamb-type [879]. Lambert [1418, 969]. laminar [1442, 1665, 427, 1767, 45, 142, 1873]. laminated [49, 694]. Lanchester [460]. Landau [1598, 1565, 1752, 1808, 371, 386, 356]. Landweber [835]. Langevin [1602, 1787]. Laplace [940, 1443, 1747, 166, 917, 782, 702, 1594, 1854]. Laplacian [1028, 451, 1332, 1286, 136, 1723]. Large [1619, 1643, 1134, 54, 1405, 304, 227, 33, 1269, 1248, 295, 454, 1111, 1529, 1395, 1209, 488, 371, 307, 1165]. large-amplitude [307]. Large-time [1643, 1111, 1529, 1209]. laser [540, 947]. lasers [661]. lateral [856, 740, 1314, 1767, 902]. lateral-inhibition [1767]. Lattice [1166, 1168, 1159, 1165, 1003, 1412, 1647, 1163, 1164, 1369, 1417, 1167, 783]. lattices [1145, 1068, 1609, 1736]. Laurent [29]. law [178, 1213, 460, 621, 1464, 365, 1835, 446, 481, 1389]. laws [857, 813, 1577, 1642, 893, 801, 1050, 371, 1295, 606]. layer [540, 216, 759, 888, 500, 69, 1213, 1194, 1404, 1819, 385, 206, 1226, 1770, 502, 144, 948, 968, 1694, 1832, 1623, 618, 1700, 795, 975, 990, 446, 1204, 723, 462, 1119, 1172, 641, 1374, 556, 941, 836, 443, 376, 1486, 1444, 520, 617]. layered [1574, 1104, 1102, 1337, 1544, 1756, 1675]. layers [625, 714, 304, 458, 1079, 1629, 58, 116, 724, 564, 24, 1093, 988, 1021, 207, 972, 1428, 580, 7, 1624, 1870, 1527]. Leading [1580]. Leading-order [1580]. leaky [1555]. Learning [1826, 1821, 1817, 1828, 1831, 854, 1621, 1825]. least [931]. least-energy [931]. Ledermann [23]. left [1344]. Lenells [1775, 1720]. length [1863, 1627, 588, 143]. lens [1518]. lenses [290]. Leone [1848]. Leslie [1057]. levelling [447]. Liapunov [763]. Liberia [1848]. Lie [822, 893, 1156, 805, 386]. Liénard [208]. Lifshitz [1565]. lifting [1150]. Lighthill [109, 1680, 1324, 1504]. Lighthill-Thwaites [1324]. like [333, 1454, 871, 444, 1506, 1448]. likelihood [1777, 1717]. Limit [1536, 777, 1522, 1871, 390, 770, 1388, 1466, 356]. limited [1864, 1453, 1817, 355]. limited-angle [1817]. Limiting [1620]. limits [333]. Line [1771, 979, 138, 1780, 1377, 1427, 1152, 1560, 1264, 586, 181, 270, 525, 950, 1618, 43]. Linear [1078, 10, 1288, 1408, 520, 473, 1322, 1121, 924, 1864, 288, 1454, 798, 1474, 66, 882, 1280, 893, 886, 14, 1019, 1030, 860, 1770, 954, 848, 594, 863, 899, 778, 1044, 1002, 1023, 240, 824, 1269, 1075, 1156, 1200, 959, 1145, 1201, 524, 684, 1004, 698, 1755, 835, 1009, 338, 776, 996, 972, 1778, 1141, 75, 246, 1502, 1100, 1029, 1547, 1206, 1310, 911, 1082, 977, 634, 963, 865, 238, 1046, 1633, 1020, 1207, 1069, 1045, 1389, 1441, 1060, 1603, 1382]. linear-hardening [594]. linearities [1097, 1348]. linearity [1611]. linearization [100]. linearly [487, 1079, 955, 454, 1341, 1657]. lines [1712]. Lining [6, 15]. Liouville [1169, 1774, 1538, 1449]. Lipschitz [663]. Liquid [1435, 1697, 1561, 1607, 1188, 1471, 1517, 912, 1650, 1634, 592, 549, 1160, 724, 1166, 1579, 1440, 1185, 1811, 872, 1172, 1205, 1742, 837, 1844, 867, 634, 1268, 375, 1632, 709, 1165]. Liquid-bridge [1435]. liquid-gas [1165]. liquid-infused [1844]. liquids [1742]. living [656]. LMFBR [1087]. load [1342]. load-driven [1342]. loaded [62, 1401, 215, 595, 1625]. loading [647, 563, 1339, 827]. Lobatto [878, 887]. Local [668, 553, 997, 82, 1491, 771, 624, 1459, 1861, 1364, 1487, 1030, 1014, 1537, 1088, 1866, 959, 1258, 1296, 1004, 1755, 1407, 686, 1082, 998, 1294, 1149]. localised [877]. Localization [1732, 1459, 1731, 1564, 1377, 613]. Localized [1733, 1730, 1097, 1104, 1729, 1728, 1464, 1758, 1727, 1105, 1724, 1637, 1736, 827]. locally [918, 1030, 1403]. Locating [1687, 1413]. location [737, 1413]. locking [1451, 592]. logarithmic [320]. logarithmically [1138]. logistic [33, 843, 1452, 1600]. Long [182, 1273, 453, 590, 1210, 1750, 361, 911, 1082, 1459, 1154, 1775, 951, 1590, 1845]. Long-time [453, 1210, 1750, 1082, 1459, 1775, 951, 1845]. Long-wave [1273, 361, 911].

Long-wavelength [590]. Longitudinal [1596, 587, 1710, 298]. Loss [342, 1553]. losses [572, 289]. lossy [1377, 725]. Lotka [954, 1410, 1056, 1421]. Love [1819]. Low [1399, 517, 649, 1384, 1457, 1289, 1830]. low-dose [1830]. Low-frequency [1399, 517]. low-Reynolds-number [1384]. low-wave-number [300]. lower [312, 340, 240, 991, 121, 1374, 1266]. lower-branch [312]. LTI [808]. lubricated [1321, 1420, 1628, 1771]. lubrication [1847]. lumped [1682]. lumped-parameter [1682]. Lyapunov [1014, 808, 1429, 1153, 1633, 885].

Macdonald [166]. Machine [1821, 1127]. Machining [199, 103, 260, 499]. Maclaurin [1841]. macro [1178]. macro-scale [1178]. Macroscale [1424, 1569, 1540]. Macroscopic [1161, 1292, 1284]. made [658, 889]. magmon [862]. magnet [844, 1010, 1300]. Magnetic [188, 1390, 19, 1188, 187, 615, 1436, 765, 1587, 1552, 602]. magnetized [629]. magneto [719]. magneto-elasticity [719]. magnetoelastic [1306]. Magnetohydrodynamic [207, 583, 1367]. magnetohydrodynamics [1084, 1317]. magnetostatic [697]. magnetostatics [721]. magnets [721]. Majda [395]. maker [1470]. malaria [1352, 1406]. Mandelbrot [1332]. Mandelbrot-type [1332]. Manifold [1621, 810, 1026]. manifolds [891, 798, 1450, 747, 811]. many [1490]. map [1721]. mapping [96, 1008]. Mappings [1822, 573, 2]. maps [132, 859, 1068, 1382]. Marangoni [1428]. Marchi [131, 167]. Marchi-Zgrablich [131, 167]. marching [195]. Marginal [1608, 1730]. Marginally [7]. marijuana [1783]. marine [1139]. marketing [94]. markets [1309, 94]. Markov [1482, 1689, 769]. Markovian [1499]. marks [1670]. martensites [1319]. mask [588]. mass [1072, 408, 1067, 1798, 1868, 1723]. massive [1799]. Matched [1384, 442, 535, 299]. Matching [232, 385, 1800, 1021]. material [871, 1106, 821, 869, 465, 895, 594, 621, 961, 1383, 1436, 468, 511, 636, 1345, 847, 1331, 889, 829]. Materials [1365, 1340, 178, 1354, 762, 1299, 513, 720, 822, 1171, 1612, 823, 481, 435, 455, 401, 1346, 578, 287]. Math [219, 278, 450, 569, 437, 37]. Mathematical [1340, 737, 1810, 1139, 794, 80, 738, 1500, 115, 1148, 296, 1518, 103, 1505, 1552, 830, 941, 1186, 358, 189, 1422, 1439, 125, 261, 1688, 883, 627, 1144, 1683, 621, 509, 202, 539, 1630, 1766, 1392, 1359, 1848, 957]. Mathematics [1821, 23, 1133, 1142, 1469, 1463, 1147, 1143, 1461]. Mathieu [250]. matrice [1413]. matrices [938, 127]. matrix [234, 734, 1557, 1017, 1320, 1196, 1549, 323, 118, 1662, 1301, 996, 1460, 1288, 1640, 1631, 1845]. matrix-functions [1460]. matter [1187, 1654, 1554]. mattress [1596, 1710]. maturation [866]. maturity [998]. Maximal [1645, 845]. maximum [1777, 1318, 1717]. Maxwell [762, 1000, 929]. Maxwell-Norton [762]. McKean [1871]. McKendrick [1418]. Mean [6, 15, 1869, 1503, 1852, 1277]. mean-field [1503, 1277]. means [1131]. measurements [1039]. measures [392]. mechanical [1507, 325, 1103, 758, 389, 1326]. Mechanics [111, 1331, 1122, 1467, 24, 61]. mechanism [549, 546]. mechanisms [1128, 372, 374]. media [1227, 1249, 1232, 19, 339, 1653, 810, 1404, 248, 923, 1705, 747, 1506, 1676, 1484, 1202, 524, 1541, 598, 1036, 322, 1357, 552, 541, 996, 793, 1485, 179, 512, 1308, 519, 247, 1667, 482]. medicine [1144, 1848]. medium [812, 1490, 1176, 692, 1744, 1222, 302, 1520, 623, 1409, 825, 1027, 659, 988, 266, 404, 1559, 1235, 1317, 237, 558, 1718, 1803, 508]. meets [1418]. Mehl [1531]. Meinhardt [1678]. melting [727, 880, 1049, 140]. membrane [883, 818, 1325, 680, 748, 1105, 608, 334, 688, 781]. membrane-regulated [1325]. membranes [429, 859, 904, 563, 205, 452, 488]. memory [1115, 1744, 1726, 1438, 1132, 628, 1704, 1045, 1382]. memory-efficient [1382]. MEMS [1430]. menisci [1596, 1095]. meniscus [1811]. merging [519]. mesa [228]. mesenchymal [1683]. mesh [658]. mesoscale [1158, 1160]. metal [1188, 503, 226]. metamaterials [1299]. metastatic [1683]. Method [964, 84, 190, 625, 397, 734, 710, 30, 1249, 1017, 946, 268, 1712, 1837, 187, 1034, 1028, 1443, 1011, 764, 782, 848, 332, 576, 1344, 780, 747, 1827, 702, 1088, 1074, 796, 1164, 689, 384, 1654, 1687, 884, 967, 1027, 34, 336, 1136, 1853, 1396, 316, 1093, 633, 1597, 986, 244, 835, 610, 644, 638, 462, 660, 1216, 92, 78, 1429, 1153, 314, 1008, 1566, 173, 1718, 1761, 1756, 781, 1026, 1060, 1165, 1382]. method-of-lines [1712]. methodology [757]. methods [200, 1823, 1781, 1276, 1685, 1769, 1163, 581, 1211, 907, 536, 1690, 1865, 711, 1255, 1661, 1545, 263, 1460, 195, 162, 1055, 1413, 1069, 1854, 1535]. MHD [1090]. Micro [975, 1049, 1741, 1465, 1159]. micro- [1159]. micro-disks [1741]. micro-organisms [1465]. Micro/nano [975]. Micro/nanoparticle [1049]. microblogs [1462]. Microbubble [1632, 1532]. microchannel [828]. microchannels [1713]. microelectrode [493]. microflows [1163]. microlocal [1817]. micromechanical [1208]. microorganisms [1585]. micropolar [1390, 641, 1394]. microscale [1569, 1424, 1540]. microstructured [1165]. Microwave [539, 679, 481, 1261, 572, 617, 1377, 601, 707, 653, 742, 756, 743]. midpoint [1538]. migration [1689]. Miller [1034]. Mindlin [1033, 1416]. miniature [844]. minimal [319]. minimizing [525]. minimum [332]. miscible [247]. misoriented [816]. mitigating [1846]. mitral [1601]. Mixed [1374, 647, 1283, 948, 1314, 1510, 1233, 1718, 1761, 885]. mixed-mode [647]. mixed-type [885]. mixing [504, 1684, 1428]. mixture [1328, 1181]. mode [647, 1474, 545, 502, 550, 1368, 661, 1494, 819]. mode-I [1494]. mode-I/mode-II [1494]. mode-II [1494]. Model [1401, 103, 771, 1261, 130, 866, 1459, 1561, 611, 1503, 870, 1810, 1405, 1571, 1796, 1688, 1118, 1139, 1244, 1604, 888, 979, 1483, 1792, 1188, 1763, 1366, 1712, 255, 350, 874, 1643, 1402, 1224, 1456, 1474, 1516, 1003, 398, 453, 1281, 676, 460, 1087, 553, 738, 1670, 1707, 1274, 860, 1482, 1352, 1650, 1591, 1599, 715, 286, 1721, 21, 1683, 1719, 1094, 1495, 472, 423, 450, 98, 272, 1537, 1479, 1174, 1758, 1488, 1006, 509, 1531, 1410, 1167, 501, 439, 1677, 932, 1325, 1678, 1530, 539, 1132, 1771, 1484, 1431, 727, 880, 1070, 395, 1245, 1388, 1430, 1056, 425]. model [1230, 758, 1239, 1407, 1059, 1254, 962, 1698, 496, 547, 985, 1473, 329, 905, 1256, 1399, 723, 1672, 1077, 1133, 404, 1387, 1485, 1339, 1349, 1345, 921, 982, 993, 1833, 1772, 389, 722, 1813, 1836, 1415, 669, 1392, 1359, 1376, 1173, 1848, 424, 1312, 1550, 1125, 209, 827, 1113, 1445, 1016, 1458, 1539, 1715, 1795, 998, 709, 1682, 730, 1307, 1657, 1798, 1294, 1323, 1175, 1406, 1600, 1525, 1137, 1794, 383, 459, 1480, 1868, 1825, 1033]. Modelling [673, 1187, 1638, 592, 1764, 756, 1284, 1302, 1851, 735, 1663, 1432, 1790, 883, 1861, 1642, 1784, 1454, 1144, 794, 831, 1158, 1532, 80, 593, 738, 1500, 1131, 721, 1849, 981, 1344, 1148, 1506, 1160, 959, 697, 633, 468, 1518, 1255, 1814, 1505, 251, 1350, 1552, 1783, 830, 941, 1181, 1492, 189, 1631, 1422, 1129, 1165, 995, 1856]. models [1340, 777, 1850, 1869, 1354, 1852, 1801, 1507, 333, 516, 1509, 1757, 997, 1749, 1826, 1728, 1336, 839, 1689, 913, 1554, 1765, 1282, 296, 1579, 1076, 1243, 792, 220, 694, 1284, 1793, 293, 1465, 1168, 1424, 1540, 632, 1178, 1426, 1626, 1277, 1190, 432, 1161, 261, 1521, 1326, 1385, 456, 1041]. Modern [1163]. modes [1741, 627, 486, 804, 422, 1520, 1403, 439, 1639, 1022, 1615, 124, 963, 1486]. modified [891, 277, 95, 1290, 633, 1214, 245, 1397, 781]. Modulated [1609, 727, 880]. modulated-temperature [727]. Modulation [977, 1575, 603, 1547, 933]. moduli [1278]. modulus [1699]. Moiré [349, 152]. moments [517, 246]. Monge [1247]. monitored [1716]. monitoring [1616]. monoclinic [895]. monodomain [1431]. monostable [1412]. Monotone [866, 1863]. monotonicity [423, 450, 1576]. Moore [775]. mooring [105]. morphogenesis [526, 389]. morphological [185, 361, 184, 391]. morphologies [1809]. Mosaic [703]. mosquito [1795]. mosquito-borne [1795]. motility [1325, 1836, 1715]. motion [1673, 700, 597, 753, 97, 535, 1080, 855, 723, 1558, 1333, 92, 1067, 53, 373]. motions [105, 616, 104]. motivated [1807, 1786]. motor [1300]. motors [844, 1010]. moulding [672, 1518]. moulds [1707]. mount [1439]. movements [1638]. moving [1215, 1213, 318, 55, 623, 169, 1705, 597, 1031, 77, 1189, 598, 1411, 39, 1363, 399, 1246]. moving-boundary [169, 598]. MR [219, 278, 450, 569, 437, 37]. MRE [1699, 1858]. MRI [1479]. muco [111]. muco-ciliary [111]. Mullins [814]. multi [1071, 1487, 1683, 1164, 1832, 1609, 876, 1548, 1544, 1326]. multi-blebbing [1326]. multi-component [1164, 1609]. multi-innovation [1548]. multi-layer [1832]. multi-layered [1544]. multi-organ [1683]. multi-shock [876]. multi-trace [1487]. multi-wavelets [1071]. multichromatic [1869, 1852]. multicompartment [1118]. multicomponent [338]. Multicyclic [1350]. Multidimensional [328, 893, 1252, 1649, 440]. multimodal [1007]. multimode [1261]. Multiphase [210, 1591, 1163, 1181]. Multiple [1598, 1194, 1476, 1665, 662, 591, 1497, 1021, 1545, 1559, 1626, 1251, 1480]. multiplicative [1797, 1496]. multiply [787, 940, 1443, 1597, 359]. multiply-connected [359]. Multipole [536, 1001]. multiregional [85]. Multiscale [1454, 1144, 1160, 1255, 1222, 1540, 1415, 1787]. multispectral [1360, 928]. multivalued [874]. multivariable [1734]. multiwavelets [1320]. mushes [1809]. mushy [88, 140]. mutual [1426]. mycelia [593].

Nagumo [1619, 406]. Namias [262]. nano [1513, 1700, 975, 1159]. nano-flow [1159]. nano-inhomogeneity [1513]. nano-slit [1700]. nanoparticle [1049]. nanoparticles [1834]. nanostructures [1366]. nanotubes [1281]. nanowires [1190]. narrow [466]. Natural [1318, 673]. Nature [1275, 1208]. Navier [1528, 1420, 385, 1274, 1138, 1391, 1457, 675, 908, 770, 1388, 693, 939]. near [731, 1215, 135, 804, 856, 1580, 1384, 740, 395, 483, 1318, 1541, 1220, 1690, 1594, 723, 1374, 355, 1870, 1625, 1632, 1444, 1302]. near-equidiffusional [483]. near-equilibrium [395]. near-field [731]. near-shore [355]. near-trapping [804]. nearest [1702]. nearly [1520, 1612, 321, 158, 382, 1669]. nearly-vertical [158]. necessary [1635]. necking [1105]. negative [1046]. neighbourhood [1731, 640]. nematic [1471, 1517, 592, 1808, 1440]. neo [847, 401]. neo-Hookean [847, 401]. nets [1225, 1858]. netting [658]. network [1702]. Networks [1305, 1619, 1651, 1829, 1643, 1402, 1822, 351, 1827, 1737, 1035, 1358, 1662, 107, 523, 902, 1288, 1644, 411, 1679, 1441, 1659]. Neuber [1042]. Neuber-based [1042]. Neumann [691, 596, 650, 931, 1865, 47, 1385]. Neumann-Kelvin [596, 650]. neural [1829, 1822, 1858, 1827, 1035, 1358, 902, 1288, 1679]. neurons [899]. neuroscience [1463]. neutral [1666, 1388]. Newell [1501]. Newtonian [1606, 1213, 1442, 1664, 1713, 1771, 1168, 837, 1184, 1371]. Nicholson [1304]. nickel [794]. nilpotent [683]. NLS [1547]. no [219, 1855, 278, 450, 569, 1839, 437, 37]. no-contrast [1855, 1839]. nodal [1017]. node [1702]. noise [1508, 367, 1419, 1745, 1349]. Noisy [1749, 1124]. Non [624, 637, 176, 1019, 984, 1759, 1664, 899, 778, 1170, 1258, 1296, 1145, 1202, 293, 1029, 10, 876, 865, 1341, 1046, 1020, 1207, 771, 1459, 1606, 487, 1061, 1322, 1861, 924, 1213, 1390, 1837, 1404, 1454, 882, 1280, 893, 886, 1097, 1553, 1030, 860, 954, 642, 1375, 1504, 863, 1537, 1044, 1088, 1002, 1007, 1866, 720, 796, 1191, 1075, 1156, 1472, 1027, 1348, 1200, 959, 1201, 691, 1771, 454, 1004, 1609, 841, 1491, 1755, 1009, 1662, 1407, 1698, 728, 1841, 972, 1778, 1141, 1168, 1502, 1100, 1547, 1611, 802, 1206, 1310, 1367, 969, 1082, 977, 963, 1738, 1184]. non [1371, 1633, 1669, 998, 1294, 1149, 1389, 1060, 1382]. non-absorbing [1404]. non-autonomous [1841]. non-backtracking [1662]. Non-classical [984]. non-commutative [1755]. non-convex [1061, 642]. non-destructive [1553]. Non-diffusive [293]. non-dispersive [1669]. Non-equilibrium [1759]. Non-existence [1046]. non-extensive [969]. Non-failure [1170]. non-Green [1837]. Non-holonomic [1202]. non-homogeneous [1027, 841, 1841, 802]. non-hypersingular [691]. non-instantaneous [1633]. non-isospectral [1191]. non-isothermal [1367]. Non-Linear [10, 1019, 899, 1145, 1029, 865, 1020, 1207, 1322, 924, 1454, 882, 1280, 893, 886, 1030, 860, 954, 863, 1044, 1002, 1075, 1156, 1200, 1201, 1004, 1755, 1009, 972, 1778, 1141, 1502, 1100, 1547, 1206, 1310, 1082, 977, 963, 1046, 1389, 1060, 1382]. non-linearities [1097, 1348]. non-linearity [1611]. non-linearly [487, 454, 1341].

Non-local [624, 1258, 1296, 771, 1459, 1861, 1030, 1537, 1088, 1866, 959, 1004, 1491, 1755, 1407, 1082, 998, 1294, 1149]. non-negative [1046]. Non-Newtonian [1664, 1606, 1213, 1771, 1168, 1184, 1371]. non-penetration [796]. Non-periodic [778]. Non-planar [637]. Non-principal [1341]. non-smooth [1472, 1698]. non-solvable [728]. Non-spectral [176]. non-standard [1375]. non-symmetrical [720]. non-topological [1738]. non-trivial [1504, 1609]. non-uniform [1390, 1007]. non-unique [1009]. Non-uniqueness [876]. non-zero [1502]. Nonclassical [473]. nonequilibrium [391]. nonexistence [1714]. nongraphical [44]. nonhomogeneous [524, 1690, 154]. Nonlinear [1606, 476, 603, 502, 185, 369, 1830, 1669, 1773, 1863, 1807, 480, 473, 178, 813, 304, 312, 69, 746, 503, 1122, 38, 1837, 1802, 1469, 831, 657, 346, 1532, 739, 1574, 420, 531, 589, 1599, 715, 917, 114, 514, 438, 490, 67, 98, 82, 281, 58, 378, 360, 417, 689, 936, 662, 445, 498, 241, 697, 97, 1560, 161, 313, 1745, 636, 1786, 121, 937, 828, 845, 660, 494, 250, 491, 580, 575, 380, 556, 1833, 561, 648, 40, 568, 838, 273, 1575, 805, 784, 179, 265, 566, 70, 1236, 1308, 163, 1753, 1548, 386, 184, 391, 709]. nonlinear [326, 5, 1709, 1525, 1794]. nonlinear-nonlocal [1236]. nonlinearities [701, 949]. Nonlinearity [1193, 538, 569, 674, 1750]. nonlinearly [818, 822, 283]. Nonlocal [414, 559, 472, 507, 1804, 935, 291, 1236]. nonmonotone [192]. nonparallel [58]. nonpenetration [907]. nonsmooth [354, 275]. nonstandard [824]. nonstratified [508]. nonsymmetric [424]. norm [1391, 332]. normal [1475, 972, 683]. Norton [762]. Note [120, 159, 16, 407, 1091, 1588, 1414, 886, 133, 138, 34, 353, 1529, 51, 321, 1309, 580, 300, 142, 571, 125, 173]. notion [1047]. novel [1861, 169, 1243, 1316]. nozzle [1398, 1037, 876]. NS [1076]. NS-alpha [1076]. NS-alpha-deconvolution [1076]. nuclear [85]. nucleate [1181]. nucleation [414]. nucleolus [1437]. null [134]. null-field [134]. number [1607, 1384, 95, 45, 300]. numbers [1200, 1248]. Numerical [861, 1232, 448, 1781, 1276, 243, 1365, 1572, 1737, 1180, 1036, 972, 982, 568, 1393, 1854, 255, 874, 309, 1685, 302, 428, 1549, 1370, 928, 1042, 1654, 1052, 409, 1835, 1622, 92, 894, 1361, 1018, 1316, 1346, 1060]. numerical-asymptotic [1370]. nutrient [632].

object [1395]. oblate [1626]. oblique [1390, 502, 1614, 550, 1566]. observation [307]. observations [79]. obstacle [736, 585, 690, 1404, 1074, 930, 1708, 1751, 971, 638, 430, 1566, 1695]. obstacles [577, 1541, 644, 793, 1615]. obtain [1807, 1786]. obtained [1780, 660]. obtaining [683]. occurring [484]. ODE [1405, 1833, 1277]. ODEs [1807, 1786]. off [723, 1183]. Ogden [1329]. oil [200, 1244]. oilwell [1846]. Omega [1841]. on-site [1399]. on-surface [625, 710]. One [282, 1232, 393, 73, 1030, 394, 930, 83, 604, 1068, 220, 1029, 845, 1714, 253, 311, 550, 1110, 605, 903, 1524]. One-dimensional [282, 393, 73, 1030, 394, 83, 604, 220, 845, 1714]. one-mode [550]. one-phase [1232, 311]. one-shell [1110]. one-sided [1029]. only [1336]. Ono [364]. onset [1860, 1608, 403]. onto [1189]. open [1788, 1403, 579]. opening [1314, 1766]. operated [772]. operating [611]. Operational [85, 426]. operations [13]. operator [736, 685, 100, 1385, 1723]. operators [777, 1774, 148, 1487, 695, 785, 1423, 1386, 1027]. opinion [1749]. optic [289]. Optical [393, 289, 1741, 627, 486, 947, 342, 1211, 1687, 1202, 1609, 329, 1727, 436, 1020]. optics [413]. Optimal [323, 1381, 1778, 1342, 737, 325, 589, 1504, 1360, 1584, 1431, 956, 1361]. Optimisation [853]. Optimization [200, 426, 397, 611, 1684, 1362, 1622, 1825]. optimize [1605]. Optimizing [1405, 1723]. Optimum [466]. option [1400, 1287, 1535]. options [1482]. Orbital [1353]. orbits [798, 1380]. order [624, 257, 165, 1648, 759, 1774, 503, 139, 1276, 1240, 59, 1580, 1538, 545, 1030, 576, 1362, 99, 832, 415, 360, 881, 936, 618, 1755, 1284, 324, 1496, 12, 1611, 1374, 405, 568, 273, 1429, 1860, 571, 493, 402, 1657, 1393, 1434, 1603]. orders [858]. ordinary [559, 274, 1147, 581, 711, 290, 1434]. ore [745]. ore-forming [745]. Oregonator [1059, 1254]. organ [1683, 1546]. organ-inspired [1546]. organisms [1465]. organizing [784]. Oriented [1365, 1346]. Origin [1727]. Orthogonal [99, 1071, 41]. Orthotropic [1339, 1053, 1559]. Oscillating [16, 1511, 1079, 352, 591, 1272, 1255, 1675]. oscillation [1670, 860, 490]. oscillations [398, 72, 1269, 671, 1398, 841, 1473, 1672, 1608, 1055, 1238]. oscillator [1472, 1419, 245, 104]. oscillators [438, 490, 838]. Oscillatory [257, 1805, 492, 1724, 1013, 838]. Oseen [347]. other [117]. outbreaks [1848]. outer [597]. outfalls [737]. output [1386]. outside-in [594]. overall [265, 566, 326]. overdetermination [1023, 1386]. overhead [270]. Overpotentials [199]. Oxidant [355]. Oxidant-limited [355]. Oxidation [464, 794, 588, 1180].

packed [699]. packet [668]. Padé [704, 20]. Painlevé [331, 448, 1467, 788, 583, 634, 1845]. paint [447]. pair [597, 922, 1114]. pairs [618]. pairwise [757]. panic [1643]. Papkovich [1042]. Parabolic [346, 266, 831, 942, 1088, 897, 1004, 475, 949, 1206, 634, 300]. parabolic-type [942]. paradox [1467]. parallel [1420, 1688, 1131, 1849, 77, 1753, 1480]. Parameter [657, 1796, 1243, 1682, 1616]. Parameterization [1297]. parameterized [60]. Parameters [190, 1396, 1080, 1449]. Parametric [1100, 531, 642, 1173]. parasitic [1833]. Parseval [791, 310]. Parseval-type [310, 791]. part [790, 1209]. partial [1802, 1062, 764, 1549, 1180, 1112, 755, 70, 238, 939]. partially [1788, 1442, 1742, 53, 970, 403]. Particle [1195, 1840, 623, 1871, 1180, 1050, 222, 399]. particles [1234, 225, 217]. particulate [21, 909]. pass [706]. Passages [648]. passive [918, 1728]. past [630, 1263, 1534, 1145, 1192, 446, 1660, 1114, 807]. paste [317]. patches [1454, 474]. path [1377]. pathway [1500]. Patrick [1726]. Pattern [1674, 1151, 506, 858, 707, 1433, 1663, 924, 1019, 521, 930, 1767, 1715, 1020]. pattern-forming [1019]. patterned [1363]. patterning [1555]. patterns [1733, 1730, 877, 228, 740, 151, 1537, 1758, 233, 1724, 1426, 1236, 1736, 1016, 1041]. PDE [1225, 1212, 1622]. PDE-constrained [1622]. PDEs [954, 848, 576, 881, 1865]. peakon [1493]. peaks [1734]. penalty [160, 174]. penalty-perturbation [160]. Pendent [1847]. pendulum [161]. penetrable [1855, 1370, 1839, 1093, 1671]. penetrant [528]. penetration [796, 550]. penetrative [961]. Percolation [86]. Peregrine [1805]. perfect [920, 94]. perfectly [759, 1378, 290, 1626, 315, 834]. perfectly-reflecting [1626]. perforated [1229, 1685, 538, 569, 1725]. performance [1405, 1021, 1548]. period [1289]. Periodic [860, 251, 1067, 557, 1441, 625, 1249, 480, 868, 1763, 798, 1516, 1412, 398, 1229, 545, 68, 1501, 778, 1489, 172, 662, 1816, 1208, 1212, 1056, 1214, 1477, 758, 1353, 1630, 543, 1234, 1349, 1361, 550, 1013, 1429, 1085, 772, 1251, 126, 147, 1710, 1862, 1723]. periodically [733, 1698, 960, 1382]. Periodicity [843]. permafrost [1392]. Permanence [1056, 456, 800]. permanent [844, 1010, 1300, 721, 613, 496]. permeabilities [1665]. permeable [1215]. perpendicular [1453]. persistence [1086, 671, 1584]. persistent [1355]. perturbation [160, 484, 204, 1084, 1747, 316, 1395, 1025, 1192, 253, 162, 389, 1286, 433, 781]. Perturbational [158]. perturbations [38, 851, 1295, 838, 1134]. Perturbed [1471, 1568, 918, 1403, 1331, 1517]. perturbing [763]. Petrov [633]. Petviashvili [331]. pH [1415]. pH-tactic [1415]. Phase [211, 1451, 526, 1663, 1232, 746, 249, 255, 333, 50, 492, 1066, 243, 286, 951, 1749, 1495, 236, 423, 450, 1728, 509, 898, 1171, 806, 1325, 413, 83, 1609, 792, 962, 1547, 1485, 311, 47, 1178, 993, 430, 13, 512, 1605, 168, 191, 261]. phase-change [286]. phase-field [1663, 1485]. phase-field-crystal [1728]. Phase-winding [211]. phased [586]. phaseless [1686]. phases [286, 525]. phenomena [735, 1200, 618, 1214, 122]. phenomenon [1442, 1311, 784]. photoacoustic [1768]. physical [1449]. physically [65]. physics [1827]. physics-informed [1827]. physiological [122]. picture [373]. Piecewise [1864, 798, 1474, 1522, 1409, 1408]. piercing [596, 650, 198]. piezoelastic [913]. piezoelectric [1627, 729, 681, 819, 1510, 1021]. pinball [1127]. pinned [1598, 1444]. pipe [622, 1442, 1401, 1639, 142, 1546, 1873]. pipes [304]. Pipkin [522]. Pitaevskii [882]. pitchfork [945]. planar [637, 73, 1509, 1242, 1842, 1558, 805, 1605]. Planck [432, 399, 1692]. Plane [1102, 487, 377, 1218, 135, 1513, 72, 918, 929, 1283, 1453, 895, 621, 335, 720, 822, 1378, 1109, 1203, 600, 635, 9, 523, 1722, 847, 836, 1297, 1342, 914, 1116, 1657, 533, 339]. Plane-strain [339]. planes [211, 1453]. planform [1151, 1520]. plankton [1586, 632, 1251]. plasma [1592, 350, 788]. plasmas [1090]. plasmid [1137]. plasmid-bearing [1137]. plasmid-free [1137]. plasmons [1637]. plastic [594, 621, 1762, 1419]. plasticity [1132, 174]. plate [234, 160, 1215, 1073, 1567, 62, 753, 879, 1762, 595, 641, 894, 1286, 35, 807, 1045, 1416, 1033]. Plateau [1618]. plates [599, 639, 716, 787, 1375, 429, 444, 49, 77, 215, 922, 1092, 1107, 1306, 87, 694, 628, 1022, 1492, 1055, 1238, 1129, 1711]. Plesset [1024, 1289]. plugging [87]. plume [1146, 181]. Plumes [381, 1242]. plurality [1576]. Point [377, 529, 1649, 777, 1850, 135, 1801, 1390, 1471, 1517, 626, 1534, 1563, 1475, 592, 42, 1212, 1594, 638, 1374, 915, 237, 566]. Point-actuated [1649]. point-circle [777]. point-contact [1850, 1801]. Point-source [377]. point-to-periodic [1212]. point-to-point [1212]. points [239, 855, 616, 402, 1798, 1382]. pointwise [832]. Poiseuille [1623, 1297, 124]. Poisson [1779, 1388, 965, 788, 938]. Pol [108, 68, 245]. Polarity [1767]. Polarity-driven [1767]. polarization [1703]. polars [1838]. pole [1010]. poles [1502]. Pollutant [381]. pollution [466]. polyconvex [1330]. polygamma [164]. polygon [782]. polygonal [1713]. polygons [1370, 1740]. polyhedra [46]. polyhedral [1039]. polymer [528, 1379, 1328, 727, 880, 941]. polymer-penetrant [528]. polymerization [1325, 496, 547]. Polynomial [1306, 719, 1303, 1362, 1191, 17, 89, 145]. polynomials [20, 99, 1843, 41, 64]. polytropic [1371]. pool [1811]. Population [981, 866, 1588, 1003, 657, 1250, 860, 472, 839, 970, 998, 1521, 1499]. Population-scale [981]. pore [1314, 1484, 222]. Poroelastic [1346, 714, 1653, 1235, 1307, 1667, 1365]. poroelasticity [484]. Porous [1372, 1227, 1744, 1232, 1653, 810, 1520, 1629, 1705, 747, 1665, 1506, 961, 294, 1140, 353, 1484, 1318, 1036, 427, 541, 988, 387, 266, 574, 404, 786, 18, 1485, 1317, 478, 237, 142, 558, 903, 519, 247, 1667, 482]. porous-media [1232]. porous-medium [558]. posed [839, 384, 530, 1830]. posedness [1438, 951, 792, 421]. positive [1843, 1239, 900, 1392, 1251, 1089]. positivity [363]. post [1430, 1783, 1101]. post-buckling [1101]. post-Initiative-502 [1783]. post-touchdown [1430]. posts [688]. potential [196, 630, 489, 1048, 721, 151, 387, 713, 222, 1110, 1219, 1001, 1523]. potentials [1869, 1852, 393, 819, 501, 1032, 1046]. power [426, 178, 1213, 621, 256, 446, 481]. power-law [178, 1213, 621, 446]. PP04 [1763, 1698]. Practical [190]. pre [646, 816, 345, 753, 879, 600, 635, 723, 1783]. pre- [1783]. pre-stress [600]. pre-stressed [646, 816, 345, 753, 879, 635, 723]. precipitation [1484, 1485, 995]. Precise [1218]. precursor [1678]. predator [551, 1174, 1245, 1212, 1514, 1239, 1407, 1256, 1724, 1426, 1016, 1057, 1041]. predator-prey [551, 1174, 1245, 1212, 1514, 1239, 1256, 1426]. Prediction [103]. predictions [1528, 270]. predictive [1124]. Preface [109]. prepolarized [729, 819]. Presence [199, 1080, 1179, 1221]. present [1145]. preservation [1351]. preserving [920]. Pressure [706, 150, 810, 1289, 1401, 366, 437, 293, 826, 1234, 1682, 1873]. pressure-driven [810, 1873]. pressure-swing [293]. pressureless [1448]. pressurized [1732]. prestressed [729, 819, 603, 1357, 937, 411]. prey [551, 1174, 1245, 1212, 1514, 1239, 1407, 1814, 1256, 1724, 1426, 1016, 1057, 1041]. Pricing [1482, 1400]. Primal [907, 796]. primal-dual [796]. Primary [374]. prime [1468]. principal [1341]. principles [11, 128, 110, 642, 696, 179, 70, 1439]. printing [1645]. prior [1817]. priori [248, 752]. prisms [1808]. Prize [1680, 1324, 1504]. Probability [534, 1456]. Problem [1519, 101, 234, 1768, 426, 1592, 397, 1815, 734, 1218, 1215, 1061, 1498, 762, 813, 1739, 407, 1782, 19, 1788, 1073, 746, 810, 1404, 350, 73, 1034, 1028, 1280, 1276, 716, 269, 341, 504, 886, 1084, 106, 302, 428, 1072, 801, 318, 458, 1699, 1858, 55, 739, 1303, 666, 192, 228, 1453, 623, 169, 594, 621, 1375, 25, 1360, 256, 236, 82, 281, 1031, 1409, 542, 741, 702, 1088, 1002, 1023, 96, 881, 796, 898, 172, 279, 384, 445, 1285, 884, 336, 241, 1746, 879, 344, 233, 1427, 1530, 691, 596, 650, 83, 88, 1070, 1111, 1290, 1529]. problem [395, 483, 535, 684, 890, 1541, 1004, 366, 437, 1025, 795, 26, 835, 1716, 766, 154, 1510, 1835, 388, 1449, 1043, 1614, 584, 547, 768, 996, 457, 1830, 973, 1806, 1237, 510, 174, 311, 802, 47, 1361, 1233, 1448, 1772, 389, 1082, 56, 358, 933, 237, 865, 28, 27, 291, 1055, 1238, 81, 140, 382, 752, 168, 343, 1064, 1566, 402, 433, 275, 223, 1397, 1515, 1761, 508, 1686, 1709, 499, 995, 1603]. problems [701, 735, 611, 864, 160, 731, 178, 196, 1121, 1213, 681, 1792, 38, 65, 204, 249, 1791, 1802, 599, 639, 485, 649, 347, 663, 1769, 919, 465, 1283, 1789, 642, 1362, 712, 780, 497, 863, 1386, 1074, 415, 824, 306, 1842, 153, 790, 1052, 1383, 1027, 1425, 938, 1054, 604, 651, 544, 42, 530, 931, 536, 598, 396, 134, 1419, 1241, 1622, 1271, 776, 1301, 1658, 943, 266, 121, 748, 1545, 638, 195, 222, 76, 875, 163, 1753, 1177, 519, 147, 1455]. procedure [1005, 1622]. procedures [611]. Process [10, 130, 1298, 1564, 1182, 1554, 170, 1177, 1060]. processes [325, 839, 745, 1381, 769, 432, 772, 1125, 1231, 247, 1382]. processor [1499]. product [1216]. Production [259, 1244, 144, 1272]. Professor [1329, 1578]. profiles [1568, 750, 446]. progress [1413]. prolate [1423, 1626]. proof [1503, 64, 275]. propagating [531]. Propagation [1866, 6, 15, 637, 812, 992, 1653, 1837, 810, 819, 572, 495, 172, 409, 1464, 674, 944, 129, 862, 532, 579, 600, 937, 505, 1235, 982, 789, 783, 1385, 1667]. properties [1459, 965, 212, 1785, 1509, 240, 1436, 763, 265, 512, 566, 140, 1026]. property [1053, 1701, 1396, 1419, 842, 751]. proposed [905]. protected [1139]. proton [883]. protruding [1710]. protrusion [1596]. Proudman [1248]. prove [1277]. Pseudo [619, 1450, 847]. pseudo-hyperbolicity [1450]. pseudo-plane [847]. Pseudo-steady-state [619]. Pseudoparabolic [976]. pulling [80, 189]. pulsating [138]. pulse [480]. pulsed [1137]. pulses [1479, 659, 1205, 1308]. Pure [1338, 1339, 1551, 5]. purely [1063]. pursuit [1820]. pursuit-evasion [1820]. PWM [1240, 1347]. PWM-in [1240]. PWM-in/PWM-out [1240]. PWM-out [1240].

quadrangles [573]. quadrant [1867]. quadratic [857, 808]. quadrature [788]. quadratures [734, 1543]. Qualitative [943, 1256, 1077, 965, 140, 1133]. quality [1826]. Quantifying [1564, 1823]. quantities [1222]. quantization [1637]. Quantum [1507, 1083]. quarter [1218, 31]. quarter-plane [1218]. quarter-space [31]. quartic [1730]. Quasi [1062, 1388, 1501, 1269, 1037, 805, 1045]. quasi- [1037]. quasi-linear [1269, 1045]. Quasi-neutral [1388]. quasi-periodic [1501]. quasi-planar [805]. Quasi-separation [1062]. quasicrystal [1155]. quasicrystalline [1279]. quasicrystals [1228]. quasilinear [801, 931, 1233, 1295]. quasilinearization [991]. Quasisteady [1757]. questions [439]. queue [1499]. quintic [1843, 1755].

radial [1760, 487, 1338, 1014, 1705, 1735, 444, 514, 1758]. Radially [904, 1741]. Radiant [93, 137]. radiated [138]. Radiating [1533, 230]. Radiation [1725, 508, 625, 710, 864, 1182, 572, 285, 410, 289, 276, 586, 607, 1575, 376, 412, 571]. radiative [189, 175]. radius [1258, 1296]. Random [1804, 851, 1628, 1782, 1647, 1482, 1355, 474, 549, 1287, 659, 129]. randomly [1598, 645, 1660]. randomly-generated [645]. randomness [674]. range [1627, 1080, 246]. Rankine [1668]. Rankine-type [1668]. Rapid [74, 647, 989]. rapidly [1194, 1456]. rarefied [1161]. rate [1478, 288, 665, 256, 1296, 174, 661, 1794]. Rates [1117, 1645, 897, 1076, 1070, 628]. ratio [304, 313, 1256, 1724, 1150, 1293, 1165]. ratio-dependent [1256, 1724]. Rational [147]. Rational-Fourier-series [147]. Ray [1329, 334]. Rayleigh [1024, 1276, 1283, 1289, 1694, 1579, 1357, 229, 1313, 1341, 1657]. Rayleigh-type [1313]. Razumikhin [873, 1153]. Razumikhin-type [873]. reactant [180, 219]. Reacting [6, 15, 381, 4, 425]. Reaction [471, 1812, 432, 1704, 383, 459, 1733, 1459, 1619, 278, 328, 1655, 288, 1643, 1516, 893, 453, 676, 565, 593, 739, 665, 974, 1591, 1721, 954, 221, 272, 8, 1369, 1417, 1804, 1258, 1296, 1734, 1750, 51, 1245, 633, 441, 532, 614, 1059, 1254, 496, 547, 896, 1630, 526, 1611, 800, 414, 1179, 477, 562, 1013, 789, 141, 669, 773, 677, 1173, 1236, 1679, 1292, 998, 1293, 1294, 1515, 1406, 1433, 1060, 670]. Reaction-diffusion [471, 432, 383, 459, 1733, 1459, 1619, 278, 328, 1643, 593, 739, 1721, 221, 8, 1369, 1417, 1258, 1296, 51, 1245, 441, 532, 496, 547, 800, 414, 477, 562, 141, 773, 677, 1236, 1292, 1294, 1433, 670]. reaction-diffusion-advection [526]. reactions [1255, 493]. Reactive [213, 193, 1227, 1712, 362, 464, 482]. reactor [85]. reactors [284, 403]. readily [127]. Real [1351, 148, 1579]. Real-time [1351]. realization [66]. receptor [1230]. receptor-based [1230]. reciprocal [1835]. Recognition [127]. Reconstruction [1703, 1298, 1691, 1355, 930, 1708, 517, 1005, 413, 1716, 1345, 1751]. Recovering [917, 1523, 1039]. Recovery [825, 1855, 1839]. rectangle [1101]. rectangular [408, 344, 1816, 922, 1587, 836, 587]. rectification [202]. rectilinear [597]. recursive [1849]. recycling [632]. redraw [1590]. Reduced [1190, 1808]. Reducing [1690]. reduction [1405, 1557, 1072, 778, 1401, 1720, 788, 691, 711, 1547, 1368, 1292]. reductions [473, 689, 1075]. reference [412]. refinements [1861]. Reflected [659]. reflecting [1626]. Reflection [224, 635, 1652, 43, 600]. Reflections [1123, 1122]. reformer [830]. refraction [996]. refractive [659, 1803]. refugee [1689]. regime [132, 1482]. regimes [1189, 1318]. region [354, 88, 1192, 1477, 28, 27, 140, 359]. Regions [84, 115, 196, 83, 680]. register [177]. Regular [227, 785, 46, 316]. regularisation [1823]. Regularity [1299, 1138, 1457, 1633]. Regularization [1088, 1613, 378, 1826, 1081]. Regularized [1430, 1805, 1565, 1386, 1716, 1332]. regulated [1325]. Reinforced [1365, 155, 1338, 733, 9, 1346]. Reissner [802, 1492, 1416, 1033]. relapse [1539, 1525]. related [731, 485, 767, 1002, 153, 439, 1419, 1573, 1050, 307, 757, 568, 100]. relation [192, 1641, 1332]. relations [1414, 1864, 524, 326]. Relationship [1386, 1427]. relative [1681, 368]. relaxation [1750, 1865, 1672, 168, 1444]. Relaxed [452, 488, 205]. released [464]. relevant [65]. reliability [771]. relief [1401]. remark [132]. Remarks [365, 63]. remodelling [1343]. remote [1513]. removal [1682]. renewal [1, 37]. repeated [1339]. Repeating [30]. representation [668, 378, 1386, 1211, 203, 1332, 1766, 1408, 1394]. representations [165, 1570]. republished [1107]. repulsion [1571, 1452, 1868]. reservoir [862]. residual [1028, 919, 825, 1526]. Resilience [670, 800]. Resistance [1234, 1668]. resistanceless [650]. resistant [1338, 1492]. Resonance [433, 438, 490, 42, 1700, 1100, 648]. Resonances [91, 1791, 531, 104]. resonant [1658, 851, 251, 1706]. resonating [1113]. resonators [1445]. respect [165, 663]. response [1796, 853, 1714, 1236, 1312]. responses [1029, 1773]. resting [1770]. restricted [1746]. result [63, 1230]. results [1705, 417, 394, 271, 550, 158, 1277, 123, 1279, 1787]. retinal [1224]. reversal [1176, 352]. reversals [591]. reverse [1128]. reversible [425]. reversing [1822, 352, 591]. Review [23, 1161]. revisited [1531]. Reynolds [1606, 1384, 1200, 1248, 491, 45, 933, 977]. Reynolds-averaged [491]. rhombic [770]. Riccati [1496]. Richards [1636]. ridge [362]. Riemann [813, 801, 1538, 1427, 1054, 1301, 1237, 1295, 1448, 1772, 1551]. Riesz [1018, 842, 751, 1393]. right [1855, 1839, 607]. right-angled [1855, 1839]. right-circular [607]. rigid [1490, 1115, 870, 1384, 276, 1220, 1128, 827, 1625, 1632]. rigid-body [1128]. rigidity [1740, 26]. rigidly [1742]. ring [899, 1736, 1302, 1020]. rings [454]. ripples [1833]. river [1826]. Rivlin [889]. rivulet [834]. rivulets [1847]. road [1284, 1772]. roasting [1591]. Robin [1739, 1685, 953, 451, 1013, 1055, 682, 459]. robotics [1422]. Robust [759, 1052]. rock [214]. rocks [745]. rod [861, 870, 609, 646, 424, 1555]. rod-and-spring [424]. rod-annular [1555]. rods [1684]. rogue [1857, 1706]. role [1224, 987, 1086, 840, 1584, 1860, 225, 217]. roles [1681]. rolling [208, 53]. roots [197]. Rossby [1607]. rotary [1846]. rotates [270]. rotating [1561, 1194, 211, 1620, 775, 1269, 968, 1742, 1486, 1184]. rotation [1198, 1319]. rotational [1868]. rough [1606, 864, 474, 1093, 1363]. roughness [910]. rubber [871]. rubber-like [871]. run [1584, 1376, 1183]. run-and-tumble [1584]. run-off [1183]. runaway [468, 511, 403, 743]. running [844, 1010, 1300, 1798]. rural [1376]. Ryaben’kii [1032].

S [1091]. S-transform [1091]. S. [23]. Sagoci [802]. sails [700]. salt [309]. sampling [1753]. sandwich [1130]. sap [1530]. Sasa [1857]. satellite [852, 754]. satisfaction [1131]. satisfying [249, 451, 1085]. Satsuma [1857]. saturated [1520]. saturation [903, 1175]. sauna [169]. savings [1437]. scalar [1557, 1577, 997, 1750, 1050, 562]. scale [981, 1530, 1484, 1230, 1337, 1178]. scale-tissue [1337]. scales [1627, 1497, 1251]. scanning [727, 880]. scattered [710]. scatterer [731]. scatterers [759, 517, 1039, 1867]. Scattering [54, 234, 625, 1788, 1791, 652, 1378, 595, 680, 1109, 198, 920, 640, 1490, 736, 1176, 797, 864, 585, 690, 731, 1040, 809, 868, 804, 927, 106, 302, 428, 529, 649, 663, 1769, 916, 1370, 1409, 542, 467, 1074, 790, 336, 277, 685, 691, 1093, 971, 1315, 1556, 1656, 750, 134, 1614, 644, 1545, 638, 793, 607, 195, 1626, 2, 56, 430, 412, 175, 343, 1566, 1718, 1695, 1761, 1803, 508, 1686, 1756, 1385]. scheme [1232, 1612, 1716, 915, 1162]. schemes [1114, 920]. Schlichting [582]. Schlichting/vortex [500]. Scholes [1535]. Schottky [1468]. Schrödinger [480, 936, 1269, 1560, 1755, 1745, 1502, 1575, 1709]. Schur [1545]. screen [538, 569, 1725]. screens [1685, 550]. sea [105]. Second [405, 1429, 493, 1648, 1276, 1240, 59, 1538, 99, 881, 501, 618, 1284, 1496, 1374, 402, 1657, 1603]. Second-order [405, 493, 1240, 59, 1538, 99, 881, 618, 1284, 1374, 402, 1657, 1603]. section [196, 1593]. sectional [1712]. sections [201]. sector [1556, 1656]. sectors [820, 822]. sedentary [970]. sedimentation [1577, 1642, 1586, 222]. seepage [1036]. Segel [932]. Segregation [317]. SEIR [1765]. Sekerka [814]. Sel′kov [553]. Selection [101, 1407, 372, 1519, 856, 1019]. Self [908, 1305, 1797, 1247, 1511, 695, 785, 322, 1716, 1558, 837, 435, 455, 693, 784, 265, 225, 217, 1389]. Self-Adaptive [1305]. self-adjoint [695, 785]. self-adjointness [1389]. self-consistent [265]. self-heating [435, 455]. self-induced [225, 217]. self-monitored [1716]. self-organizing [784]. self-oscillating [1511]. Self-similar [908, 1797, 1247, 322, 1558, 837, 693]. Semenov [468]. Semi [1304, 1129, 234, 1733, 681, 62, 495, 1804, 409, 742, 1109, 1556, 1656, 222, 634]. Semi-analytical [1304, 1129, 742]. semi-infinite [234, 681, 62, 495, 1804, 409, 1556, 1656, 222]. semi-linear [634]. semi-strong [1733]. semi-transparent [1109]. Semiconductor [947, 48, 629, 143, 220, 203, 400, 253, 661, 327]. semiconductor-device [203]. semiexplicit [378]. semilinear [514, 295, 1613]. Sensitivity [1700, 322, 1303, 1853, 907, 838]. sensor [1303]. separable [612]. separated [159]. separates [3]. Separating [308]. separation [1681, 243, 1062, 840, 806, 1626]. separation-of-variables [1626]. separations [349]. sequences [177, 1, 37]. Sequential [101]. sequestration [1629]. series [29, 1807, 622, 1300, 1849, 1442, 1096, 854, 1462, 171, 515, 1786, 1841, 147]. Serre [1696]. server [1754]. service [1437]. servomechanisms [1044]. set [398, 695, 796, 613]. sets [308, 71]. setting [1286, 1376]. settling [1712]. Several [15, 1210, 577, 1265, 17, 89, 145, 402]. SGP4 [1380]. shaded [1010]. shaded-pole [1010]. shadow [1192]. Shafranov [583]. shallow [548, 79, 984, 1269, 1199, 1318, 748, 1669]. shallow-water [79, 1199, 1669]. Shape [1684, 103, 1527, 397, 779, 1435, 294, 907, 1132, 1436, 931, 1811, 658, 1766, 1742, 682]. shapes [812, 1081, 60, 574]. Shapley [1437]. sharing [1499]. Sharp [882, 1379, 897, 925, 1016, 916, 434, 92, 575]. Sharp-type [575]. Shaw [333, 1705, 1031, 254, 1477, 372, 1134, 570]. shear [1340, 599, 716, 590, 1338, 1699, 1574, 1336, 613, 1200, 1832, 1596, 1185, 1339, 836, 1444, 464, 1692]. shear-relaxation [1444]. sheared [224, 1172]. shearing [820, 494, 401]. sheet [1856, 1590, 1367, 1851]. sheets [849, 153, 1562, 9]. shell [826, 1110]. shelled [1532]. shells [1732, 267]. shielding [921]. shift [177]. shifting [1866]. ship [208, 125]. Shock [993, 287, 1805, 1838, 1481, 876, 1162]. shocks [1295]. shooting [575]. shore [355]. Short [753, 1154, 323, 1376]. short-distance [323]. short-run [1376]. Shtrikman [512]. sided [1029]. sideways [1613]. Sierra [1848]. sign [701, 1791, 1455]. sign-changing [1791, 1455]. signal [839, 674]. signalling [1498, 395]. signals [668, 1085]. Signorini [762, 666]. Silicon [280, 188, 588, 231, 242]. similar [1797, 1247, 908, 322, 1558, 837, 693]. Similarity [1188, 337, 279, 40, 283, 386, 992, 1213, 82, 335, 427, 244, 610, 292, 482]. simple [958, 676, 1586, 821, 869, 514, 521, 858, 1349]. simplification [739]. simplified [1127, 78]. simulate [1654]. Simulation [235, 1181, 874, 1164, 1582, 1211, 1160, 1036, 129]. Simulations [1583, 1528, 1273, 1454, 738, 1166, 1159]. Simultaneous [1023]. sine [788, 1196, 1152]. sine-Gordon [1152]. sine-Poisson [788]. single [453, 844, 1066, 542, 352, 661, 296]. single-mode [661]. single-phase [1066]. single-source [542]. single-stator [844]. Singular [701, 735, 1774, 1199, 1192, 299, 484, 901, 1084, 1087, 785, 951, 1386, 1074, 1025, 991, 253, 162, 314, 433, 1026]. Singularities [229, 339, 1265, 1562, 1602, 1268, 1675, 1416]. Singularity [1011, 1385, 630, 232, 867]. singularity-free [867]. Singularly [1568]. sinh [965]. sinh-Poisson [965]. sinks [1660]. SIR [1765, 109]. SIRS [1456]. site [1399]. Sivashinsky [1504, 1777, 1368, 440]. six [306]. sixth [568]. sixth-order [568]. size [1531, 1395, 541]. sizes [1721, 1711]. skarn [745]. skin [348]. skin-depth [348]. slab [193, 1261]. slabs [707, 653]. slaving [1457]. slender [1607, 196, 1091, 980, 1065, 1816, 654, 713, 1246]. slider [1363]. sliding [732, 1474, 1476, 1616]. Slightly [150]. Slip [1593, 1420, 1011, 474, 549, 1832, 1713, 975, 142, 939, 1710]. slit [1700]. sloping [86]. sloshing [795, 251]. slot [297]. slow [1655, 1816, 1428, 35]. Slowly [1540, 193, 1674, 1328, 1799, 1424, 834]. Slugging [1776, 1617, 990]. Small [678, 1490, 1768, 1607, 1703, 1598, 652, 529, 67, 497, 390, 1687, 1200, 24, 653, 1562, 452, 371, 174, 222, 1295, 838, 13, 1150, 640, 1862]. small-amplitude [390]. small-aspect-ratio [1150]. small-strain [174]. small-time [67, 1562]. smectic [1650, 724, 872, 634]. smectic-A [1650, 872]. smelting [317]. Smoluchowski [1797]. smooth [809, 1522, 1472, 1859, 1210, 1166, 1698]. smoother [1777]. smoulder [637, 992, 548, 495, 409]. Smouldering [812]. Snaking [1738, 1736, 1726, 1735, 1732, 1734]. Sobolev [1787]. social [1604, 1521]. society [1148]. sodium [1020]. soft [234, 1673, 1335, 529, 1472, 930, 336, 1867, 214]. softening [613, 1350, 1339]. soils [786, 903]. sojourn [1499]. solid [647, 1335, 345, 1083, 495, 474, 409, 633, 1596, 1333, 586, 1313, 175, 261, 640, 1302]. solid-solid [261]. Solidification [190, 619, 761, 185, 1558, 5]. solids [487, 1673, 1099, 325, 816, 821, 869, 1285, 600, 1100, 494, 1331, 960, 1047, 1341, 1334, 1675].

Solitary

[1534, 1086, 936, 364, 1154, 1819, 1572, 1198, 1533, 1859, 1496, 1204, 1205, 1308]. solitary-wave [364]. Soliton [114, 554, 1202]. solitons [1720, 126]. Solonnikov [956]. solutal [391]. solute [464]. Solution [734, 38, 901, 382, 238, 992, 426, 364, 102, 1508, 1061, 196, 762, 746, 622, 350, 252, 1034, 1280, 398, 470, 1519, 428, 819, 1513, 1065, 347, 1769, 589, 1442, 594, 1375, 699, 1549, 98, 497, 1023, 8, 445, 1285, 1696, 1052, 277, 409, 595, 1111, 1529, 483, 51, 684, 530, 598, 1594, 1841, 469, 1545, 290, 47, 902, 575, 654, 1018, 867, 568, 399, 125, 752, 939, 223, 1798, 499, 1045, 482, 996]. Solutions [1075, 799, 703, 1304, 416, 1459, 257, 1807, 480, 57, 965, 1636, 630, 705, 1263, 849, 1213, 1528, 1797, 1188, 278, 268, 328, 227, 419, 537, 208, 1051, 406, 831, 942, 1412, 453, 599, 341, 801, 211, 553, 620, 1196, 545, 337, 817, 860, 1650, 715, 1789, 1197, 68, 114, 576, 1362, 1504, 472, 221, 319, 340, 1664, 514, 423, 450, 82, 272, 612, 675, 1031, 908, 1665, 1506, 1386, 1489, 240, 335, 513, 619, 279, 306, 1369, 1417, 662, 948, 1348, 1199, 354, 583, 697, 1140, 95, 1210, 1765, 1750, 931, 1058, 1259, 1212, 745, 742, 1092]. solutions [1107, 427, 1493, 744, 862, 203, 1477, 1009, 1835, 758, 769, 1239, 71, 229, 1745, 628, 418, 1786, 1301, 1496, 991, 12, 949, 1141, 1037, 246, 660, 251, 307, 1233, 1295, 982, 562, 1082, 836, 141, 40, 876, 960, 1429, 693, 237, 283, 814, 163, 970, 827, 1046, 1251, 1253, 402, 433, 386, 1089, 1294, 1857, 1149, 440, 1441, 1862, 1873, 292, 1603]. Solvability [771, 1314, 1033, 1744, 1565, 885]. solvable [728]. solvated [1505]. solve [780]. solvent [1328]. solvent-polymer [1328]. Solving [1802, 919, 1739, 1842, 336, 1575, 314]. Some [164, 63, 1461, 79, 248, 514, 254, 335, 417, 1336, 394, 544, 1243, 418, 847, 222, 123, 17, 184, 611, 30, 102, 1528, 1414, 212, 942, 1044, 258, 839, 515, 146, 1753, 127, 1413, 1524]. somersaults [1605]. Sound [6, 15, 54, 234, 1768, 135, 117, 144, 259, 538, 569, 930, 535, 680, 922, 1867, 543, 805, 334, 1725]. sound-soft [930, 1867]. source [473, 377, 135, 1823, 1066, 529, 1298, 1564, 138, 21, 542, 1023, 1386, 967, 282, 1452, 1264, 1594, 1630, 638, 586, 181, 1082, 950, 125]. sources [1074, 83, 475]. Space [923, 1752, 407, 268, 419, 341, 210, 649, 1574, 79, 1453, 1775, 31, 1123, 1140, 1210, 1745, 635, 75, 253, 802, 1316, 264, 315, 1393]. Space-charge [923, 419, 341]. space-inhomogeneous [75]. Space-Time [1752]. spaced [950]. spaces [732, 148, 1819, 938, 605, 1787]. Spacewise [604, 1298, 967]. spacewise-dependent [1298, 967]. Spatial [1731, 681, 824, 973, 1426, 1406, 1003, 1412, 1094, 1375, 551, 4, 1714, 1016, 1795, 1535, 1041]. spatially [703, 278, 221, 455, 300]. Spatio [1537, 1498]. Spatio-temporal [1537, 1498]. spatiotemporal [671, 1426]. Special [57, 196, 1329, 84, 1142, 1324, 1821, 1578, 1177, 1122, 139, 212, 1469, 1680, 1098]. species [1483, 1514, 1820]. spectra [1066, 444, 1704]. Spectral [1423, 1447, 561, 1017, 176, 1785, 589, 1775, 1449, 1526]. spectrally [1685]. spectroscopy [1743]. Spectrum [100, 1275, 1487, 1747, 1347]. speed [1768, 1321, 1420, 732, 706, 477, 970, 783]. speeds [1003, 1278, 1635, 562]. spermatozoan [1582]. sphere [529, 55, 1701, 536, 765, 1594, 1110, 412, 175]. spheres [230, 1747, 597, 323, 855, 159, 1302]. Spherical [1465, 1169, 731, 252, 652, 1674, 1252, 1384, 923, 267, 449, 971, 1049, 418, 1234, 493, 1445]. spherical-wave [731]. spheroidal [1423, 1594, 1234, 1626]. spheroidal-coordinate [1626]. Spike [618, 932, 1678]. spin [457]. spin-up [457]. Spinning [375, 1850, 1801, 1170, 669]. spiral [1737]. spleen [738]. Spline [1285]. Spline-interpolation [1285]. spot [601]. spread [926, 1257, 1683]. Spreading [1003, 970, 709]. spring [1401, 1067, 424, 1798, 1723]. spring-loaded [1401]. spring-mass [1723]. springs [1136]. square [393, 460, 1151, 1520]. square-law [460]. square-well [393]. squeeze [1372]. squeeze-film [1372]. squirmer [1583]. squirmers [1465]. squirming [1585]. Stability [799, 480, 1852, 1099, 1154, 516, 492, 1754, 617, 1705, 1096, 1044, 1832, 1859, 1514, 1050, 457, 10, 1349, 1206, 195, 803, 827, 1253, 1618, 1798, 247, 1060, 363, 1478, 312, 1730, 69, 1435, 278, 379, 1402, 252, 871, 1567, 1098, 663, 531, 929, 1226, 954, 1103, 221, 1585, 961, 1666, 968, 439, 932, 83, 873, 1421, 1542, 368, 1407, 584, 1353, 1630, 1727, 1037, 1105, 845, 786, 719, 47, 372, 1288, 556, 911, 757, 1640, 889, 763, 124, 963, 1312, 842, 903, 1633, 1610, 5, 1020, 520, 1515, 1069, 1057, 1869]. stabilization [1635, 1358, 698, 1306, 768]. Stabilizing [1504, 73]. Stable [108, 1289, 1678, 1475, 808, 92, 1570, 7]. stacked [1699]. stacks [883]. stage [1003, 676]. stages [1480]. stagnant [540]. Stagnation [939, 1390, 1374]. stagnation-point [1390]. stall [775]. standard [1375]. standby [1754]. star [987]. started [1623]. state [705, 1051, 1402, 470, 609, 646, 553, 715, 895, 256, 619, 434, 1037, 499]. statement [596, 650]. states [1731, 1869, 1571, 1852, 1763, 819, 1729, 1634, 1728, 1799, 1693, 392, 1630, 1727, 455, 1836, 525, 772, 1715]. static [1811, 1766]. Stationary [50, 1734, 1724, 1869, 1852, 1563, 423, 450, 1005, 552, 239, 1239, 1055]. Statistically [1365, 1346]. stator [844, 1300]. steadily [834]. Steady [705, 84, 470, 1489, 1713, 1693, 636, 404, 1715, 812, 1731, 1748, 1028, 553, 919, 105, 715, 1442, 256, 1847, 619, 1140, 604, 1009, 392, 616, 1630, 1037, 455, 1836, 499]. Steady-state [705, 470, 553, 715, 256, 1037, 499]. stealth [1824]. steam [1180]. steaming [169]. Steel [112, 1670, 1707, 509, 1531, 261]. Stefan [19, 746, 333, 73, 192, 279, 1530, 83, 88, 311, 47, 168, 223, 995]. Stefan-like [333]. Stefan-type [223]. stents [1225]. Step [6, 1863, 1111, 1387]. Steric [974]. Stewartson [817, 1063, 1311, 925]. sticky [1050]. stiff [9]. stiffened [62]. stiffness [1017, 1126, 1618]. Stiglitz [1376]. stirred [180, 219]. stirring [1684]. Stochastic [1147, 1814, 1783, 1862, 1400, 1565, 210, 1446, 1549, 1871, 1410, 1804, 1005, 1421, 1514, 1621, 1745, 610, 1288, 1277, 1292, 1089]. stochasticity [1224]. stock [1323]. stocking [1112]. stoichiometry [686]. Stokes [1274, 1388, 1091, 1528, 1276, 1195, 1012, 1048, 1580, 385, 1011, 1138, 1391, 1457, 675, 908, 1423, 77, 1743, 770, 1108, 171, 986, 1641, 855, 299, 462, 76, 693, 1394, 939]. stone [1682]. Stopping [519]. storage [94]. straightening [820]. strain [895, 1344, 720, 822, 613, 962, 434, 174, 1102, 1342, 339]. strained [924, 1136, 1396]. strains [958, 1337]. stratabound [745]. strategies [1553, 1160, 1846]. strategy [852, 1368]. Stratification [1834]. stratifications [520]. stratified [812, 363, 717, 72, 577, 963, 508]. stream [1262, 1694, 297]. streamline [296, 1570]. streams [874]. strength [614]. Stress [339, 613, 57, 1157, 609, 1278, 621, 825, 823, 434, 600, 1766, 9, 1466, 1339, 1526, 836, 374]. stress-dependent [1278]. stress-free [609, 374]. stress-softening [1339]. stress-strain [434]. stressed [646, 816, 345, 753, 879, 635, 723, 1313]. Stresses [214, 779]. stretch [1864]. stretch-limited [1864]. stretched [87]. stretches [24]. stretching [216, 1367, 1317]. strings [1864]. strip [1427, 943, 53, 76, 718]. strips [543]. Strong [1565, 698, 1230, 1733, 1053, 1836]. strong-property-fluctuation [1053]. strongly [705, 69, 501, 838, 391, 1515]. structural [1345]. structurally [1570]. structure [1003, 1222, 398, 945, 871, 420, 733, 1601, 997, 172, 758, 1241, 1209, 1727, 1295, 1640, 1551, 1539, 287, 1521, 326]. structured [866, 278, 221, 1600, 1525, 1794]. structures [176, 927, 1126, 1289, 1694, 1464, 1700, 1203, 1102, 373, 1439, 1162, 1193]. strut [259]. students [854]. studies [448, 243, 972]. Study [717, 806, 478, 627, 1837, 309, 1034, 1808, 1582, 1180, 745, 855, 1178, 1186, 1848, 1162, 1026]. Sturm [1774, 1449]. subclass [1054]. subcritical [1734, 316]. subcriticality [1738]. subdiffusion [1704, 1060]. subdiffusive [1719]. Subharmonic [104]. subharmonics [108]. subject [1340, 1012, 1112, 1102]. subjected [563, 826]. submerged [138, 1764]. suboptimal [668]. subsequences [177]. subsonic [722]. subspaces [1156]. substance [5]. substrate [966, 834]. substrates [1693]. subsurface [226]. Suction [1317, 948, 556, 1134]. Suction-induced [1317]. sufficiently [295]. sugar [1488]. super [528]. super-case [528]. supercell [1785]. superconductivity [516]. superconvergence [240]. Supercooling [191, 209]. supercritical [4, 1745]. superheating [83, 191, 209]. superimposed [609, 646]. superplastic [672]. superposed [531, 1722]. supersaturated [252]. support [161]. support-excited [161]. supported [1320, 966, 688]. suppressed [1836, 1715]. supremum [1153]. Surface [980, 1574, 285, 966, 1116, 1544, 193, 1227, 625, 1850, 710, 864, 312, 630, 90, 1861, 1801, 989, 1263, 1015, 910, 73, 517, 1072, 1534, 345, 206, 531, 895, 916, 1328, 1189, 1038, 474, 549, 294, 410, 913, 596, 650, 171, 1264, 1315, 1049, 198, 988, 1573, 1209, 644, 631, 574, 1119, 1742, 18, 1374, 607, 1637, 1363, 1313, 982, 1317, 755, 525, 1110, 412, 571, 1246, 1494]. surface-catalysed [982]. surface-defining [90]. surface-impedance [895]. surface-piercing [596, 650, 198]. Surface-tension-driven [980, 1116]. surfaces [503, 1364, 560, 1211, 149, 1166, 1818, 1255, 107, 1844, 1626, 1150, 1165]. surfaces-how [1818]. Surfactant [1200, 1172, 1860, 709]. surfactant-covered [1860]. Surfactant-induced [1172]. surfactants [1184]. surge [775]. surgery [738]. surges [466]. Surging [357]. surrounded [1697, 1273]. surrounding [1221]. survey [1413]. susceptibility [1539]. susceptible [1662]. susceptible-infected-susceptible [1662]. suspended [93, 137]. suspension [398, 1599, 1585, 662, 501, 1447, 758]. suspensions [1583]. SVIR [1458, 1539]. swarm [1422]. Swelling [828, 786, 903]. Swelling-induced [828]. Swift [1730]. swimmer [1384]. swimming [1585, 1582, 1465]. swing [293]. swirling [124]. switched [1044, 1381]. switching [1482, 1754, 757]. Sylvester [1549]. Symbolic [1026]. Symbolic-computation [1026]. symbolism [934]. symbols [1557, 1385]. Symmetric [424, 705, 148, 1775, 290]. symmetrical [720]. symmetries [559, 624, 274, 893, 923, 780, 463, 1123, 1156, 728, 935, 1543]. Symmetry [1434, 473, 514, 881, 581, 1075, 766, 1049, 728, 518, 660, 851]. symmetry- [581]. symmetry-breaking [514]. synchronies [1495]. Synchronization [1035, 1619, 1068, 1679]. Synchronous [844, 1010, 1300]. Synchrony [1644]. synergy [1463]. synthesis [1044]. system [771, 1459, 833, 1790, 852, 630, 1805, 1642, 1784, 333, 831, 1569, 593, 1567, 817, 1063, 665, 589, 1117, 954, 1754, 951, 551, 1096, 1620, 1866, 506, 521, 775, 858, 390, 1258, 1296, 806, 1734, 1259, 1311, 1452, 1212, 698, 1112, 1381, 1514, 1820, 1493, 441, 579, 614, 686, 1622, 996, 12, 1237, 1872, 290, 1067, 1448, 925, 803, 330, 829, 1251, 1253, 1089, 1515, 1057, 94, 440, 1045, 1455, 1433]. systematic [212, 938]. Systems [1305, 180, 219, 1733, 1619, 964, 3, 857, 1322, 1405, 1655, 1643, 798, 853, 942, 893, 14, 346, 565, 801, 1019, 59, 1564, 528, 576, 767, 1871, 1044, 258, 1369, 1417, 1348, 1464, 4, 1058, 873, 1243, 1421, 1542, 1447, 744, 808, 532, 846, 471, 610, 1698, 896, 526, 271, 851, 1206, 800, 1310, 1337, 1724, 1233, 1295, 477, 757, 648, 568, 1429, 525, 677, 1439, 1846, 606, 1548, 1633, 1610, 225, 217, 1706, 1704, 1149, 1441, 1862, 1723, 533, 670]. systems-characterization [1723]. Szydlowski [1323].

tactic [1415]. Tait [1579].

Tait-compressible [1579]. Takens [1731]. taking [1132]. tandem [650]. Tangential [1818, 1364, 1713, 1740]. tanks [1712, 466, 1183]. Tanner [1245, 1016]. tapered [1707]. target [413, 1600]. target-cell [1600]. targeting [852]. targets [1436]. Taylor [420, 370, 374]. TE [963]. technique [1842, 685, 598, 1124]. techniques [1243]. temperature [193, 1478, 269, 1356, 495, 549, 1005, 727, 880, 962, 481, 892, 1773]. temperature-induced [962]. temperature-rate-dependent [1478]. temperatures [305, 1757]. tempered [210]. Temporal [1760, 1498, 1537, 1037, 1426]. ten [119]. tensegrity [1126]. tensile [871, 1171]. tension [958, 73, 980, 1049, 1573, 1742, 1116, 1618, 1494]. tensor [895, 1818]. tensors [1703]. term [1496, 1448, 1082, 1523]. terms [1438, 819, 917, 497, 1023, 1630, 1251]. territory [1174, 1239]. testing [1553]. tethered [852]. tetrahedron [887]. th [1611]. th-order [1611]. their [1789, 1665, 839, 790, 530, 971, 1641, 1834, 1337, 1413]. theorem [791, 408, 423, 450, 467, 1470, 1413, 310]. theorems [29, 106, 873, 313, 338, 75, 523, 1626, 343].

Theoretical [928, 1528, 1468, 21, 161, 784, 1113, 1445]. theories [1416]. Theory [892, 1176, 964, 194, 480, 160, 888, 1091, 1653, 548, 204, 379, 489, 208, 66, 1048, 626, 1053, 1030, 1747, 151, 1375, 319, 507, 49, 362, 1146, 534, 241, 202, 1330, 296, 651, 515, 154, 1587, 1698, 985, 786, 1216, 1102, 911, 45, 123, 124, 237, 334, 1343, 957, 184, 275, 1001, 1667]. therapy [1796, 1600]. Thermal [362, 32, 743, 225, 217, 156, 180, 219, 746, 503, 305, 590, 987, 747, 319, 1167, 1623, 975, 468, 511, 756, 39, 435, 661, 403]. thermistor [416, 269]. Thermo [1629]. Thermo-convective [1629]. thermoacoustic [1608]. thermocapillary [1710]. Thermodynamic [61]. thermoelastic [1073, 647, 1117, 898, 1092, 1107, 1473, 894, 960, 1616]. Thermoelasticity [983, 1478, 726, 897, 766, 1043, 584, 494, 667, 892]. thermomechanical [647, 1707]. thermophoresis [225, 217]. Thermoviscoelastic [1038, 861, 628]. thermoviscoelasticity [845]. thick [267]. thick-walled [267]. thickness [1040, 1079, 826]. thin [1261, 625, 1606, 1157, 714, 759, 979, 1404, 639, 603, 974, 1079, 1770, 1732, 153, 1700, 1693, 1109, 943, 421, 162, 405, 1492, 966, 1238, 1268, 1008, 1870, 1101, 743, 709]. thin-film [1606, 979, 966]. thin-film/substrate [966]. thin-walled [1870]. thinning [1185]. third [1140]. thirty [113]. Thixotropy [1466]. thread [1273, 1119]. threads [867]. Three [675, 1189, 1081, 915, 566, 1266, 1326, 196, 304, 1154, 798, 1802, 1808, 1746, 742, 1514, 765, 792, 1493, 1241, 1095, 1872, 556, 814, 512, 1709, 1706]. three-component [1493, 1709]. three-coupled [1154]. three-degree-of-freedom [1872]. Three-dimensional [675, 1081, 1266, 196, 304, 798, 1802, 1808, 742, 765, 1095, 556, 814]. three-phase [792, 512]. Three-point [915, 566]. three-species [1514]. three-vortex [1746]. three-wave [1706]. Threshold [1525, 882, 1765, 260, 1795]. thrombin [1500]. throughflow [961]. thrust [1420]. Thwaites [1680, 1324, 1504]. tidal [426]. tilting [1681]. Time [540, 1176, 787, 1391, 1752, 1431, 1675, 416, 1459, 1619, 710, 22, 1810, 1703, 1797, 1643, 1516, 453, 302, 59, 1699, 1858, 1298, 1299, 1849, 1355, 1646, 1719, 1094, 1775, 951, 67, 1088, 1174, 1378, 1123, 277, 1287, 1635, 1291, 854, 1210, 1462, 1111, 1529, 1750, 1381, 1514, 1865, 808, 1745, 1347, 1812, 1209, 1562, 1349, 195, 1082, 1415, 829, 1268, 1351, 1134, 1251, 1294, 1323, 508, 1175, 1406, 1393, 1441, 1845, 1499]. time-delay [1441]. time-delayed [1810, 1294, 1406]. Time-dependent [540, 787, 416, 1176, 1703, 1699, 1298, 1646]. Time-harmonic [1675, 710, 302, 1299, 1378, 277, 508]. time-periodic [1516, 1349]. time-varying [1415]. time-wise [22]. times [392]. Timoshenko [774, 1635, 1141, 751, 1033]. tip [1302]. tip-vortex [1302]. tipping [1124]. tissue [738, 1336, 1793, 1350, 1337]. tissues [1104]. titanium [794]. Tollmien [500, 582]. Tollmien-Schlichting [500, 582]. Tollmien-Schlichting/vortex [500]. tomography [1768, 1360, 928, 1687, 1830, 1817]. Tool [199, 1468]. Topic [1462]. topographically [1022]. topography [1263, 645]. Topological [1377, 1785, 1853, 1738]. topologies [1570]. tori [770]. toroidal [724]. Torsion [1335, 155, 1002, 955, 401]. Torsional [1599, 1740, 1559, 1846]. torus [204, 1451, 1114]. Total [363, 754]. touchdown [1430]. towers [105]. trace [1487]. tracer [1583]. Tracers [303]. trachomatis [1186]. track [1380]. tract [1266]. traction [1208]. traditional [1848]. traffic [1284, 1772, 730]. train [480, 840]. trains [706, 1407, 1254]. trajectories [1195, 1587]. trajectory [1607]. transaction [1400]. transcendental [197]. transducer [1627, 1113, 1445, 1546]. transducers [1432, 1021]. transfer [1182, 408, 236, 445, 39, 1367, 830, 175, 93, 137]. transform [1091, 1443, 626, 166, 764, 782, 848, 1780, 1842, 1597, 1595, 999]. transform-based [1842]. transformation [131, 167, 1048, 1811]. Transformations [859, 583, 697, 962, 518]. transforming [1171]. transforms [791, 30, 262, 310, 1854]. Transient [992, 230, 769, 1559, 1356, 824, 253, 53]. transients [706]. transition [1226, 898, 392, 477, 1527]. transitional [1203]. transitions [1763, 1683, 1749, 509, 724, 191, 261]. translation [107]. Transmission [922, 766, 1275, 1073, 1404, 1791, 716, 1769, 1377, 233, 1541, 1043, 768, 1652, 270, 1238, 43, 1803, 1524]. transonic [876]. transparency [138]. transparent [1109, 93, 137]. transport [111, 210, 1583, 21, 923, 623, 745, 75, 1660, 1361, 1161, 399, 482]. transportation [94]. transported [1222]. transverse [599, 716]. transversely [1453, 77, 889, 1341]. Trapped [422, 1022, 1615, 1015, 1403, 549]. trapping [804]. Traveling [1636, 1799, 1480]. Travelling [1412, 593, 1197, 1094, 1369, 1417, 1058, 1259, 1212, 425, 1059, 1294, 1149, 328, 516, 1003, 817, 472, 1031, 1070, 744, 579, 846, 496, 985, 632, 380, 1013, 1297, 677, 970, 1253, 1610, 533]. travelling-wave [1031]. trawl [658]. treadmilling [1384]. treatment [90, 639, 1182, 685]. Tree [1570, 997, 1530]. trees [1, 37]. tri [1485]. tri-phase [1485]. triality [642]. triangle [878]. triangular [1522]. triangularizable [757]. tridiagonal [1631]. tridimensional [1194]. Triple [149]. Triple-deck [149]. trivial [1504, 1609]. Truncated [788]. truncation [1785]. truncations [29]. Trust [1823]. tube [1338, 603, 818, 1840, 594, 536, 1234, 982, 1870, 718, 1444]. tubes [155, 1087, 1182, 1086, 1105]. tumble [1584]. tumbling [1584]. tumour [1796, 1719]. tunnel [706, 840]. tunnelling [393, 289, 329]. turbulence [117]. turbulent [144, 259, 399]. Turing [1734, 1245, 632, 1704]. Turning [42, 1422, 402]. TV [1716]. TV-regularized [1716]. twisted [1618]. Two [29, 485, 1641, 1090, 1135, 2, 1267, 381, 709, 1232, 1483, 746, 176, 672, 1516, 1412, 676, 50, 492, 458, 347, 79, 531, 151, 700, 643, 1572, 916, 1705, 811, 323, 1006, 498, 1200, 1325, 1530, 596, 650, 316, 83, 454, 680, 1068, 1220, 653, 1214, 1820, 792, 129, 795, 1230, 198, 1254, 694, 990, 1204, 723, 462, 1547, 159, 1428, 719, 47, 1178, 911, 1114, 443, 915, 13, 297, 237, 27, 1624, 605, 520, 1756, 892, 1026]. two- [792]. two-component [1214]. Two-dimensional [1641, 1090, 2, 1267, 709, 176, 1412, 458, 347, 151, 700, 1572, 498, 596, 650, 316, 454, 653, 129, 694, 723, 462, 719, 1114, 27]. two-fluid [1006, 1200, 911]. two-layer [795, 990, 1204, 443, 520]. two-layered [1756]. two-phase [1232, 746, 492, 1325, 83, 1547, 47, 13]. two-point [915, 237]. two-scale [1530, 1230]. two-singular [1026]. two-species [1483, 1820]. two-stage [676]. two-temperature-generalized [892]. two-variable [1068, 1254]. Twofold [1319]. type [791, 1619, 888, 1073, 1247, 942, 1028, 1668, 1670, 951, 1476, 1549, 906, 1031, 1417, 790, 1027, 879, 1396, 1425, 873, 1092, 1107, 1043, 1332, 935, 896, 1100, 802, 575, 1313, 477, 1286, 1089, 885, 583, 223, 310]. types [490, 761].

ultrasonic

[1432, 1627, 853, 1113, 1445, 1546]. ultrasound [1021]. unbounded [1791, 1358]. uncertain [774, 1288]. unconstrained [726]. undamped [1567]. underlying [1819]. undulations [1650]. unexpected [1336]. unfaithful [854]. unfolded [945]. unfolding [1229]. Uniaxial [643, 1101]. uniaxially [1129]. unidirectional [1660, 81]. unified [90, 1829, 1780, 1162]. Uniform [609, 646, 1563, 203, 381, 1262, 1390, 1513, 1007, 841, 330]. uniformity [779]. uniformly [211]. unifying [1733]. unilateral [870, 1136]. unique [1519, 1386, 1009]. Uniqueness [364, 504, 106, 1634, 343, 1803, 1686, 1459, 886, 408, 467, 741, 172, 1108, 971, 776, 469, 290, 876, 1376, 1610, 1089, 499, 1603]. unit [1208]. universality [1829]. unknown [917, 1002, 1080, 644]. unloading [1339]. unreliable [1754]. unsteadily [597]. Unsteady [1242, 116, 18, 35, 622, 1509, 385, 700, 1743, 1694, 604, 544, 684, 915]. upon [945, 1766, 404]. Upper [304, 121, 929, 340, 240, 1328, 582, 991]. Upper-branch [304, 582]. Upscaling [1485]. upstream [1213, 1401, 443]. urban [1376]. urban-rural [1376]. urinary [1266]. use [1807, 30, 57, 712, 1336, 633, 1021, 1786, 1783, 466]. Using [1131, 1605, 1555, 625, 710, 979, 1071, 268, 1474, 347, 919, 545, 1344, 1014, 1504, 1871, 95, 413, 296, 1700, 1594, 143, 1800, 1698, 1284, 60, 607, 1181, 386]. Utilization [934].

vaccination [1077, 1133, 1539]. Vacuum [672]. Vajda [23]. validated [1707]. valuation [611]. value [701, 1815, 611, 1121, 1028, 599, 639, 465, 1789, 780, 82, 497, 863, 1088, 1386, 1437, 1842, 938, 1070, 1111, 1290, 1529, 536, 396, 547, 943, 748, 1806, 1233, 865, 291, 163, 382, 275, 147, 1709]. valve [1601, 1401, 1766]. Vandermonde [1595]. vanishingly [24]. vapour [1020]. variable [1276, 893, 331, 1705, 906, 1489, 1174, 1191, 1314, 32, 1529, 1068, 1239, 1254, 235, 324, 518, 1367, 1772, 562, 298, 842, 156]. variable-coefficient [331, 1191]. variable-order [1276]. variable-territory [1174, 1239]. variables [712, 1210, 1265, 1626, 17, 89, 145]. Variational [11, 148, 639, 696, 179, 128, 110, 1823, 642, 340, 415, 967, 1093, 218, 1510, 70, 1825]. Variational-iterative [148]. varies [1604]. various [201]. varying [193, 1741, 1040, 1712, 1799, 1022, 1424, 1540, 1415, 283, 1593, 834]. Vector [797, 721, 1655, 1522, 1028, 1560, 1175]. vegetation [1758]. velocities [919, 298]. velocity [1028, 1442, 823, 311, 1374, 270]. ventricle [1344]. version [891]. versus [1097, 1102]. vertex [595]. vertical [1215, 1257, 321, 198, 988, 641, 158, 382, 1207]. vertically [825, 1357]. vertically-inhomogeneous [1357]. very [1246]. Vesicle [1270]. via [1703, 1320, 1802, 517, 1780, 702, 1662, 1547, 1679]. Vibration [688, 538, 569, 1872, 1033]. vibrational [853]. vibrations [1136, 161, 1846]. vicinity [434]. view [626]. viewed [983]. virial [501]. virus [1094]. visco [1231]. visco-elastic [1231]. Viscoelastic [768, 1697, 1561, 870, 339, 712, 1317, 1286, 287]. viscoelasticity [1220, 1100, 1466]. viscoplastic [1354]. viscoplasticity [394]. viscosity [1705, 1119, 1367, 520]. Viscous [1486, 1760, 304, 1273, 325, 1065, 1242, 1189, 1832, 1291, 1623, 1816, 1573, 1119, 1428, 1860, 81, 1632, 1207]. visions [1142]. Vlasov [1000, 1871]. void [223]. voids [1762]. Voight [1274]. volatility [1400, 1364, 1482]. Voltage [1066, 844, 143, 1347, 327]. voltage-fed [844]. Volterra [164, 888, 954, 1410, 1056, 1421]. Volume [23, 760]. vortex [500, 1194, 1748, 1534, 1746, 1562, 1575, 693, 1302]. vortices [1598, 1194, 420, 58, 380]. vorticity [612, 1219, 557]. vote [1576]. Vries [1111, 933, 1290, 1529, 396, 554]. vs [1225].

W. [23]. Waals [1281, 1481, 1544]. Wake [278]. walk [1355]. walks [1647]. wall [870, 216, 1213, 1384, 77, 294, 353, 24, 574, 1374, 1615, 478, 124]. wall-modes [124]. walled [267, 1870, 1444]. walls [502, 427, 1109]. wanted [1351]. Washington [1783]. wastewater [737]. Water [927, 916, 1272, 16, 1183, 79, 984, 1572, 1826, 153, 1199, 321, 134, 198, 224, 1411, 158, 53, 78, 807, 52, 557, 382, 1246, 1267, 1669, 275]. water-wave [153, 134]. water-waves [158]. watershed [655]. Wave [1478, 1040, 868, 122, 902, 1235, 16, 866, 364, 731, 1636, 1078, 1263, 1273, 1154, 1653, 328, 1837, 1454, 804, 927, 652, 1280, 1412, 1668, 603, 1278, 817, 1283, 1197, 502, 916, 472, 1031, 1743, 172, 936, 840, 1369, 1417, 31, 153, 439, 1635, 753, 1272, 1108, 202, 944, 1093, 1058, 1259, 1612, 1212, 1109, 129, 744, 134, 1407, 1254, 1241, 776, 985, 1867, 1658, 937, 543, 1722, 505, 361, 983, 911, 1135, 1013, 78, 300, 1470, 1351, 970, 1246, 783, 1294, 1857, 356, 1706, 1385, 1149, 1480, 1667]. wave-bearing [944, 543]. wave-body [776]. Wave-Energy [16]. wave-free [1263]. Wave-front [902]. wave-fronts [866, 985]. wave-like [1454]. wave-maker [1470]. waveform [68, 1865]. wavefront [562]. Wavefronts [998, 813, 1070, 896]. waveguide [717, 1788, 510, 920]. waveguides [1007, 1403, 289, 1615]. Wavelength [856, 590, 1019]. waveless [596]. Wavelet [1211, 668]. wavelets [1071]. wavenumber [476]. Waves [315, 363, 1176, 625, 1218, 182, 1861, 849, 1154, 1805, 548, 516, 1569, 1003, 302, 428, 1195, 609, 646, 871, 593, 1819, 1574, 816, 1534, 1106, 345, 726, 531, 664, 138, 1086, 1094, 1572, 1198, 1533, 899, 467, 1737, 215, 1378, 1171, 1666, 1038, 913, 1145, 1859, 1762, 595, 1212, 1609, 322, 425, 171, 1315, 750, 1357, 321, 515, 198, 224, 441, 532, 579, 686, 846, 1059, 471, 369, 582, 1353, 496, 1411, 600, 635, 972, 1204, 371, 1652, 1206, 632, 894, 380, 1313, 1179, 158, 477, 838, 977, 1297, 805, 966, 1860, 1470, 1725, 807, 557, 382]. waves [1341, 687, 1267, 1669, 655, 126, 275, 1657, 1207, 508, 1706, 533]. wavetrain [1547]. wavy [149]. Weak [599, 1871, 1506, 1308, 341, 590, 1650, 1117, 1634, 1481, 1023, 1136, 1493, 1141]. Weakly [312, 924, 1201, 1483, 901, 1585, 648, 805, 184, 1692]. weakly-squirming [1585]. Wear [103, 1616]. weathering [1856, 1851]. Wedge [761, 1855, 1481, 619, 725, 1839, 595, 1671]. Weierstrass [1332]. weight [655]. weight-function [655]. weighted [1826, 1787]. weld [1757]. welding [959]. Well [1438, 951, 792, 180, 219, 393, 839, 1765, 421, 482]. well-posed [839]. Well-posedness [1438, 951, 792, 421]. well-stirred [180, 219]. wet [507]. wetting [1164, 834]. Wheatley [1766]. which [3, 439, 755, 270]. while [1604]. whispering [1741]. white [1419, 1745]. Whittaker [664]. whose [324]. widely [1624]. widths [1772]. Wiener [127, 234, 527, 704, 790, 118, 1427, 1216]. Wiley [23]. Will [1463, 1143]. winding [211]. window [840]. wing [722]. wings [577]. Winkler [1770]. winners [1324]. wire [421]. wireless [1702]. wise [22]. within [39, 181, 1317, 1486]. without [1785, 1237, 1738]. WKB [826, 173]. Woods [1726]. work [394]. wrinkling [523].

years [1461, 119, 113]. Yield [560, 1810, 1463, 859, 1466].

zero [1668, 1502, 1219, 356]. zerodissipation [356]. Zeros [664, 971, 324]. zeta [1551]. Zgrablich [131, 167]. ZND [322]. zone [1226, 286, 1590]. zones [974].

References

Rands:1981:ATR

[1] B. M. I. Rands and D. J. A. Welsh. Animals, trees and renewal sequences. IMA Journal of Applied Mathematics, 27(1):1–17, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See corrigendum [37].

Sleeman:1981:TDI

[2] B. D. Sleeman. Two-dimensional inverse scattering and conformal mappings. IMA Journal of Applied Mathematics, 27(1):19–31, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arscott:1981:CCS

[3] F. M. Arscott and A. Darai. Curvilinear coordinate systems in which the Helmholtz equation separates. IMA Journal of Applied Mathematics, 27 (1):33–70, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1981:SDS

[4] A. A. Lacey. The spatial dependence of supercritical reacting systems. IMA Journal of Applied Mathematics, 27(1):71–84, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wollkind:1981:NSA

[5] David J. Wollkind and Ronald D. Notestine. A nonlinear stability analysis of the solidification of a pure substance. IMA Journal of Applied Mathematics, 27(1):85–104, ???? 1981.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Nilsson:1981:PSCa

[6] Börje Nilsson and Olle Brander. The propagation of sound in cylindrical ducts with mean flow and bulk-reacting lining. 3. step discontinuities. IMA Journal of Applied Mathematics, 27 (1):105–131, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stewartson:1981:MSI

[7] K. Stewartson. Marginally stable inviscid flows with critical layers. IMA Journal of Applied Mathematics, 27 (2):133–175, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1981:SRD

[8] J. M. Hill. On the solution of reactiondiffusion equations. IMA Journal of Applied Mathematics, 27(2):177–194, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pipkin:1981:FPS

[9] A. C. Pipkin. Finite plane stress of stiff fibre-reinforced sheets. IMA Journal of Applied Mathematics, 27 (2):195–209, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rasmussen:1981:NLS

[10] H. Rasmussen and J. A. McGeough. On the non-linear stability of the electroforming process. IMA Journal of Applied Mathematics, 27(2):211–220, ????

1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Atanackovic:1981:VPC

[11] T. M. Atanacković. Variational principles for column buckling. IMA Journal of Applied Mathematics, 27 (2):221–228, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pagan:1981:BSF

[12] G. Pagan and P. J. Richards. A bound for solutions of a fourth-order dynamical system. IMA Journal of Applied Mathematics, 27(2):229–235, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sobey:1981:ESD

[13] I. J. Sobey. The effect of small dispersion on two-phase operations. IMA Journal of Applied Mathematics, 27 (2):237–249, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Clarke:1981:EAD

[14] B. M. N. Clarke. Eigenvalue assignment by delayed boundary feedback for a class of linear hyperbolic systems. IMA Journal of Applied Mathematics, 27(2):251–262, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Nilsson:1981:PSCb

[15] Börje Nilsson and Olle Brander. The propagation of sound in cylindrical ducts with mean flow and bulk-reacting lining. 4. several interacting discontinuities. IMA Journal of Applied Mathematics, 27(3):263–289, ???? 1981.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Srokosz:1981:NAO

[16] M. A. Srokosz. A note on arrays of oscillating water column wave-energy devices. IMA Journal of Applied Mathematics, 27(3):291–297, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Srivastava:1981:SPE

[17] H. M. Srivastava. Some polynomial expansions for functions of several variables. IMA Journal of Applied Mathematics, 27(3):299–306, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rasmussen:1981:UPF

[18] H. Rasmussen and D. Salhani. Unsteady porous flow with a free surface. IMA Journal of Applied Mathematics, 27(3):307–319, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bossavit:1981:HSP

[19] A. Bossavit and A. Damlamian. Homogenization of the Stefan problem and application to magnetic composite media. IMA Journal of Applied Mathematics, 27(3):319–334, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Baumel:1981:AFH

[20] R. T. Baumel, J. L. Gammel, and J. Nuttall. Asymptotic form of Hermite–Padé polynomials. IMA Journal of Applied Mathematics, 27(3):335– 357, ???? 1981. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic).

Fisher:1981:TMP

[21] B. E. A. Fisher and J. F. Macqueen. A theoretical model for particulate transport from an elevated source in the atmosphere. IMA Journal of Applied Mathematics, 27(3):359–371, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arkeryd:1981:TWA

[22] L. Arkeryd. A time-wise approximated Boltzmann equation. IMA Journal of Applied Mathematics, 27 (3):373–383, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Morris:1981:BRH

[23] A. O. Morris. Book review: Handbook of Applicable Mathematics. Volume 1. Algebra, edited by W. Ledermann and S. Vajda, Wiley, 1990. IMA Journal of Applied Mathematics, 27(3):385–386, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kuiken:1981:BLF

[24] H. K. Kuiken. On boundary layers in fluid mechanics that decay algebraically along stretches of wall that are not vanishingly small. IMA Journal of Applied Mathematics, 27 (4):387–405, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gladwell:1981:IPE

[25] G. M. L. Gladwell and S. Coen. An inverse problem in elastostatics. IMA Journal of Applied Mathematics, 27 (4):407–421, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mansfield:1981:CCR

[26] Anthony Mansfield. On the computational complexity of a rigidity problem. IMA Journal of Applied Mathematics, 27(4):423–429, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sundelius:1981:TDC

[27] B. Sundelius. A two-dimensional contact problem with given region of adhesion. IMA Journal of Applied Mathematics, 27(4):431–454, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sundelius:1981:ACP

[28] B. Sundelius. An axisymmetric contact problem with given region of adhesion. IMA Journal of Applied Mathematics, 27(4):455–475, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Abian:1981:TTT

[29] Alexander Abian. Two theorems on truncations of the Laurent series. IMA Journal of Applied Mathematics, 27 (4):477–479, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Apelblat:1981:RUI

[30] Alexander Apelblat. Repeating use of integral transforms—a new method for evaluation of some infinite integrals. IMA Journal of Applied Mathematics, 27(4):481–496, ???? 1981. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hudson:1982:DAW

[31] J. A. Hudson. The diffraction of an acoustic wave by an embedded quarter-space. IMA Journal of Applied Mathematics, 28(1):1–21, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1982:TIV

[32] A. A. Lacey and G. C. Wake. Thermal ignition with variable thermal conductivity. IMA Journal of Applied Mathematics, 28(1):23–39, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Fowler:1982:AAD

[33] A. C. Fowler. An asymptotic analysis of the delayed logistic equation when the delay is large. IMA Journal of Applied Mathematics, 28(1):41–49, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1982:NGM

[34] D. S. Jones. A note on Galerkin’s method. IMA Journal of Applied Mathematics, 28(1):51–73, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tayler:1982:USF

[35] A. B. Tayler and M. O. Nicholas. Unsteady slow flows over a cooled flat plate. IMA Journal of Applied Mathematics, 28(1):75–91, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Fletcher:1982:BFE

[36] R. Fletcher and P. J. Harley. Basis functions for the exact control of the heat equation. IMA Journal of Applied Mathematics, 28(1):93–105, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rands:1982:CAT

[37] B. M. I. Rands and D. J. A. Welsh. Corrigendum: “Animals, trees and renewal sequences” [IMA J. Appl. Math. 27 (1981), no. 1, 1–17; MR 82j:05049]. IMA Journal of Applied Mathematics, 28(1):107, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [1].

Budden:1982:SBN

[38] Peter Budden and John Norbury. Solution branches for nonlinear equilibrium problems—bifurcation and domain perturbations. IMA Journal of Applied Mathematics, 28(2):109–129, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rubinstein:1982:HTF

[39] L. Rubinstein, H. Geiman, and M. Shachaf. Heat transfer with a free boundary moving within a concentrated thermal capacity. IMA Journal of Applied Mathematics, 28(2):131– 147, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1982:SSN

[40] Ronald Smith. Similarity solutions of a nonlinear diffusion equation. IMA Journal of Applied Mathematics, 28 (2):149–160, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1982:AOP

[41] Ronald Smith. An abundance of orthogonal polynomials. IMA Journal of Applied Mathematics, 28(2):161–167, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lewis:1982:TPP

[42] Gilbert N. Lewis. Turning point problems and resonance. IMA Journal of Applied Mathematics, 28(2):169–183, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1982:RDT

[43] W. E. Williams and A. Chakrabarti. Reflection at a discontinuity in a transmission line. IMA Journal of Applied Mathematics, 28(2):185–195, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Nonweiler:1982:NAF

[44] Terence R. F. Nonweiler. A nongraphical approach to flatness calibration. IMA Journal of Applied Mathematics, 28(2):197–206, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1982:HRN

[45] F. T. Smith. On the high Reynolds number theory of laminar flows. IMA Journal of Applied Mathematics, 28 (3):207–281, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Duggan:1982:CRP

[46] R. C. Duggan and P. J. Godwin. On the capacity of regular polyhedra. IMA Journal of Applied Mathematics, 28 (3):283–286, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rubinstein:1982:GSN

[47] L. Rubinstein. Global stability of the Neumann solution of the two-phase Stefan problem. IMA Journal of Applied Mathematics, 28(3):287–299, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Please:1982:ASJ

[48] C. P. Please. An analysis of semiconductor P − N-junctions. IMA Journal of Applied Mathematics, 28 (3):301–318, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Green:1982:TLC

[49] A. E. Green and P. M. Naghdi. A theory of laminated composite plates. IMA Journal of Applied Mathematics, 29(1):1–23, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cooke:1982:SPT

[50] J. C. Cooke. Stationary phase in two dimensions. IMA Journal of Applied Mathematics, 29(1):25–37, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lee:1982:NSR

[51] Alexander I. Lee and James M. Hill. A note on the solution of reactiondiffusion equations with convection. IMA Journal of Applied Mathematics, 29(1):39–43, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tuck:1982:AWJ

[52] E. O. Tuck. Annular water jets. IMA Journal of Applied Mathematics, 29(1):45–58, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1982:TMP

[53] C. M. Smith. The transient motion of a partially immersed rolling strip in water. IMA Journal of Applied Mathematics, 29(1):59–77, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Abrahams:1982:SSL

[54] I. D. Abrahams. Scattering of sound by large finite geometries. IMA Journal of Applied Mathematics, 29(1):79–97, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Davis:1982:MBP

[55] Gregory B. Davis and James M. Hill. A moving boundary problem for the sphere. IMA Journal of Applied Mathematics, 29(1):99–111, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sleeman:1982:IPA

[56] B. D. Sleeman. The inverse problem of acoustic scattering. IMA Journal of Applied Mathematics, 29 (2):113–142, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Atkinson:1982:SAS

[57] C. Atkinson and J. Avila. Special analytical solutions for use in debond stress analysis. IMA Journal of Applied Mathematics, 29(2):143–171, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hall:1982:NEG

[58] P. Hall. On the nonlinear evolution of Görtler vortices in nonparallel boundary layers. IMA Journal of Applied Mathematics, 29(2):173–196, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dassios:1982:FTE

[59] George Dassios. Finite time equipartition for second-order hyperbolic systems. IMA Journal of Applied Mathematics, 29(2):197–202, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mullineux:1982:ASU

[60] G. Mullineux. Approximating shapes using parameterized curves. IMA Journal of Applied Mathematics, 29 (2):203–220, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Woods:1982:TIC

[61] L. C. Woods. Thermodynamic inequalities in continuum mechanics. IMA Journal of Applied Mathematics, 29 (3):221–246, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Eatwell:1982:EFL

[62] G. P. Eatwell and J. R. Willis. The excitation of a fluid-loaded plate stiffened by a semi-infinite array of beams. IMA Journal of Applied Mathematics, 29(3):247–270, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Barnes:1982:SRR

[63] David C. Barnes. Some remarks on a result of Atanacković. IMA Journal of Applied Mathematics, 29 (3):271–273, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Srivastava:1982:EPB

[64] H. M. Srivastava. An elementary proof of Bailey’s bilinear generating function for Jacobi polynomials and of its q-analogue. IMA Journal of Applied Mathematics, 29(3):275–280, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cade:1982:CPR

[65] R. Cade. A class of physically relevant insoluble Dirichlet problems. IMA Journal of Applied Mathematics, 29 (3):281–286, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chapman:1982:IDL

[66] M. J. Chapman and A. J. Pritchard. Infinite-dimensional linear realization theory. IMA Journal of Applied Mathematics, 29(3):287–308, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grundy:1982:STE

[67] R. E. Grundy. The small-time evolution of a nonlinear diffusion equation. IMA Journal of Applied Mathematics, 29(3):309–319, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Fincham:1982:FPS

[68] W. F. Fincham and C. L. Nikias. Further periodic solutions of the van der Pol equations and its application to waveform generation. IMA Journal of Applied Mathematics, 29 (3):321–332, ???? 1982. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bodonyi:1983:ADS

[69] R. J. Bodonyi, F. T. Smith, and J. Gajjar. Amplitude-dependent stability of boundary-layer flow with a strongly nonlinear critical layer. IMA Journal of Applied Mathematics, 30(1):1–19, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Thangaraj:1983:DVP

[70] D. Thangaraj and S. N. Venkatarangan. Dual variational principles for a class of nonlinear partial differential equations. IMA Journal of Applied Mathematics, 30(1):21–26, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Millar:1983:CSS

[71] Robert F. Millar. On the completeness of sets of solutions to the Helmholtz equation. IMA Journal of Applied Mathematics, 30(1):27–38, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dore:1983:FOI

[72] B. D. Dore. Forced oscillations of an inclined plane in an exponentiallystratified fluid. IMA Journal of Applied Mathematics, 30(1):39–55, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chadam:1983:SES

[73] J. Chadam and P. Ortoleva. The stabilizing effect of surface tension on the development of the free boundary in a planar, one-dimensional, Cauchy-Stefan problem. IMA Journal of Applied Mathematics, 30(1):57–66, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Alexiades:1983:RFD

[74] Vasilios Alexiades. Rapid freezing of dilute alloys. IMA Journal of Applied Mathematics, 30(1):67–79, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pettersson:1983:ETL

[75] Rolf Pettersson. Existence theorems for the linear, space-inhomogeneous transport equation. IMA Journal of Applied Mathematics, 30(1):81–105, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Spence:1983:CBE

[76] D. A. Spence. A class of biharmonic end-strip problems arising in elasticity and Stokes flow. IMA Journal of Applied Mathematics, 30(2):107–139, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Harper:1983:SFB

[77] J. F. Harper and G. C. Wake. Stokes flow between parallel plates due to a transversely moving end wall. IMA Journal of Applied Mathematics, 30 (2):141–149, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Srokosz:1983:SMC

[78] M. A. Srokosz. A simplified method for calculating the highest wave on deep water. IMA Journal of Applied Mathematics, 30(2):151–159, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Edwards:1983:SOB

[79] N. A. Edwards, C. P. Please, and R. W. Preston. Some observations on boundary conditions for the shallowwater equations in two space dimensions. IMA Journal of Applied Mathematics, 30(2):161–172, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Crowley:1983:MMH

[80] A. B. Crowley. Mathematical modelling of heat flow in Czochralski crystal pulling. IMA Journal of Applied Mathematics, 30(2):173–189, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tuck:1983:FBP

[81] E. O. Tuck, M. Bentwich, and J. van der Hoek. The free-boundary problem for gravity-driven unidirectional viscous flows. IMA Journal of Applied Mathematics, 30(2):191–208, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grundy:1983:LSS

[82] R. E. Grundy. Local similarity solutions for the initial value problem in nonlinear diffusion. IMA Journal of Applied Mathematics, 30(2):209–214, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1983:ESR

[83] A. A. Lacey and M. Shillor. The existence and stability of regions with superheating in the classical two-phase one-dimensional Stefan problem with heat sources. IMA Journal of Applied Mathematics, 30(2):215–230, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Caulk:1983:ASH

[84] D. A. Caulk. Analysis of steady heat-conduction in regions with circular holes by a special boundary-integral method. IMA Journal of Applied Mathematics, 30(3):231–246, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Haidar:1983:OAM

[85] Nassar H. S. Haidar. Operational analysis of multiregional nuclear reactor kinetics. IMA Journal of Applied Mathematics, 30(3):247–267, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Aitchison:1983:PGS

[86] J. M. Aitchison, C. M. Elliott, and J. R. Ockendon. Percolation in gently sloping beaches. IMA Journal of Applied Mathematics, 30(3):269–287, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Moodie:1983:IPS

[87] T. Bryant Moodie and D. W. Barclay. Instantaneous plugging of stretched elastic plates. IMA Journal of Applied Mathematics, 30(3):289–301, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1983:MRS

[88] A. A. Lacey and A. B. Tayler. A mushy region in a Stefan problem. IMA Journal of Applied Mathematics, 30(3):303–313, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Srivastava:1983:CPE

[89] H. M. Srivastava. Certain q-polynomial expansions for functions of several variables. IMA Journal of Applied Mathematics, 30(3):315–323, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bez:1983:UTI

[90] H. E. Bez. A unified treatment of invariance for curve- and surface-defining algorithms. IMA Journal of Applied Mathematics, 30(3):325–343, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Moore:1983:RID

[91] D. W. Moore. Resonances introduced by discretization. IMA Journal of Applied Mathematics, 31(1):1–11, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Roberts:1983:SAN

[92] A. J. Roberts. A stable and accurate numerical method to calculate the motion of a sharp interface between fluids. IMA Journal of Applied Mathematics, 31(1):13–35, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1983:RHT

[93] M. M. R. Williams. Radiant heat transfer through an aerosol suspended in a transparent gas. IMA Journal of Applied Mathematics, 31(1):37–50, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See addendum [137].

Zemanian:1983:DMS

[94] A. H. Zemanian. A dynamic marketing, storage and transportation system with perfect competition in each of its markets. IMA Journal of Applied Mathematics, 31(1):51–78, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kleinman:1983:CNI

[95] R. E. Kleinman and R. Kress. On the condition number of integral equations in acoustics using modified fundamental solutions. IMA Journal of Applied Mathematics, 31(1):79–90, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hayman:1983:CMP

[96] W. K. Hayman. A conformal mapping problem arising in elasticity. II. IMA Journal of Applied Mathematics, 31(2):91–111, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1983:IMF

[97] A. A. Lacey. Initial motion of the free boundary for a nonlinear diffusion equation. IMA Journal of Applied Mathematics, 31(2):113–119, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grundy:1983:ASM

[98] R. E. Grundy. Asymptotic solution of a model nonlinear convective diffusion equation. IMA Journal of Applied Mathematics, 31(2):121–137, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grunbaum:1983:OPS

[99] F. Alberto Grünbaum. Orthogonal polynomials and second-order differential equations. IMA Journal of Applied Mathematics, 31(2):139–144, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Veling:1983:SED

[100] E. J. M. Veling. Spectrum and eigenfunctions of a differential operator arising by linearization of the Fisher and related equations. IMA Journal of Applied Mathematics, 31(2):145–160, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Fang:1983:SAI

[101] S. C. Fang. A sequential algorithm for an inventory selection problem. IMA Journal of Applied Mathematics, 31 (2):161–168, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Atkinson:1983:ASS

[102] C. Atkinson and F. G. Leppington. The asymptotic solution of some integral equations. IMA Journal of Applied Mathematics, 31(3):169–182, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Parker:1983:MMP

[103] D. F. Parker. A mathematical model for the prediction of shape wear during electrodischarge machining. IMA Journal of Applied Mathematics, 31 (3):183–205, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Thompson:1983:SRC

[104] J. M. T. Thompson, A. R. Bokaian, and R. Ghaffari. Subharmonic resonances and chaotic motions of a bilinear oscillator. IMA Journal of Applied Mathematics, 31(3):207–234, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Elvey:1983:EDM

[105] J. S. N. Elvey. On the elimination of destabilizing motions of articulated mooring towers under steady sea conditions. IMA Journal of Applied Mathematics, 31(3):235–252, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Colton:1983:UTI

[106] David Colton and B. D. Sleeman. Uniqueness theorems for the inverse problem of acoustic scattering. IMA Journal of Applied Mathematics, 31 (3):253–259, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pipkin:1983:IND

[107] A. C. Pipkin. Inextensible networks deformed as translation surfaces. IMA Journal of Applied Mathematics, 31 (3):261–267, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

El-Abbasy:1983:SSF

[108] E. M. El-Abbasy and Eleanor M. James. Stable subharmonics of the forced van der Pol equation. IMA Journal of Applied Mathematics, 31 (3):269–279, ???? 1983. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Anonymous:1984:PSJ

[109] Anonymous. Preface: Sir James Lighthill, FRS. IMA Journal of Applied Mathematics, 32(1–3):1, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Benjamin:1984:IFF

[110] T. Brooke Benjamin. Impulse, flow force and variational principles. IMA Journal of Applied Mathematics, 32 (1–3):3–68, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See errata [128].

Blake:1984:MMC

[111] J. R. Blake. Mechanics of mucociliary transport. IMA Journal of Applied Mathematics, 32(1–3):69–87, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bland:1984:FCC

[112] D. R. Bland. Flux and the continuous casting of steel. IMA Journal of Applied Mathematics, 32(1–3):89–112, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1984:AAT

[113] J. E. Ffowcs Williams. The acoustic analogy — thirty years on. IMA Journal of Applied Mathematics, 32 (1–3):113–124, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Freeman:1984:SSN

[114] N. C. Freeman. Soliton solutions of nonlinear evolution equations. IMA Journal of Applied Mathematics, 32 (1–3):125–145, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Goldsworthy:1984:MAE

[115] F. A. Goldsworthy. A mathematical approach to the evolution of HII regions. IMA Journal of Applied Mathematics, 32(1–3):147–173, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hancock:1984:UBL

[116] G. J. Hancock. Unsteady boundary layers. IMA Journal of Applied Mathematics, 32(1–3):175–185, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Howe:1984:AST

[117] M. S. Howe. On the absorption of sound by turbulence and other hydrodynamic flows. IMA Journal of Applied Mathematics, 32(1–3):187–209, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1984:FWH

[118] D. S. Jones. Factorization of a Wiener– Hopf matrix. IMA Journal of Applied Mathematics, 32(1–3):211–220, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mees:1984:DFT

[119] A. I. Mees. Describing functions: ten years on. IMA Journal of Applied Mathematics, 32(1–3):221–233, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Meyer:1984:NEC

[120] R. E. Meyer. Note on evaporation in capillaries. IMA Journal of Applied Mathematics, 32(1–3):235–252, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pack:1984:ULB

[121] D. C. Pack, R. J. Cole, and J. Mika. Upper and lower bounds of bilinear functionals in nonlinear problems. IMA Journal of Applied Mathematics, 32 (1–3):253–266, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pedley:1984:WPP

[122] T. J. Pedley. Wave phenomena in physiological flows. IMA Journal of Applied Mathematics, 32(1–3):267–287, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Spence:1984:SNR

[123] D. A. Spence. Some new results for the integro-differential equation of jet-flap theory. IMA Journal of Applied Mathematics, 32(1–3):289–309, ???? 1984.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stewartson:1984:ICM

[124] K. Stewartson and S. N. Brown. Inviscid centre-modes and wall-modes in the stability theory of swirling Poiseuille flow. IMA Journal of Applied Mathematics, 32(1–3):311–333, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ursell:1984:MNF

[125] F. Ursell. Mathematical note on the fundamental solution (Kelvin source) in ship hydrodynamics. IMA Journal of Applied Mathematics, 32(1– 3):335–351, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Whitham:1984:CPW

[126] G. B. Whitham. Comments on periodic waves and solitons. IMA Journal of Applied Mathematics, 32(1– 3):353–366, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1984:RSR

[127] W. E. Williams. Recognition of some readily “Wiener–Hopf” factorizable matrices. IMA Journal of Applied Mathematics, 32(1–3):367–378, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Benjamin:1984:EIF

[128] T. B. Benjamin. Errata: “Impulse, flow force and variational principles”. IMA Journal of Applied Mathematics, 33(1):i, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [110].

Macaskill:1984:CST

[129] C. Macaskill and T. E. Ewart. Computer simulation of two-dimensional random wave propagation. IMA Journal of Applied Mathematics, 33(1):1– 15, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Aitchison:1984:MEP

[130] J. M. Aitchison, A. A. Lacey, and M. Shillor. A model for an electropaint process. IMA Journal of Applied Mathematics, 33(1):17–31, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bhosale:1984:MZT

[131] S. D. Bhosale and S. V. More. On Marchi-Zgrablich transformation of generalized functions. IMA Journal of Applied Mathematics, 33(1):33–42, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arkeryd:1984:RAF

[132] Leif Arkeryd. A remark about the final aperiodic regime for maps on the interval. IMA Journal of Applied Mathematics, 33(1):43–48, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Evans:1984:NCC

[133] D. V. Evans. A note on the cooling of a cylinder entering a fluid. IMA Journal of Applied Mathematics, 33(1):49–54, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Martin:1984:NFE

[134] P. A. Martin. On the null-field equations for water-wave scattering problems. IMA Journal of Applied Mathematics, 33(1):55–69, July 1984. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Balasubramanyam:1984:ASD

[135] R. Balasubramanyam. Aerodynamic sound due to a point source near a half-plane. IMA Journal of Applied Mathematics, 33(1):71–81, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zayed:1984:ELE

[136] E. M. E. Zayed. Eigenvalues of the Laplacian: an extension to higher dimensions. IMA Journal of Applied Mathematics, 33(1):83–99, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1984:RHT

[137] M. M. R. Williams. Radiant heat transfer through an aerosol suspended in a transparent gas: addendum. IMA Journal of Applied Mathematics, 33 (1):101–103, July 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [93].

Evans:1984:NTS

[138] D. V. Evans. A note on the transparency of a submerged circular cylinder to the waves radiated by a pulsating line source. IMA Journal of Applied Mathematics, 33(2):105–107, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Campos:1984:CDC

[139] L. M. B. C. Campos. On a concept of derivative of complex order with applications to special functions. IMA Journal of Applied Mathematics, 33(2): 109–133, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ughi:1984:MPM

[140] Maura Ughi. A melting problem with a mushy region: qualitative properties. IMA Journal of Applied Mathematics, 33(2):135–152, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sleeman:1984:ESC

[141] B. D. Sleeman and E. Tuma. On exact solutions of a class of reaction-diffusion equations. IMA Journal of Applied Mathematics, 33(2):153–168, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Terrill:1984:NLF

[142] R. M. Terrill. A note on laminar flow through a porous pipe with slip. IMA Journal of Applied Mathematics, 33(2): 169–174, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Markowich:1984:CCV

[143] Peter A. Markowich, Christian A. Ringhofer, and Alois Steindl. Computation of current-voltage characteristics in a semiconductor device using arc-length continuation. IMA Journal of Applied Mathematics, 33(2): 175–187, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Howe:1984:PST

[144] M. S. Howe. On the production of sound by turbulent boundary layer flow over a compliant coating. IMA Journal of Applied Mathematics, 33(2): 189–203, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Srivastava:1984:CPE

[145] H. M. Srivastava. Certain q-polynomial expansions for functions of several variables. II. IMA Journal of Applied Mathematics, 33(2):205–209, September 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Porter:1984:SIE

[146] D. Porter. On some integral equations with a Hankel function kernel. IMA Journal of Applied Mathematics, 33(3):211–228, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wood:1984:RFS

[147] David Wood. Rational-Fourier-series approximations for periodic boundary value problems. IMA Journal of Applied Mathematics, 33(3):229–244, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Burrows:1984:VIC

[148] B. L. Burrows and P. W. Core. Variational-iterative calculations for symmetric operators on real and complex spaces. IMA Journal of Applied Mathematics, 33(3):245–254, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kachoyan:1984:TDF

[149] Philip J. Kachoyan. Triple-deck flows over wavy surfaces. IMA Journal of Applied Mathematics, 33(3):255–271, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bland:1984:SIE

[150] D. R. Bland. Slightly irregular elastic bodies under pressure. IMA Journal of Applied Mathematics, 33(3):273–289, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Firby:1984:IPT

[151] P. A. Firby and D. J. Stone. Interference patterns and two-dimensional potential theory. IMA Journal of Applied Mathematics, 33(3):291–301, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Firby:1984:ADM

[152] P. A. Firby. Angles of dominance for moiré fringes. IMA Journal of Applied Mathematics, 33(3):303–321, ???? 1984. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hurley:1985:DEC

[153] D. G. Hurley and P. F. Siew. The decay of eddy-currents in thin sheets and related water-wave problems. IMA Journal of Applied Mathematics, 34

(1):1–21, ????  1985.  CODEN

IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mason:1985:NHB

[154] D. P. Mason and R. E. Robertson. On a nonhomogeneous Hilbert boundary problem in the kinetic theory of ionized gases. IMA Journal of Applied Mathematics, 34(1):23–47, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arend:1985:IBT

[155] I. J. Arend and D. F. Parker. Inflation, bending and torsion of reinforced elastic tubes. IMA Journal of Applied Mathematics, 34(1):49–71, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zaturska:1985:TEV

[156] Maria B. Zaturska and W. H. H. Banks. Thermal explosion with variable thermal conductivity: Criticality and its disappearance. IMA Journal of Applied Mathematics, 34(1):73–82, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1985:BIE

[157] D. S. Jones. Boundary integrals in elastodynamics. IMA Journal of Applied Mathematics, 34(1):83–97, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Shaw:1985:PRD

[158] D. C. Shaw. Perturbational results for diffraction of water-waves by nearlyvertical barriers. IMA Journal of Applied Mathematics, 34(1):99–117, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rawlins:1985:NCT

[159] A. D. Rawlins. Note on the capacitance of two closely separated spheres. IMA Journal of Applied Mathematics, 34(1):119–120, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Assiff:1985:PPT

[160] Thomas C. Assiff and David H. Y. Yen. On a penalty-perturbation theory for plate problems. IMA Journal of Applied Mathematics, 34(2):121–136, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mazzilli:1985:TEA

[161] Carlos Eduardo Nigro Mazzilli. A theoretical and experimental analysis of support-excited nonlinear vibrations of the extensible pendulum. IMA Journal of Applied Mathematics, 34 (2):137–154, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Schmeiser:1985:FDT

[162] Christian Schmeiser. Finite deformations of thin beams. Asymptotic analysis by singular perturbation methods. IMA Journal of Applied Mathematics, 34(2):155–164, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Varma:1985:ASN

[163] Arvind Varma and William Strieder. Approximate solutions of nonlinear boundary value problems. IMA Journal of Applied Mathematics, 34(2):165– 171, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Apelblat:1985:SIG

[164] Alexander Apelblat. Some integrals of gamma, polygamma and Volterra functions. IMA Journal of Applied Mathematics, 34(2):173–186, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Apelblat:1985:IRD

[165] Alexander Apelblat and Naftali Kravitsky. Integral representations of derivatives and integrals with respect to the order of the Bessel functions J_{v}(t), I_{v}(t), the Anger function J_{v}(t) and the integral Bessel function Ji_{v}(t). IMA Journal of Applied Mathematics, 34(2):187–210, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Exton:1985:LTM

[166] Harold Exton. The Laplace transform of the Macdonald function of argument x^{3/2} and the Airy function Ai(x). IMA Journal of Applied Mathematics, 34(2):211–212, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bhosale:1985:DMZ

[167] S. D. Bhosale and S. V. More. On distributional Marchi-Zgrablich transformation. IMA Journal of Applied Mathematics, 34(2):213–223, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Visintin:1985:SPP

[168] Augusto Visintin. Stefan problem with phase relaxation. IMA Journal of Applied Mathematics, 34(3):225–245, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Furzeland:1985:SGE

[169] R. M. Furzeland and Erik B. Hansen. The steaming up of glasses on entering a sauna — a novel movingboundary problem. IMA Journal of Applied Mathematics, 34(3):247–257, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1985:DCE

[170] A. A. Lacey. Design of a cathode for an electromachining process. IMA Journal of Applied Mathematics, 34 (3):259–267, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Longuet-Higgins:1985:ABC

[171] M. S. Longuet-Higgins. The asymptotic behaviour of the coefficients in Stokes’s series for surface gravity waves. IMA Journal of Applied Mathematics, 34 (3):269–277, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hood:1985:UPW

[172] M. J. Hood. A uniqueness problem for wave propagation in a periodic structure. IMA Journal of Applied Mathematics, 34(3):279–294, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wilmott:1985:NWM

[173] Paul Wilmott. A note on the WKB method for difference equations. IMA Journal of Applied Mathematics, 34 (3):295–302, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Reddy:1985:PAR

[174] B. D. Reddy, T. B. Griffin, and M. J. Marais. A penalty approach to the rate problem in small-strain plasticity. IMA Journal of Applied Mathematics, 34(3):303–321, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Thynell:1985:IFE

[175] S. T. Thynell and M. N. Oezisik. The integral form of the equation of radiative transfer for an inhomogeneous, anisotropically scattering, solid sphere. IMA Journal of Applied Mathematics, 34(3):323–328, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Carpentier:1985:NSC

[176] M. P. Carpentier and A. F. Dos Santos. Non-spectral complete field expansion in two-dimensional structures. IMA Journal of Applied Mathematics, 35(1):1–12, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bourne:1985:FSS

[177] Christine Bourne and Fred Piper. Finite subsequences of shift register sequences. IMA Journal of Applied Mathematics, 35(1):13–22, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Atkinson:1985:BIE

[178] C. Atkinson and C. R. Champion. A boundary integral equation formulation for problems involving nonlinear power-law materials. IMA Journal of Applied Mathematics, 35(1):23–38, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Talbot:1985:VPI

[179] D. R. S. Talbot and J. R. Willis. Variational principles for inhomogeneous nonlinear media. IMA Journal of Applied Mathematics, 35(1):39–54, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Adler:1985:CCT

[180] J. Adler and D. M. Herbert. A criticality criterion for thermal explosions in well-stirred systems with reactant consumption. IMA Journal of Applied Mathematics, 35(1):55–70, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See addendum [219].

Shaw:1985:ABC

[181] D. C. Shaw. The asymptotic behaviour of a curved line source plume within an enclosure. IMA Journal of Applied Mathematics, 35(1):71–89, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Benjamin:1985:LGC

[182] T. Brooke Benjamin and J. Graham-Eagle. Long gravity-capillary waves with edge constraints. IMA Journal of Applied Mathematics, 35(1):91–114, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kuiken:1985:EEC

[183] H. K. Kuiken. Edge effects in crystal growth under intermediate diffusivekinetic control. IMA Journal of Applied Mathematics, 35(2):117–129, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). Special issue: IMA conference on crystal growth (Oxford, 1985).

Wheeler:1985:SAW

[184] A. A. Wheeler. Some aspects of the weakly nonlinear theory of the morphological instability. IMA Journal of Applied Mathematics, 35(2): 131–144, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). Special issue: IMA conference on crystal growth (Oxford, 1985).

Jenkins:1985:NIM

[185] D. R. Jenkins. Nonlinear interaction of morphological and convective instabilities during solidification of a dilute binary alloy. IMA Journal of Applied Mathematics, 35(2):145–157, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). Special issue: IMA conference on crystal growth (Oxford, 1985).

Fowler:1985:FFB

[186] A. C. Fowler. The formation of freckles in binary alloys. IMA Journal of Applied Mathematics, 35(2): 159–174, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cartwright:1985:EAM

[187] R. A. Cartwright, N. El-Kaddah, and J. Szekely. The effect of an axial magnetic field at the interface of a crystal grown by the Czochralski method. IMA Journal of Applied Mathematics, 35(2): 175–194, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Series:1985:EDC

[188] R. W. Series, D. T. J. Hurle, and K. G. Barraclough. Effective distribution coefficient of silicon dopants during magnetic Czochralski growth. IMA Journal of Applied Mathematics, 35(2): 195–203, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stern:1985:MMR

[189] Ellen J. Stern. Mathematical modelling of radiative heat exchanges in Czochralski crystal pulling. IMA Journal of Applied Mathematics, 35(2): 205–222, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Voller:1985:HBI

[190] V. R. Voller. A heat balance integral method for estimating practical solidification parameters. IMA Journal of Applied Mathematics, 35(2): 223–232, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Visintin:1985:SSE

[191] A. Visintin. Supercooling and superheating effects in phase transitions. IMA Journal of Applied Mathematics, 35(2):233–256, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). Special issue: IMA conference on crystal growth (Oxford, 1985).

Elliott:1985:SPN

[192] Charles M. Elliott. The Stefan problem with a nonmonotone constitutive relation. IMA Journal of Applied Mathematics, 35(2):257–264, September 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). Special issue: IMA conference on crystal growth (Oxford, 1985).

Adler:1985:CCR

[193] J. Adler and P. T. Nhamburo. The critical conditions in a reactive slab with slowly varying surface temperature. IMA Journal of Applied Mathematics, 35(3):265–272, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arnold:1985:GGT

[194] J. M. Arnold. A global geometrical theory of diffraction. IMA Journal of Applied Mathematics, 35(3):273–295, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rynne:1985:SCT

[195] B. P. Rynne. Stability and convergence of time marching methods in scattering problems. IMA Journal of Applied Mathematics, 35(3):297–310, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Barone:1985:SBI

[196] M. R. Barone and D. A. Caulk. Special boundary integral equations for approximate solution of potential problems in three-dimensional regions with slender cavities of circular crosssection. IMA Journal of Applied Mathematics, 35(3):311–325, ???? 1985.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Atkinson:1985:EFR

[197] C. Atkinson. On explicit formulae for the roots of a class of transcendental equations. IMA Journal of Applied Mathematics, 35(3):327–338, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McIver:1985:SWW

[198] P. McIver. Scattering of water waves by two surface-piercing vertical barriers. IMA Journal of Applied Mathematics, 35(3):339–355, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1985:TDE

[199] A. A. Lacey. Tool design for electrochemical machining in the presence of overpotentials. IMA Journal of Applied Mathematics, 35(3):357–364, ???? 1985. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Beale:1986:OMO

[200] E. M. L. Beale. Optimization methods in oil and gas exploration. IMA Journal of Applied Mathematics, 36(1):1–10, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Baker:1986:IAC

[201] P. C. Baker, G. Poots, and G. G. Rodgers. Ice accretion on cables of various cross sections. IMA Journal of Applied Mathematics, 36(1):11–28, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kriegsmann:1986:MTF

[202] Gregory A. Kriegsmann. A mathematical theory of full-wave rectification. IMA Journal of Applied Mathematics, 36(1):29–41, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Markowich:1986:UAR

[203] Peter A. Markowich and Christian Schmeiser. Uniform asymptotic representation of solutions of the basic semiconductor-device equations. IMA Journal of Applied Mathematics, 36 (1):43–57, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cade:1986:IEP

[204] R. Cade. Integral-equation perturbation theory for electrostatic torus problems with applied fields. IMA Journal of Applied Mathematics, 36 (1):59–84, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pipkin:1986:RED

[205] Allen C. Pipkin. The relaxed energy density for isotropic elastic membranes. IMA Journal of Applied Mathematics, 36(1):85–99, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Eagles:1986:FSB

[206] P. M. Eagles and P. G. Daniels. Freesurface boundary-layer flow on a curved bed. IMA Journal of Applied Mathematics, 36(1):101–116, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Organ:1986:MBL

[207] A. E. Organ and N. Riley. Magnetohydrodynamic boundary layers in Czochralski crystal growth. IMA Journal of Applied Mathematics, 36 (2):117–128, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Caldeira-Saraiva:1986:BSL

[208] F. Caldeira-Saraiva. The boundedness of solutions of a Liénard equation arising in the theory of ship rolling. IMA Journal of Applied Mathematics, 36(2):129–139, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Visintin:1986:NMS

[209] A. Visintin. A new model for supercooling and superheating effects. IMA Journal of Applied Mathematics, 36 (2):141–157, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cushman:1986:MTS

[210] John H. Cushman. Multiphase transport in the space of stochastic tempered distributions. IMA Journal of Applied Mathematics, 36(2):159–175, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Daniels:1986:PWS

[211] P. G. Daniels and A. Golbabai. Phasewinding solutions for axisymmetric convection between rotating planes uniformly heated from below. IMA Journal of Applied Mathematics, 36 (2):177–189, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Campos:1986:SAS

[212] L. M. B. C. Campos. On a systematic approach to some properties of special functions. IMA Journal of Applied Mathematics, 36(2):191–206, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chadam:1986:RII

[213] J. Chadam, D. Hoff, E. Merino, P. Ortoleva, and A. Sen. Reactive infiltration instabilities. IMA Journal of Applied Mathematics, 36(3):207–221, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Reed:1986:SDA

[214] M. B. Reed. Stresses and displacements around a cylindrical cavity in soft rock. IMA Journal of Applied Mathematics, 36(3):223–245, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Howe:1986:ADB

[215] M. S. Howe. Attenuation and diffraction of bending waves at gaps in fluid loaded plates. IMA Journal of Applied Mathematics, 36(3):247–262, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Banks:1986:EBL

[216] W. H. H. Banks and M. B. Zaturska. Eigensolutions in boundary-layer flow adjacent to a stretching wall. IMA Journal of Applied Mathematics, 36 (3):263–273, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1986:TCB

[217] M. M. R. Williams. Thermal conduction in bispherical systems and its role in self-induced thermophoresis of aerosol particles. IMA Journal of Applied Mathematics, 36(3):275–292, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See corrigendum [225].

Mackie:1986:CVI

[218] A. G. Mackie. Complementary variational inequalities. IMA Journal of Applied Mathematics, 36(3):293–305, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Adler:1986:ACC

[219] J. Adler and D. M. Herbert. Addendum: “A criticality criterion for thermal explosions in well-stirred systems with reactant consumption” [IMA J. Appl. Math. 35 (1985), no. 1, 55– 70; MR 87c:80019]. IMA Journal of Applied Mathematics, 36(3):307, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [180].

Markowich:1986:AAO

[220] P. A. Markowich, C. A. Ringhofer, and C. Schmeiser. An asymptotic analysis of one-dimensional models of semiconductor devices. IMA Journal of Applied Mathematics, 37(1):1–24, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Graham-Eagle:1986:ESS

[221] J. G. Graham-Eagle, B. F. Gray, and G. C. Wake. On the existence and stability of spatially structured solutions of the reaction-diffusion equations. IMA Journal of Applied Mathematics, 37(1):25–35, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See addendum [278].

Shail:1986:SPP

[222] R. Shail and B. A. Packham. Some potential problems associated with the sedimentation of a small particle into a semi-infinite fluid-filled pore. IMA Journal of Applied Mathematics, 37 (1):37–66, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wilson:1986:STP

[223] D. G. Wilson and A. D. Solomon. A Stefan-type problem with void formation and its explicit solution. IMA Journal of Applied Mathematics, 37 (1):67–76, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McKee:1986:RWW

[224] W. D. McKee. Reflection of water waves from an exponentially sheared current. IMA Journal of Applied Mathematics, 37(1):77–90, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Williams:1986:CTC

[225] M. M. R. Williams. Corrigendum: “Thermal conduction in bispherical systems and its role in self-induced thermophoresis of aerosol particles”. IMA Journal of Applied Mathematics, 37(1):91, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [217].

McIver:1986:EFA

[226] M. McIver. The electromagnetic field around a subsurface flaw in a metal. IMA Journal of Applied Mathematics, 37(2):93–111, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Byatt-Smith:1986:RCS

[227] J. G. Byatt-Smith. Regular and chaotic solutions of Duffing’s equation for large forcing. IMA Journal of Applied Mathematics, 37(2):113–145, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Elliott:1986:MPD

[228] C. M. Elliott, M. A. Herrero, J. R. King, and J. R. Ockendon. The mesa problem: diffusion patterns for u_{t} = ∇ · (u^{m}∇u) as m → +∞. IMA Journal of Applied Mathematics, 37 (2):147–154, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Millar:1986:SRH

[229] Robert F. Millar. Singularities and the Rayleigh hypothesis for solutions to the Helmholtz equation. IMA Journal of Applied Mathematics, 37 (2):155–171, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bestman:1986:TFR

[230] A. R. Bestman. Transient flow of a radiating gas between concentric spheres. IMA Journal of Applied Mathematics, 37(2):173–183, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

King:1986:DDC

[231] J. R. King and C. P. Please. Diffusion of dopant in crystalline silicon: an asymptotic analysis. IMA Journal of Applied Mathematics, 37 (3):185–197, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ockendon:1986:MSD

[232] J. R. Ockendon and P. Wilmott. Matching and singularity distributions in inviscid flow. IMA Journal of Applied Mathematics, 37(3):199–211, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kirsch:1986:DFF

[233] Andreas Kirsch. The denseness of the far field patterns for the transmission problem. IMA Journal of Applied Mathematics, 37(3):213–225, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Abrahams:1986:SSS

[234] I. D. Abrahams. Scattering of sound by a semi-infinite elastic plate with a soft backing; a matrix Wiener–Hopf problem. IMA Journal of Applied Mathematics, 37(3):227–245, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Miketinac:1986:SJF

[235] M. J. Miketinac and R. L. Sani. Simulation of jet flows in ideal fluids by variable finite elements. IMA Journal of Applied Mathematics, 37 (3):247–259, ???? 1986. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Greenberg:1987:HHT

[236] J. M. Greenberg. A hyperbolic heat transfer problem with phase changes. IMA Journal of Applied Mathematics, 38(1):1–21, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stuart:1987:EST

[237] A. M. Stuart. Existence of solutions of a two-point free-boundary problem arising in the theory of porous medium combustion. IMA Journal of Applied Mathematics, 38(1):23–34, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Walker:1987:SCC

[238] Glen R. Walker. Solution to a class of coupled linear partial differential equations. IMA Journal of Applied Mathematics, 38(1):35–48, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McClure:1987:AED

[239] J. P. McClure and R. Wong. Asymptotic expansion of a double integral with a curve of stationary points. IMA Journal of Applied Mathematics, 38(1):49–59, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hebeker:1987:ASP

[240] F.-K. Hebeker, J. Mika, and D. C. Pack. Application of the superconvergence properties of the Galerkin approximation to the calculation of upper and lower bounds for linear functionals of solutions of integral equations. IMA Journal of Applied Mathematics, 38(1):61–70, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kachoyan:1987:FDP

[241] B. J. Kachoyan. The finite Dean problem: nonlinear theory. IMA Journal of Applied Mathematics, 38(1):71–86, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

King:1987:HCA

[242] J. R. King. High concentration arsenic diffusion in crystalline silicon: An asymptotic analysis. IMA Journal of Applied Mathematics, 38(2):87–95, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Elliott:1987:NSC

[243] Charles M. Elliott and Donald A. French. Numerical studies of the Cahn– Hilliard equation for phase separation. IMA Journal of Applied Mathematics, 38(2):97–128, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Maeda:1987:SMD

[244] Shigeru Maeda. The similarity method for difference equations. IMA Journal of Applied Mathematics, 38(2):129– 134, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Robinson:1987:MVP

[245] F. N. H. Robinson. The modified van der Pol oscillator. IMA Journal of Applied Mathematics, 38(2):135–150, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pettersson:1987:SHM

[246] Rolf Pettersson. On solutions and higher moments for the linear Boltzmann equation with infinite-range forces. IMA Journal of Applied Mathematics, 38(2):151–166, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Yortsos:1987:SCC

[247] Y. C. Yortsos. Stability of a certain class of miscible displacement processes in porous media. IMA Journal of Applied Mathematics, 38(3):167–179, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Fradkin:1987:SPB

[248] L. Ju. Fradkin and R. A. Dawe. Some a priori bounds for flows through heterogeneous media. IMA Journal of Applied Mathematics, 38(3):181–193, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Caginalp:1987:EPI

[249] G. Caginalp and P. C. Fife. Elliptic problems involving phase boundaries satisfying a curvature condition. IMA Journal of Applied Mathematics, 38(3):195–217, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Robinson:1987:EIN

[250] F. N. H. Robinson. An experimental investigation of a nonlinear Mathieu equation. IMA Journal of Applied Mathematics, 38(3):219–242, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Reynolds:1987:PSI

[251] D. W. Reynolds and E. A. Cox. Periodic solutions of an integrodifferential equation modelling resonant sloshing. IMA Journal of Applied Mathematics, 38(3):243–253, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chadam:1987:ESS

[252] J. Chadam, S. D. Howison, and P. Ortoleva. Existence and stability for spherical crystals growing in a supersaturated solution. IMA Journal of Applied Mathematics, 39(1):1–15, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ringhofer:1987:SPA

[253] Christian A. Ringhofer. A singular perturbation analysis for the transient semiconductor device equations in one space dimension. IMA Journal of Applied Mathematics, 39(1):17–32, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gustafsson:1987:SII

[254] Björn Gustafsson and Jacqueline Mossino. Some isoperimetric inequalities in electrochemistry and Hele Shaw flows. IMA Journal of Applied Mathematics, 39(1):33–49, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Caginalp:1987:NAA

[255] Gunduz Caginalp and Jian-Tong Lin. A numerical analysis of an anisotropic phase field model. IMA Journal of Applied Mathematics, 39(1):51–66, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Graham-Eagle:1987:SSD

[256] J. Graham-Eagle and I. Stakgold. A steady-state diffusion problem with fractional power absorption rate. IMA Journal of Applied Mathematics, 39 (1):67–73, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Angelova:1987:OAB

[257] D. Ts. Angelova and D. D. Baĭnov. Oscillatory and asymptotic behaviour of the solutions of first-order functionaldifferential equations. IMA Journal of Applied Mathematics, 39(1):75–89, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Holmes:1987:DSC

[258] Philip Holmes. Dynamical systems in chaos: some recent books. IMA Journal of Applied Mathematics, 39 (1):91–98, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Howe:1987:PST

[259] M. S. Howe. Production of sound by turbulent flow over an embedded strut. IMA Journal of Applied Mathematics, 39(2):99–120, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1987:EED

[260] A. A. Lacey and M. Shillor. Electrochemical and electro-discharge machining with a threshold current. IMA Journal of Applied Mathematics, 39 (2):121–142, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Visintin:1987:MMS

[261] A. Visintin. Mathematical models of solid-solid phase transitions in steel. IMA Journal of Applied Mathematics, 39(2):143–157, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McBride:1987:NFF

[262] A. C. McBride and F. H. Kerr. On Namias’s fractional Fourier transforms. IMA Journal of Applied Mathematics, 39(2):159–175, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Robinson:1987:BMH

[263] P. D. Robinson and P. K. Yuen. Bivariational methods for Hammerstein integral equations. IMA Journal of Applied Mathematics, 39(2):177–188, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1987:CHS

[264] S. A. Smith. Congruent harmonic and space charge electrostatic fields. IMA Journal of Applied Mathematics, 39(3):189–214, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Talbot:1987:BSC

[265] D. R. S. Talbot and J. R. Willis. Bounds and self-consistent estimates for the overall properties of nonlinear composites. IMA Journal of Applied Mathematics, 39(3):215–240, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Norbury:1987:PFB

[266] J. Norbury and A. M. Stuart. Parabolic free boundary problems arising in porous medium combustion. IMA Journal of Applied Mathematics, 39 (3):241–257, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Haughton:1987:IBT

[267] D. M. Haughton. Inflation and bifurcation of thick-walled compressible elastic spherical shells. IMA Journal of Applied Mathematics, 39(3):259–272, ???? 1987. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Budd:1988:ESS

[268] C. J. Budd and A. A. Wheeler. Exact solutions of the space charge equation using the hodograph method. IMA Journal of Applied Mathematics, 40(1):1–14, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cimatti:1988:BTT

[269] Giovanni Cimatti. A bound for the temperature in the thermistor problem. IMA Journal of Applied Mathematics, 40(1):15–22, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Skelton:1988:API

[270] P. L. I. Skelton and G. Poots. Approximate predictions of ice accretion on an overhead transmission line which rotates with constant angular velocity. IMA Journal of Applied Mathematics, 40(1):23–35, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pushpavanam:1988:CRI

[271] S. Pushpavanam and R. Narayanan. Comparison results for ignition in conjugate systems. IMA Journal of Applied Mathematics, 40(1):37–51, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grundy:1988:ASM

[272] R. E. Grundy. Asymptotic solutions of a model diffusion-reaction equation. IMA Journal of Applied Mathematics, 40(1):53–72, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smyth:1988:HON

[273] N. F. Smyth and J. M. Hill. Highorder nonlinear diffusion. IMA Journal of Applied Mathematics, 40(2):73–86, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bluman:1988:NSO

[274] George W. Bluman and Gregory J. Reid. New symmetries for ordinary differential equations. IMA Journal of Applied Mathematics, 40(2):87–94, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wienert:1988:EPB

[275] Lutz Wienert. An existence proof for a boundary value problem with nonsmooth boundary from the theory of water waves. IMA Journal of Applied Mathematics, 40(2):95–112, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lawrie:1988:ARF

[276] J. B. Lawrie. Axisymmetric radiation from a finite gap in an infinite, rigid, circular duct. IMA Journal of Applied Mathematics, 40(2):113–128, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jost:1988:IEM

[277] Gabriele Jost. Integral equations with modified fundamental solution in timeharmonic electromagnetic scattering. IMA Journal of Applied Mathematics, 40(2):129–143, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Brindley:1988:AES

[278] J. Brindley and G. C. Wake. Addendum to: “On the existence and stability of spatially structured solutions of the reaction-diffusion equations” [IMA J. Appl. Math. 37 (1986), no. 1, 25– 35; MR 90a:35119 ] by J. G. Graham-Eagle, B. F. Gray and Wake. IMA Journal of Applied Mathematics, 40 (2):145–146, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [221].

Howison:1988:SSS

[279] S. D. Howison. Similarity solutions to the Stefan problem and the binary alloy problem. IMA Journal of Applied Mathematics, 40(3):147–161, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

King:1988:EHC

[280] J. R. King. Extremely high concentration dopant diffusion in silicon. IMA Journal of Applied Mathematics, 40 (3):163–181, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grundy:1988:CPN

[281] R. E. Grundy. The Cauchy problem for a nonlinear diffusion equation with absorption and convection. IMA Journal of Applied Mathematics, 40 (3):183–204, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kenmochi:1988:ODH

[282] Nobuyuki Kenmochi and Mario Primicerio. One-dimensional heat conduction with a class of automatic heatsource controls. IMA Journal of Applied Mathematics, 40(3):205–216, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stuart:1988:SSH

[283] Andrew Stuart. Similarity solutions of a heat equation with nonlinearly varying heat capacity. IMA Journal of Applied Mathematics, 40(3):217–234, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1988:EEC

[284] Ronald Smith. Entry and exit conditions for flow reactors. IMA Journal of Applied Mathematics, 41(1):1–20, ????

1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1988:SRC

[285] D. S. Jones. Surface radiation conditions. IMA Journal of Applied Mathematics, 41(1):21–30, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Fasano:1988:PCM

[286] Antonio Fasano and Mario Primicerio. A phase-change model with a zone of coexistence of phases. IMA Journal of Applied Mathematics, 41 (1):31–46, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Warhola:1988:SSV

[287] G. T. Warhola and A. C. Pipkin. Shock structure in viscoelastic materials. IMA Journal of Applied Mathematics, 41(1):47–66, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cannon:1988:DRC

[288] John R. Cannon, Joseph V. Koebbe, and Rita M. Schmidt. The determination of rate constants from known linear combinations of concentrations in a chemical reaction. IMA Journal of Applied Mathematics, 41(1):67–84, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kath:1988:OTR

[289] William L. Kath and Gregory A. Kriegsmann. Optical tunnelling: radiation losses in bent fibre-optic waveguides. IMA Journal of Applied Mathematics, 41(2):85–103, ???? 1988.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rogers:1988:EUC

[290] Joel C. W. Rogers. Existence, uniqueness, and construction of the solution of a system of ordinary functionaldifferential equations, with application to the design of perfectly focusing symmetric lenses. IMA Journal of Applied Mathematics, 41(2):105–134, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tan:1988:BHI

[291] Yong Ji Tan. Boundary homogenization of an inverse nonlocal elliptic boundary value problem. IMA Journal of Applied Mathematics, 41 (2):135–146, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

vanDuijn:1988:CSS

[292] C. J. van Duijn, S. M. Gomes, and Hong Fei Zhang. On a class of similarity solutions of the equation u_{t} = (|u|^{m−1}u_{x})_{x} with m > −1. IMA Journal of Applied Mathematics, 41 (2):147–163, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Parker:1988:NDM

[293] D. F. Parker and A. D. Fitt. Nondiffusive models for pressure-swing adsorption. IMA Journal of Applied Mathematics, 41(3):165–192, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jenkins:1988:CFS

[294] D. R. Jenkins and N. G. Barton. Computation of the free-surface shape of an inviscid jet incident on a porous wall. IMA Journal of Applied Mathematics, 41(3):193–206, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lacey:1988:CBS

[295] A. A. Lacey and D. Tzanetis. Complete blow-up for a semilinear diffusion equation with a sufficiently large initial condition. IMA Journal of Applied Mathematics, 41(3):207–215, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Larcombe:1988:MMI

[296] P. J. Larcombe, G. Poots, P. L. I. Skelton, and M. Shillor. Mathematical models for ice accretion on conductors using free streamline theory. I. Single conductor. IMA Journal of Applied Mathematics, 41(3):217–236, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stojanovic:1988:IIF

[297] Srdjan Stojanovic. Injection of ideal fluid from a slot into a stream: two free boundaries. IMA Journal of Applied Mathematics, 41(3):237–253, ???? 1988. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1989:IVC

[298] S. H. Smith. The influence of variable convective velocities on longitudinal diffusion. IMA Journal of Applied Mathematics, 42(1):1–11, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Phillips:1989:SME

[299] Timothy N. Phillips. Singular matched eigenfunction expansions for Stokes flow around a corner. IMA Journal of Applied Mathematics, 42(1):13–26, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stuart:1989:NHL

[300] Andrew Stuart. A note on high/lowwave-number interactions in spatially discrete parabolic equations. IMA Journal of Applied Mathematics, 42 (1):27–42, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hammersley:1989:DEC

[301] J. M. Hammersley and G. Mazzarino. A differential equation connected with the dendritic growth of crystals. IMA Journal of Applied Mathematics, 42 (1):43–75, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Colton:1989:ISP

[302] D. L. Colton and P. B. Monk. The inverse scattering problem for timeharmonic acoustic waves in an inhomogeneous medium: numerical experiments. IMA Journal of Applied Mathematics, 42(1):77–95, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chatwin:1989:IAD

[303] P. C. Chatwin, N. T. Hajian, N. Mole, and C. D. Jones. Investigations on the atmospheric dispersion of clouds containing charged tracers. IMA Journal of Applied Mathematics, 42(1):97–117, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bassom:1989:UBI

[304] Andrew P. Bassom and J. S. B. Gajjar. Upper-branch instability of flow in pipes of large aspect ratio with threedimensional nonlinear viscous critical layers. IMA Journal of Applied Mathematics, 42(2):119–145, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Burnell:1989:DCA

[305] J. G. Burnell, J. G. Graham-Eagle, B. F. Gray, and G. C. Wake. Determination of critical ambient temperatures for thermal ignition. IMA Journal of Applied Mathematics, 42 (2):147–154, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Howison:1989:ESS

[306] S. D. Howison and J. R. King. Explicit solutions to six free-boundary problems in fluid flow and diffusion. IMA Journal of Applied Mathematics, 42(2):155–175, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Robinson:1989:EOL

[307] F. N. H. Robinson. Experimental observation of the large-amplitude solutions of Duffing’s and related equations. IMA Journal of Applied Mathematics, 42(2):177–201, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Firby:1989:SSE

[308] P. A. Firby. Separating sets as envelopes. IMA Journal of Applied Mathematics, 42(2):203–208, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

ChanHong:1989:IBF

[309] J. R. Chan Hong, C. J. van Duijn, D. Hilhorst, and J. van Kester. The interface between fresh and salt groundwater: a numerical study. IMA Journal of Applied Mathematics, 42 (3):209–240, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Yurekli:1989:PTT

[310] Osman Yürekli. A Parseval-type theorem applied to certain integral transforms. IMA Journal of Applied Mathematics, 42(3):241–249, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rodrigues:1989:ACO

[311] José-Francisco Rodrigues and Lisa Santos. Asymptotic convergences in a onephase continuous-casting Stefan problem with high extraction velocity. IMA Journal of Applied Mathematics, 42 (3):251–267, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bassom:1989:WNL

[312] Andrew P. Bassom. Weakly nonlinear lower-branch stability of fully developed and developing free-surface flows. IMA Journal of Applied Mathematics, 42(3):269–301, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McNabb:1989:FRT

[313] A. McNabb and L. Bass. Flux-ratio theorems for nonlinear equations of generalized diffusion. IMA Journal of Applied Mathematics, 43(1):1–9, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Varley:1989:MSS

[314] E. Varley and J. D. A. Walker. A method for solving singular integrodifferential equations. IMA Journal of Applied Mathematics, 43(1):11–45, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1989:WPC

[315] William V. Smith. Waves in a perfectly conducting fluid filling a half-space. IMA Journal of Applied Mathematics, 43(1):47–69, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kwok:1989:RPM

[316] Yue-Kuen Kwok. A regular perturbation method for subcritical flow over a two-dimensional airfoil. IMA Journal of Applied Mathematics, 43(1):71–81, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bergstrom:1989:SCP

[317] T. Bergstrøm, S. Cowley, A. C. Fowler, and P. E. Seward. Segregation of carbon paste in a smelting electrode. IMA Journal of Applied Mathematics, 43(1):83–99, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

DalPasso:1989:MBP

[318] R. Dal Passo and P. de Mottoni. On a moving boundary problem arising in fluidized bed combustion. IMA Journal of Applied Mathematics, 43 (2):101–126, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Graham-Eagle:1989:MBS

[319] James Graham-Eagle. The minimal branch of solutions of an equation arising in thermal combustion theory. IMA Journal of Applied Mathematics, 43 (2):127–131, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Estrada:1989:IEL

[320] R. Estrada and R. P. Kanwal. Integral equations with logarithmic kernels. IMA Journal of Applied Mathematics, 43(2):133–155, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mandal:1989:NDW

[321] B. N. Mandal and A. Chakrabarti. A note on diffraction of water waves by a nearly vertical barrier. IMA Journal of Applied Mathematics, 43 (2):157–165, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Logan:1989:SSS

[322] J. David Logan and Edwin L. Woerner. Sensitivity of self-similar ZND waves in condensed media. IMA Journal of Applied Mathematics, 43(2):167–184, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Guiasu:1989:OAE

[323] Ileana Guiaşu and Hans Raszillier. Optimal approximation of the electrostatic capacity matrix of two conducting spheres by a short-distance asymptotic expansion. IMA Journal of Applied Mathematics, 43(2):185–193, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Nalesso:1989:ZCB

[324] G. F. Nalesso. On the zeros of a class of Bessel functions whose argument and order are functions of a complex variable. IMA Journal of Applied Mathematics, 43(3):195–217, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Brokate:1989:EOC

[325] M. Brokate and J. Sprekels. Existence and optimal control of mechanical processes with hysteresis in viscous solids. IMA Journal of Applied Mathematics, 43(3):219–229, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Willis:1989:SOC

[326] J. R. Willis. The structure of overall constitutive relations for a class of nonlinear composites. IMA Journal of Applied Mathematics, 43(3):231–242, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Steinruck:1989:AAC

[327] Herbert Steinrück. Asymptotic analysis of the current-voltage curve of a pnpn semiconductor device. IMA Journal of Applied Mathematics, 43 (3):243–259, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Buonincontri:1989:MTW

[328] Steve Buonincontri and Thomas Hagstrom. Multidimensional travelling wave solutions to reaction-diffusion equations. IMA Journal of Applied Mathematics, 43(3):261–271, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Paris:1989:MEO

[329] R. B. Paris and A. D. Wood. A model equation for optical tunnelling. IMA Journal of Applied Mathematics, 43(3):273–284, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Soni:1990:BSA

[330] K. Soni and N. M. Temme. On a biorthogonal system associated with uniform asymptotic expansions. IMA Journal of Applied Mathematics, 44 (1):1–25, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Clarkson:1990:PAC

[331] Peter A. Clarkson. Painlevé analysis and the complete integrability of a generalized variablecoefficient Kadomtsev–Petviashvili equation. IMA Journal of Applied Mathematics, 44(1):27–53, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Freeden:1990:MNI

[332] W. Freeden, T. Gervens, and J. C. Mason. A minimum norm interpolation method for determining the displacement field of a homogeneous isotropic elastic body from discrete data. IMA Journal of Applied Mathematics, 44 (1):55–76, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Caginalp:1990:DCP

[333] G. Caginalp. The dynamics of a conserved phase field system: Stefan-like, Hele–Shaw, and Cahn–Hilliard models as asymptotic limits. IMA Journal of Applied Mathematics, 44(1):77–94, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tew:1990:RTD

[334] R. H. Tew. Ray theory of the diffraction of sound by an inhomogeneous membrane. IMA Journal of Applied Mathematics, 44(2):95–110, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1990:SIE

[335] James M. Hill. Some integrable and exact similarity solutions for plane finite elastic deformations. IMA Journal of Applied Mathematics, 44(2):111–126, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1990:MSI

[336] D. S. Jones and X. Q. Mao. A method for solving the inverse problem in soft acoustic scattering. IMA Journal of Applied Mathematics, 44 (2):127–143, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Doyle:1990:SSG

[337] J. Doyle and M. J. Englefield. Similarity solutions of a generalized Burgers equation. IMA Journal of Applied Mathematics, 44(2):145–153, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McNabb:1990:FTL

[338] A. McNabb and L. Bass. Flux theorems for linear multicomponent diffusion. IMA Journal of Applied Mathematics, 44(2):155–161, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bourne:1990:SSV

[339] J. P. Bourne and C. Atkinson. Stress singularities in viscoelastic media. II. Plane-strain stress singularities at corners. IMA Journal of Applied Mathematics, 44(2):163–180, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Graham-Eagle:1990:VAU

[340] James Graham-Eagle. A variational approach to upper and lower solutions. IMA Journal of Applied Mathematics, 44(2):181–184, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cimatti:1990:EWS

[341] Giovanni Cimatti. Existence of weak solutions for the space-charge problem. IMA Journal of Applied Mathematics, 44(2):185–195, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hobbs:1990:LBA

[342] Ann Kahlow Hobbs and William L. Kath. Loss and birefringence for arbitrarily bent optical fibres. IMA Journal of Applied Mathematics, 44 (3):197–219, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wall:1990:UTI

[343] David J. N. Wall. Uniqueness theorems for the inverse problem of elastodynamic boundary scattering. IMA Journal of Applied Mathematics, 44 (3):221–241, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Keady:1990:DAR

[344] G. Keady. The Dupuit approximation for the rectangular dam problem. IMA Journal of Applied Mathematics, 44(3):243–260, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dowaikh:1990:SWD

[345] M. A. Dowaikh and R. W. Ogden. On surface waves and deformations in a pre-stressed incompressible elastic solid. IMA Journal of Applied Mathematics, 44(3):261–284, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cosner:1990:PSN

[346] C. Cosner, S. Lenhart, and V. Protopopescu. Parabolic systems with nonlinear competitive interactions. IMA Journal of Applied Mathematics, 44(3):285–298, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dennis:1990:STD

[347] S. C. R. Dennis and S. Kocabiyik. The solution of two-dimensional Oseen flow problems using integral conditions. IMA Journal of Applied Mathematics, 45(1):1–31, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mestel:1990:MAS

[348] A. J. Mestel. More accurate skin-depth approximations. IMA Journal of Applied Mathematics, 45(1):33–48, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Firby:1990:MS

[349] P. A. Firby and S. J. Hubbard. Moiré separations. IMA Journal of Applied Mathematics, 45(1):49–80, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Caloz:1990:MPP

[350] G. Caloz. The model plasma problem: existence of a solution branch. IMA Journal of Applied Mathematics, 45(1):81–98, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Green:1990:DDE

[351] W. A. Green and Jingyu Shi. Deformations of discrete elastic networks. IMA Journal of Applied Mathematics, 45(2):99–113, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kay:1990:ADR

[352] Anthony Kay. Advection-diffusion in reversing and oscillating flows. 1. The effect of a single reversal. IMA Journal of Applied Mathematics, 45 (2):115–137, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

King:1990:NIJ

[353] A. C. King. A note on the impact of a jet on a porous wall. IMA Journal of Applied Mathematics, 45 (2):139–146, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ke:1990:ABS

[354] Chen Ke and S. Amini. Analytical behaviour of solutions of boundary integral equations for a nonsmooth region. IMA Journal of Applied Mathematics, 45(2):147–157, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1990:OLD

[355] Ronald Smith. Oxidant-limited depletion for a near-shore discharge. IMA Journal of Applied Mathematics, 45 (2):159–173, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Yang:1990:CDZ

[356] Yisong Yang. Computation, dimensionality, and zero-dissipation limit of the Ginzburg–Landau wave equation. IMA Journal of Applied Mathematics, 45(2):175–194, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Greenberg:1990:SGF

[357] J. M. Greenberg and W. Shyong. Surging glacial flows. IMA Journal of Applied Mathematics, 45(3):195–223, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1990:MAI

[358] W. R. Smith and G. C. Wake. Mathematical analysis: an inverse problem arising in convective-diffusive flow. IMA Journal of Applied Mathematics, 45(3):225–231, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zayed:1990:HEA

[359] E. M. E. Zayed. Heat equation of an arbitrary multiply-connected region in R^{2} with impedance boundary conditions. IMA Journal of Applied Mathematics, 45(3):233–241, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1990:HON

[360] James M. Hill and Desmond L. Hill. High-order nonlinear evolution equations. IMA Journal of Applied Mathematics, 45(3):243–265, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Riley:1990:LWI

[361] D. S. Riley and S. H. Davis. Long-wave interactions in morphological and convective instabilities. IMA Journal of Applied Mathematics, 45(3):267–285, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Herbert:1990:TET

[362] D. M. Herbert. Thermal explosion theory for a reactive ridge. IMA Journal of Applied Mathematics, 45 (3):287–298, ???? 1990. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Albert:1991:TPS

[363] J. P. Albert and J. L. Bona. Total positivity and the stability of internal waves in stratified fluids of finite depth. IMA Journal of Applied Mathematics, 46(1–2):1–19, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Amick:1991:UBS

[364] C. J. Amick and J. F. Toland. Uniqueness of Benjamin’s solitary-wave solution of the Benjamin–Ono equation. IMA Journal of Applied Mathematics, 46(1–2):21–28, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Lions:1991:RDL

[365] J.-L. Lions. Remarks on Darcy’s law. IMA Journal of Applied Mathematics, 46(1–2):29–38, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Lodge:1991:HPP

[366] A. S. Lodge, W. G. Pritchard, and L. R. Scott. The hole-pressure problem. IMA Journal of Applied Mathematics, 46(1–2):39–66, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See corrigendum [437]. The Brooke Benjamin special issue (University Park, PA, 1989).

Longuet-Higgins:1991:BN

[367] Michael S. Longuet-Higgins. Bubble noise. IMA Journal of Applied Mathematics, 46(1–2):67–70, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Maddocks:1991:SRE

[368] John H. Maddocks. On the stability of relative equilibria. IMA Journal of Applied Mathematics, 46(1–2):71–99, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Miles:1991:NAE

[369] John Miles. Nonlinear asymmetric excitation of edge waves. IMA Journal of Applied Mathematics, 46(1–2):101–108, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Mullin:1991:FDD

[370] T. Mullin. Finite-dimensional dynamics in Taylor–Couette flow. IMA Journal of Applied Mathematics, 46 (1–2):109–119, ???? 1991. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Promislow:1991:AIL

[371] Keith Promislow and Roger Temam. Approximate interaction laws for small and large waves in the Ginzburg– Landau equation. IMA Journal of Applied Mathematics, 46(1–2):121–136, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Saffman:1991:SMS

[372] P. G. Saffman. Selection mechanisms and stability of fingers and bubbles in Hele–Shaw cells. IMA Journal of Applied Mathematics, 46(1–2):137–145, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Stuart:1991:LPF

[373] J. T. Stuart. The Lagrangian picture of fluid motion and its implication for flow structures. IMA Journal of Applied Mathematics, 46(1–2):147–163, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Tavener:1991:PFE

[374] S. J. Tavener and K. A. Cliffe. Primary flow exchange mechanisms in the Taylor apparatus applying impermeable stress-free boundary conditions. IMA Journal of Applied Mathematics, 46(1–2):165–199, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). The Brooke Benjamin special issue (University Park, PA, 1989).

Wang:1991:SLF

[375] C. Y. Wang, L. T. Watson, and K. A. Alexander. Spinning of a liquid film from an accelerating disc. IMA Journal of Applied Mathematics, 46 (3):201–210, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Straughan:1991:CCR

[376] Brian Straughan. Convection caused by radiation through the layer. IMA Journal of Applied Mathematics, 46 (3):211–216, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Asghar:1991:PSD

[377] S. Asghar, Bashir Ahmad, and M. Ayub. Point-source diffraction by an absorbing half-plane. IMA Journal of Applied Mathematics, 46(3):217– 224, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hanke:1991:ARR

[378] Michael Hanke. On the asymptotic representation of a regularization approach to nonlinear semiexplicit higher-index differential-algebraic equations. IMA Journal of Applied Mathematics, 46(3):225–245, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cade:1991:STC

[379] R. Cade. The stability theory of charge distributions on conductors. IMA Journal of Applied Mathematics, 46 (3):247–267, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Seddougui:1991:IGV

[380] Sharon O. Seddougui and Andrew P. Bassom. On the instability of Görtler vortices to nonlinear travelling waves.

IMA Journal of Applied Mathematics, 46(3):269–296, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wicks:1991:DTC

[381] P. J. Wicks. The diffusion of two chemically reacting pollutant discharge plumes in a uniform flow. IMA Journal of Applied Mathematics, 47(1):1–21, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

VijayaBharathi:1991:SBV

[382] L. Vijaya Bharathi and A. Chakrabarti. Solution of a boundary value problem associated with diffraction of water waves by a nearly vertical barrier. IMA Journal of Applied Mathematics, 47(1):23–32, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zhang:1991:RDM

[383] G. Zhang, J. H. Merkin, and S. K. Scott. Reaction-diffusion model for combustion with fuel consumption. I. Dirichlet boundary conditions. IMA Journal of Applied Mathematics, 47 (1):33–60, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ingham:1991:BEM

[384] D. B. Ingham, Y. Yuan, and H. Han. The boundary-element method for an improperly posed problem. IMA Journal of Applied Mathematics, 47(1):61– 79, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dennis:1991:AMC

[385] S. C. R. Dennis and S. Kocabiyik. An asymptotic matching condition for unsteady boundary-layer flows governed by the Navier–Stokes equations. IMA Journal of Applied Mathematics, 47 (1):81–98, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Webb:1991:SSN

[386] G. M. Webb and J. F. McKenzie. Similarity solutions of the nonlinear Landau damping equation using Lie group analysis. IMA Journal of Applied Mathematics, 47(1):99–108, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Mikelic:1991:PFI

[387] A. Mikelić, A. Suhadolc, and K. Veselić. On the potential flow of an ideal incompressible fluid through a porous boundary. IMA Journal of Applied Mathematics, 47(2):109–125, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McIver:1991:IPE

[388] M. McIver. An inverse problem in electromagnetic crack detection. IMA Journal of Applied Mathematics, 47 (2):127–145, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sherratt:1991:PPA

[389] Jonathan A. Sherratt. A perturbation problem arising from a mechanical model for epithelial morphogenesis. IMA Journal of Applied Mathematics, 47(2):147–162, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

James:1991:CSE

[390] E. M. James and N. G. Lloyd. A cubic system with eight small-amplitude limit cycles. IMA Journal of Applied Mathematics, 47(2):163–171, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wheeler:1991:SNA

[391] A. A. Wheeler. A strongly nonlinear analysis of the morphological instability of a freezing binary alloy: solutal convection, density change, and nonequilibrium effects. IMA Journal of Applied Mathematics, 47(2):173–192, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McNabb:1991:HCF

[392] A. McNabb and G. C. Wake. Heat conduction and finite measures for transition times between steady states. IMA Journal of Applied Mathematics, 47 (2):193–206, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Burzlaff:1991:OTO

[393] J. Burzlaff and A. D. Wood. Optical tunnelling from one-dimensional square-well potentials. IMA Journal of Applied Mathematics, 47(2):207–215, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ionescu:1991:SER

[394] Ioan R. Ionescu. Some existence results in one-dimensional dynamic viscoplasticity with work hardening. IMA Journal of Applied Mathematics, 47 (3):217–228, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ledder:1991:SPN

[395] Glenn Ledder and J. David Logan. A signalling problem for near-equilibrium flows in the Fickett-Majda model of combustion. IMA Journal of Applied Mathematics, 47(3):229–246, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Marchant:1991:IBV

[396] T. R. Marchant and N. F. Smyth. Initial-boundary value problems for the Korteweg–de Vries equation. IMA Journal of Applied Mathematics, 47 (3):247–264, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Angell:1991:CMS

[397] Thomas S. Angell and Ralph E. Kleinman. A constructive method for shape optimization: a problem in hydromechanics. IMA Journal of Applied Mathematics, 47(3):265–281, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Choi:1991:SSS

[398] Y. S. Choi, K. C. Jen, and P. J. McKenna. The structure of the solution set for periodic oscillations in a suspension bridge model. IMA Journal of Applied Mathematics, 47 (3):283–306, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tang:1991:MBS

[399] T. Tang, S. McKee, and M. W. Reeks. On a moving boundary solution to the Fokker–Planck equation for particle transport in turbulent flows with absorbing boundaries. IMA Journal of Applied Mathematics, 47(3):307–318, ???? 1991. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Reyna:1992:DDC

[400] Luis G. Reyna and Michael J. Ward. Drift-dominated current flow in forward-biased semiconductor devices. IMA Journal of Applied Mathematics, 48(1):1–21, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tao:1992:CST

[401] L. Tao, K. R. Rajagopal, and A. S. Wineman. Circular shearing and torsion of generalized neo-Hookean materials. IMA Journal of Applied Mathematics, 48(1):23–37, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wazwaz:1992:ASE

[402] Abdul-Majid Wazwaz. Asymptotic solutions of an eigenvalue problem with several second-order turning points. IMA Journal of Applied Mathematics, 48(1):39–51, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ward:1992:OTR

[403] Michael J. Ward and Eric F. Van de Velde. The onset of thermal runaway in partially insulated or cooled reactors.

IMA Journal of Applied Mathematics, 48(1):53–83, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Qin:1992:SCP

[404] Yu Qin and P. N. Kaloni. Steady convection in a porous medium based upon the Brinkman model. IMA Journal of Applied Mathematics, 48 (1):85–95, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Siew:1992:SOE

[405] P. F. Siew. Second-order effects in the induction of thin discs. IMA Journal of Applied Mathematics, 48 (1):97–106, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chen:1992:ASN

[406] Zhi Xiong Chen and Ben Yu Guo. Analytic solutions of the Nagumo equation. IMA Journal of Applied Mathematics, 48(2):107–115, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bhattacharya:1992:NSC

[407] T. Bhattacharya. A note on the space charge problem. IMA Journal of Applied Mathematics, 48(2):117–124, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gaultier:1992:EUT

[408] Maurice Gaultier and Mikel Lezaun. An existence and uniqueness theorem for the transfer of mass and heat in a rectangular cavity. IMA Journal of Applied Mathematics, 48(2):125–148, ????

1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kerr:1992:DCS

[409] Camilla E. Kerr. Diffusion-controlled smoulder propagation in a semi-infinite solid: a numerical solution. IMA Journal of Applied Mathematics, 48 (2):149–162, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1992:ISR

[410] D. S. Jones. An improved surface radiation condition. IMA Journal of Applied Mathematics, 48(2):163–193, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Steigmann:1992:EPN

[411] D. J. Steigmann. Equilibrium of prestressed networks. IMA Journal of Applied Mathematics, 48(2):195–215, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Teymur:1992:ASR

[412] M. Teymur. Assessment of the surface radiation condition by reference to acoustic scattering by a hard sphere. IMA Journal of Applied Mathematics, 48(3):217–241, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Kriegsmann:1992:ATR

[413] G. A. Kriegsmann. Acoustic target reconstruction using geometrical optics phase information. IMA Journal of Applied Mathematics, 48(3):243–247, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rubinstein:1992:NRD

[414] Jacob Rubinstein and Peter Sternberg. Nonlocal reaction-diffusion equations and nucleation. IMA Journal of Applied Mathematics, 48(3):249–264, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hilgers:1992:GCV

[415] M. G. Hilgers and A. C. Pipkin. The Graves condition for variational problems of arbitrary order. IMA Journal of Applied Mathematics, 48 (3):265–269, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Allegretto:1992:EST

[416] W. Allegretto and H. Xie. Existence of solutions for the time-dependent thermistor equations. IMA Journal of Applied Mathematics, 48(3):271–281, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1992:SER

[417] James M. Hill, Aaron J. Avagliano, and Maureen P. Edwards. Some exact results for nonlinear diffusion with absorption. IMA Journal of Applied Mathematics, 48(3):283–304, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Murphy:1992:SNC

[418] Jeremiah G. Murphy. Some new closedform solutions describing spherical inflation in compressible finite elasticity. IMA Journal of Applied Mathematics, 48(3):305–316, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cade:1992:EEF

[419] R. Cade. Electrohydrostatic equilibrium and field solutions under spacecharge conditions. IMA Journal of Applied Mathematics, 49(1):1–13, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Denier:1992:SFN

[420] James P. Denier. The structure of fully nonlinear Taylor vortices. IMA Journal of Applied Mathematics, 49 (1):15–33, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rynne:1992:WPI

[421] Bryan P. Rynne. The well-posedness of the integral equations for thin wire antennas. IMA Journal of Applied Mathematics, 49(1):35–44, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Evans:1992:TAM

[422] D. V. Evans. Trapped acoustic modes. IMA Journal of Applied Mathematics, 49(1):45–60, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grinfeld:1992:MTA

[423] M. Grinfeld, J. E. Furter, and J. C. Eilbeck. A monotonicity theorem and its application to stationary solutions of the phase field model. IMA Journal of Applied Mathematics, 49(1):61–72, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See corrigendum [450].

Tavener:1992:SNE

[424] Simon J. Tavener. Symmetric and nonsymmetric equilibria of a rod-and-spring model. IMA Journal of Applied Mathematics, 49(1):73–102, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Logan:1992:TWM

[425] J. David Logan and Steven R. Dunbar. Travelling waves in model reacting flows with reversible kinetics. IMA Journal of Applied Mathematics, 49(2):103–121, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Alderson:1992:SOC

[426] Steven G. Alderson. A solution by operational calculus to the problem of the optimization of tidal power. IMA Journal of Applied Mathematics, 49 (2):123–138, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lu:1992:ABS

[427] Chun Qing Lu, A. Dean MacGillivray, and Stuart P. Hastings. Asymptotic behaviour of solutions of a similarity equation for laminar flows in channels with porous walls. IMA Journal of Applied Mathematics, 49(2):139–162, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Colton:1992:NSI

[428] David Colton and Peter Monk. The numerical solution of an inverse scattering problem for acoustic waves. IMA Journal of Applied Mathematics, 49 (2):163–184, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gottlieb:1992:AIA

[429] H. P. W. Gottlieb. Axisymmetric isospectral annular plates and membranes. IMA Journal of Applied Mathematics, 49(2):185–192, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sleeman:1992:ABC

[430] B. D. Sleeman and Chen Hua. An analogue of Berry’s conjecture for the phase in fractal obstacle scattering. IMA Journal of Applied Mathematics, 49(3):193–202, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

OldeDaalhuis:1992:HEC

[431] A. B. Olde Daalhuis. Hyperasymptotic expansions of confluent hypergeometric functions. IMA Journal of Applied Mathematics, 49(3):203–216, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Spigler:1992:RDM

[432] Renato Spigler and Damián H. Zanette. Reaction-diffusion models from the Fokker–Planck formulation of chemical processes. IMA Journal of Applied Mathematics, 49(3):217–229, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wazwaz:1992:RAS

[433] Abdul-Majid Wazwaz. Resonance and asymptotic solutions of a singular perturbation problem. IMA Journal of Applied Mathematics, 49(3):231–244, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Movchan:1992:ABS

[434] A. B. Movchan and S. A. Nazarov. Asymptotic behaviour of stress-strain state in the vicinity of sharp defects in an elastic body. IMA Journal of Applied Mathematics, 49(3):245–272, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sisson:1992:SHD

[435] R. A. Sisson, A. Swift, G. C. Wake, and B. F. Gray. The self-heating of damp cellulosic materials. I. High thermal conductivity and diffusivity. IMA Journal of Applied Mathematics, 49 (3):273–291, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ryder:1992:CEE

[436] E. Ryder and D. F. Parker. Coupled evolution equations for axially inhomogeneous optical fibres. IMA Journal of Applied Mathematics, 49 (3):293–309, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lodge:1992:CHP

[437] A. S. Lodge, W. G. Pritchard, and L. R. Scott. Corrigendum to: “The hole-pressure problem” [IMA J. Appl. Math. 46 (1991), no. 1–2, 39–66; MR 92b:76025]. IMA Journal of Applied Mathematics, 49(3):311, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See

Gross:1993:HRF

[438] Peter Gross. On harmonic resonance in forced nonlinear oscillators exhibiting a Hopf bifurcation. IMA Journal of Applied Mathematics, 50(1):1–12, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jones:1993:CDE

[439] M. C. W. Jones. On coupled differential equations which model the evolution of interacting capillary-gravity wave modes and related questions of stability. IMA Journal of Applied Mathematics, 50(1):13–28, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zhang:1993:DSM

[440] Ling Hai Zhang. Decay of solutions of the multidimensional generalized Kuramoto–Sivashinsky system. IMA Journal of Applied Mathematics, 50(1):29–42, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Merkin:1993:CRD

[441] J. H. Merkin, D. J. Needham, and S. K. Scott. Coupled reaction-diffusion waves in an isothermal autocatalytic chemical system. IMA Journal of Applied Mathematics, 50(1):43–76, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Skinner:1993:MAE

[442] L. A. Skinner. Matched asymptotic expansions of integrals. IMA Journal of Applied Mathematics, 50(1):77–90, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1993:CUC

[443] Ronald Smith. The calculation of upstream concentrations from downstream data for a two-layer flow. IMA Journal of Applied Mathematics, 50 (2):91–105, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gottlieb:1993:IAP

[444] H. P. W. Gottlieb. Inhomogeneous annular plates with exactly beam-like radial spectra. IMA Journal of Applied Mathematics, 50(2):107–112, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ingham:1993:SNI

[445] D. B. Ingham and Y. Yuan. The solution of a nonlinear inverse problem in heat transfer. IMA Journal of Applied Mathematics, 50(2):113–132, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pakdemirli:1993:BLF

[446] Mehmet Pakdemirli. Boundary layer flow of power-law fluids past arbitrary profiles. IMA Journal of Applied Mathematics, 50(2):133–148, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wilson:1993:LPF

[447] S. K. Wilson. The levelling of paint films. IMA Journal of Applied Mathematics, 50(2):149–166, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bassom:1993:NSF

[448] A. P. Bassom, P. A. Clarkson, and A. C. Hicks. Numerical studies of the fourth Painlevé equation. IMA Journal of Applied Mathematics, 50 (2):167–193, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1993:CSI

[449] James M. Hill. Cylindrical and spherical inflation in compressible finite elasticity. IMA Journal of Applied Mathematics, 50(2):195–201, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grinfeld:1993:CMT

[450] M. Grinfeld, J. E. Furter, and J. C. Eilbeck. Corrigendum to: “A monotonicity theorem and its application to stationary solutions of the phase field model” [IMA J. Appl. Math. 49 (1992), no. 1, 61–72; MR 93j:80003]. IMA Journal of Applied Mathematics, 50(2):203–204, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [423].

Keady:1993:FCL

[451] Grant Keady and Alex McNabb. Functions with constant Laplacian satisfying homogeneous Robin boundary conditions. IMA Journal of Applied Mathematics, 50(3):205–224, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pipkin:1993:RED

[452] A. C. Pipkin. Relaxed energy densities for small deformations of membranes. IMA Journal of Applied Mathematics, 50(3):225–237, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Choi:1993:LTB

[453] Y. S. Choi and Roger Lui. Longtime behaviour of solutions of an electrophoretic model with a single reaction. IMA Journal of Applied Mathematics, 50(3):239–252, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

LanzadeCristoforis:1993:LDN

[454] Massimo Lanza de Cristoforis. The large deformation of non-linearly elastic rings in a two-dimensional compressible flow. IMA Journal of Applied Mathematics, 50(3):253–283, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sisson:1993:SHD

[455] R. A. Sisson, A. Swift, G. C. Wake, and B. F. Gray. The self-heating of damp cellulosic materials. II. On the steady states of the spatially distributed case. IMA Journal of Applied Mathematics, 50(3):285–306, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zhao:1993:PKC

[456] X.-Q. Zhao and B. D. Sleeman. Permanence in Kolmogorov competition models with diffusion. IMA Journal of Applied Mathematics, 51(1):1–11, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Otto:1993:SFA

[457] S. R. Otto. Stability of the flow around a cylinder: the spin-up problem. IMA Journal of Applied Mathematics, 51 (1):13–26, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dando:1993:CGP

[458] Andrew H. Dando and Sharon O. Seddougui. The compressible Görtler problem in two-dimensional boundary layers. IMA Journal of Applied Mathematics, 51(1):27–67, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zhang:1993:RDM

[459] G. Zhang, J. H. Merkin, and S. K. Scott. Reaction-diffusion model for combustion with fuel consumption. II. Robin boundary conditions. IMA Journal of Applied Mathematics, 51 (1):69–93, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Colegrave:1993:LSL

[460] R. K. Colegrave and J. M. Hyde. The Lanchester square-law model extended to a (2, 2) conflict. IMA Journal of Applied Mathematics, 51(2):95–109, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cimatti:1993:EEH

[461] Giovanni Cimatti and Eduard Feireisl. The equations of electrohydrodynamics with highly irregular initial data. IMA Journal of Applied Mathematics, 51(2):111–122, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Power:1993:CDL

[462] Henry Power. The completed double layer boundary integral equation method for two-dimensional Stokes flow. IMA Journal of Applied Mathematics, 51(2):123–145, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Guo:1993:HSE

[463] Ann Guo and Barbara Abraham-Shrauner. Hidden symmetries of energy-conserving differential equations. IMA Journal of Applied Mathematics, 51(2):147–153, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wicks:1993:ORS

[464] P. J. Wicks and Ronald Smith. Oxidation of reactive solute released into shear flow. IMA Journal of Applied Mathematics, 51(2):155–168, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

El-Ali:1993:BVP

[465] K. El-Ali and C. Atkinson. Boundary value problems involving Herschel– Bulkley material. IMA Journal of Applied Mathematics, 51(2):169–186, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Smith:1993:OUH

[466] Ronald Smith. Optimum use of holding tanks to avoid pollution surges in narrow estuaries. IMA Journal of Applied Mathematics, 51(2):187–199, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hahner:1993:UTI

[467] Peter Hähner. A uniqueness theorem in inverse scattering of elastic waves. IMA Journal of Applied Mathematics, 51(3):201–215, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McIntosh:1993:SAM

[468] A. C. McIntosh. A Semenov approach to the modelling of thermal runaway of damp combustible material. IMA Journal of Applied Mathematics, 51 (3):217–237, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Paullet:1993:EUS

[469] Joseph E. Paullet. Existence and uniqueness of a solution for a BVP from developmental biology. IMA Journal of Applied Mathematics, 51 (3):239–249, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Choi:1993:SSS

[470] Y. S. Choi and Xun Yu. Steadystate solution for electroplating. IMA Journal of Applied Mathematics, 51 (3):251–267, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Metcalf:1993:RDW

[471] M. J. Metcalf, J. H. Merkin, and S. K. Scott. Reaction-diffusion waves in coupled isothermal autocatalytic chemical systems. IMA Journal of Applied Mathematics, 51(3):269–298, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gourley:1993:ITW

[472] S. A. Gourley and N. F. Britton. Instability of travelling wave solutions of a population model with nonlocal effects. IMA Journal of Applied Mathematics, 51(3):299–310, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arrigo:1994:NSR

[473] D. J. Arrigo, J. M. Hill, and P. Broadbridge. Nonclassical symmetry reductions of the linear diffusion equation with a nonlinear source. IMA Journal of Applied Mathematics, 52(1):1–24, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jansons:1994:DSB

[474] Kalvis M. Jansons. Determination of the slip boundary condition between a gas and a solid surface with a random array of rough patches. IMA Journal of Applied Mathematics, 52 (1):25–29, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lowe:1994:DCP

[475] Bruce D. Lowe and William Rundell. The determination of a coefficient in a parabolic equation from input sources. IMA Journal of Applied Mathematics, 52(1):31–50, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Blennerhassett:1994:NHW

[476] P. J. Blennerhassett and Andrew P. Bassom. Nonlinear high-wavenumber Bénard convection. IMA Journal of Applied Mathematics, 52(1):51–77, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sherratt:1994:SAT

[477] Jonathan A. Sherratt. On the speed of amplitude transition waves in reactiondiffusion systems of λ-ω type. IMA Journal of Applied Mathematics, 52 (1):79–92, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Stavre:1994:SJI

[478] Ruxandra Stavre. Study of a jet incident on a porous wall in a gravity field. IMA Journal of Applied Mathematics, 52(1):93–103, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Choi:1994:EIF

[479] Y. S. Choi, Roger Lui, and Xun Yu. Electrolysis and isoelectric focusing. IMA Journal of Applied Mathematics, 52(2):105–122, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Arnold:1994:STP

[480] J. M. Arnold. Stability theory for periodic pulse train solutions of the nonlinear Schrödinger equation. IMA Journal of Applied Mathematics, 52 (2):123–140, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pincombe:1994:MHM

[481] A. H. Pincombe and N. F. Smyth. Microwave heating of materials with power law temperature dependencies.

IMA Journal of Applied Mathematics, 52(2):141–176, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

vanDuijn:1994:FRT

[482] C. J. van Duijn and Peter Knabner. Flow and reactive transport in porous media induced by well injection: similarity solution. IMA Journal of Applied Mathematics, 52(2):177–200, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Ledder:1994:ASN

[483] Glenn Ledder. Analytical solution of a near-equidiffusional flame problem. IMA Journal of Applied Mathematics, 52(3):201–220, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Atkinson:1994:SPA

[484] C. Atkinson and R. V. Craster. A singular perturbation approach to integral equations occurring in poroelasticity. IMA Journal of Applied Mathematics, 52(3):221–252, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Craster:1994:TRF

[485] R. V. Craster. Two related free boundary problems. IMA Journal of Applied Mathematics, 52(3):253–270, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Bonnet-Bendhia:1994:HFA

[486] Anne-Sophie Bonnet-Bendhia and Rabia Djellouli. High-frequency asymptotics of guided modes in optical fibres. IMA Journal of Applied Mathematics, 52(3):271–287, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Aron:1994:CPR

[487] M. Aron. On a class of plane radial deformations of compressible non-linearly elastic solids. IMA Journal of Applied Mathematics, 52(3):289–296, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Pipkin:1994:RED

[488] A. C. Pipkin. Relaxed energy densities for large deformations of membranes. IMA Journal of Applied Mathematics, 52(3):297–308, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cade:1994:IEA

[489] R. Cade. On integral equations of axisymmetric potential theory. IMA Journal of Applied Mathematics, 53 (1):1–25, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gross:1994:OTF

[490] Peter Gross. On oscillation types in forced nonlinear oscillators close to harmonic resonance. IMA Journal of Applied Mathematics, 53(1):27–43, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rothmayer:1994:ARA

[491] A. P. Rothmayer, D. W. Black, and W. Hao. Asymptotic Reynoldsaveraged closure of nonlinear ensembles. IMA Journal of Applied Mathematics, 53(1):45–73, ???? 1994.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Coward:1994:SOT

[492] Adrian V. Coward and Demetrios T. Papageorgiou. Stability of oscillatory two-phase Couette flow. IMA Journal of Applied Mathematics, 53 (1):75–93, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Tutty:1994:SOK

[493] O. R. Tutty and G. Denuault. Secondorder kinetics for EC′ reactions at a spherical microelectrode. IMA Journal of Applied Mathematics, 53(1):95–109, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rajagopal:1994:FCS

[494] K. R. Rajagopal and Yu-Ning Huang. Finite circumferential shearing of nonlinear solids in the context of thermoelasticity. IMA Journal of Applied Mathematics, 53(2):111–125, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Herbert:1994:DCS

[495] D. M. Herbert, C. E. Kerr, and J. Adler. Diffusion-controlled smoulder propagation in a semi-infinite solid: the asymptotic temperature distribution. IMA Journal of Applied Mathematics, 53(2):127–136, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Needham:1994:RDM

[496] D. J. Needham, A. C. King, and J. H. Merkin. A reaction-diffusion model for autocatalytic polymerization. I. Permanent form travelling waves. IMA Journal of Applied Mathematics, 53 (2):137–149, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grundy:1994:ASF

[497] R. E. Grundy and H. R. Allen. The asymptotic solution of a family of boundary value problems involving exponentially small terms. IMA Journal of Applied Mathematics, 53 (2):151–168, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jansons:1994:GDN

[498] Kalvis M. Jansons. A general duality for nonlinear two-dimensional conduction. IMA Journal of Applied Mathematics, 53(2):169–172, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Yuan:1994:EUC

[499] Guang Wei Yuan. Existence and uniqueness of the classical solution for the steady-state electrochemical machining problem. IMA Journal of Applied Mathematics, 53(2):173–190, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Blackaby:1994:TSV

[500] Nicholas D. Blackaby. Tollmien-Schlichting/vortex interactions in compressible boundary-layer flows. IMA Journal of Applied Mathematics, 53 (2):191–214, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jansons:1994:SVC

[501] Kalvis M. Jansons. The second virial coefficient for a strongly Brownian suspension of arbitrary convex bodies with model interaction potentials. IMA Journal of Applied Mathematics, 53 (3):215–219, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gajjar:1994:NEF

[502] J. S. B. Gajjar. Nonlinear evolution of a first mode oblique wave in a compressible boundary layer. I. Heated/cooled walls. IMA Journal of Applied Mathematics, 53(3):221–248, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Broadbridge:1994:IFO

[503] P. Broadbridge and P. Tritscher. An integrable fourth-order nonlinear evolution equation applied to thermal grooving of metal surfaces. IMA Journal of Applied Mathematics, 53 (3):249–265, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cimatti:1994:UPM

[504] Giovanni Cimatti. Uniqueness for a problem of mixing in electrohydrodynamics. IMA Journal of Applied Mathematics, 53(3):267–273, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

RegoSilva:1994:NIF

[505] J. J. Rěgo Silva, H. Power, and L. C. Wrobel. A new integral formulation for 3D elastic wave propagation. IMA Journal of Applied Mathematics, 53 (3):275–293, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1994:PFC

[506] R. Hill and J. H. Merkin. Pattern formation in a coupled cubic autocatalator system. IMA Journal of Applied Mathematics, 53(3):295–322, ???? 1994. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Graham-Eagle:1995:NEA

[507] James Graham-Eagle. A nonlocal equation arising in the theory of wet combustion. IMA Journal of Applied Mathematics, 54(1):1–8, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Xu:1995:RCS

[508] Yongzhi Xu. Radiation condition and scattering problem for time-harmonic acoustic waves in a stratified medium with a nonstratified inhomogeneity. IMA Journal of Applied Mathematics, 54(1):9–29, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Homberg:1995:MMP

[509] D. Hömberg. A mathematical model for the phase transitions in eutectoid carbon steel. IMA Journal of Applied Mathematics, 54(1):31–57, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Rawlins:1995:BCW

[510] A. D. Rawlins. A bifurcated circular waveguide problem. IMA Journal of Applied Mathematics, 54(1):59–81, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McIntosh:1995:TRC

[511] A. C. McIntosh and J. F. Griffiths. On the thermal runaway of combustible fluids in lagging material. IMA Journal of Applied Mathematics, 54(1):83–96, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Talbot:1995:IHS

[512] D. R. S. Talbot, J. R. Willis, and V. Nesi. On improving the Hashin-Shtrikman bounds for the effective properties of three-phase composite media. IMA Journal of Applied Mathematics, 54(1):97–107, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1995:NFE

[513] James M. Hill and Daniel J. Arrigo. New families of exact solutions for finitely deformed incompressible elastic materials. IMA Journal of Applied Mathematics, 54(2):109–123, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Grindrod:1995:SSE

[514] Peter Grindrod. Some semilinear elliptic equations with nonlinear boundary conditions: radial solutions and simple symmetry-breaking. IMA Journal of Applied Mathematics, 54(2):125–137, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Martin:1995:AAF

[515] P. A. Martin. Asymptotic approximations for functions defined by series, with some applications to the theory of guided waves. IMA Journal of Applied Mathematics, 54(2):139–157, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chapman:1995:STW

[516] S. J. Chapman. Stability of travelling waves in models of superconductivity. IMA Journal of Applied Mathematics, 54(2):159–169, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Charalambopoulos:1995:RSS

[517] Antonios Charalambopoulos. The reconstruction of the surface of scatterers with continuous curvature via lowfrequency moments. IMA Journal of Applied Mathematics, 54(2):171–201, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Qu:1995:ATS

[518] Chang Zheng Qu. Allowed transformations and symmetry classes of variable coefficient Burgers equations. IMA Journal of Applied Mathematics, 54 (3):203–225, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Witelski:1995:SMP

[519] Thomas P. Witelski. Stopping and merging problems for the porous media equation. IMA Journal of Applied Mathematics, 54(3):227–243, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Wu:1995:LHS

[520] Charles C. K. Wu and Yue-Kuen Kwok. Linear hydrodynamical stability of two-layer inclined flow with viscosity and density stratifications. IMA Journal of Applied Mathematics, 54 (3):245–256, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Hill:1995:ECP

[521] R. Hill and J. H. Merkin. The effects of coupling on pattern formation in a simple autocatalytic system. IMA Journal of Applied Mathematics, 54(3):257–281, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Steigmann:1995:CP

[522] David Steigmann. A. C. Pipkin. IMA Journal of Applied Mathematics, 54(3):283, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). Obituary.

Pipkin:1995:ETI

[523] A. C. Pipkin. Energy theorems for infinitesimal plane wrinkling of inextensible networks. IMA Journal of Applied Mathematics, 54(3):285–299, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lakhtakia:1995:CEC

[524] Akhlesh Lakhtakia and Werner S. Weiglhofer. On a constraint on the electromagnetic constitutive relations of nonhomogeneous linear media. IMA Journal of Applied Mathematics, 54 (3):301–306, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Steigmann:1995:EMS

[525] D. J. Steigmann and Dongqing Li. Energy-minimizing states of capillary systems with bulk, surface, and line phases. IMA Journal of Applied Mathematics, 55(1):1–17, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Perumpanani:1995:PDR

[526] Abbey J. Perumpanani, Jonathan A. Sherratt, and Philip K. Maini. Phase differences in reaction-diffusion-advection systems and applications to morphogenesis. IMA Journal of Applied Mathematics, 55(1):19–33, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Abrahams:1995:FCW

[527] I. D. Abrahams and J. B. Lawrie. On the factorization of a class of Wiener– Hopf kernels. IMA Journal of Applied Mathematics, 55(1):35–47, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Edwards:1995:ECD

[528] David A. Edwards and Donald S. Cohen. The effect of a changing diffusion coefficient in super-case II polymerpenetrant systems. IMA Journal of Applied Mathematics, 55(1):49–66, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Dassios:1995:PSE

[529] George Dassios and Gregory Kamvyssas. Point source excitation in direct and inverse scattering: the soft and the hard small sphere. IMA Journal of Applied Mathematics, 55(1):67–84, ????

1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Li:1995:IPC

[530] B. C. Li and S. Syngellakis. On improperly posed Cauchy problems and their approximate solution. IMA Journal of Applied Mathematics, 55(1):85–95, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

El-Dib:1995:SPR

[531] Yusry O. El-Dib. On stability of parametric resonances of nonlinear surface waves propagating between two superposed electrified fluids. IMA Journal of Applied Mathematics, 55 (2):97–116, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Merkin:1995:PIR

[532] J. H. Merkin, M. J. Metcalf, D. J. Needham, and S. K. Scott. The propagation of isothermal reaction-diffusion waves in coupled autocatalytic systems. IMA Journal of Applied Mathematics, 55(2):117–133, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

vanVuuren:1995:ETP

[533] Jan H. van Vuuren. The existence of travelling plane waves in a general class of competition-diffusion systems. IMA Journal of Applied Mathematics, 55(2):135–148, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jansons:1995:PDT

[534] Kalvis M. Jansons and L. C. G. Rogers. Probability and dispersion theory. IMA Journal of Applied Mathematics, 55 (2):149–162, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Leppington:1995:IBS

[535] F. G. Leppington. The interaction between sound and circulatory fluid motion as a problem in matched expansions. IMA Journal of Applied Mathematics, 55(2):163–186, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Linton:1995:MMB

[536] C. M. Linton. Multipole methods for boundary-value problems involving a sphere in a tube. IMA Journal of Applied Mathematics, 55(2):187–204, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Cade:1995:DSI

[537] R. Cade. On discontinuous solutions of the integral equations of electrostatics. IMA Journal of Applied Mathematics, 55(3):205–220, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Howe:1995:DSV

[538] M. S. Howe. Damping of sound and vibration by flow nonlinearity in the apertures of a perforated elastic screen. IMA Journal of Applied Mathematics, 55(3):221–242, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See corrigendum [569].

Kriegsmann:1995:MHC

[539] G. A. Kriegsmann and B. A. Wagner. Microwave heating of carbon-coated ceramic fibres: a mathematical model. IMA Journal of Applied Mathematics, 55(3):243–255, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Adler:1995:TDL

[540] J. Adler. Time-dependent laser ignition of a combustible stagnant boundary layer. IMA Journal of Applied Mathematics, 55(3):257–268, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

McNabb:1995:BSD

[541] A. McNabb. Block-size determination in fractured porous media. IMA Journal of Applied Mathematics, 55 (3):269–279, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gutman:1995:GCS

[542] Semion Gutman, Michael V. Klibanov, and Alexander V. Tikhonravov. Global convexity in a single-source 3-D inverse scattering problem. IMA Journal of Applied Mathematics, 55(3):281–302, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Popple:1995:SFD

[543] D. F. Popple and F. G. Leppington. The sound field due to a periodic array of forced wave-bearing strips. IMA Journal of Applied Mathematics, 55 (3):303–315, ???? 1995. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Levine:1996:SUD

[544] Harold Levine and Alex McNabb. Some unsteady diffusion and absorption problems. IMA Journal of Applied Mathematics, 56(1):1–19, February 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_01/560001.sgm.abs.html.

Donescu:1996:EDH

[545] Pompiliu Donescu and Lawrence N. Virgin. Efficient determination of higher-order periodic solutions using nmode harmonic balance. IMA Journal of Applied Mathematics, 56(1):21–32, February 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_01/560021.sgm.abs.html.

King:1996:AAC

[546] J. R. King and M. G. Meere. Asymptotic analysis of a combined dissociative and kick-out diffusion mechanism. IMA Journal of Applied Mathematics, 56(1):33–63, February 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_01/560033.sgm.abs.html.

Needham:1996:RDM

[547] D. J. Needham, A. C. King, and J. H. Merkin. A reaction-diffusion model for autocatalytic polymerization: II. The initial value problem. IMA Journal of Applied Mathematics, 56(1):65–86, February 1996. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_01/560065.sgm.abs.html.

Buckmaster:1996:TSS

[548] J. Buckmaster. A theory of shallow smoulder waves. IMA Journal of Applied Mathematics, 56(1):87–102, February 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_01/560087.sgm.abs.html.

Jansons:1996:MSB

[549] Kalvis M. Jansons. A mechanism for slip between a liquid and a random surface with trapped gas due to a temperature gradient. IMA Journal of Applied Mathematics, 56(1):103–108, February 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_01/560103.sgm.abs.html.

Scarpetta:1996:EAR

[550] Edoardo Scarpetta and Mezhlum A. Sumbatyan. Explicit analytical results for one-mode oblique penetration into a periodic array of screens. IMA Journal of Applied Mathematics, 56(2):109–120, April 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gourley:1996:IPP

[551] S. A. Gourley. Instability in a predatorprey system with delay and spatial averaging. IMA Journal of Applied Mathematics, 56(2):121–132, April 1996.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Markov:1996:SDH

[552] K. Z. Markov, D. R. S. Talbot, and J. R. Willis. On stationary diffusion in heterogeneous media. IMA Journal of Applied Mathematics, 56 (2):133–144, April 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Davidson:1996:LGB

[553] F. A. Davidson and B. P. Rynne. Local and global behaviour of steadystate solutions of the Sel′kov model. IMA Journal of Applied Mathematics, 56(2):145–155, April 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Marchant:1996:SIE

[554] T. R. Marchant and N. F. Smyth. Soliton interaction for the extended Korteweg–de Vries equation. IMA Journal of Applied Mathematics, 56 (2):157–176, April 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Jansons:1996:GDD

[555] Kalvis M. Jansons. The general drift decomposition for diffusion with advection. IMA Journal of Applied Mathematics, 56(2):177–181, April 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Seddougui:1996:ESN

[556] Sharon O. Seddougui and Andrew P. Bassom. The effects of suction on the nonlinear stability of a threedimensional compressible boundary layer. IMA Journal of Applied Mathematics, 56(2):183–206, April 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Vanden-Broeck:1996:PWC

[557] J. M. Vanden-Broeck. Periodic waves with constant vorticity in water of infinite depth. IMA Journal of Applied Mathematics, 56(3):207– 217, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560207.sgm.abs.html.

Wang:1996:BEA

[558] Shin-Hwa H. Wang. Bifurcation of an equation arising in porousmedium combustion. IMA Journal of Applied Mathematics, 56(3):219– 234, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560219.sgm.abs.html.

Abraham-Shrauner:1996:HSN

[559] Barbara Abraham-Shrauner. Hidden symmetries and nonlocal group generators for ordinary differential equations. IMA Journal of Applied Mathematics, 56(3):235–252, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560235.sgm.abs.html.

Craster:1996:YSH

[560] R. V. Craster. Yield surfaces for Herschel–Bulkley flows in complex geometries. IMA Journal of Applied Mathematics, 56(3):253–276, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560253.sgm.abs.html.

Shepherd:1996:SDN

[561] R. Shepherd and R. J. Wiltshire. Spectral decompositions in nonlinear coupled diffusion. IMA Journal of Applied Mathematics, 56(3):277– 287, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560277.sgm.abs.html.

Sherratt:1996:ADV

[562] Jonathan A. Sherratt and Ben P. Marchant. Algebraic decay and variable speeds in wavefront solutions of a scalar reaction-diffusion equation. IMA Journal of Applied Mathematics, 56(3):289–302, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560289.sgm.abs.html.

Haughton:1996:AEM

[563] D. M. Haughton. Axisymmetric elastic membranes subjected to fluid loading. IMA Journal of Applied Mathematics, 56(3):303–320, June 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_56/
Issue_03/560303.sgm.abs.html.

King:1996:AAG

[564] J. R. King and C. P. Please. Asymptotic analysis of the growth of cake layers in filters. IMA Journal of Applied Mathematics, 57(1):1–28, August 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_01/570001.sgm.abs.html.

Crompton:1996:GAF

[565] Stephen Crompton and Peter Grindrod. A geometric approach to fingering instabilities for reaction fronts in fully coupled geochemical systems. IMA Journal of Applied Mathematics, 57(1):29–40, August 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL http://www3.oup.co.
uk/imamat/hdb/Volume_57/Issue_01/
570029.sgm.abs.html.

Talbot:1996:TPB

[566] D. R. S. Talbot and J. R. Willis. Three-point bounds for the overall properties of a nonlinear composite dielectric. IMA Journal of Applied Mathematics, 57(1):41–52, August 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_01/570041.sgm.abs.html.

Galaktionov:1996:BBI

[567] V. A. Galaktionov and J. R. King. On the behaviour of blow-up interfaces for an inhomogeneous filtration equation. IMA Journal of Applied Mathematics, 57(1):53–77, August 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_57/
                                        Issue_01/570053.sgm.abs.html.

Smith:1996:NAS

[568] W. R. Smith, S. D. Howison, and D. F. Mayers. Numerical and asymptotic solution of a sixth-order nonlinear diffusion equation and related coupled systems. IMA Journal of Applied Mathematics, 57(1):79–98, August 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_57/
                                        Issue_01/570079.sgm.abs.html.

Howe:1996:CDS

[569] M. S. Howe. Corrigendum to: “Damping of sound and vibration by flow nonlinearity in the apertures of a perforated elastic screen” [IMA J. Appl. Math. 55 (1995), no. 3, 221–242; MR 97b:76114]. IMA Journal of Applied Mathematics, 57(1):99, August 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). See [538].

Xu:1996:IIF

[570] Jian-Jun Xu. Interfacial instabilities and fingering formation in Hele–Shaw flow. IMA Journal of Applied Mathematics, 57(2):101–135, October 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Teymur:1996:NHO

[571] M. Teymur. A note on higher-order surface radiation conditions. IMA Journal of Applied Mathematics, 57 (2):137–163, October 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

GarciaReimbert:1996:ERL

[572] C. Garcia Reimbert, A. A. Minzoni, and N. F. Smyth. Effect of radiation losses on hotspot formation and propagation in microwave heating. IMA Journal of Applied Mathematics, 57 (2):165–179, October 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Craster:1996:CMI

[573] R. V. Craster. Conformal mappings involving curvilinear quadrangles. IMA Journal of Applied Mathematics, 57 (2):181–191, October 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Peng:1996:FSS

[574] Weidong Peng. Free-surface shapes of a jet injecting through a porous wall. IMA Journal of Applied Mathematics, 57(2):193–210, October 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_02/570193.sgm.abs.html.

Sanchez-Garduno:1996:SAA

[575] Faustino Sánchez-Garduño, Philip K. Maini, and M. E. Kappos. A shooting argument approach to a Sharptype solution for nonlinear degenerate Fisher-KPP equations. IMA Journal of Applied Mathematics, 57(3):211–221, December 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_03/570211.sgm.abs.html.

Fusco:1996:MFE

[576] D. Fusco and N. Manganaro. A method for finding exact solutions to hyperbolic systems of first-order PDEs. IMA Journal of Applied Mathematics, 57(3):223–242, December 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_03/570223.sgm.abs.html.

Krutitskii:1996:FSF

[577] P. A. Krutitskii. Fast stratified flow over several obstacles, including wings. IMA Journal of Applied Mathematics, 57(3):243–256, December 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_03/570243.sgm.abs.html.

Truscott:1996:EDA

[578] J. E. Truscott, J. Brindley, A. McIntosh, and J. Griffiths. The effect of diffusion on the autoignition of combustible fluids in insulation materials. IMA Journal of Applied Mathematics, 57(3):257–271, December 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_03/570257.sgm.abs.html.

Merkin:1996:PTW

[579] J. H. Merkin and M. A. Sadiq. The propagation of travelling waves in an open cubic autocatalytic chemical system. IMA Journal of Applied Mathematics, 57(3):273–309, December 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_03/570273.sgm.abs.html.

Saccomandi:1996:NID

[580] Giuseppe Saccomandi. A note on inhomogeneous deformations of nonlinear elastic layers. IMA Journal of Applied Mathematics, 57(3):311–324, December 1996. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_57/
Issue_03/570311.sgm.abs.html.

Hood:1997:SAB

[581] Simon Hood. On symmetry- and ansatz-based methods for integrating ordinary differential equations. IMA Journal of Applied Mathematics, 58 (1):1–18, February 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

www3.oup.co.uk/imamat/hdb/Volume_
58/Issue_01/580001.sgm.abs.html.

Mureithi:1997:EBU

[582] Eunice W. Mureithi, James P. Denier, and Jillian A. K. Stott. The effect of buoyancy on upper-branch Tollmien-Schlichting waves. IMA Journal of Applied Mathematics, 58(1):19– 50, February 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_01/580019.sgm.abs.html.

Khater:1997:BTP

[583] A. H. Khater, R. S. Ibrahim, A. B. Shamardan, and D. K. Callebaut. Bäcklund transformations and Painlevé analysis: exact solutions for a Grad-Shafranov-type magnetohydrodynamic equilibrium. IMA Journal of Applied Mathematics, 58(1):51–69, February 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_58/
                                        Issue_01/580051.sgm.abs.html.

MunozRivera:1997:ESC

[584] Jaime E. Muñoz Rivera and Milton de Lacerda Oliveira. Exponential stability for a contact problem in thermoelasticity. IMA Journal of Applied Mathematics, 58(1):71–82, February 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_58/
                                        Issue_01/580071.sgm.abs.html.

Athanasiadis:1997:ESC

[585] Christodoulos Athanasiadis and Ioannis G. Stratis. Electromagnetic scattering by a chiral obstacle. IMA Journal of Applied Mathematics, 58(1):83– 91, February 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_58/
                                        Issue_01/580083.sgm.abs.html.

Rogoff:1997:ARP

[586] Z. M. Rogoff and R. H. Tew. Acoustic radiation from a phased line source at a fluid-solid boundary. IMA Journal of Applied Mathematics, 58(2):93– 110, April 1997. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_02/580093.sgm.abs.html.

Wang:1997:LFT

[587] C. Y. Wang. Longitudinal flow through an array of rectangular cylinders. IMA Journal of Applied Mathematics, 58(2):111–120, April 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_02/580111.sgm.abs.html.

Evans:1997:FLM

[588] J. D. Evans, A. B. Tayler, and J. R. King. Finite-length mask effects in the isolation oxidation of silicon. IMA Journal of Applied Mathematics, 58(2):121–146, April 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_02/580121.sgm.abs.html.

Elnagar:1997:OSC

[589] Gamal N. Elnagar and Ahmed Khamayseh. On the optimal spectral Chebyshev solution of a controlled nonlinear dynamical system. IMA Journal of Applied Mathematics, 58(2):147– 157, April 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_02/580147.sgm.abs.html.

Cox:1997:LWT

[590] Stephen M. Cox. Long-wavelength thermal convection in a weak shear flow. IMA Journal of Applied Mathematics, 58(2):159–184, April 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL http://www3.oup.co.
uk/imamat/hdb/Volume_58/Issue_02/
580159.sgm.abs.html.

Kay:1997:ADR

[591] Anthony Kay. Advection-diffusion in reversing and oscillating flows. 2. Flows with multiple reversals. IMA Journal of Applied Mathematics, 58(3):185– 210, June 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_03/580185.sgm.abs.html.

GuidonePeroli:1997:MCL

[592] G. Guidone Peroli and E. G. Virga. Modelling the capillary locking of point defects in nematic liquid crystals. IMA Journal of Applied Mathematics, 58(3):211–236, June 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_03/580211.sgm.abs.html.

Davidson:1997:TWR

[593] F. A. Davidson, B. D. Sleeman, and J. W. Crawford. Travelling waves in a reaction-diffusion system modelling fungal mycelia. IMA Journal of Applied Mathematics, 58(3):237– 257, June 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_03/580237.sgm.abs.html.

Gao:1997:EFD

[594] Xin-Lin Gao and Satya N. Atluri. An exact finite deformation elasto-plastic solution for the outside-in free eversion problem of a tube of elastic linear-hardening material. IMA Journal of Applied Mathematics, 58(3):259– 275, June 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_58/
Issue_03/580259.sgm.abs.html.

Lawrie:1997:SFL

[595] J. B. Lawrie and I. D. Abrahams. Scattering of fluid loaded elastic plate waves at the vertex of a wedge of arbitrary angle. I. Analytic solution. IMA Journal of Applied Mathematics, 59(1):1– 23, August 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_01/590001.sgm.abs.html.

Kuznetsov:1997:WST

[596] N. Kuznetsov and O. Motygin. On the waveless statement of the twodimensional Neumann-Kelvin problem for a surface-piercing body. IMA Journal of Applied Mathematics, 59(1):25– 42, August 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_01/590025.sgm.abs.html.

Green:1997:IOF

[597] M. R. Green and F. G. Leppington. The inner and outer flow fields due to a pair of spheres moving unsteadily through a compressible fluid: rectilinear motion. IMA Journal of Applied Mathematics, 59(1):43–69, August 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_01/590043.sgm.abs.html.

Liu:1997:CES

[598] F. Liu and D. L. S. McElwain. A computationally efficient solution technique for moving-boundary problems in finite media. IMA Journal of Applied Mathematics, 59(1):71–84, August 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_01/590071.sgm.abs.html.

Chudinovich:1997:WSI

[599] Igor Chudinovich and Christian Constanda. Weak solutions of interior boundary-value problems for plates with transverse shear deformation. IMA Journal of Applied Mathematics, 59(1):85–94, August 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_01/590085.sgm.abs.html.

Ogden:1997:EPS

[600] R. W. Ogden and D. A. Sotiropoulos. The effect of pre-stress on the propagation and reflection of plane waves in incompressible elastic solids. IMA Journal of Applied Mathematics, 59(1):95– 121, August 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_01/590095.sgm.abs.html.

Kriegsmann:1997:HSF

[601] G. A. Kriegsmann. Hot spot formation in microwave heated ceramic fibres. IMA Journal of Applied Mathematics, 59(2):123–148, October 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_02/590123.sgm.abs.html.

Siew:1997:DBM

[602] P. F. Siew. On the differentiation between the magnetic anomalies of cylinders and disks. IMA Journal of Applied Mathematics, 59(2):149–164, October 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_02/590149.sgm.abs.html.

Demiray:1997:NWM

[603] Hilmi Demiray. Nonlinear wave modulation in a prestressed thin elastic tube filled with an inviscid fluid. IMA Journal of Applied Mathematics, 59(2):165– 181, October 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_02/590165.sgm.abs.html.

Lesnic:1997:SCI

[604] D. Lesnic, L. Elliott, and D. B. Ingham. Spacewise coefficient identification in steady and unsteady one-dimensional diffusion problems. IMA Journal of Applied Mathematics, 59(2):183–192, October 1997. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_02/590183.sgm.abs.html.

Walpole:1997:IOT

[605] L. J. Walpole. An inclusion in one of two joined isotropic elastic half-spaces. IMA Journal of Applied Mathematics, 59(2):193–209, October 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_02/590193.sgm.abs.html.

Vanneste:1997:DFC

[606] J. Vanneste. On the derivation of fluxes for conservation laws in Hamiltonian systems. IMA Journal of Applied Mathematics, 59(2):211–220, October 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_02/590211.sgm.abs.html.

Roxburgh:1997:ESR

[607] D. G. Roxburgh. Electromagnetic scattering from a right-circular cylinder using a surface radiation condition. IMA Journal of Applied Mathematics, 59(2):221–230, October 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_02/590221.sgm.abs.html.

Tait:1997:EDC

[608] R. J. Tait and P. Connor. On the expansion of a deformed cylindrical elastic membrane. IMA Journal of Applied Mathematics, 59(3):231–243, December 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_03/590231.sgm.abs.html.

Dai:1997:UAA

[609] H.-H. Dai. Uniform asymptotic analysis for waves in an incompressible elastic rod. I. Disturbances superimposed on an initially stressfree state. IMA Journal of Applied Mathematics, 59(3):245–260, December 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_03/590245.sgm.abs.html.

Misawa:1997:SMS

[610] Tetsuya Misawa. The similarity method in stochastic dynamical systems. IMA Journal of Applied Mathematics, 59(3):261–272, December 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_59/
Issue_03/590261.sgm.abs.html.

Apabhai:1997:MVB

[611] M. Z. Apabhai, N. I. Georgikopoulos, D. Hasnip, R. K. D. Jamie, M. Kim, and P. Wilmott. A model for the value of a business, some optimization problems in its operating procedures and the valuation of its debt. IMA Journal of Applied Mathematics, 59(3):273–285, December 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_03/590273.sgm.abs.html.

Grundy:1997:GBS

[612] R. E. Grundy and R. McLaughlin. Global blow-up of separable solutions of the vorticity equation. IMA Journal of Applied Mathematics, 59(3):287–307, December 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_03/590287.sgm.abs.html.

Huntley:1997:SSS

[613] H. E. Huntley, A. S. Wineman, and K. R. Rajagopal. Stress softening, strain localization and permanent set in the circumferential shear of an incompressible elastomeric cylinder. IMA Journal of Applied Mathematics, 59(3):309–338, December 1997. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_59/
                                        Issue_03/590309.sgm.abs.html.

Merkin:1998:EEF

[614] J. H. Merkin, H. Ševčíková, D. Šnita, and M. Marek. The effects of an electric field on an autocatalytic ionic reaction in a system with high ionic strength. IMA Journal of Applied Mathematics, 60(1):1–31, February 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_60/
                                        Issue_01/600001.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        60/Issue_01/pdf/600001.pdf.

Jones:1998:ICM

[615] D. S. Jones. Irradiation of a cone by a magnetic dipole. IMA Journal of Applied Mathematics, 60(1):33–53, February 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_01/600033.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_01/pdf/600033.pdf.

Nordenholz:1998:SMI

[616] T. R. Nordenholz and O. M. O’Reilly. On steady motions of isotropic, elastic Cosserat points. IMA Journal of Applied Mathematics, 60(1):55–72, February 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_01/600055.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_01/pdf/600055.pdf.

Gilchrist:1998:SMH

[617] John Gilchrist, Gregory A. Kriegsmann, and Demetrios T. Papageorgiou. Stability of a microwave heated fluid Layer. IMA Journal of Applied Mathematics, 60(1):73–89, February 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_01/600073.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_01/pdf/600073.pdf.

Kelley:1998:SLP

[618] Walter Kelley. Spike layer phenomena for pairs of second-order equations. IMA Journal of Applied Mathematics, 60(1):91–108, February 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_01/600091.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_01/pdf/600091.pdf.

Hoang:1998:PSS

[619] Ha V. Hoang, James M. Hill, and J. N. Dewynne. Pseudo-steady-state solutions for solidification in a wedge. IMA Journal of Applied Mathematics, 60(2):109–121, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_02/600109.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_02/pdf/600109.pdf.

DeGodefroy:1998:BSG

[620] Akmel Dé Godefroy. Blow up of solutions of a generalized Boussinesq equation. IMA Journal of Applied Mathematics, 60(2):123–138, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_02/600123.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_02/pdf/600123.pdf.

Gao:1998:MAE

[621] Xin-Lin Gao. A mathematical analysis of the elasto-plastic plane stress problem of a power-law material. IMA Journal of Applied Mathematics, 60(2):139– 149, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_02/600139.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_02/pdf/600139.pdf.

Bujurke:1998:CES

[622] N. M. Bujurke, N. P. Pai, and G. Jayaraman. Computer extended series solution for unsteady flow in a contracting or expanding pipe. IMA Journal of Applied Mathematics, 60(2):151– 165, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_02/600151.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_02/pdf/600151.pdf.

Frosali:1998:AAP

[623] Giovanni Frosali. Asymptotic analysis for a particle transport problem in a moving medium. IMA Journal of Applied Mathematics, 60(2):167– 185, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_02/600167.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_02/pdf/600167.pdf.

Adam:1998:NLS

[624] A. A. Adam and F. M. Mahomed. Non-local symmetries of first-order equations. IMA Journal of Applied Mathematics, 60(2):187–198, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_02/600187.sgm.abs.html; http:

                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        60/Issue_02/pdf/600187.pdf.

Ammari:1998:SWT

[625] H. Ammari. Scattering of waves by thin periodic layers at high frequencies using the on-surface radiation condition method. IMA Journal of Applied Mathematics, 60(2):199– 214, April 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_60/
                                        Issue_02/600199.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        60/Issue_02/pdf/600199.pdf.

Dattoli:1998:APV

[626] G. Dattoli, A. Torre, and G. Mazzacurati. An alternative point of view to the theory of fractional Fourier transform. IMA Journal of Applied Mathematics, 60(3):215– 224, June 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_60/
                                        Issue_03/600215.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        60/Issue_03/pdf/600215.pdf.

Bonnet-BenDhia:1998:GMI

[627] A. S. Bonnet-BenDhia, G. Caloz, and F. Mahé. Guided modes of integrated optical guides. A mathematical study. IMA Journal of Applied Mathematics, 60(3):225–261, June 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_60/
                                        Issue_03/600225.sgm.abs.html; http:

//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_03/pdf/600225.pdf.

MunozRivera:1998:DRS

[628] Jaime E. Muñoz Rivera and Rioco Kamei Barreto. Decay rates of solutions to thermoviscoelastic plates with memory. IMA Journal of Applied Mathematics, 60(3):263– 283, June 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_03/600263.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_03/pdf/600263.pdf.

Krutitskii:1998:ECE

[629] P. A. Krutitskii. On the electric current from electrodes in a magnetized semiconductor film. IMA Journal of Applied Mathematics, 60(3):285– 297, June 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_03/600285.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_03/pdf/600285.pdf.

Bellamy-Knights:1998:ISS

[630] P. G. Bellamy-Knights. An image system and surface singularity solutions for potential flow past an elliptical cylinder. IMA Journal of Applied Mathematics, 60(3):299– 310, June 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_60/
Issue_03/600299.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
60/Issue_03/pdf/600299.pdf.

Peck:1998:KFS

[631] Bill Peck and Lorenz Sigurdson. On the kinematics at a free surface. IMA Journal of Applied Mathematics, 61(1):1–13, September 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_01/610001.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_01/pdf/610001.pdf.

Ruan:1998:TIT

[632] Shigui Ruan. Turing instability and travelling waves in diffusive plankton models with delayed nutrient recycling. IMA Journal of Applied Mathematics, 61(1):15–32, September 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_01/610015.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_01/pdf/610015.pdf.

Liu:1998:UMP

[633] F. Liu, D. L. S. McElwain, and E. Donskoi. The use of a modified Petrov–Galerkin method for gassolid reaction modelling. IMA Journal of Applied Mathematics, 61(1):33–46, September 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_01/610033.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_01/pdf/610033.pdf.

Stewart:1998:PAS

[634] I. W. Stewart. Painlevé analysis for a semi-linear parabolic equation arising in smectic liquid crystals. IMA Journal of Applied Mathematics, 61(1):47–60, September 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_61/
Issue_01/610047.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
61/Issue_01/pdf/610047.pdf.

Ogden:1998:RPW

[635] R. W. Ogden and D. A. Sotiropoulos. Reflection of plane waves from the boundary of a pre-stressed compressible elastic half-space. IMA Journal of Applied Mathematics, 61(1):61–90, September 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_61/
Issue_01/610061.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
61/Issue_01/pdf/610061.pdf.

Mityushev:1998:SHC

[636] V. Mityushev. Steady heat conduction of a material with an array of cylindrical holes in the nonlinear case. IMA Journal of Applied Mathematics, 61(1):91–102, September 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_61/
Issue_01/610091.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
61/Issue_01/pdf/610091.pdf.

Adler:1998:NPS

[637] J. Adler and D. M. Herbert. Nonplanar smoulder propagation governed by diffusion. IMA Journal of Applied Mathematics, 61(2):105–117, October 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_02/610105.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_02/pdf/610105.pdf.

Potthast:1998:PSM

[638] Roland Potthast. A point source method for inverse acoustic and electromagnetic obstacle scattering problems. IMA Journal of Applied Mathematics, 61(2):119–140, October 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_02/610119.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_02/pdf/610119.pdf.

Chudinovich:1998:VTE

[639] Igor Chudinovich and Christian Constanda. Variational treatment of exterior boundary-value problems for thin elastic plates. IMA Journal of Applied Mathematics, 61(2):141–153, October 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_02/610141.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_02/pdf/610141.pdf.

Williams:1998:SSD

[640] D. P. Williams and R. V. Craster. Scattering by small defects in the neighbourhood of a fluidsolid interface. IMA Journal of Applied Mathematics, 61(2):155–177, October 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_61/
Issue_02/610155.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
61/Issue_02/pdf/610155.pdf.

Rees:1998:FCB

[641] D. A. S. Rees and I. Pop. Free convection boundary-layer flow of a micropolar fluid from a vertical flat plate. IMA Journal of Applied Mathematics, 61(2):179–197, October 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_61/
Issue_02/610179.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
61/Issue_02/pdf/610179.pdf.

Gao:1998:DTC

[642] David Yang Gao. Duality, triality and complementary extremum principles in non-convex parametric variational problems with applications. IMA Journal of Applied Mathematics, 61(3):199–235, December 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_61/
Issue_03/610199.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
61/Issue_03/pdf/610199.pdf.

Frigaard:1998:UEF

[643] I. A. Frigaard and O. Scherzer. Uniaxial exchange flows of two Bingham fluids in a cylindrical duct. IMA Journal of Applied Mathematics, 61(3):237–266, December 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_03/610237.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_03/pdf/610237.pdf.

Norris:1998:DIS

[644] Andrew N. Norris. A direct inverse scattering method for imaging obstacles with unknown surface conditions. IMA Journal of Applied Mathematics, 61(3):267–290, December 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_03/610267.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_03/pdf/610267.pdf.

Penney:1998:ADG

[645] R. W. Penney. Analytic determination of gradients of intervisibility over continuous randomly-generated topography. IMA Journal of Applied Mathematics, 61(3):291–303, December 1998. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_61/
                                        Issue_03/610291.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        61/Issue_03/pdf/610291.pdf.

Dai:1999:UAA

[646] Hui-Hui Dai and Zongxi Cai. Uniform asymptotic analysis for waves in an incompressible elastic rod. II. Disturbances superimposed on a pre-stressed state. IMA Journal of Applied Mathematics, 62(1):1–29, February 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_01/620001.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_01/pdf/620001.pdf.

Brock:1999:RCG

[647] L. M. Brock. Rapid crack growth in a thermoelastic solid under mixed-mode thermomechanical loading. IMA Journal of Applied Mathematics, 62(1):31– 44, February 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_01/620031.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_01/pdf/620031.pdf.

Skinner:1999:PTR

[648] L. A. Skinner. Passages through resonance in weakly nonlinear systems. IMA Journal of Applied Mathematics, 62(1):45–59, February 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_01/620045.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_01/pdf/620045.pdf.

Dassios:1999:HSS

[649] George Dassios and Ralph Kleinman (deceased). Half space scattering problems at low frequencies. IMA Journal of Applied Mathematics, 62(1):61– 79, February 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_01/620061.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_01/pdf/620045.pdf.

Kuznetsov:1999:RST

[650] N. Kuznetsov and O. Motygin. On the resistanceless statement of the two-dimensional Neumann-Kelvin problem for a surface-piercing tandem. IMA Journal of Applied Mathematics, 62(1):81–99, February 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_01/620081.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_01/pdf/620081.pdf.

Lesnic:1999:ACI

[651] D. Lesnic, L. Elliott, and D. B. Ingham. Analysis of coefficient identification problems associated to the inverse Euler–Bernoulli beam theory. IMA Journal of Applied Mathematics, 62(2):101–116, April 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_02/620101.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_02/pdf/620101.pdf.

Charalambopoulos:1999:SSW

[652] A. Charalambopoulos and G. Dassios. Scattering of a spherical wave by a small ellipsoid. IMA Journal of Applied Mathematics, 62(2):117– 136, April 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_02/620117.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_02/pdf/620117.pdf.

Liu:1999:MHT

[653] B. Liu and T. R. Marchant. The microwave heating of two-dimensional slabs with small Arrhenius absorptivity. IMA Journal of Applied Mathematics, 62(2):137–166, April 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_02/620137.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_02/pdf/620137.pdf.

Sellier:1999:ASE

[654] A. Sellier. Asymptotic solution for the electrostatic field around a slender conducting body. IMA Journal of Applied Mathematics, 62(2):167– 193, April 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_02/620167.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_02/pdf/620167.pdf.

Watt:1999:WFA

[655] S. D. Watt, R. O. Weber, H. S. Sidhu, and G. N. Mercer. A weightfunction approach for determining watershed initial conditions for combustion waves. IMA Journal of Applied Mathematics, 62(2):195–206, April 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_02/620195.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_02/pdf/620195.pdf.

Choi:1999:EII

[656] Y. S. Choi, Diana Resasco, Jim Schaff, and Boris Slepchenko. Electrodiffusion of ions inside living cells. IMA Journal of Applied Mathematics, 62(3):207– 226, June 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_03/620207.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_03/pdf/620207.pdf.

Cho:1999:PEN

[657] Chung-Ki Cho and YongHoon Kwon. Parameter estimation in nonlinear age-dependent population dynamics. IMA Journal of Applied Mathematics, 62(3):227–244, June 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_03/620227.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_03/pdf/620227.pdf.

ONeill:1999:ATC

[658] F. G. O’Neill. Axisymmetric trawl cod-ends made from netting of a generalized mesh shape. IMA Journal of Applied Mathematics, 62(3):245– 262, June 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_03/620245.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_03/pdf/620245.pdf.

Kim:1999:RPR

[659] Jeong-Hoon Kim. Reflected pulses from a refractive random medium at grazing incidence. IMA Journal of Applied Mathematics, 62(3):263– 282, June 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_03/620263.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_03/pdf/620263.pdf.

Qu:1999:ESN

[660] Changzheng Qu. Exact solutions to nonlinear diffusion equations obtained by a generalized conditional symmetry method. IMA Journal of Applied Mathematics, 62(3):283– 302, June 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_62/
Issue_03/620283.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
62/Issue_03/pdf/620283.pdf.

Smith:1999:SMR

[661] W. R. Smith, J. R. King, B. Tuck, and J. W. Orton. The single-mode rate equations for semiconductor lasers with thermal effects. IMA Journal of Applied Mathematics, 63(1):1–36, August 1999. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_01/630001.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_01/pdf/630001.pdf.

Humphreys:1999:MPS

[662] L. D. Humphreys and P. J. McKenna. Multiple periodic solutions for a nonlinear suspension bridge equation. IMA Journal of Applied Mathematics, 63(1):37–49, August 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL http://www3.oup.co.
                                        uk/imamat/hdb/Volume_63/Issue_01/
                                        630037.sgm.abs.html; http://www3.
                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_01/pdf/630037.pdf.

Djellouli:1999:CFD

[663] Rabia Djellouli, Charbel Farhat, Jan Mandel, and Petr Vaněk. Continuous Fréchet differentiability with respect to a Lipschitz domain and a stability estimate for direct acoustic scattering problems. IMA Journal of Applied Mathematics, 63(1):51– 69, August 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_01/630051.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_01/pdf/630051.pdf.

Esparza:1999:ZWF

[664] Javier Esparza, José L. López, and Javier Sesma. Zeros of the Whittaker function associated to Coulomb waves. IMA Journal of Applied Mathematics, 63(1):71–87, August 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Edwards:1999:ERC

[665] David A. Edwards. Estimating rate constants in a convectiondiffusion system with a boundary reaction. IMA Journal of Applied Mathematics, 63(1):89–112, August 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_63/
Issue_01/630089.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
63/Issue_01/pdf/630089.pdf.

Drabla:1999:ASP

[666] Salah Drabla and Mircea Sofonea. Analysis of a Signorini problem with friction. IMA Journal of Applied Mathematics, 63(2):113–130, October 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_63/
Issue_02/630113.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
63/Issue_02/pdf/630113.pdf.

Saccomandi:1999:IDF

[667] Giuseppe Saccomandi. On inhomogeneous deformations in finite thermoelasticity. IMA Journal of Applied Mathematics, 63(2):131–148, October 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_63/
Issue_02/630131.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
63/Issue_02/pdf/630131.pdf.

Bukkapatnam:1999:LEB

[668] Satish T. S. Bukkapatnam, Soundar R. T. Kumara, and Akhlesh Lakhtakia. Local eigenfunctions based suboptimal wavelet packet representation of contaminated chaotic signals. IMA Journal of Applied Mathematics, 63(2):149– 162, October 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_02/630149.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_02/pdf/630149.pdf.

Strunin:1999:AMS

[669] D. V. Strunin. Autosoliton model of the spinning fronts of reaction. IMA Journal of Applied Mathematics, 63(2):163– 177, October 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_02/630163.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_02/pdf/630163.pdf.

vanVuuren:1999:RRD

[670] Jan H. van Vuuren. Resilience in reaction-diffusion systems. IMA Journal of Applied Mathematics, 63(2):179– 197, October 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_02/630179.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_02/pdf/630179.pdf.

Kay:1999:PSO

[671] Alison L. Kay and Jonathan A. Sherratt. On the persistence of spatiotemporal oscillations generated by invasion. IMA Journal of Applied Mathematics, 63(2):199–216, October 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_63/
Issue_02/630199.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
63/Issue_02/pdf/630199.pdf.

Chapman:1999:VMS

[672] S. J. Chapman, A. D. Fitt, and G. C. Pulos. Vacuum moulding of a superplastic in two dimensions. IMA Journal of Applied Mathematics, 63(3):217–246, December 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_63/
Issue_03/630217.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
63/Issue_03/pdf/630217.pdf.

Alvarez-Vazquez:1999:MCN

[673] L. J. Alvarez-Vázquez and A. Martínez. Modelling and control of natural convection in canned foods. IMA Journal of Applied Mathematics, 63(3):247–265, December 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL http://www3.oup.co.
uk/imamat/hdb/Volume_63/Issue_03/
630247.sgm.abs.html; http://www3.
oup.co.uk/imamat/hdb/Volume_63/
Issue_03/pdf/630247.pdf.

Komarova:1999:CBN

[674] Natalia L. Komarova and Alan C. Newell. The competition between nonlinearity, dispersion and randomness in signal propagation. IMA Journal of Applied Mathematics, 63(3):267–286, December 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_03/630267.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_03/pdf/630267.pdf.

Grundy:1999:TDB

[675] R. E. Grundy and R. McLaughlin. Three-dimensional blow-up solutions of the Navier–Stokes equations. IMA Journal of Applied Mathematics, 63(3):287–306, December 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_03/630287.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_03/pdf/630287.pdf.

Clifford:1999:SMT

[676] Michael J. Clifford and Stephen M. Cox. A simple model for a twostage chemical reaction with diffusion. IMA Journal of Applied Mathematics, 63(3):307–318, December 1999. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_63/
                                        Issue_03/630307.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        63/Issue_03/pdf/630307.pdf.

Takase:2000:GTC

[677] H. Takase and B. D. Sleeman. Generalized travelling coordinate for systems of reaction-diffusion equations. IMA Journal of Applied Mathematics, 64(1):1–22, February 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_01/640001.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_01/pdf/640001.pdf.

Fowler:2000:SEA

[678] A. C. Fowler, G. Kember, and S. G. B. O’Brien. Small exponent asymptotics. IMA Journal of Applied Mathematics, 64(1):23–38, February 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_01/640023.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_01/pdf/640023.pdf.

Pelesko:2000:MHC

[679] John A. Pelesko and Gregory A. Kriegsmann. Microwave heating of ceramic composites. IMA Journal of Applied Mathematics, 64(1):39–50, February 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_01/640039.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_01/pdf/640039.pdf.

Leppington:2000:SSI

[680] F. G. Leppington and W. M. Pang. Scattering of sound by an infinite membrane fixed on two circular regions. IMA Journal of Applied Mathematics, 64(1):51–72, February 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_01/640051.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_01/pdf/640051.pdf.

Borrelli:2000:SEE

[681] A. Borrelli and M. C. Patria. Spatial energy estimates in dynamical problems for a semi-infinite piezoelectric beam. IMA Journal of Applied Mathematics, 64(1):73–93, February 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_01/640073.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_01/pdf/640073.pdf.

Zayed:2000:HSB

[682] E. M. E. Zayed. On hearing the shape of a bounded domain with Robin boundary conditions. IMA Journal of Applied Mathematics, 64(1):95–108, February 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_01/640095.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_01/pdf/640095.pdf.

Zhang:2000:AAO

[683] Weiyi Zhang and Koncay Huseyin. An algebraic approach for obtaining nilpotent normal forms in dimension 4. IMA Journal of Applied Mathematics, 64(2):109–123, April 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/

Issue_02/640109.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_02/pdf/640109.pdf.

Levine:2000:GSU

[684] Harold Levine. A general solution for an unsteady linear diffusion problem. IMA Journal of Applied Mathematics, 64(2):125–137, April 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_02/640125.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_02/pdf/640125.pdf.

King:2000:ATS

[685] I. D. King and T. E. Hodgetts. An approximate treatment of scattering based on the delta boundary operator technique. IMA Journal of Applied Mathematics, 64(2):139– 155, April 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_02/640139.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_02/pdf/640139.pdf.

Merkin:2000:EEF

[686] J. H. Merkin, H. Sevciková, and D. Snita. The effect of an electric field on the local stoichiometry of front waves in an ionic chemical system. IMA Journal of Applied Mathematics, 64(2):157–188, April 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_02/640157.sgm.abs.html; http:

                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_02/pdf/640157.pdf.

Walker:2000:DEW

[687] E. J. Walker, G. R. Wickham, and P. A. Lewis. The diffraction of elastic waves by an imperfect fusion bond. IMA Journal of Applied Mathematics, 64(2):189– 211, April 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_02/640189.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_02/pdf/640189.pdf.

Wang:2000:VIM

[688] C. Y. Wang. Vibration of an infinite membrane supported by an array of posts. IMA Journal of Applied Mathematics, 64(3):213–221, June 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_03/640213.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_03/pdf/640213.pdf.

Hood:2000:DIR

[689] Simon Hood. On direct, implicit reductions of a nonlinear diffusion equation with an arbitrary function — generalizations of Clarkson’s and Kruskal’s method. IMA Journal of Applied Mathematics, 64(3):223– 244, June 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_03/640223.sgm.abs.html; http:

//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_03/pdf/640223.pdf.

Athanasiadis:2000:ESH

[690] Christodoulos Athanasiadis, George Costakis, and Ioannis G. Stratis. Electromagnetic scattering by a homogeneous chiral obstacle in a chiral environment. IMA Journal of Applied Mathematics, 64(3):245–258, June 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_03/640245.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_03/pdf/640245.pdf.

Krutitskii:2000:NAR

[691] P. A. Krutitskii. A new approach to reduction of the Neumann problem in acoustic scattering to a non-hypersingular integral equation. IMA Journal of Applied Mathematics, 64(3):259–269, June 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_03/640259.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
64/Issue_03/pdf/640259.pdf.

Amendola:2000:ABC

[692] Giovambattista Amendola. Asymptotic behaviour in a conducting electromagnetic medium. IMA Journal of Applied Mathematics, 64(3):271– 282, June 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_64/
Issue_03/640271.sgm.abs.html; http:

                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_03/pdf/640271.pdf.

Stein:2000:ECS

[693] Carl Fredrik Stein. On the existence of conically self-similar freevortex solutions to the Navier–Stokes equations. IMA Journal of Applied Mathematics, 64(3):283–300, June 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_64/
                                        Issue_03/640283.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        64/Issue_03/pdf/640283.pdf.

Motygin:2000:JKH

[694] O. V. Motygin and S. A. Nazarov. Justification of the Kirchhoff hypotheses and error estimation for twodimensional models of anisotropic and inhomogeneous plates, including laminated plates. IMA Journal of Applied Mathematics, 65(1):1–28, August 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/
                                        Issue_01/650001.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_01/pdf/650001.pdf.

El-Gebeily:2000:CSE

[695] M. A. El-Gebeily and K. M. Furati. On the completeness of the set of eigenvectors of a certain class of self-adjoint finite difference operators. IMA Journal of Applied Mathematics, 65(1):29– 43, August 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/

Issue_01/650029.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
65/Issue_01/pdf/650029.pdf.

Lazzari:2000:VPE

[696] Barbara Lazzari and Roberta Nibbi. Variational principles in electromagnetism. IMA Journal of Applied Mathematics, 65(1):45–95, August 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_65/
Issue_01/650045.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
65/Issue_01/pdf/650045.pdf.

Khater:2000:BTE

[697] A. H. Khater, D. K. Callebaut, and O. H. El-Kalaawy. Bäcklund transformations and exact solutions for a nonlinear elliptic equation modelling isothermal magnetostatic atmosphere. IMA Journal of Applied Mathematics, 65(1):97–108, August 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_65/
Issue_01/650097.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
65/Issue_01/pdf/650097.pdf.

Liu:2000:SSS

[698] Wei-Jiu Liu and Miroslav Krstić. Strong stabilization of the system of linear elasticity by a Dirichlet boundary feedback. IMA Journal of Applied Mathematics, 65(2):109–121, October 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_65/

                                        Issue_02/650109.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_02/pdf/650109.pdf.

Gielen:2000:GPC

[699] J. L. W. Gielen. The general packed column: an analytical solution. IMA Journal of Applied Mathematics, 65(2):123–146, October 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/
                                        Issue_02/650123.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_02/pdf/650123.pdf.

Fitt:2000:UMT

[700] A. D. Fitt and T. R. B. Lattimer. On the unsteady motion of twodimensional sails. IMA Journal of Applied Mathematics, 65(2):147–171, October 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/
                                        Issue_02/650147.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_02/pdf/650147.pdf.

Agarwal:2000:SIB

[701] Ravi P. Agarwal and Donal O’Regan. Singular initial and boundary value problems with sign changing nonlinearities. IMA Journal of Applied Mathematics, 65(2):173–198, October 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/
                                        Issue_02/650173.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_02/pdf/650173.pdf.

Hao:2000:CPL

[702] Dinh Nho Hào and D. Lesnic. The Cauchy problem for Laplace’s equation via the conjugate gradient method. IMA Journal of Applied Mathematics, 65(2):199–217, October 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_65/
Issue_02/650199.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
65/Issue_02/pdf/650199.pdf.

Abell:2000:MSE

[703] K. A. Abell, A. R. Humphries, and E. S. Van Vleck. Mosaic solutions and entropy for spatially discrete Cahn– Hilliard equations. IMA Journal of Applied Mathematics, 65(3):219–255, December 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_65/
Issue_03/650219.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
65/Issue_03/pdf/650219.pdf.

Abrahams:2000:APA

[704] I. David Abrahams. The application of Padé approximants to Wiener– Hopf factorization. IMA Journal of Applied Mathematics, 65(3):257–281, December 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_65/
Issue_03/650257.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
65/Issue_03/pdf/650257.pdf.

Billingham:2000:SSS

[705] John Billingham. Steady-state solutions for strongly exothermic ignition in symmetric geometries. IMA Journal of Applied Mathematics, 65(3):283–313, December 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/
                                        Issue_03/650283abs.pdf;   http:/
                                        /www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_03/pdf/650283.pdf.

Howe:2000:PTG

[706] M. S. Howe. Pressure transients generated when high-speed trains pass in a tunnel. IMA Journal of Applied Mathematics, 65(3):315–334, December 2000. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_65/
                                        Issue_03/650315.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        65/Issue_03/pdf/650315.pdf.

Kriegsmann:2001:PFM

[707] G. A. Kriegsmann. Pattern formation in microwave heated ceramics: cylinders and slabs. IMA Journal of Applied Mathematics, 66(1):1–32, February 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_01/660001.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_01/pdf/660001.pdf.

Cade:2001:IEA

[708] R. Cade. An integral equation approach to electromagnetic diffraction by a cylinder. IMA Journal of Applied Mathematics, 66(1):33–54, February 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_01/660033.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_01/pdf/660033.pdf.

Williams:2001:TDN

[709] H. A. R. Williams and O. E. Jensen. Two-dimensional nonlinear advection-diffusion in a model of surfactant spreading on a thin liquid film. IMA Journal of Applied Mathematics, 66(1):55–82, February 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_01/660055.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_01/pdf/660055.pdf.

Antoine:2001:FAC

[710] Xavier Antoine. Fast approximate computation of a time-harmonic scattered field using the on-surface radiation condition method. IMA Journal of Applied Mathematics, 66(1):83–110, February 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_01/660083.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_01/pdf/660083.pdf.

Muriel:2001:NMR

[711] C. Muriel and J. L. Romero. New methods of reduction for ordinary differential equations. IMA Journal of Applied Mathematics, 66(2):111–125, April 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        66/Issue_02/660111.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_66/Issue_02/
                                        pdf/660111.pdf.

Gerritsma:2001:UCV

[712] M. I. Gerritsma and T. N. Phillips. On the use of characteristic variables in viscoelastic flow problems. IMA Journal of Applied Mathematics, 66(2):127– 147, April 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_02/660127.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_02/pdf/660127.pdf.

Sellier:2001:SDB

[713] A. Sellier. A slender dielectric body embedded in an arbitrary external potential. IMA Journal of Applied Mathematics, 66(2):149–173, April 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_02/660149abs.pdf;   http:/
                                        /www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_02/pdf/660149.pdf.

Barry:2001:ABT

[714] S. I. Barry and M. Holmes. Asymptotic behaviour of thin poroelastic layers. IMA Journal of Applied Mathematics, 66(2):175–194, April 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_02/660175.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_02/pdf/660175.pdf.

Fang:2001:EMN

[715] Weifu Fang and Kazufumi Ito. An electrochemistry model with nonlinear diffusion: steady-state solutions. IMA Journal of Applied Mathematics, 66(2):195–213, April 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_02/660195.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_02/pdf/660195.pdf.

Chudinovich:2001:TPB

[716] Igor Chudinovich and Christian Constanda. The transmission problem in bending of plates with transverse shear deformation. IMA Journal of Applied Mathematics, 66(3):215– 229, June 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_03/660215.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_03/pdf/660215.pdf.

Bonnet-Bendhia:2001:SHF

[717] A. S. Bonnet-Bendhia, G. Caloz, M. Dauge, and F. Mahé. Study at high frequencies of a stratified waveguide. IMA Journal of Applied Mathematics, 66(3):231–257, June 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/

                                        Issue_03/660231.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_03/pdf/660231.pdf.

Wang:2001:FCT

[718] C. Y. Wang. Flow in a circular tube with a centred strip. IMA Journal of Applied Mathematics, 66(3):259– 267, June 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_03/660259.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_03/pdf/660259.pdf.

Rivera:2001:PST

[719] Jaime E. Muñoz Rivera and Reinhard Racke. Polynomial stability in two-dimensional magnetoelasticity. IMA Journal of Applied Mathematics, 66(3):269–283, June 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_03/660269.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_03/pdf/660269.pdf.

Hill:2001:NSP

[720] James M. Hill and Anna M. Milan. On a non-symmetrical plane strain deformation for hyperelastic materials. IMA Journal of Applied Mathematics, 66(3):285–291, June 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_03/660285.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_03/pdf/660285.pdf.

Fernandes:2001:VPF

[721] P. Fernandes and I. Perugia. Vector potential formulation for magnetostatics and modelling of permanent magnets. IMA Journal of Applied Mathematics, 66(3):293–318, June 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_03/660293.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_03/pdf/660293.pdf.

Shubov:2001:AAA

[722] Marianna A. Shubov. Asymptotic analysis of aircraft wing model in subsonic air flow. IMA Journal of Applied Mathematics, 66(4):319–356, August 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_04/660319.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_04/pdf/660319.pdf.

Pichugin:2001:TDM

[723] Aleksey V. Pichugin and Graham A. Rogerson. A two-dimensional model for extensional motion of a pre-stressed incompressible elastic layer near cutoff frequencies. IMA Journal of Applied Mathematics, 66(4):357–385, August 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_04/660357.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_04/pdf/660357.pdf.

Kidd:2001:FTC

[724] J. E. Kidd, I. W. Stewart, and C. Constanda. Freedericksz transitions in circular toroidal layers of smectic C liquid crystals. IMA Journal of Applied Mathematics, 66(4):387– 409, August 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_04/660387.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_04/pdf/660387.pdf.

Jones:2001:ILW

[725] D. S. Jones. Impedance of a lossy wedge. IMA Journal of Applied Mathematics, 66(4):411–422, August 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_04/660411.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_04/pdf/660411.pdf.

Dunwoody:2001:HWH

[726] J. Dunwoody. Harmonic waves in heat conducting bars: a comparison of constrained and unconstrained thermoelasticity. IMA Journal of Applied Mathematics, 66(5):423–447, October 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_05/660423.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_05/pdf/660423.pdf.

Lacey:2001:MPM

[727] A. A. Lacey and C. Nikolopoulos. A model for polymer melting during modulated-temperature differential scanning calorimetry. IMA Journal of Applied Mathematics, 66(5):449– 476, October 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_05/660449.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_05/pdf/660449.pdf.

Muriel:2001:CSN

[728] C. Muriel and J. L. Romero. Csymmetries and non-solvable symmetry algebras. IMA Journal of Applied Mathematics, 66(5):477–498, October 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_05/660477.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_05/pdf/660477.pdf.

Baesu:2001:AFP

[729] Eveline Baesu and Eugen Soós. Antiplane fracture in a prestressed and prepolarized piezoelectric crystal. IMA Journal of Applied Mathematics, 66(5):499–508, October 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_05/660499.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_05/pdf/660499.pdf.

Wilson:2001:AGC

[730] R. Eddie Wilson. An analysis of Gipps’s car-following model of highway traffic. IMA Journal of Applied Mathematics, 66(5):509–537, October 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_05/660509.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_05/pdf/660509.pdf.

Athanasiadis:2001:SWS

[731] C. Athanasiadis, P. A. Martin, and I. G. Stratis. On sphericalwave scattering by a spherical scatterer and related near-field inverse problems. IMA Journal of Applied Mathematics, 66(6):539–549, December 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_06/660539.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_06/pdf/660539.pdf.

Brock:2001:ISC

[732] L. M. Brock and H. G. Georgiadis. An illustration of sliding contact at any constant speed on highly elastic half-spaces. IMA Journal of Applied Mathematics, 66(6):551–566, December 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_66/
                                        Issue_06/660551.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        66/Issue_06/pdf/660551.pdf.

ElJarroudi:2001:ABC

[733] Mustapha El Jarroudi and Alain Brillard. Asymptotic behaviour of a cylindrical elastic structure periodically reinforced along identical fibres. IMA Journal of Applied Mathematics, 66(6):567–590, December 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_06/660567.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_06/pdf/660567.pdf.

Antipov:2001:SQP

[734] Y. A. Antipov. Solution by quadratures of the problem of a cylindrical crack by the method of matrix factorization. IMA Journal of Applied Mathematics, 66(6):591–619, December 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_06/660591.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_06/pdf/660591.pdf.

Agarwal:2001:SPI

[735] Ravi P. Agarwal and Donal O’Regan. Singular problems on the infinite interval modelling phenomena in draining flows. IMA Journal of Applied Mathematics, 66(6):621–635, December 2001. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_66/
Issue_06/660621.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
66/Issue_06/pdf/660621.pdf.

Alves:2002:FFO

[736] Carlos J. S. Alves and Rainer Kress. On the far-field operator in elastic obstacle scattering. IMA Journal of Applied Mathematics, 67(1):1–21, February 2002. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_01/670001.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_01/pdf/670001.pdf.

Alvarez-Vazquez:2002:MAO

[737] L. J. Alvarez-Vázquez, A. Martínez, C. Rodríguez, and M. E. Vázquez-Méndez. Mathematical analysis of the optimal location of wastewater outfalls. IMA Journal of Applied Mathematics, 67(1):23–39, February 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_01/670023.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_01/pdf/670023.pdf.

Davies:2002:MMK

[738] Penny J. Davies, Fiona J. Carter, and Alfred Cuschieri. Mathematical modelling for keyhole surgery simulations: a biomechanical model for spleen tissue. IMA Journal of Applied Mathematics, 67(1):41–67, February 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_01/670041.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_01/pdf/670041.pdf.

DePablo:2002:ASR

[739] Arturo De Pablo, Fernando Quirós, and Julio D. Rossi. Asymptotic simplification for a reaction-diffusion problem with a nonlinear boundary condition. IMA Journal of Applied Mathematics, 67(1):69–98, February 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_01/670069.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_01/pdf/670069.pdf.

Ferguson:2002:CPN

[740] V. R. Ferguson and P. G. Daniels. Convection patterns near a lateral boundary. IMA Journal of Applied Mathematics, 67(2):99–125, April 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_02/670099.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_02/pdf/670099.pdf.

Hahner:2002:UIP

[741] Peter Hähner. On uniqueness for an inverse problem in inhomogeneous elasticity. IMA Journal of Applied Mathematics, 67(2):127–143, April 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_02/670127.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_02/pdf/670127.pdf.

Liu:2002:MHT

[742] B. Liu and T. R. Marchant. The microwave heating of three-dimensional blocks: semi-analytical solutions. IMA Journal of Applied Mathematics, 67 (2):145–175, April 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_02/670145.sgm.abs.html;
                                        http://www3.oup.co.uk/imamat/hdb/
                                        Volume_67/Issue_02/pdf/670145.pdf.

Ward:2002:TRM

[743] Michael J. Ward. Thermal runaway and microwave heating in thin cylindrical domains. IMA Journal of Applied Mathematics, 67(2):177– 200, April 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_02/670177.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_02/pdf/670177.pdf.

Marchant:2002:DTW

[744] B. P. Marchant and John Norbury. Discontinuous travelling wave solutions for certain hyperbolic systems. IMA Journal of Applied Mathematics, 67(2):201– 224, April 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_02/670201.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_02/pdf/670201.pdf.

Ling:2002:ASO

[745] Qi-Cong Ling, Cong-Qiang Liu, Zheng-Yu Bao, and Fanglin Li. Analytical solutions for ore fluid transport in fractured rocks: a case study on ore-forming processes of the stratabound skarn copper deposits in China. IMA Journal of Applied Mathematics, 67(3):225–247, June 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_03/670225.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_03/pdf/670225.pdf.

Briozzo:2002:ESI

[746] Adriana C. Briozzo and Domingo A. Tarzia. An explicit solution for an instantaneous two-phase Stefan problem with nonlinear thermal coefficients. IMA Journal of Applied Mathematics, 67(3):249–261, June 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_03/670249.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_03/pdf/670249.pdf.

Goldfarb:2002:DTE

[747] Igor Goldfarb, Vladimir Gol’dshtein, and Ann Zinoviev. Delayed thermal explosion in porous media: method of invariant manifolds. IMA Journal of Applied Mathematics, 67(3):263– 280, June 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_03/670263.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_03/pdf/670263.pdf.

Palamides:2002:BVP

[748] P. K. Palamides. Boundary-value problems for shallow elastic membrane caps. IMA Journal of Applied Mathematics, 67(3):281–299, June 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/

                                        Issue_03/670281.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_03/pdf/670281.pdf.

Lopez:2002:AEG

[749] José L. López and Chelo Ferreira. Asymptotic expansions of the generalized Epstein–Hubbell integral. IMA Journal of Applied Mathematics, 67(3):301–319, June 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_03/670301.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_03/pdf/670301.pdf.

Mabrouk:2002:APA

[750] Mongi Mabrouk and Zouheir Helali. Asymptotic profiles and asymptotic distributions of energy in elastic waves scattering. IMA Journal of Applied Mathematics, 67(4):321–356, August 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_04/670321.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_04/pdf/670321.pdf.

Xu:2002:RBP

[751] Gen-Qi Xu and De-Xing Feng. The Riesz basis property of a Timoshenko beam with boundary feedback and application. IMA Journal of Applied Mathematics, 67(4):357–370, August 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_04/670357.sgm.abs.html; http:

//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_04/pdf/670357.pdf.

Villacampa:2002:PES

[752] Y. Villacampa and A. Balaguer. A priori estimates of the solution for the Dirichlet problem. IMA Journal of Applied Mathematics, 67(4):371– 382, August 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_04/670371.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_04/pdf/670371.pdf.

Kaplunov:2002:SWM

[753] J. D. Kaplunov, E. V. Nolde, and G. A. Rogerson. Short wave motion in a pre-stressed incompressible elastic plate. IMA Journal of Applied Mathematics, 67(4):383–399, August 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_04/670383.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_04/pdf/670383.pdf.

Kostolansky:2002:DTE

[754] Eduard Kostolanský. On the duration of the total eclipse of a satellite of a body. IMA Journal of Applied Mathematics, 67(4):401–410, August 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_04/670401.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_04/pdf/670401.pdf.

Site:2002:PDE

[755] L. Delle Site. A partial differential equation which describes an interatomic surface. IMA Journal of Applied Mathematics, 67(4):411–417, August 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_04/670411.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_04/pdf/670411.pdf.

Mercado:2002:MTF

[756] Gema A. Mercado, Benjamin P. Luce, and Jack Xin. Modelling thermal front dynamics in microwave heating. IMA Journal of Applied Mathematics, 67(5):419–439, October 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_05/670419.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_05/pdf/670419.pdf.

Shorten:2002:NMS

[757] Robert Shorten and Fiacre Ó Cairbre. A new methodology for the stability analysis of pairwise triangularizable and related switching systems. IMA Journal of Applied Mathematics, 67(5):441–457, October 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_67/
                                        Issue_05/670441.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        67/Issue_05/pdf/670441.pdf.

McKenna:2002:GSP

[758] P. J. McKenna and K. S. Moore. The global structure of periodic solutions to a suspension bridge mechanical model. IMA Journal of Applied Mathematics, 67(5):459–478, October 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_05/670459.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_05/pdf/670459.pdf.

Bartoli:2002:RHO

[759] Nathalie Bartoli and Abderrahmane Bendali. Robust and high-order effective boundary conditions for perfectly conducting scatterers coated by a thin dielectric layer. IMA Journal of Applied Mathematics, 67(5):479– 508, October 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_05/670479.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_05/pdf/670479.pdf.

Anonymous:2002:IV

[760] Anonymous. Index to Volume 67. IMA Journal of Applied Mathematics, 67(6):i–ii, December 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL http://www3.oup.co.
uk/imamat/hdb/Volume_67/Issue_06/
pdf/67000i.pdf.

Hoang:2002:WSD

[761] Viêt Hà Hoàng and R. V. Craster. Wedge solidification with differing types of boundary conditions. IMA Journal of Applied Mathematics, 67(6):509–524, December 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_06/670509.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_06/pdf/670509.pdf.

Barral:2002:ESS

[762] P. Barral and P. Quintela. Existence of a solution for a Signorini contact problem for Maxwell-Norton materials. IMA Journal of Applied Mathematics, 67(6):525–549, December 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_06/670525.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_06/pdf/670525.pdf.

Soliman:2002:BSP

[763] A. A. Soliman. On ∨_{0}-boundedness and ∨_{0}-stability properties with perturbing Liapunov functional. IMA Journal of Applied Mathematics, 67(6):551–557, December 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_06/670551.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_06/pdf/670551.pdf.

Fokas:2002:NTM

[764] A. S. Fokas. A new transform method for evolution partial differential equations. IMA Journal of Applied Mathematics, 67(6):559–590, December 2002. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_67/
Issue_06/670559.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
67/Issue_06/pdf/670559.pdf.

Lohofer:2003:ECS

[765] G. Lohöfer. An electrically conducting sphere in a three-dimensional, alternating magnetic field. IMA Journal of Applied Mathematics, 68(1):1– 21, February 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_01/680001.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_01/pdf/680001.pdf.

Marzocchi:2003:TPT

[766] Alfredo Marzocchi, Jaime E. Muñoz Rivera, and Maria Grazia Naso. Transmission problem in thermoelasticity with symmetry. IMA Journal of Applied Mathematics, 68(1):23–46, February 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_01/680023.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_01/pdf/680023.pdf.

Grundy:2003:ABI

[767] R. E. Grundy and A. Kay. The asymptotics of blow-up in inviscid Boussinesq flow and related systems. IMA Journal of Applied Mathematics, 68(1):47– 81, February 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_01/680047.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_01/pdf/680047.pdf.

Oquendo:2003:VBS

[768] Higidio Portillo Oquendo. Viscoelastic boundary stabilization for a transmission problem in elasticity. IMA Journal of Applied Mathematics, 68(1):83– 97, February 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_01/680083.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_01/pdf/680083.pdf.

Mederer:2003:TSM

[769] Michael Mederer. Transient solutions of Markov processes and generalized continued fractions. IMA Journal of Applied Mathematics, 68(1):99–118, February 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_01/680099.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_01/pdf/680099.pdf.

Kim:2003:BIL

[770] Sun-Chul Kim and Hisashi Okamoto. Bifurcations and inviscid limit of rhombic Navier–Stokes flows in tori. IMA Journal of Applied Mathematics, 68 (2):119–134, April 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_02/hxg007.sgm.abs.html;
                                        http://www3.oup.co.uk/imamat/hdb/
                                        Volume_68/Issue_02/pdf/hxg007.pdf.

Ackleh:2003:SNL

[771] Azmy S. Ackleh, Keng Deng, and Wei Li. Solvability of a non-local hyperbolic model arising from a reliability system. IMA Journal of Applied Mathematics, 68(2):135–148, April 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_02/hxg008.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_02/pdf/hxg008.pdf.

VanNoorden:2003:ECP

[772] T. L. Van Noorden, S. M. Verduyn Lunel, and A. Bliek. The efficient computation of periodic states of cyclically operated chemical processes. IMA Journal of Applied Mathematics, 68(2):149–166, April 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_02/hxg005.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_02/pdf/hxg005.pdf.

Szeftel:2003:ABC

[773] Jérémie Szeftel. Absorbing boundary conditions for reaction-diffusion equations. IMA Journal of Applied Mathematics, 68(2):167–184, April 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_02/hxg009.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_02/pdf/hxg009.pdf.

Hlavacek:2003:BTB

[774] Ivan Hlaváček. Buckling of a Timoshenko beam on elastic foundation with uncertain input data. IMA Journal of Applied Mathematics, 68(2):185– 204, April 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_02/hxg010.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_02/pdf/hxg010.pdf.

Hos:2003:BAS

[775] Csaba Hős, Alan Champneys, and László Kullmann. Bifurcation analysis of surge and rotating stall in the Moore–Greitzer compression system. IMA Journal of Applied Mathematics, 68(2):205–228, April 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_02/hxg012.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_02/pdf/hxg012.pdf.

Motygin:2003:UCL

[776] O. V. Motygin and P. McIver. A uniqueness criterion for linear problems of wave-body interaction. IMA Journal of Applied Mathematics, 68(3):229– 250, June 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_03/hxg015.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_03/pdf/hxg015.pdf.

Allahverdiev:2003:EDF

[777] B. P. Allahverdiev. Extensions, dilations and functional models of discrete Dirac operators in limit pointcircle cases. IMA Journal of Applied Mathematics, 68(3):251–267, June 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_03/hxg016.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_03/pdf/hxg016.pdf.

Gustafsson:2003:NPE

[778] Björn Gustafsson and Jacqueline Mossino. Non-periodic explicit homogenization and reduction of dimension: the linear case. IMA Journal of Applied Mathematics, 68(3):269– 298, June 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_03/hxg011.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_03/pdf/hxg011.pdf.

Antipov:2003:USI

[779] Y. A. Antipov and P. Schiavone. On the uniformity of stresses inside an inhomogeneity of arbitrary shape. IMA Journal of Applied Mathematics, 68(3):299– 311, June 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_03/hxg014.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_03/pdf/hxg014.pdf.

Goard:2003:ICS

[780] Joanna Goard. Iterating the classical symmetries method to solve initial value problems. IMA Journal of Applied Mathematics, 68(3):313– 328, June 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_03/hxg017.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_03/pdf/hxg017.pdf.

Yu:2003:FFD

[781] L. H. Yu. Fundamental frequency of a doubly connected membrane: a modified perturbation method. IMA Journal of Applied Mathematics, 68(4):329– 354, August 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_04/hxg021.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_04/pdf/hxg021.pdf.

Fokas:2003:TML

[782] A. S. Fokas and A. A. Kapaev. On a transform method for the Laplace equation in a polygon. IMA Journal of Applied Mathematics, 68(4):355– 408, August 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_04/hxg019.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_04/pdf/hxg019.pdf.

Weng:2003:ASP

[783] Peixuan Weng, Huaxiong Huang, and Jianhong Wu. Asymptotic speed of propagation of wave fronts in a lattice delay differential equation with global interaction. IMA Journal of Applied Mathematics, 68(4):409–439, August 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_04/hxg020.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_04/pdf/hxg020.pdf.

Sugimoto:2003:TAF

[784] Takeshi Sugimoto. A theoretical analysis of formation flight as a nonlinear self-organizing phenomenon. IMA Journal of Applied Mathematics, 68(5):441–470, October 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_05/680441.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_05/pdf/680441.pdf.

El-Gebeily:2003:RAS

[785] Mohamed A. El-Gebeily. Regular approximation of singular self-adjoint differential operators. IMA Journal of Applied Mathematics, 68(5):471–489, October 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_05/680471.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_05/pdf/680471.pdf.

Quintanilla:2003:EST

[786] R. Quintanilla. On existence and stability in the theory of swelling porous elastic soils. IMA Journal of Applied Mathematics, 68(5):491–506, October 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_05/680491.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_05/pdf/680491.pdf.

Chudinovich:2003:TDB

[787] Igor Chudinovich and Christian Constanda. Time-dependent boundary integral equations for multiply connected plates. IMA Journal of Applied Mathematics, 68(5):507–522, October 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_05/680507.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_05/pdf/680507.pdf.

Ibrahim:2003:TPE

[788] R. S. Ibrahim. Truncated Painlevé; expansion and reduction of sine-Poisson equation to a quadrature in collisionless cold plasma. IMA Journal of Applied Mathematics, 68(5):523–535, October 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_05/680523.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_05/pdf/680523.pdf.

Simon:2003:IFP

[789] Peter L. Simon, Serafim Kalliadasis, John H. Merkin, and Stephen K. Scott. Inhibition of flame propagation by an endothermic reaction. IMA Journal of Applied Mathematics, 68(5):537–562, October 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_05/680537.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_05/pdf/680537.pdf.

Idemen:2003:CFE

[790] Mithat Idemen and Ali Alkumru. On a class of functional equations of the Wiener–Hopf type and their applications in n-part scattering problems. IMA Journal of Applied Mathematics, 68(6):563–586, December 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_06/680563.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_06/pdf/680563.pdf.

Adawi:2003:PTT

[791] A. Adawi and A. Alawneh. A Parseval-type theorem applied to certain integral transforms on generalized functions. IMA Journal of Applied Mathematics, 68(6):587–593, December 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_06/680587.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_06/pdf/680587.pdf.

Louaked:2003:WPI

[792] M. Louaked, L. Hanich, and C. P. Thompson. Well-posedness of incompressible models of two- and threephase flow. IMA Journal of Applied Mathematics, 68(6):595–620, December 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_06/680595.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_06/pdf/680595.pdf.

Potthast:2003:DDS

[793] Roland Potthast and Ioannis G. Stratis. On the domain derivative for scattering by impenetrable obstacles in chiral media. IMA Journal of Applied Mathematics, 68(6):621–635, December 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_06/680621.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_06/pdf/680621.pdf.

Chen:2003:MMO

[794] C. Y. Chen, H. M. Byrne, and J. R. King. Mathematical modelling of the oxidation of nickel titanium alloys. IMA Journal of Applied Mathematics, 68(6):637–664, December 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_68/
                                        Issue_06/680637.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        68/Issue_06/pdf/680637.pdf.

Mackey:2003:DTL

[795] D. Mackey and E. A. Cox. Dynamics of a two-layer fluid sloshing problem. IMA Journal of Applied Mathematics, 68(6):665–686, December 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_68/
Issue_06/680665.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
68/Issue_06/pdf/680665.pdf.

Hintermuller:2004:PDA

[796] M. Hintermüller, V. A. Kovtunenko, and K. Kunisch. The primal-dual active set method for a crack problem with non-penetration. IMA Journal of Applied Mathematics, 69(1):1–26, February 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_69/
Issue_01/690001.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
69/Issue_01/pdf/690001.pdf.

Antipov:2004:VFD

[797] Y. A. Antipov and V. V. Silvestrov. Vector functional-difference equation in electromagnetic scattering. IMA Journal of Applied Mathematics, 69(1):27– 69, February 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_69/
Issue_01/690027.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
69/Issue_01/pdf/690027.pdf.

Carmona:2004:IMP

[798] Victoriano Carmona, Emilio Freire, Enrique Ponce, and Francisco Torres. Invariant manifolds of periodic orbits for piecewise linear threedimensional systems. IMA Journal of Applied Mathematics, 69(1):71–91, February 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_69/
                                        Issue_01/690071.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        69/Issue_01/pdf/690071.pdf.

Aassila:2004:SAB

[799] Mohammed Aassila. Stability and asymptotic behaviour of solutions of the heat equation. IMA Journal of Applied Mathematics, 69(1):93–109, February 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_69/
                                        Issue_01/690093.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        69/Issue_01/pdf/690093.pdf.

Rowell:2004:CPR

[800] Jonathan T. Rowell. Coexistence, permanence and resilience in biological reaction-diffusion systems. IMA Journal of Applied Mathematics, 69(2):111– 129, April 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_69/
                                        Issue_02/690111.sgm.abs.html; http:
                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        69/Issue_02/pdf/690111.pdf.

Dai:2004:BSG

[801] Hui-Hui Dai and De-Xing Kong. Blowup of solutions to generalized Riemann problem for quasilinear hyperbolic systems of conservation laws. IMA Journal of Applied Mathematics, 69(2):131–158, April 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

                                        oup.co.uk/imamat/hdb/Volume_69/
                                        Issue_02/690131.sgm.abs.html; http:

//www3.oup.co.uk/imamat/hdb/Volume_
69/Issue_02/pdf/690131.pdf.

Rokne:2004:RST

[802] J. Rokne, B. M. Singh, R. S. Dhaliwal, and J. Vrbik. The Reissner–Sagoci type problem for a non-homogeneous elastic cylinder embedded in an elastic nonhomogeneous half-space. IMA Journal of Applied Mathematics, 69(2):159– 173, April 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_69/
Issue_02/690159.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
69/Issue_02/pdf/690159.pdf.

Simon:2004:SFE

[803] Peter L. Simon, Serafim Kalliadasis, John H. Merkin, and Stephen K. Scott. Stability of flames in an exothermicendothermic system. IMA Journal of Applied Mathematics, 69(2):175– 203, April 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_69/
Issue_02/690175.sgm.abs.html; http:
//www3.oup.co.uk/imamat/hdb/Volume_
69/Issue_02/pdf/690175.pdf.

Chamberlain:2004:EEW

[804] P. G. Chamberlain. The effect of evanescent wave modes on scattering and near-trapping. IMA Journal of Applied Mathematics, 69(2):205– 218, April 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://www3.

oup.co.uk/imamat/hdb/Volume_69/
Issue_02/690205.sgm.abs.html; http:

                                        //www3.oup.co.uk/imamat/hdb/Volume_
                                        69/Issue_02/pdf/690205.pdf.

Soh:2004:LGI

[805] C. Wafo Soh and Ebrahim Momoniat. A Lie group investigation of exact weakly nonlinear quasi-planar sound waves. IMA Journal of Applied Mathematics, 69(3):219–232, June 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL http://imamat.oupjournals.
                                        org/cgi/content/abstract/69/3/219;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/69/3/219.

Kessler:2004:SSI

[806] Daniel Kessler, Jean-François Scheid, Giulio Schimperna, and Ulisse Stefanelli. Study of a system for the isothermal separation of components in a binary alloy with change of phase. IMA Journal of Applied Mathematics, 69(3):233–257, June 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL http://imamat.oupjournals.
                                        org/cgi/content/abstract/69/3/233;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/69/3/233.

Tooley:2004:CWP

[807] S. Tooley and J.-M. Vanden-Broeck. Capillary waves past a flat plate in water of finite depth. IMA Journal of Applied Mathematics, 69(3):259– 269, June 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oupjournals.org/cgi/content/abstract/
                                        69/3/259;   http://imamat.oupjournals.
                                        org/cgi/reprint/69/3/259.

Mason:2004:CQL

[808] Oliver Mason and Robert Shorten. On common quadratic Lyapunov functions for stable discrete-time LTI systems. IMA Journal of Applied Mathematics, 69(3):271–283, June 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/3/271;
http://imamat.oupjournals.org/
cgi/reprint/69/3/271.

Bernard:2004:ESS

[809] J. M. L. Bernard and M. A. Lyalinov. Electromagnetic scattering by a smooth convex impedance cone. IMA Journal of Applied Mathematics, 69(3):285–333, June 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/3/285;
http://imamat.oupjournals.org/
cgi/reprint/69/3/285.

Bykov:2004:IMA

[810] Viatcheslav Bykov, Igor Goldfarb, and Vladimir Gol’dshtein. On the integral manifold approach to a flame propagation problem: pressure-driven flames in porous media. IMA Journal of Applied Mathematics, 69(4):335–352, August 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/4/335;
http://imamat.oupjournals.org/
cgi/reprint/69/4/335.

Goldfarb:2004:CAT

[811] Igor Goldfarb, Vladimir Gol’dshtein, and Ulrich Maas. Comparative analysis of two asymptotic approaches based on integral manifolds. IMA Journal of Applied Mathematics, 69(4):353–374, August 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/4/353;
 http://imamat.oupjournals.org/
cgi/reprint/69/4/353.

Adler:2004:SFS

[812] J. Adler and D. M. Herbert. Smouldering front shapes for steady propagation in a stratified medium. IMA Journal of Applied Mathematics, 69(4):375–390, August 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/4/375;
 http://imamat.oupjournals.org/
cgi/reprint/69/4/375.

Baskar:2004:RPK

[813] S. Baskar and Phoolan Prasad. Riemann problem for kinematical conservation laws and geometrical features of nonlinear wavefronts. IMA Journal of Applied Mathematics, 69(4):391–419, August 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/4/391;
 http://imamat.oupjournals.org/
cgi/reprint/69/4/391.

Su:2004:ATD

[814] Jianzhong Su and Bao Loc Tran. Axisymmetric three-dimensional finger solutions in Mullins–Sekerka equation. IMA Journal of Applied Mathematics, 69(4):421–435, August 2004. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/69/4/421;
 http://imamat.oupjournals.org/
cgi/reprint/69/4/421.

Scott:2005:E

[815] Nigel Scott. Editorial. IMA Journal of Applied Mathematics, 70(1):1– 2, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oupjournals.org/cgi/reprint/70/
1/1.

Destrade:2005:IWM

[816] M. Destrade. On interface waves in misoriented pre-stressed incompressible elastic solids. IMA Journal of Applied Mathematics, 70(1):3–14, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oupjournals.org/cgi/content/abstract/
70/1/3;   http://imamat.oupjournals.
org/cgi/reprint/70/1/3.

Eden:2005:TWS

[817] Alp Eden and Saadet Erbay. On travelling wave solutions of a generalized Davey–Stewartson system. IMA Journal of Applied Mathematics, 70(1):15–24, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/15;
 http://imamat.oupjournals.org/
cgi/reprint/70/1/15.

Erbay:2005:DEC

[818] H. A. Erbay and V. H. Tüzel. Dynamic extension of a compressible nonlinearly elastic membrane tube. IMA Journal of Applied Mathematics, 70(1):25–38, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL http://imamat.oupjournals.
                                        org/cgi/content/abstract/70/1/25;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/70/1/25.

Craciun:2005:GST

[819] Eduard-Marius Craciun, Eveline Baesu, and Eugen Soós. General solution in terms of complex potentials for incremental antiplane states in prestressed and prepolarized piezoelectric crystals: application to Mode III fracture propagation. IMA Journal of Applied Mathematics, 70(1):39–52, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL http://imamat.oupjournals.
                                        org/cgi/content/abstract/70/1/39;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/70/1/39.

Aron:2005:CAS

[820] M. Aron. Combined axial shearing, extension, and straightening of elastic annular cylindrical sectors. IMA Journal of Applied Mathematics, 70(1):53–63, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL http://imamat.oupjournals.
                                        org/cgi/content/abstract/70/1/53;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/70/1/53.

Dunwoody:2005:CCDa

[821] J. Dunwoody. On a class of controllable deformations of isotropic incompressible elastic solids with simple material inhomogeneity. IMA Journal of Applied Mathematics, 70(1):64–79, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oupjournals.org/cgi/content/abstract/
70/1/64;   http://imamat.oupjournals.
org/cgi/reprint/70/1/64.

Horgan:2005:LGA

[822] Cornelius O. Horgan and Jeremiah G. Murphy. Lie group analysis and plane strain bending of cylindrical sectors for compressible nonlinearly elastic materials. IMA Journal of Applied Mathematics, 70(1):80–91, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/80;
http://imamat.oupjournals.org/
cgi/reprint/70/1/80.

McCue:2005:NSV

[823] Scott W. McCue, I. Kenneth Johnpillai, and James M. Hill. New stress and velocity fields for highly frictional granular materials. IMA Journal of Applied Mathematics, 70(1):92–118, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/92;
http://imamat.oupjournals.org/
cgi/reprint/70/1/92.

Horgan:2005:SDT

[824] C. O. Horgan and R. Quintanilla. Spatial decay of transient end effects for nonstandard linear diffusion problems. IMA Journal of Applied Mathematics, 70(1):119–128, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/119;
 http://imamat.oupjournals.org/
cgi/reprint/70/1/119.

Ivanov:2005:RRS

[825] Sergei A. Ivanov, Chi-Sing Man, and Gen Nakamura. Recovery of residual stress in a vertically heterogeneous elastic medium. IMA Journal of Applied Mathematics, 70(1):129–146, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oupjournals.org/cgi/content/abstract/
70/1/129;   http://imamat.oupjournals.
org/cgi/reprint/70/1/129.

Pour:2005:WAB

[826] M. Sanjarani Pour. A WKB analysis of the buckling condition for a cylindrical shell of arbitrary thickness subjected to an external pressure. IMA Journal of Applied Mathematics, 70(1):147–161, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/147;
 http://imamat.oupjournals.org/
cgi/reprint/70/1/147.

Wadee:2005:SLS

[827] M. Khurram Wadee. Stability of localized solutions under rigid loading in a heuristic buckling model. IMA Journal of Applied Mathematics, 70(1):162–172, February 2005.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/162;
 http://imamat.oupjournals.org/
cgi/reprint/70/1/162.

Pence:2005:SIM

[828] Thomas J. Pence and Hungyu Tsai. Swelling-induced microchannel formation in nonlinear elasticity. IMA Journal of Applied Mathematics, 70(1):173–189, February 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/1/173;
 http://imamat.oupjournals.org/
cgi/reprint/70/1/173.

Tang:2005:GTD

[829] Yanbin Tang and Li Zhou. Great time delay in a system with material cycling and delayed biomass growth. IMA Journal of Applied Mathematics, 70(2):191–200, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/2/191;
 http://imamat.oupjournals.org/
cgi/reprint/70/2/191.

Selsil:2005:MMH

[830] A. Selsil, N. V. Movchan, A. B. Movchan, S. T. Kolaczkowski, and S. Awdry. Mathematical modelling of heat transfer in a catalytic reformer. IMA Journal of Applied Mathematics, 70(2):201–220, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.

                                        org/cgi/content/abstract/70/2/201;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/70/2/201.

Chen:2005:GEB

[831] Hua Chen and Xin-Hua Zhong. Global existence and blow-up for the solutions to nonlinear parabolic– elliptic system modelling chemotaxis. IMA Journal of Applied Mathematics, 70(2):221–240, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL http://imamat.oupjournals.
                                        org/cgi/content/abstract/70/2/221;
                                         http://imamat.oupjournals.org/
                                        cgi/reprint/70/2/221.

He:2005:PEO

[832] Wen-Ming He and Jun-Zhi Cui. A pointwise estimate on the 1-order approximation of G_{x_{0}}^{ϵ}. IMA Journal of Applied Mathematics, 70(2):241– 269, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oupjournals.org/cgi/content/abstract/
                                        70/2/241;   http://imamat.oupjournals.
                                        org/cgi/reprint/70/2/241.

Amendola:2005:ABE

[833] Giovambattista Amendola. On the asymptotic behaviour for an electromagnetic system with a dissipative boundary condition. IMA Journal of Applied Mathematics, 70(2):270– 292, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oupjournals.org/cgi/content/abstract/
                                        70/2/270;   http://imamat.oupjournals.
                                        org/cgi/reprint/70/2/270.

Wilson:2005:RPW

[834] S. K. Wilson and B. R. Duffy. A rivulet of perfectly wetting fluid draining steadily down a slowly varying substrate. IMA Journal of Applied Mathematics, 70(2):293–322, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/2/293;
http://imamat.oupjournals.org/
cgi/reprint/70/2/293.

Marin:2005:BEL

[835] Liviu Marin and Daniel Lesnic. Boundary element-Landweber method for the Cauchy problem in linear elasticity. IMA Journal of Applied Mathematics, 70(2):323–340, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/2/323;
http://imamat.oupjournals.org/
cgi/reprint/70/2/323.

Singh:2005:CFS

[836] B. M. Singh, J. Rokne, and R. S. Dhaliwal. Closed-form solutions for a rectangular layer with central crack under anti-plane shear stress. IMA Journal of Applied Mathematics, 70(2):341–352, April 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oupjournals.
org/cgi/content/abstract/70/2/341;
http://imamat.oupjournals.org/
cgi/reprint/70/2/341.

Renardy:2005:CSS

[837] Michael Renardy. A comment on selfsimilar breakup for inertialess Newtonian liquid jets. IMA Journal of Applied Mathematics, 70(3):353– 358, June 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/3/353;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/3/353.

Smith:2005:SSN

[838] Warren R. Smith. On the sensitivity of strongly nonlinear autonomous oscillators and oscillatory waves to small perturbations. IMA Journal of Applied Mathematics, 70(3):359– 385, June 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/3/359;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/3/359.

Horstmann:2005:SCD

[839] Dirk Horstmann. On some crossdiffusion models in population dynamics and their connections to well-posed filters in signal enhancement processes. IMA Journal of Applied Mathematics, 70(3):386–399, June 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/3/386;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/3/386.

Howe:2005:RSC

[840] M. S. Howe. On the role of separation in compression wave generation by a train entering a tunnel hood with a window. IMA Journal of Applied Mathematics, 70(3):400– 418, June 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/3/400;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/3/400.

Lobo:2005:EOC

[841] Miguel Lobo, Serguei A. Nazarov, and Eugenia Perez. Eigen-oscillations of contrasting non-homogeneous elastic bodies: asymptotic and uniform estimates for eigenvalues. IMA Journal of Applied Mathematics, 70(3):419– 458, June 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/3/419;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/3/419.

Wang:2005:RBP

[842] Jun-Min Wang, Gen-Qi Xu, and Siu-Pang Yung. Riesz basis property, exponential stability of variable coefficient Euler–Bernoulli beams with indefinite damping. IMA Journal of Applied Mathematics, 70(3):459– 477, June 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/3/459;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/3/459.

Hui:2005:PIL

[843] Jing Hui and Lansun Chen. Periodicity in an impulsive logistic equation with a distributed delay. IMA Journal of Applied Mathematics, 70(4):479– 487, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/4/479;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/4/479.

Collins:2005:SRV

[844] W. D. Collins. Synchronous running of voltage-fed miniature single-stator permanent magnet motors. IMA Journal of Applied Mathematics, 70(4):488– 508, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/4/488;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/4/488.

Qin:2005:ESM

[845] Yuming Qin. Exponential stability and maximal attractors for a one-dimensional nonlinear thermoviscoelasticity. IMA Journal of Applied Mathematics, 70(4):509–526, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/4/509;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/4/509.

Merkin:2005:ECA

[846] J. H. Merkin and H. Sevcíková. The effects of a complexing agent on travelling waves in autocatalytic systems with applied electric fields. IMA Journal of Applied Mathematics, 70(4):527– 549, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/4/527;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/4/527.

Saccomandi:2005:SGP

[847] Giuseppe Saccomandi. Some generalized pseudo-plane deformations for the neo-Hookean material. IMA Journal of Applied Mathematics, 70(4):550– 563, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/4/550;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/4/550.

Fokas:2005:TML

[848] A. S. Fokas and B. Pelloni. A transform method for linear evolution PDEs on a finite interval. IMA Journal of Applied Mathematics, 70(4):564– 587, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/4/564;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/4/564.

Blyth:2005:NSC

[849] M. G. Blyth and J.-M. Vanden-Broeck. New solutions for capillary waves on curved sheets of fluid. IMA Journal of Applied Mathematics, 70(4):588– 601, August 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/4/588;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/4/588.

Fenwick:2005:F

[850] Alan J. Fenwick, Marian Wiercigroch, and Alan R. Champneys. Foreword. IMA Journal of Applied Mathematics, 70(5):603–604, October 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

                                        //imamat.oxfordjournals.org/cgi/
                                        reprint/70/5/603.

Ramakrishnan:2005:RPR

[851] N. Ramakrishnan and N. Sri Namachchivaya. Random perturbations of 1: 1-resonant systems with SO(2) symmetry. IMA Journal of Applied Mathematics, 70(5):605–625, October 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        70/5/605.

Barkow:2005:TSD

[852] B. Barkow, A. Steindl, and H. Troger. A targeting strategy for the deployment of a tethered satellite system. IMA Journal of Applied Mathematics, 70(5):626–644, October 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

                                        //imamat.oxfordjournals.org/cgi/
                                        reprint/70/5/626.

Cartmell:2005:OVR

[853] M. P. Cartmell, F. C. N. Lim, A. Cardoni, and M. Lucas. Optimisation of the vibrational response of ultrasonic cutting systems. IMA Journal of Applied Mathematics, 70(5): 645–656, October 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
reprint/70/5/645.

Kobayashi:2005:FUS

[854] Ryota Kobayashi, Youichi Miyazaki, and Shigeru Shinomoto. Faithful and unfaithful students in time series learning. IMA Journal of Applied Mathematics, 70(5):657–665, October 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/reprint/
70/5/657.

Mullin:2005:ESF

[855] T. Mullin, Y. Li, C. Del Pino, and J. Ashmore. An experimental study of fixed points and chaos in the motion of spheres in a Stokes flow. IMA Journal of Applied Mathematics, 70 (5):666–676, October 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
reprint/70/5/666.

Daniels:2005:WSC

[856] P. G. Daniels and V. R. Ferguson. Wavelength selection in convection near a lateral boundary. IMA Journal of Applied Mathematics, 70(6):677–699, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/70/6/677;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/677.

Asakura:2005:GHC

[857] Fumioki Asakura and Mitsuru Yamazaki. Geometry of Hugoniot curves in 2 × 2 systems of hyperbolic conservation laws with quadratic flux functions. IMA Journal of Applied Mathematics, 70(6):700–722, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/700;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/700.

Hubbard:2005:PFS

[858] M. E. Hubbard, J. A. Leach, and J. C. Wei. Pattern formation in a 2D simple chemical system with general orders of autocatalysis and decay. IMA Journal of Applied Mathematics, 70(6):723–747, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/723;   http://imamat.
oxfordjournals.org/cgi/content/
full/hxh076/DC1;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/723.

Gottlieb:2005:TBI

[859] H. P. W. Gottlieb. Transformations between isospectral membranes yield conformal maps. IMA Journal of Applied Mathematics, 70(6):748–752, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/748;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/748.

Elabbasy:2005:PSO

[860] E. M. Elabbasy and S. H. Saker.

Periodic solutions and oscillation of discrete non-linear delay population dynamics model with external force. IMA Journal of Applied Mathematics, 70(6):753–767, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/753;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/753.

Andrews:2005:NAD

[861] K. T. Andrews, J. R. Fernández, and M. Shillor. Numerical analysis of dynamic thermoviscoelastic contact with damage of a rod. IMA Journal of Applied Mathematics, 70(6):768–795, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/768;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/768.

Marchant:2005:ASM

[862] T. R. Marchant and N. F. Smyth. Approximate solutions for magmon propagation from a reservoir. IMA Journal of Applied Mathematics, 70(6):796–813, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/796;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/796.

Grundy:2005:AHI

[863] R. E. Grundy. The application of Hermite interpolation to the analysis of non-linear diffusive initial-boundary value problems. IMA Journal of Applied Mathematics, 70(6):814–838, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/6/814;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/6/814.

Arens:2005:RCR

[864] Tilo Arens and Thorsten Hohage. On radiation conditions for rough surface scattering problems. IMA Journal of Applied Mathematics, 70(6):839–847, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/6/839;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/6/839.

Sun:2005:NLF

[865] Wujun Sun, Hongmin Chen, and Donghua Cui. Non-linear functional boundary value problem for discontinuous functional differential equations. IMA Journal of Applied Mathematics, 70(6):848–857, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/6/848;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/6/848.

Al-Omari:2005:MWF

[866] J. F. M. Al-Omari and S. A. Gourley. Monotone wave-fronts in a structured population model with distributed maturation delay. IMA Journal of Applied Mathematics, 70(6):858–879, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/858;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/858.

Shikhmurzaev:2005:CBL

[867] Yulii D. Shikhmurzaev. Capillary breakup of liquid threads: a singularity-free solution. IMA Journal of Applied Mathematics, 70(6): 880–907, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/70/6/880;   http:
//imamat.oxfordjournals.org/cgi/
reprint/70/6/880.

Biggs:2005:WSA

[868] N. R. T. Biggs and D. Porter. Wave scattering by an array of periodic barriers. IMA Journal of Applied Mathematics, 70(6):908–936, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/70/6/908;   http://imamat.
oxfordjournals.org/cgi/reprint/
70/6/908.

Dunwoody:2005:CCDb

[869] J. Dunwoody. On a class of controllable deformations of isotropic incompressible elastic solids with simple material inhomogeneity: II. IMA Journal of Applied Mathematics, 70(6):937–953, December 2005. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/70/6/937;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        70/6/937.

Atanackovic:2006:MVR

[870] Teodor M. Atanackovic, Ljubica Oparnica, and Stevan Pilipovic. On a model of viscoelastic rod in unilateral contact with a rigid wall. IMA Journal of Applied Mathematics, 71 (1):1–13, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

                                        //imamat.oxfordjournals.org/cgi/
                                        content/abstract/71/1/1;   http:
                                        //imamat.oxfordjournals.org/cgi/
                                        reprint/71/1/1.

Dai:2006:GSS

[871] Hui-Hui Dai and De-Xing Kong. Global structure stability of impactinduced tensile waves in a rubberlike material. IMA Journal of Applied Mathematics, 71(1):14–33, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/1/14;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/1/14.

Napoli:2006:SLC

[872] Gaetano Napoli. On smectic-A liquid crystals in an electrostatic field. IMA Journal of Applied Mathematics, 71(1):34–46, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/71/1/34;   http://imamat.

oxfordjournals.org/cgi/reprint/
71/1/34.

Liu:2006:RTT

[873] Bin Liu, Xinzhi Liu, Kok Lay Teo, and Qing Wang. Razumikhin-type theorems on exponential stability of impulsive delay systems. IMA Journal of Applied Mathematics, 71(1):47–61, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/1/47;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/1/47.

Calvo:2006:CMM

[874] Nati Calvo, José Durany, Ana I. Muñoz, Emanuele Schiavi, and Carlos Vázquez. A coupled multivalued model for ice streams and its numerical simulation. IMA Journal of Applied Mathematics, 71(1):62–91, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/1/62;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/1/62.

Stewart:2006:CCP

[875] David E. Stewart. Convolution complementarity problems with application to impact problems. IMA Journal of Applied Mathematics, 71(1):92–119, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/1/92;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/1/92.

Smith:2006:NUM

[876] D. H. Smith. Non-uniqueness and multi-shock solutions for transonic nozzle flows. IMA Journal of Applied Mathematics, 71(1):120–132, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/1/120;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/1/120.

Coman:2006:ILB

[877] Ciprian D. Coman. Inhomogeneities and localised buckling patterns. IMA Journal of Applied Mathematics, 71(1):133–152, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/71/1/133;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/1/133.

Blyth:2006:LIG

[878] M. G. Blyth and C. Pozrikidis. A Lobatto interpolation grid over the triangle. IMA Journal of Applied Mathematics, 71(1):153–169, February 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/1/153;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/1/153.

Kaplunov:2006:LTP

[879] J. D. Kaplunov, A. V. Pichugin, and G. A. Rogerson. On a Lamb-type problem for a bi-axially pre-stressed incompressible elastic plate. IMA Journal of Applied Mathematics, 71(2):171– 185, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/2/171;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/2/171.

Lacey:2006:MPM

[880] A. A. Lacey and C. V. Nikolopoulos. A 1D model for polymer melting during modulated temperature differential scanning calorimetry. IMA Journal of Applied Mathematics, 71(2):186– 209, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/2/186;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/2/186.

Hillgarter:2006:CSC

[881] Erik Hillgarter. A contribution to the symmetry classification problem for second-order PDEs in z(x, y). IMA Journal of Applied Mathematics, 71 (2):210–231, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/71/2/210;   http:
//imamat.oxfordjournals.org/cgi/
reprint/71/2/210.

Chen:2006:STG

[882] Guanggan Chen and Jian Zhang. Sharp threshold of global existence for non-linear Gross–Pitaevskii equation in N. IMA Journal of Applied Mathematics, 71(2):232–240, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/2/232;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/2/232.

Berg:2006:ECP

[883] Peter Berg, Atife Caglar, Keith Promislow, Jean St-Pierre, and Brian Wetton. Electrical coupling in proton exchange membrane fuel cell stacks: mathematical and computational modelling. IMA Journal of Applied Mathematics, 71(2):241– 261, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/2/241;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/2/241.

Johansson:2006:IMC

[884] Tomas Johansson. An iterative method for a Cauchy problem for the heat equation. IMA Journal of Applied Mathematics, 71(2):262– 286, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/2/262;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/2/262.

Xu:2006:SMT

[885] Shufang Xu and Mingsong Cheng. On the solvability for the mixed-type Lyapunov equation. IMA Journal of Applied Mathematics, 71(2):287– 294, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/2/287;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/2/287.

Cimatti:2006:NEU

[886] Giovanni Cimatti. A note on existence, uniqueness and continuous dependence for a non-linear heat conduction problem. IMA Journal of Applied Mathematics, 71(2):295–297, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/2/295;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/2/295.

Luo:2006:LIG

[887] H. Luo and C. Pozrikidis. A Lobatto interpolation grid in the tetrahedron. IMA Journal of Applied Mathematics, 71(2):298–313, April 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/2/298;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/2/298.

Bijura:2006:ILT

[888] Angelina M. Bijura. Initial-layer theory and model equations of Volterra type. IMA Journal of Applied Mathematics, 71(3):315–331, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/3/315;   http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        71/3/315.

Soldatos:2006:SCR

[889] Kostas P. Soldatos. On the stability of a compressible Rivlin’s cube made of transversely isotropic material. IMA Journal of Applied Mathematics, 71(3):332–353, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/71/3/332;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/3/332.

Levine:2006:CDP

[890] Harold Levine. A classical diffusion problem. IMA Journal of Applied Mathematics, 71(3):354–358, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        71/3/354.

Bykov:2006:MVI

[891] Viatcheslav Bykov, Igor Goldfarb, Vladimir Gol’dshtein, and Ulrich Maas. On a modified version of ILDM approach: asymptotic analysis based on integral manifolds. IMA Journal of Applied Mathematics, 71(3):359– 382, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/3/359;   http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        full/hxh100/DC1;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/3/359.

Youssef:2006:TTT

[892] H. M. Youssef. Theory of twotemperature-generalized thermoelasticity. IMA Journal of Applied Mathematics, 71(3):383–390, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/3/383;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/3/383.

Cherniha:2006:LSC

[893] Roman Cherniha and John R. King. Lie symmetries and conservation laws of non-linear multidimensional reaction–diffusion systems with variable diffusivities. IMA Journal of Applied Mathematics, 71(3):391–408, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/3/391;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/3/391.

Salnikov:2006:ANA

[894] Valeriy Salnikov and Nigel Scott. Asymptotic and numerical analysis of thermoelastic waves in an isotropic plate. IMA Journal of Applied Mathematics, 71(3):409–433, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/3/409;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/3/409.

Fu:2006:EES

[895] Y. B. Fu and D. W. Brookes. An explicit expression for the surfaceimpedance tensor of a compressible monoclinic material in a state of plane strain. IMA Journal of Applied Mathematics, 71(3):434–445, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/3/434;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/3/434.

Needham:2006:FAR

[896] D. J. Needham. On the formation of acceleration and reaction–diffusion wavefronts in autocatalytic-type reaction– diffusion systems. IMA Journal of Applied Mathematics, 71(3):446– 458, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/3/446;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/3/446.

Irmscher:2006:SDR

[897] Tilman Irmscher and Reinhard Racke. Sharp decay rates in parabolic and hyperbolic thermoelasticity. IMA Journal of Applied Mathematics, 71(3):459– 478, June 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/3/459;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/3/459.

Homberg:2006:TCP

[898] Dietmar Hömberg and Alexander Khludnev. A thermoelastic contact problem with a phase transition. IMA Journal of Applied Mathematics, 71(4):479–495, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/4/479;   http://imamat.
oxfordjournals.org/cgi/content/
full/hxl003/DC1;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/4/479.

Guo:2006:NLW

[899] Shangjiang Guo and Lihong Huang. Non-linear waves in a ring of neurons. IMA Journal of Applied Mathematics, 71(4):496–518, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/4/496;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/4/496.

Simon:2006:CPC

[900] Peter L. Simon. Classification of positive convex functions according to focal equivalence. IMA Journal of Applied Mathematics, 71(4):519–533, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/4/519;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/4/519.

Chakrabarti:2006:SCW

[901] A. Chakrabarti. Solution of certain weakly singular integral equations. IMA Journal of Applied Mathematics, 71(4):534–543, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/71/4/534;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/4/534.

Ruktamatakul:2006:WFS

[902] Sittipong Ruktamatakul, Jonathan Bell, and Yongwimon Lenbury. Wavefront solution behaviour for continuous neural networks with lateral inhibition. IMA Journal of Applied Mathematics, 71(4):544–564, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/4/544;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/4/544.

Wang:2006:SSP

[903] Jun-Min Wang and Bao-Zhu Guo. On the stability of swelling porous elastic soils with fluid saturation by one internal damping. IMA Journal of Applied Mathematics, 71(4):565–582, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/71/4/565;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        71/4/565.

Gottlieb:2006:RIA

[904] H. P. W. Gottlieb. Radially isospectral annular membranes. IMA Journal of Applied Mathematics, 71(4):583– 589, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/4/583;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/4/583.

Payne:2006:PMH

[905] L. E. Payne and J. C. Song. On a proposed model for heat conduction. IMA Journal of Applied Mathematics, 71(4):590–599, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/4/590;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/4/590.

Grajales:2006:IBT

[906] Juan Carlos Muñoz Grajales and André Nachbin. Improved Boussinesqtype equations for highly variable depth. IMA Journal of Applied Mathematics, 71(4):600–633, August 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/4/600;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/4/600.

Kovtunenko:2006:PDM

[907] V. A. Kovtunenko. Primal–dual methods of shape sensitivity analysis for curvilinear cracks with nonpenetration. IMA Journal of Applied Mathematics, 71(5):635–657, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/635;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/635.

Guo:2006:SSS

[908] Zhenhua Guo and Song Jiang. Selfsimilar solutions to the isothermal compressible Navier–Stokes equations. IMA Journal of Applied Mathematics, 71(5):658–669, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/658;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/658.

MacKay:2006:DCG

[909] Cameron MacKay, Sean McKee, and Anthony J. Mulholland. Diffusion and convection of gaseous and fine particulate from a chimney. IMA Journal of Applied Mathematics, 71(5):670– 691, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/670;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/670.

BuAli:2006:GES

[910] Q. BuAli, L. E. Johns, and R. Narayanan. Geometric effects on surface roughness in electrodeposition. IMA Journal of Applied Mathematics, 71(5):692– 714, October 2006. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/692;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/692.

Segin:2006:LWL

[911] Tetyana M. Segin, Lou Kondic, and Burt S. Tilley. Long-wave linear stability theory for two-fluid channel flow including compressibility effects. IMA Journal of Applied Mathematics, 71(5):715–739, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/715;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/715.

Decent:2006:HAD

[912] S. P. Decent. Hydrodynamic assist and the dynamic contact angle in the coalescence of liquid drops. IMA Journal of Applied Mathematics, 71(5):740– 767, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/740;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/740.

Kaplunov:2006:EME

[913] J. Kaplunov, A. Zakharov, and D. Prikazchikov. Explicit models for elastic and piezoelastic surface waves. IMA Journal of Applied Mathematics, 71(5):768–782, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/768;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/768.

Sudak:2006:GFP

[914] L. J. Sudak and X. Wang. Green’s function in plane anisotropic bimaterials with imperfect interface. IMA Journal of Applied Mathematics, 71(5):783– 794, October 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/5/783;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/5/783.

Smith:2006:TPS

[915] Ronald Smith. Three-point scheme and two-point boundary formulations for unsteady diffusion with flow. IMA Journal of Applied Mathematics, 71(6):795–810, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/6/795;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/795.

Gayen:2006:WWS

[916] Rupanwita Gayen, B. N. Mandal, and A. Chakrabarti. Water wave scattering by two sharp discontinuities in the surface boundary conditions. IMA Journal of Applied Mathematics, 71(6):811–831, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/

abstract/71/6/811;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/811.

Fasino:2006:RUT

[917] Dario Fasino and Gabriele Inglese. Recovering unknown terms in a nonlinear boundary condition for Laplace’s equation. IMA Journal of Applied Mathematics, 71(6):832–852, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/6/832;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/832.

Duran:2006:HEL

[918] Mario Durán, Ignacio Muga, and Jean-Claude Nédélec. The Helmholtz equation in a locally perturbed half-plane with passive boundary. IMA Journal of Applied Mathematics, 71(6):853–876, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/6/853;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/853.

Donaldson:2006:SSI

[919] Roger D. Donaldson and Brian R. Wetton. Solving steady interface problems using residual velocities. IMA Journal of Applied Mathematics, 71(6):877–897, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/6/877;   http://imamat.

oxfordjournals.org/cgi/reprint/
71/6/877.

Volkov:2006:SPC

[920] D. Volkov and G. A. Kriegsmann. Scattering by a perfect conductor in a waveguide: energy-preserving schemes for integral equations. IMA Journal of Applied Mathematics, 71(6):898–923, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/6/898;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/898.

Ryabenkii:2006:ASM

[921] Victor S. Ryaben’kii and Sergei V. Utyuzhnikov. Active shielding model for hyperbolic equations. IMA Journal of Applied Mathematics, 71(6):924–939, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/71/6/924;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/924.

Leppington:2006:TST

[922] F. G. Leppington, E. G. Broadbent, and G. F. Butler. Transmission of sound through a pair of rectangular elastic plates. IMA Journal of Applied Mathematics, 71(6):940–955, December 2006. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/71/6/940;   http://imamat.
oxfordjournals.org/cgi/reprint/
71/6/940.

Fromhold:2007:SCE

[923] A. T. Fromhold. Space-charge effects on current transport in media having cylindrical and spherical symmetries. IMA Journal of Applied Mathematics, 72(1):1–8, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/72/1/1;   http:
//imamat.oxfordjournals.org/cgi/
reprint/72/1/1.

Blanariu:2007:WNL

[924] Mihaela Blanariu and Brian J. Spencer. Weakly non-linear bifurcation analysis of pattern formation in strained alloy film growth. IMA Journal of Applied Mathematics, 72(1): 9–35, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/72/1/9;   http:
//imamat.oxfordjournals.org/cgi/
reprint/72/1/9.

Shu:2007:SCG

[925] Ji Shu and Jian Zhang. Sharp conditions of global existence for the generalized Davey–Stewartson system. IMA Journal of Applied Mathematics, 72(1):36–42, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/72/1/36;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/1/36.

Chapman:2007:KCS

[926] S. J. Chapman. The Kelly criterion for spread bets. IMA Journal of Applied Mathematics, 72(1):43–51, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/1/43;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/1/43.

Chamberlain:2007:WWS

[927] P. G. Chamberlain. Water wave scattering by finite arrays of circular structures. IMA Journal of Applied Mathematics, 72(1):52–66, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/1/52;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/1/52.

Han:2007:TNA

[928] Weimin Han and Ge Wang. Theoretical and numerical analysis on multispectral bioluminescence tomography. IMA Journal of Applied Mathematics, 72(1):67–85, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/72/1/67;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/1/67.

Eldabe:2007:HSP

[929] Nabil T. M. Eldabe, M. F. El-Sabbagh, and M. A.-S. El-Sayed (Hajjaj). Hydromagnetic stability of plane Couette flow of an upper convected Maxwell fluid. IMA Journal of Applied Mathematics, 72(1):86–95, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/1/86;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/1/86.

Johansson:2007:RAS

[930] Tomas Johansson and Brian D. Sleeman. Reconstruction of an acoustically sound-soft obstacle from one incident field and the far-field pattern. IMA Journal of Applied Mathematics, 72(1):96–112, February 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/1/96;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/1/96.

Li:2007:SLE

[931] Yi Li and Chunshan Zhao. On the shape of least-energy solutions for a class of quasilinear elliptic Neumann problems. IMA Journal of Applied Mathematics, 72(2):113–139, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/2/113;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/2/113.

Kang:2007:SDS

[932] K. Kang, T. Kolokolnikov, and M. J. Ward. The stability and dynamics of a spike in the 1D Keller– Segel model. IMA Journal of Applied Mathematics, 72(2):140–162, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/2/140;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/2/140.

Smith:2007:EME

[933] Warren R. Smith. Explicit modulation equations, Reynolds averaging and the closure problem for the Korteweg–de Vries–Burgers equation. IMA Journal of Applied Mathematics, 72(2):163– 179, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/2/163;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/2/163.

Fabrikant:2007:UDI

[934] V. I. Fabrikant. Utilization of divergent integrals and a new symbolism in contact and crack analysis. IMA Journal of Applied Mathematics, 72(2):180– 190, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/2/180;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/2/180.

Muriel:2007:SNS

[935] C. Muriel and J. L. Romero. C^{∞}- symmetries and nonlocal symmetries of exponential type. IMA Journal of Applied Mathematics, 72(2):191– 205, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/2/191;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/2/191.

Hoseini:2007:SWI

[936] S. M. Hoseini and T. R. Marchant. Solitary wave interaction for a higherorder nonlinear Schrödinger equation. IMA Journal of Applied Mathematics, 72(2):206–222, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/72/2/206;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/2/206.

Parnell:2007:EWP

[937] W. J. Parnell. Effective wave propagation in a prestressed nonlinear elastic composite bar. IMA Journal of Applied Mathematics, 72(2):223– 244, April 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/2/223;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/2/223.

Kirchner:2007:SDG

[938] Gerhard Kirchner. A systematic derivation of Green’s and Poisson’s matrices for boundary-value problems in elastostatic half-spaces. IMA Journal of Applied Mathematics, 72(3):245– 270, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/3/245;   http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        72/3/245.

Wang:2007:SFC

[939] C. Y. Wang. Stagnation flow on a cylinder with partial slip—an exact solution of the Navier–Stokes equations. IMA Journal of Applied Mathematics, 72(3):271–277, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/3/271;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/3/271.

Crowdy:2007:GFL

[940] Darren Crowdy and Jonathan Marshall. Green’s functions for Laplace’s equation in multiply connected domains. IMA Journal of Applied Mathematics, 72(3):278–301, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/72/3/278;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/3/278.

Shah:2007:MMC

[941] A. A. Shah, Gwang-Soo Kim, and K. Promislow. Mathematical modelling of the catalyst layer of a polymer electrolyte fuel cell. IMA Journal of Applied Mathematics, 72(3):302– 330, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/3/302;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/3/302.

Chen:2007:ESS

[942] Hua Chen and Shaohua Wu. On existence of solutions for some hyperbolic–parabolic-type chemotaxis systems. IMA Journal of Applied Mathematics, 72(3):331–347, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/3/331;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/3/331.

Nolde:2007:QAI

[943] E. Nolde. Qualitative analysis of initial-value problems for a thin elastic strip. IMA Journal of Applied Mathematics, 72(3):348–375, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/3/348;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/3/348.

Lawrie:2007:EEA

[944] Jane B. Lawrie. On eigenfunction expansions associated with wave propagation along ducts with wavebearing boundaries. IMA Journal of Applied Mathematics, 72(3):376– 394, June 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/3/376;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/3/376.

Clarke:2007:BEU

[945] Sam Clarke and Peter Grindrod. A Bayesian estimation of unfolded pitchfork bifurcation structure based upon experimental data. IMA Journal of Applied Mathematics, 72(4):395–404, August 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/4/395;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/4/395.

Bondar:2007:IBF

[946] M. L. Bondar, R. M. M. Mattheij, and J. H. M. ten Thije Boonkkamp. Investigation of Bunsen flame dynamics by the method of characteristics. IMA Journal of Applied Mathematics, 72(4):405– 419, August 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/4/405;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/4/405.

Hek:2007:SLF

[947] Geertje Hek and Vivi Rottschäfer. Semiconductor laser with filtered optical feedback: from optical injection to conventional feedback. IMA Journal of Applied Mathematics, 72(4):420– 450, August 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/4/420;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/4/420.

Ishak:2007:DSM

[948] A. Ishak, R. Nazar, and I. Pop. Dual solutions in mixed convection boundary-layer flow with suction or injection. IMA Journal of Applied Mathematics, 72(4):451–463, August 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/4/451;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/4/451.

Pan:2007:ABN

[949] Jiaqing Pan. The approximating behavior on nonlinearities of solutions of a degenerate parabolic equation. IMA Journal of Applied Mathematics, 72(4):464–475, August 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/4/464;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/4/464.

Thompson:2007:ECS

[950] I. Thompson and C. M. Linton. On the excitation of a closely spaced array by a line source. IMA Journal of Applied Mathematics, 72(4):476– 497, August 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/4/476;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/4/476.

Gilardi:2007:WPL

[951] Gianni Gilardi and Elisabetta Rocca. Well-posedness and long-time behaviour for a singular phase field system of conserved type. IMA Journal of Applied Mathematics, 72(4):498– 530, August 2007. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/4/498;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/4/498.

Quintanilla:2007:E

[952] Ramon Quintanilla and Giuseppe Saccomandi. Editorial. IMA Journal of Applied Mathematics, 72(5):531, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        72/5/531.

Hayes:2007:FRK

[953] Michael Hayes. Foreword: Robin Knops. IMA Journal of Applied Mathematics, 72(5):532–539, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        72/5/532.

Flavin:2007:CDI

[954] J. N. Flavin and S. Rionero. Crossdiffusion influence on the non-linear L^{2}-stability analysis for a Lotka– Volterra reaction–diffusion system of PDEs. IMA Journal of Applied Mathematics, 72(5):540–555, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/5/540;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/5/540.

Horgan:2007:TFG

[955] Cornelius O. Horgan. On the torsion of functionally graded anisotropic linearly elastic bars. IMA Journal of Applied Mathematics, 72(5):556–562, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/556;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/556.

Payne:2007:BOC

[956] L. E. Payne. A bound for the optimal constant in an inequality of Ladyzhenskaya and Solonnikov. IMA Journal of Applied Mathematics, 72(5):563–569, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/563;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/563.

Villaggio:2007:MTC

[957] Piero Villaggio. A mathematical theory of climbing. IMA Journal of Applied Mathematics, 72(5):570–576, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/570;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/570.

Boulanger:2007:BES

[958] PH. Boulanger and M. Hayes. Bounds on the extensional strains in elastic homogeneous crystals under simple tension. IMA Journal of Applied Mathematics, 72(5):577–596, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/577;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/577.

Kavallaris:2007:BNL

[959] N. I. Kavallaris, A. A. Lacey, C. V. Nikolopoulos, and C. Voong. Behaviour of a non-local equation modelling linear friction welding. IMA Journal of Applied Mathematics, 72(5):597–616, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/597;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/597.

Spencer:2007:CES

[960] A. J. M. Spencer. A class of exact solutions for periodically inhomogeneous thermoelastic solids. IMA Journal of Applied Mathematics, 72(5):617–634, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/617;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/617.

Hill:2007:GSP

[961] A. A. Hill, S. Rionero, and B. Straughan. Global stability for penetrative convection with throughflow in a porous material. IMA Journal of Applied Mathematics, 72(5):635–643, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/635;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/635.

Mielke:2007:MTI

[962] Alexander Mielke. A model for temperature-induced phase transformations in finite-strain elasticity. IMA Journal of Applied Mathematics, 72(5):644–658, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/644-a;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/644-a.

Stuart:2007:EST

[963] C. A. Stuart. Existence and stability of TE modes in a stratified nonlinear dielectric. IMA Journal of Applied Mathematics, 72(5):659–679, October 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/5/659;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/5/659.

Antipov:2007:MIE

[964] Y. A. Antipov and V. V. Silvestrov. Method of integral equations for systems of difference equations in diffraction theory. IMA Journal of Applied Mathematics, 72(6):681–705, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/681;   http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        72/6/681.

Bartolucci:2007:EQP

[965] Daniele Bartolucci and Angela Pistoia. Existence and qualitative properties of concentrating solutions for the sinh-Poisson equation. IMA Journal of Applied Mathematics, 72(6):706–729, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/6/706;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/6/706.

Steigmann:2007:SWS

[966] D. J. Steigmann and R. W. Ogden. Surface waves supported by thin-film/ substrate interactions. IMA Journal of Applied Mathematics, 72(6):730–747, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/6/730;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/6/730.

Johansson:2007:VMI

[967] B. Tomas Johansson and Daniel Lesnic. A variational method for identifying a spacewise-dependent heat source. IMA Journal of Applied Mathematics, 72(6):748–760, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/6/748;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/6/748.

John:2007:SBL

[968] Jo-Anne L. John and Sharon O. Stephen. The stability of the boundary layer on a compliant rotating disc. IMA Journal of Applied Mathematics, 72(6):761–784, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/761;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/761.

Shafee:2007:LFN

[969] Fariel Shafee. Lambert function and a new non-extensive form of entropy. IMA Journal of Applied Mathematics, 72(6):785–800, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/785;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/785.

Volkov:2007:SST

[970] Darko Volkov and Roger Lui. Spreading speed and travelling wave solutions of a partially sedentary population. IMA Journal of Applied Mathematics, 72(6):801–816, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/801;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/801.

Liu:2007:ZBS

[971] Hongyu Liu and Jun Zou. Zeros of the Bessel and spherical Bessel functions and their applications for uniqueness in inverse acoustic obstacle scattering. IMA Journal of Applied Mathematics, 72(6):817–831, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/817;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/817.

Papageorgiou:2007:NAS

[972] Demetrios T. Papageorgiou and J.-M. Vanden-Broeck. Numerical and analytical studies of non-linear gravity– capillary waves in fluid layers under normal electric fields. IMA Journal of Applied Mathematics, 72(6):832–853, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/832;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/832.

Payne:2007:SDC

[973] L. E. Payne and J. C. Song. Spatial decay in a cross-diffusion problem. IMA Journal of Applied Mathematics, 72(6):854–864, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/72/6/854;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/854.

Edwards:2007:SHE

[974] David A. Edwards. Steric hindrance effects in thin reaction zones: applications to BIAcore. IMA Journal of Applied Mathematics, 72(6):865–893, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/865;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/865.

Matthews:2007:MNT

[975] Miccal T. Matthews and James M. Hill. Micro/nano thermal boundary layer equations with slip–creep–jump boundary conditions. IMA Journal of Applied Mathematics, 72(6):894–911, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/72/6/894;   http://imamat.
oxfordjournals.org/cgi/reprint/
72/6/894.

Ptashnyk:2007:PEC

[976] Mariya Ptashnyk. Pseudoparabolic equations with convection. IMA Journal of Applied Mathematics, 72(6): 912–922, December 2007. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/72/6/912;   http:
//imamat.oxfordjournals.org/cgi/
reprint/72/6/912.

Smith:2007:MER

[977] Warren R. Smith. Modulation equations and Reynolds averaging for finiteamplitude non-linear waves in an incompressible fluid. IMA Journal of Applied Mathematics, 72(6):923–945, December 2007. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/72/6/923;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        72/6/923.

Needham:2008:FAK

[978] D. J. Needham and J. Billingham. Foreword: Andy King. IMA Journal of Applied Mathematics, 73(1):1– 3, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        73/1/1.

Billingham:2008:GDT

[979] J. Billingham. Gravity-driven thin-film flow using a new contact line model. IMA Journal of Applied Mathematics, 73(1):4–36, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

                                        //imamat.oxfordjournals.org/cgi/
                                        content/abstract/73/1/4;   http:
                                        //imamat.oxfordjournals.org/cgi/
                                        reprint/73/1/4.

Decent:2008:STD

[980] S. P. Decent and A. C. King. Surface-tension-driven flow in a slender cone. IMA Journal of Applied Mathematics, 73(1):37–68, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/1/37;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/1/37.

Fozard:2008:PSM

[981] J. A. Fozard and J. R. King. Population-scale modelling of cellular chemotaxis and aggregation. IMA Journal of Applied Mathematics, 73 (1):69–106, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/1/69;   http:
//imamat.oxfordjournals.org/cgi/
reprint/73/1/69.

Sharpe:2008:NSM

[982] G. J. Sharpe, S. A. E. G. Falle, and J. Billingham. Numerical solutions of a model for the propagation of a surface-catalysed flame in a tube. IMA Journal of Applied Mathematics, 73(1):107–122, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/107;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/107.

Scott:2008:TGT

[983] N. H. Scott. Thermoelasticity and generalized thermoelasticity viewed as wave hierarchies. IMA Journal of Applied Mathematics, 73(1):123–136, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/123;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/123.

Edwards:2008:NCS

[984] Carina M. Edwards, S. D. Howison, H. Ockendon, and J. R. Ockendon. Non-classical shallow water flows. IMA Journal of Applied Mathematics, 73(1):137–157, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/137;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/137.

Needham:2008:ETW

[985] D. J. Needham and J. A. Leach. The evolution of travelling wave-fronts in a hyperbolic Fisher model. I. The travelling wave theory. IMA Journal of Applied Mathematics, 73(1):158–198, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/158;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/158.

Lukyanov:2008:CBF

[986] A. V. Lukyanov, Y. D. Shikhmurzaev, and A. C. King. A combined BIE-FE method for the Stokes equations. IMA Journal of Applied Mathematics, 73(1):199–224, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/199;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/199.

Falle:2008:RTI

[987] S. A. E. G. Falle. The role of thermal instability in star formation. IMA Journal of Applied Mathematics, 73(1):225–230, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/225;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/225.

Mealey:2008:FCB

[988] L. Mealey and J. H. Merkin. Free convection boundary layers on a vertical surface in a heat-generating porous medium. IMA Journal of Applied Mathematics, 73(1):231–253, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/231;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/231.

Binder:2008:IRC

[989] B. J. Binder, F. Dias, and J.-M. Vanden-Broeck. Influence of rapid changes in a channel bottom on free-surface flows. IMA Journal of Applied Mathematics, 73(1):254–273, February 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/1/254;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/1/254.

Needham:2008:DST

[990] D. J. Needham, J. Billingham, R. M. S. M. Schulkes, and A. C. King. The development of slugging in twolayer hydraulic flows. IMA Journal of Applied Mathematics, 73(1):274–322, February 2008. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/1/274;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/1/274.

ORegan:2008:EUL

[991] Donal O’Regan and Mohamed El-Gebeily. Existence, upper and lower solutions and quasilinearization for singular differential equations. IMA Journal of Applied Mathematics, 73(2):323– 344, April 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/2/323;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/2/323.

Adler:2008:TDC

[992] J. Adler and D. M. Herbert. Transient diffusion-controlled smoulder propagation: a similarity solution approach. IMA Journal of Applied Mathematics, 73(2):345–360, April 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/2/345;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/2/345.

Sharpe:2008:SIE

[993] G. J. Sharpe, V. Gorchkov, and M. Short. Shock initiation of explosives: the idealized condensedphase model. IMA Journal of Applied Mathematics, 73(2):361–373, April 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/2/361;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/2/361.

Kohr:2008:GFB

[994] Mirela Kohr, G. P. Raja Sekhar, and John R. Blake. Green’s function of the Brinkman equation in a 2D anisotropic case. IMA Journal of Applied Mathematics, 73(2):374– 392, April 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/2/374;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/2/374.

vanNoorden:2008:SPM

[995] T. L. van Noorden and I. S. Pop. A Stefan problem modelling crystal dissolution and precipitation. IMA Journal of Applied Mathematics, 73(2):393– 411, April 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/2/393;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/2/393.

Ortner:2008:FMS

[996] Norbert Ortner and Peter Wagner. The fundamental matrix of the system of linear elastodynamics in hexagonal media. Solution to the problem of conical refraction. IMA Journal of Applied Mathematics, 73(2):412– 447, April 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/

                                        abstract/73/2/412;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/2/412.

Fiedler:2008:LIS

[997] Bernold Fiedler and Andreas Schuppert. Local identification of scalar hybrid models with tree structure. IMA Journal of Applied Mathematics, 73(3):449–476, June 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/73/3/449;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/3/449.

Weng:2008:WNL

[998] Peixuan Weng and Jianhong Wu. Wavefronts for a non-local reaction– diffusion population model with general distributive maturity. IMA Journal of Applied Mathematics, 73(3):477– 495, June 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/3/477;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/3/477.

Sedelnikov:2008:DDT

[999] I. Sedelnikov, A. Averbuch, and Y. Shkolnisky. The discrete diffraction transform. IMA Journal of Applied Mathematics, 73(3):496–538, June 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/3/496;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        73/3/496.

Bostan:2008:HVM

[1000] M. Bostan. Homogenization of the 1D Vlasov–Maxwell equations. IMA Journal of Applied Mathematics, 73(3):539– 555, June 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/3/539;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/3/539.

Wilson:2008:PMT

[1001] Miles Wilson, John R. Blake, and Peter M. Haese. A potential multipole theory for the hydrodynamics of bubble clouds. IMA Journal of Applied Mathematics, 73(3):556– 577, June 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/3/556;   http://imamat.
oxfordjournals.org/cgi/reprint/
73/3/556.

Hasanov:2008:DUC

[1002] Alemdar Hasanov and Arzu Erdem.

Determination of unknown coefficient in a non-linear elliptic problem related to the elastoplastic torsion of a bar. IMA Journal of Applied Mathematics, 73(4):579–591, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/4/579.

Cheng:2008:SST

[1003] Cui-Ping Cheng, Wan-Tong Li, and Zhi-Cheng Wang. Spreading speeds and travelling waves in a delayed population model with stage structure on a 2D spatial lattice. IMA Journal of Applied Mathematics, 73 (4):592–618, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/4/592.

Litvinov:2008:DEC

[1004] W. G. Litvinov. Dynamics of electrorheological clutch and a problem for non-linear parabolic equation with non-local boundary conditions. IMA Journal of Applied Mathematics, 73 (4):619–640, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/4/619.

Johansson:2008:PRS

[1005] B. Tomas Johansson. A procedure for the reconstruction of a stochastic stationary temperature field. IMA Journal of Applied Mathematics, 73 (4):641–650, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/4/641.

Holmaas:2008:AIT

[1006] H. Holmås, T. Sira, M. Nordsveen, H. P. Langtangen, and R. Schulkes. Analysis of a 1D incompressible twofluid model including artificial diffusion. IMA Journal of Applied Mathematics, 73(4):651–667, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/73/4/651.

Hazard:2008:IMA

[1007] Christophe Hazard and Eric Lunéville.

An improved multimodal approach for non-uniform acoustic waveguides. IMA Journal of Applied Mathematics, 73 (4):668–690, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/4/668.

Visser:2008:MCM

[1008] F. C. Visser. On the method of conformal mapping applied to the flow around thin cylindrically curved aerofoils. IMA Journal of Applied Mathematics, 73(4):691–697, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/4/691.

Marston:2008:EEN

[1009] J. O. Marston, S. P. Decent, and M. J. H. Simmons. Experimental evidence of non-unique solutions to a steady non-linear coating flow. IMA Journal of Applied Mathematics, 73 (4):698–702, August 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/4/698.

Collins:2008:SRS

[1010] W. D. Collins and R. Brown. Synchronous running of shaded-pole permanent magnet motors. IMA Journal of Applied Mathematics, 73(5): 703–723, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

                                        //imamat.oxfordjournals.org/cgi/
                                        content/abstract/73/5/703.

Elasmi:2008:SMS

[1011] Lassaad Elasmi. Singularity method for Stokes flow with slip boundary condition. IMA Journal of Applied Mathematics, 73(5):724–739, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/5/724.

Crowdy:2008:CDS

[1012] Darren Crowdy. Conducting drops subject to electric fields in 2D Stokes flows. IMA Journal of Applied Mathematics, 73(5):740–758, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

                                        //imamat.oxfordjournals.org/cgi/
                                        content/abstract/73/5/740.

Sherratt:2008:CPT

[1013] Jonathan A. Sherratt. A comparison of periodic travelling wave generation by Robin and Dirichlet boundary conditions in oscillatory reaction– diffusion equations. IMA Journal of Applied Mathematics, 73(5):759–781, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/73/5/759.

Giesl:2008:CLG

[1014] Peter Giesl. Construction of a local and global Lyapunov function using radial basis functions. IMA Journal of Applied Mathematics, 73(5):

782–802, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/5/782.

Blyth:2008:FSF

[1015] M. G. Blyth and J.-M. Vanden-Broeck. Free-surface flow over a trapped bubble. IMA Journal of Applied Mathematics, 73(5):803–814, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/5/803.

Wang:2008:SSP

[1016] Mingxin Wang, Peter Y. H. Pang, and Wenyan Chen. Sharp spatial patterns of the diffusive Holling– Tanner prey–predator model in heterogeneous environment. IMA Journal of Applied Mathematics, 73(5): 815–835, October 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/5/815.

Bittencourt:2008:NSS

[1017] Marco L. Bittencourt and Thais G.

Vazquez. A nodal spectral stiffness matrix for the finite-element method. IMA Journal of Applied Mathematics, 73 (6):837–849, December 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/6/837.

Shen:2008:FSN

[1018] S. Shen, F. Liu, V. Anh, and I. Turner.

The fundamental solution and numerical solution of the Riesz fractional advection–dispersion equation. IMA Journal of Applied Mathematics, 73 (6):850–872, December 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/6/850.

Daniels:2008:NLW

[1019] P. G. Daniels, D. Ho, and A. C. Skeldon. Non-linear wavelength selection for pattern-forming systems in channels. IMA Journal of Applied Mathematics, 73(6):873–901, December 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/6/873.

Wollkind:2008:NLS

[1020] D. J. Wollkind, F. J. Alvarado, and D. E. Edmeade. Non-linear stability analyses of optical pattern formation in an atomic sodium vapour ring cavity. IMA Journal of Applied Mathematics, 73(6):902–935, December 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/6/902.

Mulholland:2008:EPP

[1021] Anthony J. Mulholland, Nishal Ramadas, Richard L. O’Leary, Agnes C. S. Parr, Gordon Hayward, Alexandre Troge, and Richard A. Pethrick. Enhancing the performance of piezoelectric ultrasound transducers by the use of multiple matching layers. IMA Journal of Applied Mathematics, 73 (6):936–949, December 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/73/6/936.

Postnova:2008:TMT

[1022] J. Postnova and R. V. Craster. Trapped modes in 3D topographically varying plates. IMA Journal of Applied Mathematics, 73(6):950–963, December 2008. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/73/6/950.

Hasanov:2009:SDS

[1023] Alemdar Hasanov. Simultaneous determination of the source terms in a linear hyperbolic problem from the final overdetermination: weak solution approach. IMA Journal of Applied Mathematics, 74(1):1–19, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Barlow:2009:ARP

[1024] Euan Barlow, Anthony J. Mulholland, Anthony Gachagan, Alison Nordon, and Kenneth MacPherson. Analysis of the Rayleigh–Plesset equation with chirp excitation. IMA Journal of Applied Mathematics, 74(1):20–34, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Lopez:2009:SPP

[1025] José L. López and Ester Pérez Sinusía. A singular perturbation problem with discontinuous data in a cuboid. IMA Journal of Applied Mathematics, 74(1):35–45, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Zhang:2009:SCS

[1026] Hai-Qiang Zhang, Bo Tian, Juan Li, Tao Xu, and Ya-Xing Zhang. Symboliccomputation study of integrable properties for the (2+1)-dimensional Gardner equation with the two-singular manifold method. IMA Journal of Applied Mathematics, 74(1):46–61, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Johansson:2009:AMC

[1027] B. Tomas Johansson and Vladimir A. Kozlov. An alternating method for Cauchy problems for Helmholtztype operators in non-homogeneous medium. IMA Journal of Applied Mathematics, 74(1):62–73, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chen:2009:RVM

[1028] Wan Chen and Brian Wetton. The residual velocity method applied to a steady free boundary-value problem of vector Laplacian type. IMA Journal of Applied Mathematics, 74 (1):74–84, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Qian:2009:NLR

[1029] Q. Qian and L. Wang. Non-linear responses of a one-sided constrained beam with base excitation. IMA Journal of Applied Mathematics, 74 (1):85–96, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Duruk:2009:HOB

[1030] N. Duruk, A. Erkip, and H. A. Erbay. A higher-order Boussinesq equation in locally non-linear theory of onedimensional non-local elasticity. IMA Journal of Applied Mathematics, 74 (1):97–106, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Gunther:2009:TWS

[1031] Matthias Günther and Georg Prokert.

On travelling-wave solutions for a moving boundary problem of Hele–Shaw type. IMA Journal of Applied Mathematics, 74(1):107–127, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Utyuzhnikov:2009:GCR

[1032] Sergei V. Utyuzhnikov. Generalized Calderón–Ryaben’kii’s potentials. IMA Journal of Applied Mathematics, 74 (1):128–148, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

vanRensburg:2009:SRM

[1033] N. F. J. van Rensburg, L. Zietsman, and A. J. van der Merwe. Solvability of a Reissner–Mindlin–Timoshenko plate–beam vibration model. IMA Journal of Applied Mathematics, 74 (1):149–162, February 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Chen:2009:NSB

[1034] Ke Chen, Jin Cheng, and Paul J. Harris. A new study of the Burton and Miller method for the solution of a 3D Helmholtz problem. IMA Journal of Applied Mathematics, 74(2):163– 177, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/163;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/163.

Li:2009:SDC

[1035] Chun-Hsien Li and Suh-Yuh Yang. Synchronization in delayed Cohen– Grossberg neural networks with bounded external inputs. IMA Journal of Applied Mathematics, 74(2):178– 200, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/178;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/178.

Liu:2009:NSS

[1036] Q. Liu, F. Liu, I. Turner, and V. Anh. Numerical simulation for the 3D seepage flow with fractional derivatives in porous media. IMA Journal of Applied Mathematics, 74(2):201– 229, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/201;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/201.

Pasinlioglu:2009:TSS

[1037] Senay Pasinlioglu, Can F. Delale, and Günter H. Schnerr. On the temporal stability of steady-state quasi-1D bubbly cavitating nozzle flow solutions. IMA Journal of Applied Mathematics, 74(2):230–249, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/230;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/230.

Ivanov:2009:TSW

[1038] Ts. P. Ivanov and R. Savova. Thermoviscoelastic surface waves of an assigned frequency. IMA Journal of Applied Mathematics, 74(2):250– 263, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/250;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/250.

Liu:2009:RPS

[1039] Hongyu Liu. On recovering polyhedral scatterers with acoustic farfield measurements. IMA Journal of Applied Mathematics, 74(2):264– 272, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/264;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/264.

Bennetts:2009:WSA

[1040] Luke G. Bennetts, Nicholas R. T.

Biggs, and David Porter. Wave scattering by an axisymmetric ice floe of varying thickness. IMA Journal of Applied Mathematics, 74(2):273– 295, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/2/273;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/273.

Zou:2009:ISP

[1041] Henghui Zou. Inhomogeneous spatial patterns for predator–prey models: bifurcation at a higher eigenvalue. IMA Journal of Applied Mathematics, 74(2):296–323, April 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/74/2/296;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/2/296.

Hintermuller:2009:PNB

[1042] M. Hintermüller, V. A. Kovtunenko, and K. Kunisch. A Papkovich–Neuberbased numerical approach to cracks with contact in 3D. IMA Journal of Applied Mathematics, 74(3):325– 343, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/3/325;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/3/325.

Messaoudi:2009:EDT

[1043] Salim A. Messaoudi and Belkacem Said-Houari. Energy decay in a transmission problem in thermoelasticity of type III. IMA Journal of Applied Mathematics, 74(3):344– 360, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/

abstract/74/3/344;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/3/344.

Halanay:2009:SES

[1044] A. Halanay and I. Ursu. Stability of equilibria of some switched non-linear systems with applications to control synthesis for electrohydraulic servomechanisms. IMA Journal of Applied Mathematics, 74(3):361– 373, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/3/361;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/3/361.

Zhang:2009:GSQ

[1045] Qiong Zhang. Global solution for a quasi-linear plate system with boundary memory damping. IMA Journal of Applied Mathematics, 74(3):374– 391, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/3/374;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/3/374.

Wang:2009:NEG

[1046] Yanjin Wang. Non-existence of global solutions of a class of coupled non-linear Klein–Gordon equations with non-negative potentials and arbitrary initial energy. IMA Journal of Applied Mathematics, 74(3):392– 415, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/3/392;   http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        74/3/392.

Trimarco:2009:ENE

[1047] Carmine Trimarco. An extended notion of enthalpy. Electromagnetic solids. IMA Journal of Applied Mathematics, 74(3):416–426, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/cgi/content/
                                        abstract/74/3/416;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/3/416.

Dassios:2009:KTP

[1048] George Dassios. The Kelvin transformation in potential theory and Stokes flow. IMA Journal of Applied Mathematics, 74(3):427–438, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/3/427;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/3/427.

McCue:2009:MNM

[1049] Scott W. McCue, Bisheng Wu, and James M. Hill. Micro/nanoparticle melting with spherical symmetry and surface tension. IMA Journal of Applied Mathematics, 74(3):439–457, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/3/439;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/3/439.

Moutsinga:2009:SSP

[1050] Octave Moutsinga. Stability of sticky particle dynamics and related scalar conservation laws. IMA Journal of Applied Mathematics, 74(3):458– 467, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/3/458;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/3/458.

Cao:2009:ABG

[1051] Daomin Cao, Shuangjie Peng, and Shusen Yan. Asymptotic behaviour of ground state solutions for the Hénon equation. IMA Journal of Applied Mathematics, 74(3):468–480, June 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/3/468;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/3/468.

Jaoua:2009:RNA

[1052] Mohamed Jaoua, Juliette Leblond, Moncef Mahjoub, and Jonathan R. Partington. Robust numerical algorithms based on analytic approximation for the solution of inverse problems in annular domains. IMA Journal of Applied Mathematics, 74(4):481– 506, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/4/481;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/4/481.

Duncan:2009:HOE

[1053] Andrew J. Duncan, Tom G. Mackay, and Akhlesh Lakhtakia. On the homogenization of orthotropic elastic composites by the strong-property-fluctuation theory. IMA Journal of Applied Mathematics, 74(4):507–532, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/4/507;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/4/507.

Kucerovsky:2009:ASR

[1054] Dan Kucerovsky and Amir T. Payandeh Najafabadi. An approximation for a subclass of the Riemann– Hilbert problems. IMA Journal of Applied Mathematics, 74(4):533–547, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/4/533;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/4/533.

Thomson:2009:IEM

[1055] G. R. Thomson and C. Constanda. Integral equation methods for the Robin problem in stationary oscillations of elastic plates. IMA Journal of Applied Mathematics, 74(4):548– 558, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/4/548;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/4/548.

Liu:2009:PEI

[1056] Zhijun Liu, Jianhua Wu, and Ronghua Tan. Permanence and extinction of an impulsive delay competitive Lotka–Volterra model with periodic coefficients. IMA Journal of Applied Mathematics, 74(4):559–573, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/4/559;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/4/559.

Yuan:2009:BSA

[1057] Sanling Yuan and Yongli Song. Bifurcation and stability analysis for a delayed Leslie–Gower predator– prey system. IMA Journal of Applied Mathematics, 74(4):574–603, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/4/574;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/4/574.

Li:2009:TWS

[1058] Kun Li and Xiong Li. Travelling wave solutions in diffusive and competition–cooperation systems with delays. IMA Journal of Applied Mathematics, 74(4):604–621, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/4/604;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/4/604.

Merkin:2009:TWO

[1059] J. H. Merkin. Travelling waves in the Oregonator model for the BZ reaction. IMA Journal of Applied Mathematics, 74(4):622–643, August 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/4/622;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/4/622.

Zhuang:2009:SCI

[1060] P. Zhuang, F. Liu, V. Anh, and I. Turner. Stability and convergence of an implicit numerical method for the non-linear fractional reaction– subdiffusion process. IMA Journal of Applied Mathematics, 74(5):645–667, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/5/645;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/5/645.

Baganis:2009:ASE

[1061] G. Baganis and M. Hadjinicolaou. Analytic solution of an exterior Dirichlet problem in a non-convex domain. IMA Journal of Applied Mathematics, 74(5):668–684, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/74/5/668;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        74/5/668.

Everitt:2009:QSB

[1062] W. N. Everitt, B. T. Johansson, L. L.

Littlejohn, and C. Markett. Quasiseparation of the biharmonic partial differential equation. IMA Journal of Applied Mathematics, 74(5):685–709, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/5/685;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/5/685.

Eden:2009:FDC

[1063] A. Eden, T. B. Gürel, and E. Kuz.

Focusing and defocusing cases of the purely elliptic generalized Davey– Stewartson system. IMA Journal of Applied Mathematics, 74(5):710–725, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/5/710;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/5/710.

Wang:2009:CPC

[1064] Shubin Wang and Meiling Li. The Cauchy problem for coupled IMBq equations. IMA Journal of Applied Mathematics, 74(5):726–740, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/5/726;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/5/726.

Decent:2009:ASS

[1065] S. P. Decent. Asymptotic solution of slender viscous jet breakup. IMA Journal of Applied Mathematics, 74(5):741–781, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/5/741;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/5/741.

Cox:2009:VCS

[1066] Stephen M. Cox. Voltage and current spectra for a single-phase voltage source inverter. IMA Journal of Applied Mathematics, 74(5):782–805, October 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/5/782;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/5/782.

Shearer:2009:PMM

[1067] Michael Shearer, Pierre Gremaud, and Kristoph Kleiner. Periodic motion of a mass–spring system. IMA Journal of Applied Mathematics, 74(6):807–826, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/807;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/807.

Lin:2009:CSL

[1068] Wen-Wei Lin, Chen-Chang Peng, and Yi-Qian Wang. Chaotic synchronization in lattices of twovariable maps coupled with one variable. IMA Journal of Applied Mathematics, 74(6):827–850, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/827;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/827.

Xu:2009:ADD

[1069] Yang Xu and Jing jun Zhao. Analysis of delay-dependent stability of linear θ-methods for linear delay-integro-differential equations. IMA Journal of Applied Mathematics, 74(6):851–869, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/851;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/851.

Leach:2009:ETW

[1070] J. A. Leach and D. J. Needham.

The evolution of travelling wavefronts in a hyperbolic Fisher model. III. The initial-value problem when the initial data has exponential decay rates. IMA Journal of Applied Mathematics, 74(6):870–903, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/870;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/870.

Bouboulis:2009:COM

[1071] P. Bouboulis. Construction of orthogonal multi-wavelets using generalizedaffine fractal interpolation functions. IMA Journal of Applied Mathematics, 74(6):904–933, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/904;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/904.

DAmbrosio:2009:SEA

[1072] P. D’Ambrosio, D. De Tommasi, L. Granieri, and F. Maddalena. A surface energy approach to the mass reduction problem for elastic bodies. IMA Journal of Applied Mathematics, 74(6):934–949, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/934;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/934.

Bravo:2009:TPT

[1073] Juan C. Vila Bravo and Jaime E. Muñoz Rivera. The transmission problem to thermoelastic plate of hyperbolic type. IMA Journal of Applied Mathematics, 74(6):950–962, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/74/6/950;   http://imamat.
oxfordjournals.org/cgi/reprint/
74/6/950.

Hassen:2010:SSM

[1074] M. F. Ben Hassen, K. Erhard, and R. Potthast. The singular sources method for 3D inverse acoustic obstacle scattering problems. IMA Journal of Applied Mathematics, 75 (1):1–16, February 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/75/1/1;   http:
//imamat.oxfordjournals.org/cgi/
reprint/75/1/1.

Ji:2010:SSR

[1075] Lina Ji, Changzheng Qu, and Yaojun Ye. Solutions and symmetry reductions of the n-dimensional non-linear convection–diffusion equations. IMA Journal of Applied Mathematics, 75(1):17–55, February 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/75/1/17;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/1/17.

Layton:2010:EHD

[1076] William Layton, Leo Rebholz, and Myron Sussman. Energy and helicity dissipation rates of the NSalpha and NS-alpha-deconvolution models. IMA Journal of Applied Mathematics, 75(1):56–74, February 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/75/1/56;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/1/56.

Podder:2010:QDV

[1077] Chandra N. Podder and Abba B.

Gumel. Qualitative dynamics of a vaccination model for HSV-2. IMA Journal of Applied Mathematics, 75 (1):75–107, February 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/75/1/75;   http:
//imamat.oxfordjournals.org/cgi/
reprint/75/1/75. See corrigendum

[1133].

Bennetts:2010:LWF

[1078] Luke G. Bennetts and Vernon A. Squire. Linear wave forcing of an array of axisymmetric ice floes. IMA Journal of Applied Mathematics, 75(1):108–138, February 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/75/1/108;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/1/108.

ElJarroudi:2010:ABT

[1079] Mustapha El Jarroudi. Asymptotic behaviour of thin linearly elastic layers of oscillating thickness. IMA Journal of Applied Mathematics, 75(1):139–163, February 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/1/139;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/1/139.

Ma:2010:FER

[1080] Lili Ma, Chengyu Cao, Naira Hovakimyan, Craig Woolsey, and Warren E. Dixon. Fast estimation for range identification in the presence of unknown motion parameters. IMA Journal of Applied Mathematics, 75(2):165– 189, April 2010. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/2/165;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/2/165.

Mastroberardino:2010:TDE

[1081] Antonio Mastroberardino and Brian J.

Spencer. Three-dimensional equilibrium crystal shapes with corner energy regularization. IMA Journal of Applied Mathematics, 75(2):190– 205, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/2/190;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/2/190.

Shi:2010:LTB

[1082] Peihu Shi and Mingxin Wang. Longtime behaviour of solutions of a nonlinear diffusion problem with nonlocal source term. IMA Journal of Applied Mathematics, 75(2):206– 221, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/2/206;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/2/206.

Han:2010:CEQ

[1083] Xiaoying Han and Brian Spencer.

Composition evolution of quantum dots during the growth of solid films. IMA Journal of Applied Mathematics, 75(2):222–239, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/2/222;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/2/222.

Cimatti:2010:SPP

[1084] Giovanni Cimatti. A singular perturbation problem in magnetohydrodynamics. IMA Journal of Applied Mathematics, 75(2):240–245, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/2/240;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/2/240.

Tan:2010:CPA

[1085] Lihui Tan, Lihua Yang, and Daren Huang. Construction of periodic analytic signals satisfying the circular Bedrosian identity. IMA Journal of Applied Mathematics, 75(2):246– 256, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/2/246;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/2/246.

Fu:2010:SWF

[1086] Y. B. Fu and A. T. Il’ichev. Solitary waves in fluid-filled elastic tubes: existence, persistence, and the role of axial displacement. IMA Journal of Applied Mathematics, 75(2):257– 268, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/2/257;   http://imamat.

oxfordjournals.org/cgi/reprint/
75/2/257.

Cuminato:2010:SID

[1087] J. A. Cuminato, A. D. Fitt, M. J. S.

Mphaka, and A. Nagamine. A singular integro-differential equation model for dryout in LMFBR boiler tubes. IMA Journal of Applied Mathematics, 75(2):269–290, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/cgi/content/
abstract/75/2/269;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/2/269.

Hao:2010:RPE

[1088] Dinh Nho Hào, Nguyen Van Duc, and D. Lesnic. Regularization of parabolic equations backward in time by a non-local boundary value problem method. IMA Journal of Applied Mathematics, 75(2):291–315, April 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/2/291;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/2/291.

Wu:2010:EUG

[1089] Fuke Wu and Yangzi Hu. Existence and uniqueness of global positive solutions to the stochastic functional Kolmogorov-type system. IMA Journal of Applied Mathematics, 75(3):317– 332, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/3/317;   http://imamat.

                                        oxfordjournals.org/cgi/reprint/
                                        75/3/317.

Nunez:2010:TDM

[1090] Manuel Núñez. Two-dimensional MHD equilibria in cold gravitating plasmas. IMA Journal of Applied Mathematics, 75(3):333–342, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/3/333;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/3/333.

Blake:2010:NTS

[1091] J. R. Blake, E. O. Tuck, and P. W. Wakeley. A note on the S-transform and slender body theory in Stokes flow. IMA Journal of Applied Mathematics, 75(3):343– 355, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/3/343;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/3/343.

Liu:2010:ASTa

[1092] Zhuangyi Liu and Ramón Quintanilla. Analyticity of solutions in type III thermoelastic plates. IMA Journal of Applied Mathematics, 75(3):356– 365, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/3/356;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/

75/3/356. See republication in [1107].

Lechleiter:2010:VMW

[1093] Armin Lechleiter and Sebastian Ritterbusch. A variational method for wave scattering from penetrable rough layers. IMA Journal of Applied Mathematics, 75(3):366– 391, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/3/366;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/3/366.

Gan:2010:TWH

[1094] Qintao Gan, Rui Xu, Pinghua Yang, and Zheng Wu. Travelling waves of a hepatitis B virus infection model with spatial diffusion and time delay. IMA Journal of Applied Mathematics, 75(3):392– 417, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/3/392;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/3/392.

Pozrikidis:2010:CTD

[1095] C. Pozrikidis. Computation of threedimensional hydrostatic menisci. IMA Journal of Applied Mathematics, 75 (3):418–438, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
content/abstract/75/3/418;   http:
//imamat.oxfordjournals.org/cgi/
reprint/75/3/418.

Guo:2010:SAN

[1096] Lina Guo, Houbao Xu, Chao Gao, and Guangtian Zhu. Stability analysis of a new kind series system. IMA Journal of Applied Mathematics, 75(3):439– 460, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/3/439;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/3/439.

Coman:2010:LEB

[1097] Ciprian D. Coman. Localized elastic buckling: non-linearities versus inhomogeneities. IMA Journal of Applied Mathematics, 75(3):461–474, June 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/3/461;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/3/461.

Destrade:2010:ISI

[1098] Michel Destrade and Giuseppe Saccomandi. Introduction to special issue on stability under finite deformation. IMA Journal of Applied Mathematics, 75(4):475–478, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http:

//imamat.oxfordjournals.org/cgi/
reprint/75/4/475.

Biscari:2010:SGK

[1099] Paolo Biscari and Claudio Omati. Stability of generalized Knowles solids. IMA Journal of Applied Mathematics, 75(4):479–491, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/4/479;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/4/479.

Pucci:2010:PRN

[1100] Edvige Pucci and Giuseppe Saccomandi. Parametric resonance in non-linear viscoelasticity: solids of differential type. IMA Journal of Applied Mathematics, 75(4):492–505, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/4/492;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/4/492.

Wang:2010:ABA

[1101] Fan-Fan Wang and Hui-Hui Dai.

Asymptotic bifurcation analysis and post-buckling for uniaxial compression of a thin incompressible hyperelastic rectangle. IMA Journal of Applied Mathematics, 75(4):506–524, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/4/506;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/4/506.

Roccabianca:2010:PSB

[1102] S. Roccabianca, M. Gei, and D. Bigoni.

Plane strain bifurcations of elastic layered structures subject to finite bending: theory versus experiments. IMA Journal of Applied Mathematics, 75(4):525–548, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/4/525;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/4/525.

Goriely:2010:MSG

[1103] Alain Goriely and Rebecca Vandiver. On the mechanical stability of growing arteries. IMA Journal of Applied Mathematics, 75(4):549–570, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/4/549;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/4/549.

Dervaux:2010:LGL

[1104] Julien Dervaux and Martine Ben Amar. Localized growth of layered tissues. IMA Journal of Applied Mathematics, 75(4):571–580, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/4/571;   http://imamat.
                                        oxfordjournals.org/cgi/reprint/
                                        75/4/571.

Pearce:2010:CSL

[1105] S. P. Pearce and Y. B. Fu. Characterization and stability of localized bulging/necking in inflated membrane tubes. IMA Journal of Applied Mathematics, 75(4):581–602, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/cgi/content/
                                        abstract/75/4/581;   http://imamat.

oxfordjournals.org/cgi/reprint/
75/4/581.

Dorfmann:2010:EWF

[1106] A. Dorfmann and R. W. Ogden. Electroelastic waves in a finitely deformed electroactive material. IMA Journal of Applied Mathematics, 75(4):603– 636, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/4/603;   http://imamat.
oxfordjournals.org/cgi/reprint/
75/4/603.

Liu:2010:ASTb

[1107] Zhuangyi Liu and Ramón Quintanilla. Analyticity of solutions in type III thermoelastic plates [republished article]. IMA Journal of Applied Mathematics, 75(4):637–646, August 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/cgi/content/
abstract/75/4/637;   http://imamat.
oxfordjournals.org/cgi/reprint/

75/4/637. See original publication in [1092].

Kobayashi:2010:GUS

[1108] Kenta Kobayashi. On the global uniqueness of Stokes’ wave of extreme form. IMA Journal of Applied Mathematics, 75(5):647–675, October 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/75/5/
647.full.pdf+html.

Lyalinov:2010:SPE

[1109] M. A. Lyalinov, N. Y. Zhu, and V. P. Smyshlyaev. Scattering of a plane electromagnetic wave by a hollow circular cone with thin semitransparent walls. IMA Journal of Applied Mathematics, 75(5):676–719, October 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/75/5/
                                        676.full.pdf+html.

Sten:2010:IDS

[1110] Johan C.-E. Sten and George Dassios. Image distribution and surface potential of a dipole in a one-shell conducting sphere. IMA Journal of Applied Mathematics, 75(5):720–731, October 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/75/5/
                                        720.full.pdf+html.

Leach:2010:LTD

[1111] J. A. Leach. The large-time development of the solution to an initialvalue problem for the Korteweg– de Vries–Burgers equation: I. Initial data has a discontinuous compressive step. IMA Journal of Applied Mathematics, 75(5):732–776, October 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/75/5/
                                        732.full.pdf+html.

Liu:2010:CPE

[1112] Zhijun Liu, Jianhua Wu, and Robert A. Cheke. Coexistence and partial extinction in a delay competitive system subject to impulsive harvesting and stocking. IMA Journal of Applied Mathematics, 75(5):777–795, October 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/75/5/
777.full.pdf+html.

Walker:2010:TME

[1113] Alan J. Walker and Anthony J. Mulholland. A theoretical model of an electrostatic ultrasonic transducer incorporating resonating conduits. IMA Journal of Applied Mathematics, 75(5):796– 810, October 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/75/5/
796.full.pdf+html.

Semenko:2010:FTC

[1114] Evgeny V. Semenko. Flows in torus classification: schemes of twodimensional flows past a pair of aerofoils. IMA Journal of Applied Mathematics, 75(6):811–832, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/75/6/
811.full.pdf+html.

Amendola:2010:FER

[1115] Giovambattista Amendola, Mauro Fabrizio, and John Murrough Golden. Free energies for a rigid heat conductor with memory. IMA Journal of Applied Mathematics, 75(6):833–856, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/75/6/
                                        833.full.pdf+html.

Thompson:2010:STD

[1116] Alice B. Thompson, John Billingham, and Richard H. Tew. Surfacetension-driven flow in a half-plane. IMA Journal of Applied Mathematics, 75(6):857–880, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/content/75/6/
                                        857.full.pdf+html.

Fatori:2010:RDW

[1117] Luci Harue Fatori and Jaime E. Muñoz Rivera. Rates of decay to weak thermoelastic Bresse system. IMA Journal of Applied Mathematics, 75(6): 881–904, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        imamat.oxfordjournals.org/content/
                                        75/6/881.full.pdf+html.

Begg:2010:MAD

[1118] Ronald Begg, David J. N. Wall, and Graeme C. Wake. On a multicompartment age-distribution model of cell growth. IMA Journal of Applied Mathematics, 75(6):905–931, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/75/6/
                                        905.full.pdf+html.

Pozrikidis:2010:ESV

[1119] C. Pozrikidis and Vikram Juneja. Effect of surface viscosity on the capillary instability of an annular layer or viscous thread. IMA Journal of Applied Mathematics, 75(6):932–950, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/75/6/
932.full.pdf+html.

Pavliotis:2010:AAG

[1120] G. A. Pavliotis. Asymptotic analysis of the Green–Kubo formula. IMA Journal of Applied Mathematics, 75(6):951–967, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/75/6/
951.full.pdf+html.

Biondini:2010:IBV

[1121] Gino Biondini and Danhua Wang.

Initial-boundary-value problems for discrete linear evolution equations. IMA Journal of Applied Mathematics, 75(6):968–997, December 2010. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/75/6/
968.full.pdf+html.

Budd:2011:ISI

[1122] Chris Budd, Alan Champneys, Gabriel Lord, and Lawrie Virgin. Introduction to special issue on reflections in nonlinear mechanics. IMA Journal of Applied Mathematics, 76(1):1, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/1/
1.full.pdf+html.

Hunt:2011:RSS

[1123] Giles W. Hunt. Reflections and symmetries in space and time. IMA Journal of Applied Mathematics, 76(1):2– 26, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        2.full.pdf+html.

Thompson:2011:CTN

[1124] J. M. T. Thompson and Jan Sieber. Climate tipping as a noisy bifurcation: a predictive technique. IMA Journal of Applied Mathematics, 76(1):27–46, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        27.full.pdf+html.

Varkonyi:2011:GMC

[1125] P. L. Várkonyi and G. Domokos. A general model for collision-based abrasion processes. IMA Journal of Applied Mathematics, 76(1):47–56, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        47.full.pdf+html.

Guest:2011:STS

[1126] S. D. Guest. The stiffness of tensegrity structures. IMA Journal of Applied Mathematics, 76(1):57–66, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        57.full.pdf+html.

Pring:2011:DSP

[1127] S. R. Pring and C. J. Budd. The dynamics of a simplified pinball machine. IMA Journal of Applied Mathematics, 76(1):67–84, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/76/1/
67.full.pdf+html.

Nordmark:2011:FIR

[1128] Arne Nordmark, Harry Dankowicz, and Alan Champneys. Friction-induced reverse chatter in rigid-body mechanisms with impacts. IMA Journal of Applied Mathematics, 76(1):85–119, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/1/
85.full.pdf+html.

Wadee:2011:SAM

[1129] M. Ahmer Wadee and Christina Völlmecke. Semi-analytical modelling of buckling driven delamination in uniaxially compressed damaged plates. IMA Journal of Applied Mathematics, 76(1):120–145, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/1/
120.full.pdf+html.

Yiatros:2011:IBS

[1130] Stylianos Yiatros and M. Ahmer Wadee. Interactive buckling in sandwich beam-columns. IMA Journal of Applied Mathematics, 76(1):146–168, February 2011. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        146.full.pdf+html.

Edmunds:2011:UCS

[1131] R. Edmunds, B. J. Hicks, and G. Mullineux. Using constraint satisfaction as a means for modelling parallel folding evolution. IMA Journal of Applied Mathematics, 76(1):169–192, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        169.full.pdf+html.

Kruzik:2011:MSM

[1132] Martin Kružík and Johannes Zimmer. A model of shape memory alloys taking into account plasticity. IMA Journal of Applied Mathematics, 76(1):193–216, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        193.full.pdf+html.

Podder:2011:CQD

[1133] Chandra N. Podder and Abba B. Gumel. Corrigendum to ‘Qualitative dynamics of a vaccination model for HSV-2, IMA J. Applied Mathematics, 75(1)(2010): 75–107’. IMA Journal of Applied Mathematics, 76(1):217–218, February 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/1/
                                        217.full.pdf+html. See [1133].

Vondenhoff:2011:LTB

[1134] E. Vondenhoff. Large time behaviour of Hele–Shaw flow with injection or suction for perturbations of balls in R^{N}. IMA Journal of Applied Mathematics, 76(2):219–241, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
219.full.pdf+html.

Shermenev:2011:TWI

[1135] Alexander Shermenev. Two wave interaction in a cylinder. IMA Journal of Applied Mathematics, 76(2):242– 250, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
242.full.pdf+html.

Junca:2011:MSC

[1136] Stéphane Junca and Bernard Rousselet. The method of strained coordinates for vibrations with weak unilateral springs. IMA Journal of Applied Mathematics, 76(2):251– 276, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
251.full.pdf+html.

Yuan:2011:BAM

[1137] Sanling Yuan, Weiguo Zhang, and Yu Zhao. Bifurcation analysis of a model of plasmid-bearing, plasmidfree competition in a pulsed chemostat with an internal inhibitor. IMA Journal of Applied Mathematics, 76(2):277– 297, April 2011. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
277.full.pdf+html.

Fan:2011:LIR

[1138] Jishan Fan, Song Jiang, and Gen Nakamura. On logarithmically improved regularity criteria for the Navier– Stokes equations in R^{n}. IMA Journal of Applied Mathematics, 76(2):298– 311, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
298.full.pdf+html.

Berezansky:2011:MMM

[1139] L. Berezansky, L. Idels, and M. Kipnis. Mathematical model of marine protected areas. IMA Journal of Applied Mathematics, 76(2):312– 325, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
312.full.pdf+html.

Kimiaeifar:2011:ECF

[1140] A. Kimiaeifar, O. T. Thomsen, and E. Lund. Evaluation of closedform solutions based on the Hpm and Ham approaches for the steady flow of third grade fluids in a porous half space. IMA Journal of Applied Mathematics, 76(2):326– 339, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
326.full.pdf+html.

Park:2011:EDS

[1141] Jin-Han Park and Jum-Ran Kang.

Energy decay of solutions for Timoshenko beam with a weak nonlinear dissipation. IMA Journal of Applied Mathematics, 76(2):340–350, April 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/2/
340.full.pdf+html.

Champneys:2011:SIV

[1142] Alan Champneys and Fordyce Davidson. Special issue: 2020 visions of applied mathematics. IMA Journal of Applied Mathematics, 76(3):351– 352, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/3/
351.full.pdf+html.

Jones:2011:WCC

[1143] Christopher K. R. T. Jones. Will climate change mathematics (?). IMA Journal of Applied Mathematics, 76 (3):353–370, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

imamat.oxfordjournals.org/content/
76/3/353.full.pdf+html.

Chaplain:2011:MMM

[1144] Mark A. J. Chaplain. Multiscale mathematical modelling in biology and medicine. IMA Journal of Applied Mathematics, 76(3):371–388, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/3/
371.full.pdf+html.

Kevrekidis:2011:NLW

[1145] P. G. Kevrekidis. Non-linear waves in lattices: past, present, future. IMA Journal of Applied Mathematics, 76(3):389–423, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/content/76/3/
                                        389.full.pdf+html.

Hunt:2011:CPT

[1146] G. R. Hunt and T. S. van den Bremer. Classical plume theory: 1937– 2010 and beyond. IMA Journal of Applied Mathematics, 76(3):424– 448, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/3/
                                        424.full.pdf+html.

Higham:2011:SOD

[1147] Desmond J. Higham. Stochastic ordinary differential equations in applied and computational mathematics. IMA Journal of Applied Mathematics, 76(3):449–474, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/content/76/3/
                                        449.full.pdf+html.

Grindrod:2011:MMD

[1148] Peter Grindrod, CBE. Mathematical modelling for the digital society. IMA Journal of Applied Mathematics, 76(3):475–492, June 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.

oxfordjournals.org/content/76/3/
475.full.pdf+html.

Yu:2011:TWS

[1149] Zhi-Xian Yu and Rong Yuan. Travelling wave solutions in non-local convolution diffusive competitive– cooperative systems. IMA Journal of Applied Mathematics, 76(4):493– 513, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/4/
493.full.pdf+html.

Wang:2011:ASA

[1150] Q. X. Wang. Analysis of smallaspect-ratio lifting surfaces in extreme curved-ground effect. IMA Journal of Applied Mathematics, 76(4):514– 523, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/4/
514.full.pdf+html.

Daniels:2011:PFS

[1151] P. G. Daniels and E. Jhugroo. Pattern formation in square planform containers. IMA Journal of Applied Mathematics, 76(4):524–553, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/4/
524.full.pdf+html.

Lenells:2011:IGS

[1152] Jonatan Lenells. An integrable generalization of the sine-Gordon equation on the half-line. IMA Journal of Applied Mathematics, 76(4):554– 572, August 2011. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/4/
                                        554.full.pdf+html.

Stamova:2011:LRM

[1153] Ivanka M. Stamova. Lyapunov– Razumikhin method for impulsive differential equations with ‘supremum’. IMA Journal of Applied Mathematics, 76(4):573–581, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/content/76/4/
                                        573.full.pdf+html.

Borluk:2011:SSW

[1154] Handan Borluk and Saadet Erbay. Stability of solitary waves for threecoupled long wave–short wave interaction equations. IMA Journal of Applied Mathematics, 76(4):582–598, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/4/
                                        582.full.pdf+html.

Gao:2011:DCQ

[1155] Yang Gao. Decay conditions for 1D quasicrystal beams. IMA Journal of Applied Mathematics, 76(4):599–609, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/4/
                                        599.full.pdf+html.

Ji:2011:CLB

[1156] Lina Ji and Changzheng Qu. Conditional Lie–Bäcklund symmetries and invariant subspaces to non-linear diffusion equations. IMA Journal of Applied Mathematics, 76(4):610–632, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/4/
610.full.pdf+html.

Ballarini:2011:ESD

[1157] Roberto Ballarini and Piero Villaggio. Elastic stress diffusion around a thin corrugated inclusion. IMA Journal of Applied Mathematics, 76(4):633– 641, August 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/4/
633.full.pdf+html.

Collins:2011:EIM

[1158] Michael W. Collins and Duncan A.

Lockerby. Editorial introduction: modelling the mesoscale. IMA Journal of Applied Mathematics, 76(5):643–649, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/5/
643.full.pdf+html.

Shan:2011:LBM

[1159] Xiaowen Shan. Lattice Boltzmann in micro- and nano-flow simulations. IMA Journal of Applied Mathematics, 76(5):650–660, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/5/
650.full.pdf+html.

Kalweit:2011:MSS

[1160] Marco Kalweit and Dimitris Drikakis. Multiscale simulation strategies and mesoscale modelling of gas and liquid flows. IMA Journal of Applied Mathematics, 76(5):661–671, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/5/
                                        661.full.pdf+html.

Struchtrup:2011:MTM

[1161] Henning Struchtrup and Peyman Taheri. Macroscopic transport models for rarefied gas flows: a brief review. IMA Journal of Applied Mathematics, 76(5):672–697, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/5/
                                        672.full.pdf+html.

Xu:2011:IUG

[1162] Kun Xu and Juan-Chen Huang. An improved unified gas-kinetic scheme and the study of shock structures. IMA Journal of Applied Mathematics, 76(5):698–711, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/5/
                                        698.full.pdf+html.

Falcucci:2011:MLB

[1163] G. Falcucci, S. Ubertini, D. Chiappini, and S. Succi. Modern lattice Boltzmann methods for multiphase microflows. IMA Journal of Applied Mathematics, 76(5):712–725, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/5/
712.full.pdf+html.

Hollis:2011:DWB

[1164] A. P. Hollis, T. J. Spencer, I. Halliday, and C. M. Care. Dynamic wetting boundary condition for continuum hydrodynamics with multicomponent lattice Boltzmann equation simulation method. IMA Journal of Applied Mathematics, 76(5):726–742, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/5/
726.full.pdf+html.

Zu:2011:LBM

[1165] Y. Q. Zu and Y. Y. Yan. Lattice Boltzmann method for modelling droplets on chemically heterogeneous and microstructured surfaces with large liquid-gas density ratio. IMA Journal of Applied Mathematics, 76(5):743– 760, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/5/
743.full.pdf+html.

Kunert:2011:LBS

[1166] Christian Kunert and Jens Harting. Lattice Boltzmann simulations of liquid film drainage between smooth surfaces. IMA Journal of Applied Mathematics, 76(5):761–773, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/5/
                                        761.full.pdf+html.

Hung:2011:CLB

[1167] Li-Hsin Hung and Jaw-Yen Yang. A coupled lattice Boltzmann model for thermal flows. IMA Journal of Applied Mathematics, 76(5):774–789, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/5/
                                        774.full.pdf+html.

Phillips:2011:LBM

[1168] T. N. Phillips and G. W. Roberts. Lattice Boltzmann models for non-Newtonian flows. IMA Journal of Applied Mathematics, 76(5):790–816, October 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/5/
                                        790.full.pdf+html.

Adler:2011:SLA

[1169] J. Adler. The spherical Liouville and associated differential equations. IMA Journal of Applied Mathematics, 76(6):817–833, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/76/6/
                                        817.full.pdf+html.

Hagen:2011:NFF

[1170] Thomas Hagen and Michael Renardy. Non-failure of filaments and global existence for the equations of fiber spinning. IMA Journal of Applied Mathematics, 76(6):834–846, December 2011. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
834.full.pdf+html.

Huang:2011:IIT

[1171] Shou-Jun Huang. Impact-induced tensile waves in a kind of phase transforming materials. IMA Journal of Applied Mathematics, 76(6):847–858, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
847.full.pdf+html.

Pozrikidis:2011:SII

[1172] C. Pozrikidis and A. I. Hill. Surfactantinduced instability of a sheared liquid layer. IMA Journal of Applied Mathematics, 76(6):859–875, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
859.full.pdf+html.

Tang:2011:DPA

[1173] Yilei Tang, Deqing Huang, and Weinian Zhang. Direct parametric analysis of an enzyme-catalyzed reaction model. IMA Journal of Applied Mathematics, 76(6):876–898, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
876.full.pdf+html.

He:2011:BCB

[1174] Zhimin He and Xiaowei Jiang. Bifurcation and chaotic behaviour of a discrete-time variable-territory predator-prey model. IMA Journal of Applied Mathematics, 76(6):899–918, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
899.full.pdf+html.

Xu:2011:GDV

[1175] Rui Xu and Zhien Ma. Global dynamics of a vector disease model with saturation incidence and time delay. IMA Journal of Applied Mathematics, 76(6):919–937, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
919.full.pdf+html.

Amar:2011:TRS

[1176] Chokri Ben Amar and Christophe Hazard. Time reversal and scattering theory for time-dependent acoustic waves in a homogeneous medium. IMA Journal of Applied Mathematics, 76(6):938–955, December 2011. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/76/6/
938.full.pdf+html.

Wilson:2012:SIF

[1177] Mark Wilson, John Brindley, Michael Fairweather, Andy McIntosh, Gary Sharpe, Martin Braithwaite, Stephen Decent, and David Youdan. Special issue: fluid problems in process engineering. IMA Journal of Applied Mathematics, 77(1):1, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/1/
1.full.pdf+html.

Schwendeman:2012:CST

[1178] Donald W. Schwendeman, Ashwani K.

Kapila, and William D. Henshaw. A comparative study of two macroscale models of condensed-phase explosives. IMA Journal of Applied Mathematics, 77(1):2–17, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/77/1/
2.full.pdf+html.

Sharples:2012:ACW

[1179] J. J. Sharples, H. S. Sidhu, A. C. McIntosh, J. Brindley, and V. V. Gubernov. Analysis of combustion waves arising in the presence of a competitive endothermic reaction. IMA Journal of Applied Mathematics, 77(1):18– 31, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/1/
18.full.pdf+html.

Kestel:2012:NSP

[1180] M. Kestel, P. Nikrityuk, O. Hennig, and C. Hasse. Numerical study of the partial oxidation of a coal particle in steam and dry air atmospheres. IMA Journal of Applied Mathematics, 77(1):32–46, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/1/
32.full.pdf+html.

Shala:2012:SNB

[1181] Mehmet Shala. Simulation of nucleate boiling flow using a multiphase mixture modelling approach. IMA Journal of Applied Mathematics, 77(1):47– 58, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/1/
                                        47.full.pdf+html.

Elmabrouk:2012:EHT

[1182] E. M. Elmabrouk and Y. Wu. Enhancing the heat transfer in a heat treatment furnace through improving the combustion process in the radiation tubes. IMA Journal of Applied Mathematics, 77(1):59–71, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/1/
                                        59.full.pdf+html.

Thompson:2012:WRA

[1183] Alice B. Thompson and Joseph K.- W. Lam. Water run-off in aircraft fuel tanks. IMA Journal of Applied Mathematics, 77(1):72–85, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/1/
                                        72.full.pdf+html.

Uddin:2012:DFR

[1184] Jamal Uddin and Stephen P. Decent. Drop formation in rotating non-Newtonian jets with surfactants. IMA Journal of Applied Mathematics, 77(1):86–96, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/77/1/
86.full.pdf+html.

Mohsin:2012:BDF

[1185] Muhammad Mohsin, Jamal Uddin, Stephen P. Decent, and Mark J. H. Simmons. Break-up and droplet formation in shear thinning compound liquid jets. IMA Journal of Applied Mathematics, 77(1):97–108, February 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/1/
97.full.pdf+html.

Sharomi:2012:MSH

[1186] O. Sharomi and A. B. Gumel. Mathematical study of in-host dynamics of Chlamydia trachomatis. IMA Journal of Applied Mathematics, 77(2):109– 139, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/2/
109.full.pdf+html.

Colombo:2012:MDG

[1187] Rinaldo M. Colombo, Graziano Guerra, and Francesca Monti. Modelling the dynamics of granular matter. IMA Journal of Applied Mathematics, 77(2):140–156, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/77/2/
140.full.pdf+html.

Brighi:2012:SSM

[1188] Bernard Brighi and Je-Chiang Tsai. Similarity solutions for a model arising from high frequency excitation of liquid metal in an antisymmetric magnetic field. IMA Journal of Applied Mathematics, 77(2):157– 195, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/2/
                                        157.full.pdf+html.

Hlod:2012:TFR

[1189] A. Hlod, A. C. T. Aarts, A. A. F. van de Ven, and M. A. Peletier. Three flow regimes of viscous jet falling onto a moving surface. IMA Journal of Applied Mathematics, 77(2):196– 219, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/2/
                                        196.full.pdf+html.

Slastikov:2012:RMF

[1190] Valeriy V. Slastikov and Charles Sonnenberg. Reduced models for ferromagnetic nanowires. IMA Journal of Applied Mathematics, 77(2):220– 235, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/2/
                                        220.full.pdf+html.

Hon:2012:BBP

[1191] Y. C. Hon and Engui Fan. Binary Bell polynomial approach to the non-isospectral and variable-coefficient KP equations. IMA Journal of Applied Mathematics, 77(2):236–251, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/2/
236.full.pdf+html.

Lynch:2012:SPA

[1192] Sean Lynch and Charles Knessl. Singular perturbation analysis of driftdiffusion past a circle: shadow region. IMA Journal of Applied Mathematics, 77(2):252–278, April 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/77/2/
252.full.pdf+html.

Hone:2012:ICN

[1193] Andrew Hone, Beatrice Pelloni, and Demetrios Papageorgiou. IMA Conference on Nonlinearity and Coherent Structures. IMA Journal of Applied Mathematics, 77(3):279–281, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
279.full.pdf+html.

Caillol:2012:MVI

[1194] Philippe Caillol. Multiple vortices induced by a tridimensional critical layer in a rapidly rotating vortex. IMA Journal of Applied Mathematics, 77(3):282– 292, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
282.full.pdf+html.

Constantin:2012:PTE

[1195] Adrian Constantin. Particle trajectories in extreme Stokes waves. IMA Journal of Applied Mathematics, 77 (3):293–307, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        imamat.oxfordjournals.org/content/
                                        77/3/293.full.pdf+html.

Demontis:2012:CFS

[1196] Francesco Demontis and Cornelis van der Mee. Closed form solutions to the matrix sine–Gordon equation. IMA Journal of Applied Mathematics, 77(3):308–315, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/3/
                                        308.full.pdf+html.

Feng:2012:TWS

[1197] Zhaosheng Feng, Jing Tian, Shenzhou Zheng, and Hanfang Lu. Travelling wave solutions of the Burgers– Huxley equation. IMA Journal of Applied Mathematics, 77(3):316–325, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/3/
                                        316.full.pdf+html.

Grimshaw:2012:ERI

[1198] Roger Grimshaw and Karl Helfrich. The effect of rotation on internal solitary waves. IMA Journal of Applied Mathematics, 77(3):326–339, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
326.full.pdf+html.

Kalisch:2012:SSS

[1199] Henrik Kalisch and Darko Mitrovic.

Singular solutions for the shallowwater equations. IMA Journal of Applied Mathematics, 77(3):340–350, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
340.full.pdf+html.

Kalogirou:2012:SDN

[1200] A. Kalogirou, D. T. Papageorgiou, and Y.-S. Smyrlis. Surfactant destabilization and non-linear phenomena in two-fluid shear flows at small Reynolds numbers. IMA Journal of Applied Mathematics, 77(3):351– 360, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
351.full.pdf+html.

Khusnutdinova:2012:WNL

[1201] K. R. Khusnutdinova and K. R.

Moore. Weakly non-linear extension of d’Alembert’s formula. IMA Journal of Applied Mathematics, 77(3):361– 381, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
361.full.pdf+html.

Kundu:2012:NHD

[1202] Anjan Kundu. Non-holonomic deformation of the DNLS equation for controlling optical soliton in doped fibre media. IMA Journal of Applied Mathematics, 77(3):382– 390, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/3/
                                        382.full.pdf+html.

Manneville:2012:CST

[1203] Paul Manneville. Coherent structures in transitional plane Couette flow. IMA Journal of Applied Mathematics, 77(3):391–398, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/3/
                                        391.full.pdf+html.

Parau:2012:GSW

[1204] Emilian I. Parau and Hugh C. Woolfenden. Gap solitary waves in two-layer fluids. IMA Journal of Applied Mathematics, 77(3):399– 407, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/3/
                                        399.full.pdf+html.

Pradas:2012:BIS

[1205] Marc Pradas, Serafim Kalliadasis, and Dmitri Tseluiko. Binary interactions of solitary pulses in falling liquid films. IMA Journal of Applied Mathematics, 77(3):408–419, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/3/
                                        408.full.pdf+html.

Rottmann-Matthes:2012:SFW

[1206] Jens Rottmann-Matthes. Stability and freezing of waves in non-linear hyperbolic-parabolic systems. IMA Journal of Applied Mathematics, 77 (3):420–429, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

imamat.oxfordjournals.org/content/
77/3/420.full.pdf+html.

Wray:2012:NLW

[1207] A. W. Wray, O. Matar, and D. T.

Papageorgiou. Non-linear waves in electrified viscous film flow down a vertical cylinder. IMA Journal of Applied Mathematics, 77(3):430– 440, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/3/
430.full.pdf+html.

Li:2012:NPT

[1208] Shuguang Li. On the nature of periodic traction boundary conditions in micromechanical Fe analyses of unit cells. IMA Journal of Applied Mathematics, 77(4):441– 450, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
441.full.pdf+html.

Needham:2012:IDJ

[1209] D. J. Needham. The initial development of a jet caused by fluid, body and free surface interaction. Part 4. The large-time structure. IMA Journal of Applied Mathematics, 77(4):451– 472, June 2012. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
451.full.pdf+html.

Kong:2012:LTB

[1210] De-Xing Kong. Long-time behaviour of smooth solutions to the compressible Euler equations with damping in several space variables. IMA Journal of Applied Mathematics, 77(4):473– 494, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
473.full.pdf+html.

Jester:2012:WMR

[1211] Philipp Jester, Christoph Menke, and Karsten Urban. Wavelet methods for the representation, analysis and simulation of optical surfaces. IMA Journal of Applied Mathematics, 77(4):495– 515, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
495.full.pdf+html.

Liang:2012:TWS

[1212] Dong Liang, Peixuan Weng, and Jianhong Wu. Travelling wave solutions in a delayed predator-prey diffusion PDE system point-to-periodic and point-to-point waves. IMA Journal of Applied Mathematics, 77(4):516– 545, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
516.full.pdf+html.

Bognar:2012:SSB

[1213] Gabriella Bognár. On similarity solutions of boundary layer problems with upstream moving wall in non-Newtonian power-law fluids. IMA Journal of Applied Mathematics, 77(4):546–562, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/77/4/
546.full.pdf+html.

Liu:2012:BPM

[1214] Jingjing Liu and Zhaoyang Yin. On the blow-up phenomena for a modified periodic two-component Camassa– Holm equation. IMA Journal of Applied Mathematics, 77(4):563–577, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
563.full.pdf+html.

Bachok:2012:IPF

[1215] Norfifah Bachok, Anuar Ishak, Roslinda Nazar, and Ioan Pop. Ingham problem for free convection near a continuously moving vertical permeable plate. IMA Journal of Applied Mathematics, 77(4):578–589, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/4/
578.full.pdf+html.

Rawlins:2012:MFP

[1216] Anthony D. Rawlins. The method of finite-product extraction and an application to Wiener–Hopf theory. IMA Journal of Applied Mathematics, 77(4):590–602, June 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.
                                        oxfordjournals.org/content/77/4/
                                        590.full.pdf+html.

Champneys:2012:E

[1217] Alan R. Champneys. Editorial. IMA Journal of Applied Mathematics, 77 (5):603–604, October 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        imamat.oxfordjournals.org/content/
                                        77/5/603.full.pdf+html.

Assier:2012:PDD

[1218] R. C. Assier and N. Peake. Precise description of the different far fields encountered in the problem of diffraction of acoustic waves by a quarter-plane. IMA Journal of Applied Mathematics, 77(5):605–625, October 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/5/
                                        605.full.pdf+html.

Stewart:2012:CEC

[1219] A. L. Stewart and P. J. Dellar. Crossequatorial channel flow with zero potential vorticity under the complete Coriolis force. IMA Journal of Applied Mathematics, 77(5):626–651, October 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/5/
                                        626.full.pdf+html.

Lind:2012:EVD

[1220] S. J. Lind and T. N. Phillips. The effect of viscoelasticity on the dynamics of two gas bubbles near a rigid boundary. IMA Journal of Applied Mathematics, 77(5):652–677, October 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/5/
652.full.pdf+html.

Thompson:2012:ICP

[1221] Alice B. Thompson and John Billingham. Inviscid coalescence in the presence of a surrounding fluid. IMA Journal of Applied Mathematics, 77(5):678– 696, October 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/5/
678.full.pdf+html.

Chernyavsky:2012:CMS

[1222] Igor L. Chernyavsky, Ian L. Dryden, and Oliver E. Jensen. Characterizing the multiscale structure of fluctuations of transported quantities in a disordered medium. IMA Journal of Applied Mathematics, 77(5):697–725, October 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/5/
697.full.pdf+html.

VanDuijn:2012:E

[1223] C. J. Van Duijn, M. A. Peletier, I. S. Pop, S. M. Hassanizadeh, A. Mikelić, and B. Schweizer. Editorial. IMA Journal of Applied Mathematics, 77(6):727–728, December 2012. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/6/
                                        727.full.pdf+html.

Capasso:2012:RSM

[1224] Vincenzo Capasso, Daniela Morale, and Giuseppe Facchetti. The role of stochasticity in a model of retinal angiogenesis. IMA Journal of Applied Mathematics, 77(6):729–747, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/6/
                                        729.full.pdf+html.

Canic:2012:CSP

[1225] Suncica Canić and Josip Tambaca. Cardiovascular stents as PDE nets: 1D vs. 3D. IMA Journal of Applied Mathematics, 77(6):748–770, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/6/
                                        748.full.pdf+html.

Elenius:2012:ECT

[1226] Maria T. Elenius, Jan M. Nordbotten, and Henrik Kalisch. Effects of a capillary transition zone on the stability of a diffusive boundary layer. IMA Journal of Applied Mathematics, 77(6):771–787, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/77/6/
                                        771.full.pdf+html.

Allaire:2012:HRF

[1227] G. Allaire and H. Hutridurga. Homogenization of reactive flows in porous media and competition between bulk and surface diffusion. IMA Journal of Applied Mathematics, 77(6):788–815, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/6/
788.full.pdf+html.

Braides:2012:IEQ

[1228] Andrea Braides, Andrea Causin, and Margherita Solci. Interfacial energies on quasicrystals. IMA Journal of Applied Mathematics, 77(6):816–836, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/6/
816.full.pdf+html.

Cioranescu:2012:PUF

[1229] D. Cioranescu, G. Griso, and A. Damlamian.

The periodic unfolding for a Fredholm alternative in perforated domains. IMA Journal of Applied Mathematics, 77(6):837–854, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/6/
837.full.pdf+html.

Marciniak-Czochra:2012:STS

[1230] Anna Marciniak-Czochra. Strong twoscale convergence and corrector result for a receptor-based model of intercellular communication. IMA Journal of Applied Mathematics, 77(6):855–868, December 2012. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/6/
855.full.pdf+html.

Visintin:2012:HVE

[1231] Augusto Visintin. On the homogenization of visco-elastic processes. IMA Journal of Applied Mathematics, 77(6):869–886, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/6/
869.full.pdf+html.

Baber:2012:NSC

[1232] Katherina Baber, Klaus Mosthaf, Bernd Flemisch, Rainer Helmig, Steffen Müthing, and Barbara Wohlmuth. Numerical scheme for coupling twophase compositional porous-media flow and one-phase compositional free flow. IMA Journal of Applied Mathematics, 77(6):887–909, December 2012. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/77/6/
887.full.pdf+html.

Shao:2013:ABG

[1233] Zhi-Qiang Shao. Asymptotic behaviour of global classical solutions to the mixed initial-boundary value problem for diagonalizable quasilinear hyperbolic systems. IMA Journal of Applied Mathematics, 78(1):1– 31, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/1/
1.full.pdf+html.

Pozrikidis:2013:RPC

[1234] C. Pozrikidis and J. M. Davis.

Resistance and pressure coefficients for a periodic array of spherical, spheroidal, and cylindrical particles inside a circular tube. IMA Journal of Applied Mathematics, 78(1):32–58, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/1/
32.full.pdf+html.

Sharma:2013:WPD

[1235] M. D. Sharma. Wave propagation in a dissipative poroelastic medium. IMA Journal of Applied Mathematics, 78(1):59–69, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/78/1/
59.full.pdf+html.

Tian:2013:APR

[1236] Yanling Tian and Peixuan Weng.

Asymptotic patterns of a reactiondiffusion equation with nonlinearnonlocal functional response. IMA Journal of Applied Mathematics, 78 (1):70–101, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

imamat.oxfordjournals.org/content/
78/1/70.full.pdf+html.

Peng:2013:RPB

[1237] Yue-Jun Peng and Jérémy Ruiz.

Riemann problem for the Born– Infeld system without differential constraints. IMA Journal of Applied Mathematics, 78(1):102–131, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/1/
                                        102.full.pdf+html.

Thomson:2013:TPH

[1238] Gavin R. Thomson and Christian Constanda. The transmission problem for harmonic oscillations of thin plates. IMA Journal of Applied Mathematics, 78(1):132–145, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/1/
                                        132.full.pdf+html.

Mei:2013:EPS

[1239] Linfeng Mei and Xiaoyan Zhang. Existence of positive stationary solutions for a diffusive variable-territory predator-prey model. IMA Journal of Applied Mathematics, 78(1):146–158, February 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/1/
                                        146.full.pdf+html.

Cox:2013:SOP

[1240] S. M. Cox, C. K. Lam, and M. T. Tan. A second-order PWM-in/PWM-out class-D audio amplifier. IMA Journal of Applied Mathematics, 78(2):159– 180, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/2/
                                        159.full.pdf+html.

Mondal:2013:EFW

[1241] R. Mondal, S. K. Mohanty, and T. Sahoo. Expansion formulae for wave structure interaction problems in three dimensions. IMA Journal of Applied Mathematics, 78(2):181– 205, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/2/
181.full.pdf+html.

Forbes:2013:UPP

[1242] Lawrence K. Forbes and Graeme C.

Hocking. Unsteady plumes in planar flow of viscous and inviscid fluids. IMA Journal of Applied Mathematics, 78(2):206–234, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/78/2/
206.full.pdf+html.

Liu:2013:SNT

[1243] F. Liu, K. Burrage, and N. A. Hamilton. Some novel techniques of parameter estimation for dynamical models in biological systems. IMA Journal of Applied Mathematics, 78(2):235– 260, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/2/
235.full.pdf+html.

Berg:2013:EEO

[1244] Peter Berg, Paul Hanz, and Ian Milton. An energy-economic oil production model. IMA Journal of Applied Mathematics, 78(2):261–286, April 2013. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/2/
                                        261.full.pdf+html.

Li:2013:HBT

[1245] Xin Li, Weihua Jiang, and Junping Shi. Hopf bifurcation and Turing instability in the reactiondiffusion Holling–Tanner predatorprey model. IMA Journal of Applied Mathematics, 78(2):287–306, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/2/
                                        287.full.pdf+html.

Wang:2013:ASB

[1246] Q. X. Wang. The aerodynamics of a slender body moving very close to a water wave surface. IMA Journal of Applied Mathematics, 78(2):307– 337, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/2/
                                        307.full.pdf+html.

Budd:2013:SSB

[1247] C. J. Budd and V. A. Galaktionov. On self-similar blow-up in evolution equations of Monge– Ampère type. IMA Journal of Applied Mathematics, 78(2):338–378, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/2/
                                        338.full.pdf+html.

Kim:2013:GPJ

[1248] Sun-Chul Kim and Hisashi Okamoto.

The generalized Proudman–Johnson equation at large Reynolds numbers. IMA Journal of Applied Mathematics, 78(2):379–403, April 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/78/2/
379.full.pdf+html.

Arens:2013:AEM

[1249] Tilo Arens, Kai Sandfort, Susanne Schmitt, and Armin Lechleiter. Analysing Ewald’s method for the evaluation of Green’s functions for periodic media. IMA Journal of Applied Mathematics, 78(3):405– 431, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
405.full.pdf+html.

Dauhoo:2013:DID

[1250] M. Z. Dauhoo, B. S. N. Korimboccus, and S. B. Issack. On the dynamics of illicit drug consumption in a given population. IMA Journal of Applied Mathematics, 78(3):432– 448, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
432.full.pdf+html.

Wang:2013:FPP

[1251] Dongshu Wang. Four positive periodic solutions of a delayed plankton allelopathy system on time scales with multiple exploited (or harvesting) terms. IMA Journal of Applied Mathematics, 78(3):449–473, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
449.full.pdf+html.

Choudhury:2013:DHM

[1252] Bejoy K. Choudhury. Diffusion of heat in multidimensional composite spherical body. IMA Journal of Applied Mathematics, 78(3):474– 493, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
474.full.pdf+html.

Wang:2013:STF

[1253] Lina Wang, Hailei Shao, and Yaping Wu. Stability of travelling front solutions for a forest dynamical system with cross-diffusion. IMA Journal of Applied Mathematics, 78(3):494– 512, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
494.full.pdf+html.

Merkin:2013:WTA

[1254] J. H. Merkin. On wave trains arising in the two-variable Oregonator model for the BZ reaction. IMA Journal of Applied Mathematics, 78(3):513– 536, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
513.full.pdf+html.

ONaraigh:2013:MMM

[1255] Lennon Ó Náraigh and Khalid Kamhawi. Multiscale methods and modelling for chemical reactions on oscillating surfaces. IMA Journal of Applied Mathematics, 78(3):537– 565, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
537.full.pdf+html.

Peng:2013:QAD

[1256] Rui Peng. Qualitative analysis on a diffusive and ratio-dependent predator-prey model. IMA Journal of Applied Mathematics, 78(3):566– 586, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
566.full.pdf+html.

Fidler:2013:DVF

[1257] D. A. Fidler and A. A. Lacey. Dynamics of vertical fire spread. IMA Journal of Applied Mathematics, 78(3):587– 613, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
587.full.pdf+html.

Kavallaris:2013:NLR

[1258] Nikos I. Kavallaris and Takashi Suzuki.

Non-local reaction-diffusion system involved by reaction radius I. IMA Journal of Applied Mathematics, 78(3):614– 632, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/3/
614.full.pdf+html.

Li:2013:TWS

[1259] Kun Li and Xiong Li. Travelling wave solutions in integro-difference competition system. IMA Journal of Applied Mathematics, 78(3):633– 650, June 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/3/
                                        633.full.pdf+html.

Milewski:2013:E

[1260] Paul Milewski and Demetrios Papageorgiou. Editorial. IMA Journal of Applied Mathematics, 78(4):651, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/4/
                                        651.full.pdf+html.

Agrawal:2013:MMH

[1261] S. Agrawal and G. A. Kriegsmann. A model for the microwave heating of a thin ceramic slab in a multimode cavity. IMA Journal of Applied Mathematics, 78(4):652–664, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/4/
                                        652.full.pdf+html.

Blyth:2013:DEC

[1262] M. G. Blyth and E. I. Parau. Deformation of an elastic cell in a uniform stream and in a circulatory flow. IMA Journal of Applied Mathematics, 78(4):665–684, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
665.full.pdf+html.

Binder:2013:FSF

[1263] Benjamin J. Binder, M. G. Blyth, and Scott W. McCue. Free-surface flow past arbitrary topography and an inverse approach for wave-free solutions. IMA Journal of Applied Mathematics, 78(4):685–696, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
685.full.pdf+html.

Lustri:2013:EAF

[1264] Christopher J. Lustri, Scott W. Mc-Cue, and S. Jonathan Chapman. Exponential asymptotics of free surface flow due to a line source. IMA Journal of Applied Mathematics, 78(4):697– 713, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
697.full.pdf+html.

Malakuti:2013:DCS

[1265] Kamyar Malakuti, Russel E. Caflisch, Michael Siegel, and Alex Virodov. Detection of complex singularities for a function of several variables. IMA Journal of Applied Mathematics, 78(4):714–728, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/78/4/
714.full.pdf+html.

Tziannaros:2013:TDE

[1266] Marios Tziannaros, Stephen E. Glavin, and Frank T. Smith. Threedimensional effects in the lower urinary tract. IMA Journal of Applied Mathematics, 78(4):729–749, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
729.full.pdf+html.

Wang:2013:TDF

[1267] Z. Wang, J-M. Vanden-Broeck, and P. A. Milewski. Two-dimensional flexural-gravity waves of finite amplitude in deep water. IMA Journal of Applied Mathematics, 78(4):750– 761, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
750.full.pdf+html.

Tseluiko:2013:HCA

[1268] Dmitri Tseluiko, Josh Baxter, and Uwe Thiele. A homotopy continuation approach for analysing finite-time singularities in thin liquid films. IMA Journal of Applied Mathematics, 78(4):762– 776, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
762.full.pdf+html.

Hunter:2013:QLS

[1269] John K. Hunter and Mihaela Ifrim. A quasi-linear Schrödinger equation for large amplitude inertial oscillations in a rotating shallow fluid. IMA Journal of Applied Mathematics, 78(4):777– 796, August 2013. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
777.full.pdf+html.

McConnell:2013:VED

[1270] Lane C. McConnell, Michael J. Miksis, and Petia M. Vlahovska. Vesicle electrohydrodynamics in DC electric fields. IMA Journal of Applied Mathematics, 78(4):797–817, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
797.full.pdf+html.

Moore:2013:ACI

[1271] M. R. Moore, J. R. Ockendon, and J. M. Oliver. Air-cushioning in impact problems. IMA Journal of Applied Mathematics, 78(4):818–838, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
818.full.pdf+html.

Keller:2013:WWP

[1272] Joseph B. Keller. Water wave production by oscillating bodies. IMA Journal of Applied Mathematics, 78(4):839– 850, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
839.full.pdf+html.

Booty:2013:LWE

[1273] M. R. Booty, D. T. Papageorgiou, M. Siegel, and Q. Wang. Longwave equations and direct simulations for the breakup of a viscous fluid thread surrounded by an immiscible viscous fluid. IMA Journal of Applied Mathematics, 78(4):851–867, August 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/4/
851.full.pdf+html.

Ebrahimi:2013:NSV

[1274] M. A. Ebrahimi and Evelyn Lunasin. The Navier–Stokes–Voight model for image inpainting. IMA Journal of Applied Mathematics, 78(5):869–894, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
869.full.pdf+html.

Bellis:2013:NTE

[1275] Cédric Bellis, Fioralba Cakoni, and Bojan B. Guzina. Nature of the transmission eigenvalue spectrum for elastic bodies. IMA Journal of Applied Mathematics, 78(5):895–923, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
895.full.pdf+html.

Chen:2013:NMV

[1276] Chang-Ming Chen, F. Liu, K. Burrage, and Y. Chen. Numerical methods of the variable-order Rayleigh– Stokes problem for a heated generalized second grade fluid with fractional derivative. IMA Journal of Applied Mathematics, 78(5):924–944, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
924.full.pdf+html.

Simon:2013:ESM

[1277] Péter L. Simon and Istvan Z. Kiss.

From exact stochastic to mean-field ODE models: a new approach to prove convergence results. IMA Journal of Applied Mathematics, 78(5):945–964, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
945.full.pdf+html.

Destrade:2013:SDE

[1278] M. Destrade and R. W. Ogden. On stress-dependent elastic moduli and wave speeds. IMA Journal of Applied Mathematics, 78(5):965–997, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
965.full.pdf+html.

Wang:2013:ECK

[1279] Xu Wang and Peter Schiavone. Extension of certain key results from classical elasticity to decagonal quasicrystalline composites. IMA Journal of Applied Mathematics, 78(5):998–1014, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
998.full.pdf+html.

Chen:2013:GES

[1280] Guowang Chen and Xianjun Han.

Global existence of solution of Cauchy problem for a non-linear wave equation. IMA Journal of Applied Mathematics, 78(5):1015–1031, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
1015.full.pdf+html.

Claeyssen:2013:EEB

[1281] Julio R. Claeyssen, Teresa Tsukazan, and Rosemaira D. Copetti. Eigenanalysis of an Euler–Bernoulli model coupled with van der Waals forces for carbon nanotubes. IMA Journal of Applied Mathematics, 78(5):1032–1050, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
1032.full.pdf+html.

Kuehn:2013:DAE

[1282] Christian Kuehn, Stefan Siegmund, and Thilo Gross. Dynamical analysis of evolution equations in generalized models. IMA Journal of Applied Mathematics, 78(5):1051–1077, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
1051.full.pdf+html.

Erbas:2013:RWF

[1283] Baris Erbas, Julius Kaplunov, and Danila A. Prikazchikov. The Rayleigh wave field in mixed problems for a half-plane. IMA Journal of Applied Mathematics, 78(5):1078–1086, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
1078.full.pdf+html.

Moutari:2013:MRT

[1284] S. Moutari, M. Herty, A. Klein, M. Oeser, B. Steinauer, and V. Schleper. Modelling road traffic accidents using macroscopic second-order models of traffic flow. IMA Journal of Applied Mathematics, 78(5):1087–1108, October 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/5/
1087.full.pdf+html.

Ivanshin:2013:SIS

[1285] P. N. Ivanshin and E. A. Shirokova.

Spline-interpolation solution of 3D Dirichlet problem for a certain class of solids. IMA Journal of Applied Mathematics, 78(6):1109–1129, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1109.full.pdf+html.

Silva:2013:VPE

[1286] M. A. Jorge Silva and T. F. Ma. On a viscoelastic plate equation with history setting and perturbation of p-Laplacian type. IMA Journal of Applied Mathematics, 78(6):1130–1146, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1130.full.pdf+html.

Jun:2013:DBO

[1287] Doobae Jun and Hyejin Ku. Digital barrier option contract with exponential random time. IMA Journal of Applied Mathematics, 78(6):1147–1155, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1147.full.pdf+html.

Sakthivel:2013:LMI

[1288] R. Sakthivel, R. Raja, and S. Marshal Anthoni. Linear matrix inequality approach to stochastic stability of uncertain delayed BAM neural networks. IMA Journal of Applied Mathematics, 78(6):1156–1178, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1156.full.pdf+html.

Hegedus:2013:SPS

[1289] F. Hegedüs, C. Hös, and L. Kullmann. Stable period 1, 2 and 3 structures of the harmonically excited Rayleigh– Plesset equation applying low ambient pressure. IMA Journal of Applied Mathematics, 78(6):1179–1195, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1179.full.pdf+html.

Leach:2013:IVP

[1290] J. A. Leach. An initial-value problem for the modified Korteweg–de Vries equation. IMA Journal of Applied Mathematics, 78(6):1196–1213, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1196.full.pdf+html.

Kay:2013:VBD

[1291] Anthony Kay. A viscous buoyancydriven flow exhibiting finite-time blow-up. IMA Journal of Applied Mathematics, 78(6):1214–1236, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1214.full.pdf+html.

Wang:2013:MRS

[1292] W. Wang and A. J. Roberts. Macroscopic reduction for stochastic reactiondiffusion equations. IMA Journal of Applied Mathematics, 78(6):1237–1264, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1237.full.pdf+html.

Wu:2013:EIR

[1293] Jianhua Wu, Cui Ma, and Gaihui Guo.

The effect of interaction ratio in a chemical reaction. IMA Journal of Applied Mathematics, 78(6):1265–1289, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/78/6/
1265.full.pdf+html.

Wu:2013:TWS

[1294] Chufen Wu and Dongmei Xiao. Travelling wave solutions in a non-local and time-delayed reaction-diffusion model. IMA Journal of Applied Mathematics, 78(6):1290–1317, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/6/
                                        1290.full.pdf+html.

Shao:2013:GSI

[1295] Zhi-Qiang Shao. Global structure instability of Riemann solutions for general quasilinear hyperbolic systems of conservation laws under small BV perturbations of the initial data: shocks and contact discontinuities. IMA Journal of Applied Mathematics, 78(6):1318–1355, December 2013. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/78/6/
                                        1318.full.pdf+html.

Kavallaris:2014:NLR

[1296] Nikos I. Kavallaris and Takashi Suzuki. Non-local reaction-diffusion system involving reaction radius II: rate of convergence. IMA Journal of Applied Mathematics, 79(1):1–21, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/1/
                                        1.full.pdf+html.

Smith:2014:PTW

[1297] Warren R. Smith and Jan G. Wissink. Parameterization of travelling waves in plane Poiseuille flow. IMA Journal of Applied Mathematics, 79(1):22–32, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/1/
                                        22.full.pdf+html.

Dhaeyer:2014:RSD

[1298] Sam D’haeyer, B. Tomas Johansson, and Marián Slodicka. Reconstruction of a spacewise-dependent heat source in a time-dependent heat diffusion process. IMA Journal of Applied Mathematics, 79(1):33–53, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
33.full.pdf+html.

Fernandes:2014:RTH

[1299] Paolo Fernandes, Marina Ottonello, and Mirco Raffetto. Regularity of time-harmonic electromagnetic fields in the interior of bianisotropic materials and metamaterials. IMA Journal of Applied Mathematics, 79(1):54– 93, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
54.full.pdf+html.

Collins:2014:SRD

[1300] W. D. Collins. Synchronous running of a double-stator permanent magnet motor with a series capacitor. IMA Journal of Applied Mathematics, 79(1):94–108, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
94.full.pdf+html.

Najafabadi:2014:ESC

[1301] Amir T. Payandeh Najafabadi and Dan Kucerovsky. Exact solutions for a class of matrix Riemann– Hilbert problems. IMA Journal of Applied Mathematics, 79(1):109–123, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
109.full.pdf+html.

Wilson:2014:MSS

[1302] Miles Wilson, David M. Leppinen, Peter M. Haese, and John R. Blake. Modelling of solid spheres and bubbles near a tip-vortex and a vortex ring. IMA Journal of Applied Mathematics, 79(1):124–162, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
124.full.pdf+html.

DeStaelen:2014:SSC

[1303] Rob H. De Staelen and Guillaume Crevecoeur. A sensor sensitivity and correlation analysis through polynomial chaos in the EEG problem. IMA Journal of Applied Mathematics, 79(1):163–174, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
163.full.pdf+html.

Alfifi:2014:SAS

[1304] H. Y. Alfifi, T. R. Marchant, and M. I. Nelson. Semi-analytical solutions for the 1- and 2-D diffusive Nicholson’s blowflies equation. IMA Journal of Applied Mathematics, 79(1):175–199, February 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/1/
175.full.pdf+html.

Rodriguez:2014:NSA

[1305] Julio Rodriguez and Max-Olivier Hongler. Networks of self-adaptive dynamical systems. IMA Journal of Applied Mathematics, 79(2):201–240, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/2/
201.full.pdf+html.

Ma:2014:PSM

[1306] To Fu Ma, Jaime Edilberto Muñoz Rivera, Higidio Portillo Oquendo, and Fredy Maglorio Sobrado Suárez. Polynomial stabilization of magnetoelastic plates. IMA Journal of Applied Mathematics, 79(2):241–253, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/2/
241.full.pdf+html.

Wirth:2014:CCP

[1307] B. Wirth, I. Sobey, and A. Eisenträger.

Conditions for choking in a poroelastic flow model. IMA Journal of Applied Mathematics, 79(2):254– 273, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/2/
254.full.pdf+html.

Tseluiko:2014:WIS

[1308] Dmitri Tseluiko and Serafim Kalliadasis. Weak interaction of solitary pulses in active dispersive-dissipative nonlinear media. IMA Journal of Applied Mathematics, 79(2):274–299, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/2/
                                        274.full.pdf+html.

Mijatovic:2014:NDH

[1309] Aleksandar Mijatović and Mikhail Urusov. A note on delta hedging in markets with jumps. IMA Journal of Applied Mathematics, 79(2):300– 312, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/2/
                                        300.full.pdf+html.

Ruan:2014:CDA

[1310] Ning Ruan and D. Y. Gao. Canonical duality approach for non-linear dynamical systems. IMA Journal of Applied Mathematics, 79(2):313– 325, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/2/
                                        313.full.pdf+html.

Li:2014:BPG

[1311] Xiaoguang Li, Jian Zhang, Shaoyong Lai, and Yonghong Wu. On the blow-up phenomenon for a generalized Davey–Stewartson system. IMA Journal of Applied Mathematics, 79(2):326– 346, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/2/
                                        326.full.pdf+html.

Tian:2014:GSD

[1312] Xiaohong Tian and Rui Xu. Global stability of a delayed HIV-1 infection model with absorption and CTL immune response. IMA Journal of Applied Mathematics, 79(2):347– 359, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/2/
347.full.pdf+html.

Shams:2014:RTS

[1313] M. Shams and R. W. Ogden. On Rayleigh-type surface waves in an initially stressed incompressible elastic solid. IMA Journal of Applied Mathematics, 79(2):360–376, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/2/
360.full.pdf+html.

Johansson:2014:SAH

[1314] B. Tomas Johansson and Vladimir A.

Kozlov. Solvability and asymptotics of the heat equation with mixed variable lateral conditions and applications in the opening of the exocytotic fusion pore in cells. IMA Journal of Applied Mathematics, 79(2):377– 392, April 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/2/
377.full.pdf+html.

Lyalinov:2014:ESC

[1315] Mikhail A. Lyalinov. Electromagnetic scattering by a circular impedance cone: diffraction coefficients and surface waves. IMA Journal of Applied Mathematics, 79(3):393–430, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
393.full.pdf+html.

Shen:2014:NNA

[1316] S. Shen, F. Liu, V. Anh, I. Turner, and J. Chen. A novel numerical approximation for the space fractional advectiondispersion equation. IMA Journal of Applied Mathematics, 79(3):431– 444, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
431.full.pdf+html.

Siddheshwar:2014:SIM

[1317] P. G. Siddheshwar, A. Chan, and U. S. Mahabaleswar. Suction-induced magnetohydrodynamics of a viscoelastic fluid over a stretching surface within a porous medium. IMA Journal of Applied Mathematics, 79(3):445– 458, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
445.full.pdf+html.

Leppinen:2014:NCS

[1318] D. M. Leppinen and D. A. S. Rees. Natural convection in shallow porous cavities near the density maximum: the conduction and intermediate regimes. IMA Journal of Applied Mathematics, 79(3):459– 493, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
459.full.pdf+html.

Kasinathan:2014:TRB

[1319] Sakthivel Kasinathan. Twofold rotation between kinematically compatible martensites. IMA Journal of Applied Mathematics, 79(3):494– 501, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
494.full.pdf+html.

Cen:2014:ECC

[1320] Yi-Gang Cen, Rui-Zhen Zhao, Li-Hui Cen, Zhen-Jiang Miao, and Xiao-Fang Chen. Explicit construction of compactly supported biorthogonal multiwavelets via matrix extension. IMA Journal of Applied Mathematics, 79(3):502–534, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
502.full.pdf+html.

Bailey:2014:DHS

[1321] N. Y. Bailey, K. A. Cliffe, S. Hibberd, and H. Power. On the dynamics of a high-speed coned fluidlubricated bearing. IMA Journal of Applied Mathematics, 79(3):535– 561, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
535.full.pdf+html.

Balachandran:2014:CNL

[1322] K. Balachandran and J. Kokila. Controllability of non-linear implicit fractional dynamical systems. IMA Journal of Applied Mathematics, 79(3):562– 570, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
562.full.pdf+html.

Wu:2014:KSM

[1323] Xiaoqin P. Wu and Liancheng Wang. A Krawiec–Szydlowski model of business cycles with a time delay in capital stock. IMA Journal of Applied Mathematics, 79(3):571–599, June 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/3/
571.full.pdf+html.

Fu:2014:FSI

[1324] Yibin Fu. Foreword to special issue for winners of the 2nd IMA Lighthill-Thwaites Prize. IMA Journal of Applied Mathematics, 79(4):601–602, August 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/4/
601.

Kimpton:2014:EMR

[1325] Laura S. Kimpton, Jonathan P. Whiteley, Sarah L. Waters, and James M. Oliver. The effect of membrane-regulated actin polymerization on a two-phase flow model for cell motility. IMA Journal of Applied Mathematics, 79(4):603–635, August 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/4/
603.

Woolley:2014:TMM

[1326] Thomas E. Woolley, Eamonn A.

Gaffney, Sarah L. Waters, James M. Oliver, Ruth E. Baker, and Alain Goriely. Three mechanical models for blebbing and multi-blebbing. IMA Journal of Applied Mathematics, 79 (4):636–660, August 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

imamat.oxfordjournals.org/content/
79/4/636.

Moore:2014:ACA

[1327] M. R. Moore and J. M. Oliver.

On air cushioning in axisymmetric impacts. IMA Journal of Applied Mathematics, 79(4):661–680, August 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/4/
661.

Hennessy:2014:DSE

[1328] Matthew G. Hennessy and Andreas Münch. Dynamics of a slowly evaporating solvent-polymer mixture with a deformable upper surface. IMA Journal of Applied Mathematics, 79(4):681– 720, August 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/4/
681.

Bustamante:2014:FSI

[1329] Roger Bustamante, Jose Merodio, and David Steigmann. Foreword to special issue celebrating Professor Ray Ogden’s 70th birthday. IMA Journal of Applied Mathematics, 79(5):721, October 2014. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
721.

Krishnan:2014:PFI

[1330] Jyothi Krishnan and David J. Steigmann. A polyconvex formulation of isotropic elastoplasticity theory. IMA Journal of Applied Mathematics, 79(5):722–738, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
722.

Selvadurai:2014:MCB

[1331] A. P. S. Selvadurai. Mechanics of contact between bi-material elastic solids perturbed by a flexible interface. IMA Journal of Applied Mathematics, 79(5):739–752, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
739.

Michelitsch:2014:RRF

[1332] Thomas M. Michelitsch, Gérard A. Maugin, Shahram Derogar, and Mujibur Rahman. A regularized representation of the fractional Laplacian in n dimensions and its relation to Weierstrass–Mandelbrot-type fractal functions. IMA Journal of Applied Mathematics, 79(5):753–777, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
753.

Quintanilla:2014:AEG

[1333] R. Quintanilla and K. R. Rajagopal.

Analysis of the equations governing the motion of a degrading elastic solid due to diffusion of a fluid. IMA Journal of Applied Mathematics, 79(5):778– 789, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
778.

Wang:2014:FDH

[1334] Xu Wang and Peter Schiavone. Finite deformation of harmonic solids with cusp cracks. IMA Journal of Applied Mathematics, 79(5):790–803, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
790.

Ciarletta:2014:TIS

[1335] Pasquale Ciarletta and Michel Destrade. Torsion instability of soft solid cylinders. IMA Journal of Applied Mathematics, 79(5):804–819, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
804.

Horgan:2014:SUB

[1336] C. O. Horgan and J. G. Murphy. Some unexpected behaviour in shear for elasticity models of arterial tissue that only use the I_{1}, I_{4}, I_{6} invariants. IMA Journal of Applied Mathematics, 79(5):820– 829, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        820.

Rudykh:2014:AEF

[1337] Stephan Rudykh and Mary C. Boyce. Analysis of elasmoid fish imbricated layered scale-tissue systems and their bio-inspired analogues at finite strains and bending. IMA Journal of Applied Mathematics, 79(5):830–847, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        830.

Dagher:2014:PAS

[1338] M. A. Dagher and K. P. Soldatos. Pure azimuthal shear deformation of an incompressible tube reinforced by radial fibres resistant in bending. IMA Journal of Applied Mathematics, 79(5):848– 868, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        848.

Rickaby:2014:OCS

[1339] Stephen R. Rickaby and Nigel H. Scott. Orthotropic cyclic stress-softening model for pure shear during repeated loading and unloading. IMA Journal of Applied Mathematics, 79(5):869– 888, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        869.

Adler:2014:MCM

[1340] J. H. Adler, L. Dorfmann, D. Han, S. MacLachlan, and C. Paetsch. Mathematical and computational models of incompressible materials subject to shear. IMA Journal of Applied Mathematics, 79(5):889–914, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
889.

Vinh:2014:NPR

[1341] Pham Chi Vinh, Jose Merodio, Trinh Thi Thanh Hue, and Nguyen Thi Nam. Non-principal Rayleigh waves in deformed transversely isotropic incompressible non-linearly elastic solids. IMA Journal of Applied Mathematics, 79(5):915–928, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
915.

Springhetti:2014:OEH

[1342] Roberta Springhetti, Eliana Bortot, Gal deBotton, and Massimiliano Gei. Optimal energy-harvesting cycles for load-driven dielectric generators in plane strain. IMA Journal of Applied Mathematics, 79(5):929–946, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/
929.

Topol:2014:TDD

[1343] Heiko Topol, Hasan Demirkoparan, Thomas J. Pence, and Alan Wineman. A theory for deformation dependent evolution of continuous fibre distribution applicable to collagen remodelling. IMA Journal of Applied Mathematics, 79(5):947–977, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        947.

Gao:2014:DFS

[1344] Hao Gao, David Carrick, Colin Berry, Boyce E. Griffith, and Xiaoyu Luo. Dynamic finite-strain modelling of the human left ventricle in health and disease using an immersed boundaryfinite element method. IMA Journal of Applied Mathematics, 79(5):978–1010, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        978.

Roy:2014:FEA

[1345] David Roy, Gerhard A. Holzapfel, Claude Kauffmann, and Gilles Soulez. Finite element analysis of abdominal aortic aneurysms: geometrical and structural reconstruction with application of an anisotropic material model. IMA Journal of Applied Mathematics, 79(5):1011–1026, October 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/5/
                                        1011.

Tomic:2014:PMR

[1346] Aleksandar Tomic, Alfio Grillo, and Salvatore Federico. Poroelastic materials reinforced by statistically oriented fibres — numerical implementation and application to articular cartilage. IMA Journal of Applied Mathematics, 79(5):1027–1059, October 2014. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/5/

1027. See erratum [1365].

Moore:2014:DTE

[1347] David C. Moore, Milijana Odavic, and Stephen M. Cox. Dead-time effects on the voltage spectrum of a PWM inverter. IMA Journal of Applied Mathematics, 79(6):1061–1076, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/6/
1061.

Kadem:2014:BSS

[1348] A. Kadem, M. Kirane, C. M. Kirk, and W. E. Olmstead. Blowing-up solutions to systems of fractional differential and integral equations with exponential non-linearities. IMA Journal of Applied Mathematics, 79(6):1077–1088, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/6/
1077.

Roberts:2014:STP

[1349] Melanie E. Roberts and Andrew P.

Bassom. Stability of a time-periodic flow incorporating a simple model for noise. IMA Journal of Applied Mathematics, 79(6):1089–1106, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/6/

Rickaby:2014:MMS

[1350] Stephen R. Rickaby and Nigel H. Scott. Multicyclic modelling of softening in biological tissue. IMA Journal of Applied Mathematics, 79(6):1107–1125, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/6/
                                        1107.

Utyuzhnikov:2014:RTA

[1351] Sergey V. Utyuzhnikov. Real-time active wave control with preservation of wanted field. IMA Journal of Applied Mathematics, 79(6):1126–1138, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/6/
                                        1126.

Elsheikh:2014:AMM

[1352] Sara M.-A. S. Elsheikh, Kailash C. Patidar, and Rachid Ouifki. Analysis of a malaria model with a distributed delay. IMA Journal of Applied Mathematics, 79(6):1139–1160, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/79/6/
                                        1139.

Natali:2014:OSP

[1353] Fábio Natali and Aloisio Neves. Orbital stability of periodic waves. IMA Journal of Applied Mathematics, 79(6):1161–1179, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://imamat.

oxfordjournals.org/content/79/6/
1161.

Barboteu:2014:HDC

[1354] Mikael Barboteu, Flavius Patrulescu, Ahmad Ramadan, and Mircea Sofonea. History-dependent contact models for viscoplastic materials. IMA Journal of Applied Mathematics, 79(6):1180–1200, December 2014. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/79/6/
1180.

Fok:2015:RPR

[1355] Pak-Wing Fok, Qunhui Han, and Tom Chou. Reconstruction of a persistent random walk from exit time distributions. IMA Journal of Applied Mathematics, 80(1):1–23, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
1.

Hao:2015:DAT

[1356] Dinh Nho Hào, Phan Xuan Thanh, and D. Lesnic. Determination of the ambient temperature in transient heat conduction. IMA Journal of Applied Mathematics, 80(1):24–46, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
24.

Man:2015:DRW

[1357] Chi-Sing Man, Gen Nakamura, Kazumi Tanuma, and Shengzhang Wang. Dispersion of Rayleigh waves in vertically-inhomogeneous prestressed elastic media. IMA Journal of Applied Mathematics, 80(1):47–84, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
47.

Li:2015:GES

[1358] Xiaodi Li, Donal O’Regan, and Haydar Akca. Global exponential stabilization of impulsive neural networks with unbounded continuously distributed delays. IMA Journal of Applied Mathematics, 80(1):85–99, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
85.

Sun:2015:MMK

[1359] Chengjun Sun, Rudolf Leube, Reinhard Windoffer, and Stéphanie Portet. A mathematical model for the keratin cycle of assembly and disassembly. IMA Journal of Applied Mathematics, 80(1):100–114, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
100.

Gong:2015:OFE

[1360] Rongfang Gong and Xiaoliang Cheng. An optimal finite element error estimate for an inverse problem in multispectral bioluminescence tomography. IMA Journal of Applied Mathematics, 80(1):115–134, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
115.

Saumier:2015:ENA

[1361] Louis-Philippe Saumier, Martial Agueh, and Boualem Khouider. An efficient numerical algorithm for the L^{2} optimal transport problem with periodic densities. IMA Journal of Applied Mathematics, 80(1):135–157, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
135.

Gao:2015:CSC

[1362] Timothy K. Gao. Complete solutions to a class of eighth-order polynomial optimization problems. IMA Journal of Applied Mathematics, 80(1):158–176, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
158.

Schwendeman:2015:HAC

[1363] Donald W. Schwendeman, Colin P.

Please, Burt S. Tilley, and Ferdinand Hendriks. A homogenization analysis of the compressible flow between a slider and a moving patterned rough surface. IMA Journal of Applied Mathematics, 80(1):177–211, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
177.

Cezaro:2015:TCC

[1364] Adriano De Cezaro and Jorge P.

Zubelli. The tangential cone condition for the iterative calibration of local volatility surfaces. IMA Journal of Applied Mathematics, 80(1):212–232, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/
212.

Federico:2015:EPM

[1365] Salvatore Federico and Alfio Grillo. Erratum to: Poroelastic Materials Reinforced by Statistically Oriented Fibres — Numerical Implementation and Application to Articular Cartilage. IMA Journal of Applied Mathematics, 80(1):233–234, February 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/1/

233. See [1346].

Bunder:2015:GMD

[1366] J. E. Bunder and James M. Hill. General model for diffusive growth of nanostructures. IMA Journal of Applied Mathematics, 80(2):235– 254, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
235.

Salem:2015:EGM

[1367] A. M. Salem. Entropy generation for magnetohydrodynamic heat transfer over a non-isothermal stretching sheet with variable viscosity. IMA Journal of Applied Mathematics, 80(2):255– 272, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
255.

Schmuck:2015:NMR

[1368] M. Schmuck, M. Pradas, G. A.

Pavliotis, and S. Kalliadasis. A new mode reduction strategy for the generalized Kuramoto–Sivashinsky equation. IMA Journal of Applied Mathematics, 80(2):273–301, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
273.

Hsu:2015:TWSa

[1369] Cheng-Hsiung Hsu, Jian-Jhong Lin, and Tzi-Sheng Yang. Travelling wave solutions for delayed lattice reactiondiffusion systems. IMA Journal of Applied Mathematics, 80(2):302–323, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
302.

Groth:2015:HNA

[1370] Samuel P. Groth, David P. Hewett, and Stephen Langdon. Hybrid numericalasymptotic approximation for highfrequency scattering by penetrable convex polygons. IMA Journal of Applied Mathematics, 80(2):324–353, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
324.

Wang:2015:IIN

[1371] Yulan Wang and Zhaoyin Xiang. The interfaces of an inhomogeneous non-Newtonian polytropic filtration equation with convection. IMA Journal of Applied Mathematics, 80(2):354– 375, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
354.

Knox:2015:PSF

[1372] D. J. Knox, S. K. Wilson, B. R. Duffy, and S. McKee. Porous squeeze-film flow. IMA Journal of Applied Mathematics, 80(2):376–409, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
376.

Miranville:2015:GAC

[1373] Alain Miranville and Ramon Quintanilla. A generalization of the Allen–Cahn equation. IMA Journal of Applied Mathematics, 80(2):410– 430, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
410.

Rosca:2015:MCB

[1374] Natalia C. Rosca, Alin V. Rosca, John H. Merkin, and Ioan Pop. Mixed convection boundary-layer flow near the lower stagnation point of a horizontal circular cylinder with a secondorder wall velocity condition and a constant surface heat flux. IMA Journal of Applied Mathematics, 80(2):431– 451, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
431.

Ghiba:2015:SES

[1375] Ionel-Dumitrel Ghiba and Emilian Bulgariu. On spatial evolution of the solution of a non-standard problem in the bending theory of elastic plates. IMA Journal of Applied Mathematics, 80(2):452–473, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
452.

Tabata:2015:EUS

[1376] Minoru Tabata, Nobuoki Eshima, Yuichi Kiyonari, and Ichiro Takagi. The existence and uniqueness of short-run equilibrium of the Dixit–Stiglitz–Krugman model in an urban-rural setting. IMA Journal of Applied Mathematics, 80(2):474– 493, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
474.

Gulgowski:2015:TAT

[1377] Jacek Gulgowski and Jerzy Julian Michalski. Topological attitude towards path following, applied to localization of complex dispersion characteristics for a lossy microwave, ferritecoupled transmission line. IMA Journal of Applied Mathematics, 80(2):494– 507, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/

Hu:2015:STH

[1378] Guanghui Hu and Andreas Rathsfeld. Scattering of time-harmonic electromagnetic plane waves by perfectly conducting diffraction gratings. IMA Journal of Applied Mathematics, 80(2):508– 532, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/2/
                                        508.

Edwards:2015:SFF

[1379] D. A. Edwards. Sharp Fickian fronts in conjugated polymer films. IMA Journal of Applied Mathematics, 80(2):533– 553, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/2/
                                        533.

Easthope:2015:ESA

[1380] P. F. Easthope. Examination of SGP4 along-track errors for initially circular orbits. IMA Journal of Applied Mathematics, 80(2):554–568, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/2/
                                        554.

Liu:2015:OCS

[1381] Chongyang Liu. Optimal control of a switched autonomous system with time delay arising in fedbatch processes. IMA Journal of Applied Mathematics, 80(2):569–584, April 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/2/
569.

vandeRotten:2015:MEB

[1382] B. A. van de Rotten and S. M. Verduyn Lunel. A memory-efficient Broyden method to compute fixed points of non-linear maps arising in periodically forced processes. IMA Journal of Applied Mathematics, 80(3):585– 607, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
585.

Jia:2015:EIA

[1383] Yue Jia, Cosmin Anitescu, Seyed Shahram Ghorashi, and Timon Rabczuk. Extended isogeometric analysis for material interface problems. IMA Journal of Applied Mathematics, 80(3):608– 633, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
608.

Davis:2015:MAS

[1384] Anthony M. J. Davis and Darren G. Crowdy. Matched asymptotics for a spherical low-Reynolds-number treadmilling swimmer near a rigid wall. IMA Journal of Applied Mathematics, 80(3):634–650, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
634.

Yoon:2015:SSD

[1385] Ji-Hun Yoon, Jeong-Hoon Kim, and Jinnam Jo. Singularity of scattering and Dirichlet-to-Neumann operator symbols in elliptic wave propagation models. IMA Journal of Applied Mathematics, 80(3):651–675, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
651.

Hasanov:2015:RBR

[1386] Alemdar Hasanov and Balgaisha Mukanova. Relationship between representation formulas for unique regularized solutions of inverse source problems with final overdetermination and singular value decomposition of input-output operators. IMA Journal of Applied Mathematics, 80(3):676– 696, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
676.

Ratz:2015:NDI

[1387] Andreas Rätz. A new diffuseinterface model for step flow in epitaxial growth. IMA Journal of Applied Mathematics, 80(3):697–711, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
697.

Li:2015:QNL

[1388] Ye-Ping Li and Wen-An Yong. Quasineutral limit in a 3D compressible Navier–Stokes–Poisson–Korteweg model. IMA Journal of Applied Mathematics, 80(3):712–727, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/80/3/
712.

Zhang:2015:DAN

[1389] Zhi-Yong Zhang and Yu-Fu Chen.

Determination of approximate nonlinear self-adjointness and approximate conservation law. IMA Journal of Applied Mathematics, 80(3):728– 746, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
728.

Borrelli:2015:INU

[1390] A. Borrelli, G. Giantesio, and M. C.

Patria. Influence of a non-uniform external magnetic field on the oblique stagnation-point flow of a micropolar fluid. IMA Journal of Applied Mathematics, 80(3):747–765, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
747.

Foias:2015:TAH

[1391] Ciprian Foias, Michael S. Jolly, Ruomeng Lan, Rishika Rupam, Yong Yang, and Bingsheng Zhang. Time analyticity with higher norm estimates for the 2D Navier–Stokes equations. IMA Journal of Applied Mathematics, 80(3):766–810, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
766.

Sudakov:2015:MMP

[1392] Ivan Sudakov and Sergey A. Vakulenko. A mathematical model for a positive permafrost carbon-climate feedback. IMA Journal of Applied Mathematics, 80(3):811–824, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/3/
                                        811.

Ye:2015:NAT

[1393] H. Ye, F. Liu, V. Anh, and I. Turner. Numerical analysis for the time distributed-order and Riesz space fractional diffusions on bounded domains. IMA Journal of Applied Mathematics, 80(3):825–838, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/3/
                                        825.

Vafeas:2015:AID

[1394] Panayiotis Vafeas, Polycarpos K. Papadopoulos, and Pavlos M. Hatzikonstantinou. Analytical integro-differential representation of flow fields for the micropolar Stokes flow of a conducting ferrofluid. IMA Journal of Applied Mathematics, 80(3):839– 864, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/3/
                                        839.

Ledger:2015:PEF

[1395] Paul D. Ledger and William R. B. Lionheart. The perturbation of electromagnetic fields at distances that are large compared with the object’s size. IMA Journal of Applied Mathematics, 80(3):865–892, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
865.

Karjanto:2015:MSP

[1396] Natanael Karjanto. On the method of strained parameters for a KdV type of equation with exact dispersion property. IMA Journal of Applied Mathematics, 80(3):893–905, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
893.

Wu:2015:CPM

[1397] Xinglong Wu and Boling Guo. The Cauchy problem of the modified CH and DP equations. IMA Journal of Applied Mathematics, 80(3):906– 930, June 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/3/
906.

Kim:2015:COE

[1398] Hyeon Jeong Kim, Marco A. Fontelos, and Hyung Ju Hwang. Capillary oscillations at the exit of a nozzle. IMA Journal of Applied Mathematics, 80(4):931–962, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
931.

Perrusson:2015:LFS

[1399] G. Perrusson, P. Vafeas, I. K. Chatjigeorgiou, and D. Lesselier. Low-frequency on-site identification of a highly conductive body buried in earth from a model ellipsoid. IMA Journal of Applied Mathematics, 80(4):963–980, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        963.

Caflisch:2015:EOP

[1400] R. E. Caflisch, G. Gambino, M. Sammartino, and C. Sgarra. European option pricing with transaction costs and stochastic volatility: an asymptotic analysis. IMA Journal of Applied Mathematics, 80(4):981–1008, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        981.

Ho:2015:MRD

[1401] Csaba Hős, Csaba Bazsó, and Alan Champneys. Model reduction of a direct spring-loaded pressure relief valve with upstream pipe. IMA Journal of Applied Mathematics, 80(4):1009– 1024, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        1009.

Cao:2015:GSE

[1402] Jinde Cao, Yi Wang, Abdulaziz Alofi, Abdullah Al-Mazrooei, and Ahmed Elaiw. Global stability of an epidemic model with carrier state in heterogeneous networks. IMA Journal of Applied Mathematics, 80(4):1025– 1048, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1025.

Hazard:2015:ATM

[1403] Christophe Hazard. On the absence of trapped modes in locally perturbed open waveguides. IMA Journal of Applied Mathematics, 80(4):1049–1062, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1049.

Cakoni:2015:AAT

[1404] Fioralba Cakoni, Nicolas Chaulet, and Houssem Haddar. Asymptotic analysis of the transmission eigenvalue problem for a Dirichlet obstacle coated by a thin layer of nonabsorbing media. IMA Journal of Applied Mathematics, 80(4):1063–1098, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1063.

Barannyk:2015:OPD

[1405] Lyudmyla L. Barannyk and Alexander Panchenko. Optimizing performance of the deconvolution model reduction for large ODE systems. IMA Journal of Applied Mathematics, 80(4):1099– 1123, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1099.

Xu:2015:SDT

[1406] Zhiting Xu and Yiyi Zhang. Spatial dynamics of a time-delayed reaction and diffusion malaria model. IMA Journal of Applied Mathematics, 80(4):1124– 1154, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        1124.

Merchant:2015:SSW

[1407] Sandra M. Merchant and Wayne Nagata. Selection and stability of wave trains behind predator invasions in a model with non-local prey competition. IMA Journal of Applied Mathematics, 80(4):1155–1177, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        1155.

Sakurai:2015:LCR

[1408] Tomoaki Sakurai and Toshiaki Murofushi. Linear complementarity representation of piecewise linear functions. IMA Journal of Applied Mathematics, 80(4):1178–1198, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        1178.

Guo:2015:ISP

[1409] Jun Guo, Qinghua Wu, and Guozheng Yan. The inverse scattering problem by a crack buried in a piecewise homogeneous medium. IMA Journal of Applied Mathematics, 80(4):1199– 1218, August 2015. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1199.

Hu:2015:EES

[1410] Yangzi Hu and Fuke Wu. Exponential extinction of a stochastic Lotka– Volterra model with expectations in coefficients. IMA Journal of Applied Mathematics, 80(4):1219–1234, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1219.

Nersisyan:2015:GWW

[1411] Hayk Nersisyan, Denys Dutykh, and Enrique Zuazua. Generation of 2D water waves by moving bottom disturbances. IMA Journal of Applied Mathematics, 80(4):1235–1253, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1235.

Cheng:2015:TWS

[1412] Cui-Ping Cheng, Wan-Tong Li, and Guo Lin. Travelling wave solutions in periodic monostable equations on a two-dimensional spatial lattice. IMA Journal of Applied Mathematics, 80(4):1254–1272, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/4/
1254.

Wu:2015:SPL

[1413] Junliang Wu and Jianguo Zhao. A survey of the progress of locating methods of complex matrices’ eigenvalues and some new location theorem and their applications. IMA Journal of Applied Mathematics, 80(4):1273–1285, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/4/
                                        1273.

Bustamante:2015:NSN

[1414] R. Bustamante and K. R. Rajagopal. A note on some new classes of constitutive relations for elastic bodies. IMA Journal of Applied Mathematics, 80(5):1287–1299, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/5/
                                        1287.

Stinner:2015:MMP

[1415] Christian Stinner, Christina Surulescu, and Gülnihal Meral. A multiscale model for pH-tactic invasion with time-varying carrying capacities. IMA Journal of Applied Mathematics, 80(5):1300–1321, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/5/
                                        1300.

Zhou:2015:CSC

[1416] Mingru Zhou and Li Sun. On the corner singularities in the classical and Reissner–Mindlin plate theories. IMA Journal of Applied Mathematics, 80(5):1322–1335, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1322.

Hsu:2015:TWSb

[1417] Cheng-Hsiung Hsu, Jian-Jhong Lin, and Ting-Hui Yang. Travelling wave solutions for Kolmogorov-type delayed lattice reaction-diffusion systems. IMA Journal of Applied Mathematics, 80(5):1336–1367, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1336.

Pakes:2015:LMK

[1418] Anthony G. Pakes. Lambert’s W meets Kermack–McKendrick epidemics. IMA Journal of Applied Mathematics, 80(5):1368–1386, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1368.

Mertz:2015:DDP

[1419] Laurent Mertz and Alain Bensoussan. Degenerate Dirichlet problems related to the ergodic property of an elasto-plastic oscillator excited by a filtered white noise. IMA Journal of Applied Mathematics, 80(5):1387–1408, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1387.

Bailey:2015:DPH

[1420] N. Y. Bailey, K. A. Cliffe, S. Hibberd, and H. Power. Dynamics of a parallel, high-speed, lubricated thrust bearing with Navier slip boundary conditions. IMA Journal of Applied Mathematics, 80(5):1409–1430, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/5/
                                        1409.

Liu:2015:GAS

[1421] Meng Liu. Global asymptotic stability of stochastic Lotka–Volterra systems with infinite delays. IMA Journal of Applied Mathematics, 80(5):1431–1453, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/5/
                                        1431.

Taylor-King:2015:MMT

[1422] Jake P. Taylor-King, Benjamin Franz, Christian A. Yates, and Radek Erban. Mathematical modelling of turning delays in swarm robotics. IMA Journal of Applied Mathematics, 80(5):1454–1474, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/80/5/
                                        1454.

Hadjinicolaou:2015:SDS

[1423] M. Hadjinicolaou and E. Protopapas. Spectral decomposition of the Stokes flow operators in the inverted prolate spheroidal coordinates. IMA Journal of Applied Mathematics, 80(5):1475–1491, October 2015. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1475.

Roberts:2015:MSV

[1424] A. J. Roberts. Macroscale, slowly varying, models emerge from the microscale dynamics. IMA Journal of Applied Mathematics, 80(5):1492–1518, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1492.

Kerimov:2015:DIP

[1425] Nazim B. Kerimov and Mansur I. Ismailov. Direct and inverse problems for the heat equation with a dynamic-type boundary condition. IMA Journal of Applied Mathematics, 80(5):1519–1533, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1519.

Shi:2015:STS

[1426] Hong-Bo Shi and Shigui Ruan. Spatial, temporal and spatiotemporal patterns of diffusive predator-prey models with mutual interference. IMA Journal of Applied Mathematics, 80(5):1534–1568, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1534.

Kisil:2015:RBS

[1427] Anastasia V. Kisil. The relationship between a strip Wiener–Hopf problem and a line Riemann–Hilbert problem. IMA Journal of Applied Mathematics, 80(5):1569–1581, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1569.

Rickett:2015:IMM

[1428] Lydia M. Rickett, Robert Penfold, Mark G. Blyth, Richard Purvis, and Mark J. Cooker. Incipient mixing by Marangoni effects in slow viscous flow of two immiscible fluid layers. IMA Journal of Applied Mathematics, 80(5):1582–1618, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1582.

Stamov:2015:SML

[1429] Gani Stamov and Ivanka Stamova. Second method of Lyapunov and almost periodic solutions for impulsive differential systems of fractional order. IMA Journal of Applied Mathematics, 80(5):1619–1633, October 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/5/
1619.

Lindsay:2015:RMP

[1430] A. E. Lindsay, J. Lega, and K. B. Glasner. Regularized model of posttouchdown configurations in electrostatic MEMS: interface dynamics. IMA Journal of Applied Mathematics, 80(6):1635–1663, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/80/6/
1635.

Kunisch:2015:TOC

[1431] Karl Kunisch and Armin Rund.

Time optimal control of the monodomain model in cardiac electrophysiology. IMA Journal of Applied Mathematics, 80(6):1664–1683, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1664.

Algehyne:2015:FEA

[1432] Ebrahem A. Algehyne and Anthony J.

Mulholland. A finite element approach to modelling fractal ultrasonic transducers. IMA Journal of Applied Mathematics, 80(6):1684–1702, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1684.

Zhou:2015:PFG

[1433] Jun Zhou and Junping Shi. Pattern formation in a general glycolysis reactiondiffusion system. IMA Journal of Applied Mathematics, 80(6):1703–1738, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1703.

daSilva:2015:SAC

[1434] Priscila Leal da Silva and Igor Leite Freire. Symmetry analysis of a class of autonomous even-order ordinary differential equations. IMA Journal of Applied Mathematics, 80(6):1739–1758, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1739.

Bostwick:2015:LBS

[1435] J. B. Bostwick and P. H. Steen. Liquid-bridge shape stability by energy bounding. IMA Journal of Applied Mathematics, 80(6):1759–1775, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1759.

Ledger:2015:CSM

[1436] Paul D. Ledger and William R. B. Lionheart. Characterizing the shape and material properties of hidden targets from magnetic induction data. IMA Journal of Applied Mathematics, 80(6):1776–1798, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1776.

Hou:2015:SVN

[1437] Dongshuang Hou, Theo Driessen, and Hao Sun. The Shapley value and the nucleolus of service cost savings games as an application of 1convexity. IMA Journal of Applied Mathematics, 80(6):1799–1807, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1799.

Chen:2015:WPC

[1438] Hua Chen and Gongwei Liu. Wellposedness for a class of Kirchhoff equations with damping and memory terms. IMA Journal of Applied Mathematics, 80(6):1808–1836, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1808.

Thornby:2015:MPD

[1439] John A. Thornby, Robert S. MacKay, and Mark A. Williams. Mathematical principles for the design of isostatic mount systems for dynamic structures. IMA Journal of Applied Mathematics, 80(6):1837–1854, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1837.

Liu:2015:BCN

[1440] Qiao Liu. On blow-up criteria for the 3D nematic liquid crystal flows. IMA Journal of Applied Mathematics, 80(6):1855–1870, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1855.

Zhang:2015:PSC

[1441] Xinhong Zhang, Wenxue Li, and Ke Wang. Periodic solutions of coupled systems on networks with both time-delay and linear coupling. IMA Journal of Applied Mathematics, 80(6):1871–1889, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1871.

Fullard:2015:ASS

[1442] Luke A. Fullard and Graeme C. Wake. An analytical series solution to the steady laminar flow of a Newtonian fluid in a partially filled pipe, including the velocity distribution and the dip phenomenon. IMA Journal of Applied Mathematics, 80(6):1890–1901, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1890.

Crowdy:2015:TML

[1443] Darren Crowdy. A transform method for Laplace’s equation in multiply connected circular domains. IMA Journal of Applied Mathematics, 80(6):1902–1931, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/80/6/
1902.

Whittaker:2015:SRB

[1444] Robert J. Whittaker. A shearrelaxation boundary layer near the pinned ends of a buckled elasticwalled tube. IMA Journal of Applied Mathematics, 80(6):1932–1967, December 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/80/6/
1932.

Walker:2016:TMU

[1445] Alan J. Walker and Anthony J. Mulholland. A theoretical model of an ultrasonic transducer incorporating spherical resonators. IMA Journal of Applied Mathematics, 81(1):1–25, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/1/
1.

Djordjevic:2016:CBD

[1446] Jasmina Djordjević. On a class of backward doubly stochastic differential equations with continuous coefficients. IMA Journal of Applied Mathematics, 81(1):26–41, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/1/
26.

Malik:2016:SAC

[1447] Josef Malík. Spectral analysis connected with suspension bridge systems. IMA Journal of Applied Mathematics, 81(1):42–75, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/81/1/
42.

Shen:2016:RPP

[1448] Chun Shen. The Riemann problem for the pressureless Euler system with the Coulomb-like friction term. IMA Journal of Applied Mathematics, 81(1):76– 99, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/1/
                                        76.

Meneses:2016:SLP

[1449] Rodrigo Meneses and Oscar Orellana. On a Sturm–Liouville problem with spectral and physical parameters in boundary conditions. IMA Journal of Applied Mathematics, 81(1):100–131, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/1/
                                        100.

Du:2016:AIM

[1450] Zhengdong Du, Weinian Zhang, and Shengfu Deng. Approximation to invariant manifolds under pseudohyperbolicity. IMA Journal of Applied Mathematics, 81(1):132–151, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/1/
                                        132.

Grindrod:2016:PLT

[1451] Peter Grindrod and Ebrahim L. Patel. Phase locking to the ntorus. IMA Journal of Applied Mathematics, 81(1):152–164, February 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/1/
                                        152.

Li:2016:ARC

[1452] Xie Li and Zhaoyin Xiang. On an attraction-repulsion chemotaxis system with a logistic source. IMA Journal of Applied Mathematics, 81(1):165–198, February 2016. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/1/
165.

Fabrikant:2016:CPT

[1453] V. I. Fabrikant. Contact problem for a transversely isotropic half-space limited by a plane perpendicular to its planes of isotropy. IMA Journal of Applied Mathematics, 81(2):199– 227, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
199.

Cao:2016:MMC

[1454] Meng Cao and A. J. Roberts. Multiscale modelling couples patches of nonlinear wave-like simulations. IMA Journal of Applied Mathematics, 81(2):228– 254, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
228.

Zhao:2016:FBP

[1455] Yonggang Zhao and Mingxin Wang.

Free boundary problems for the diffusive competition system in higher dimension with sign-changing coefficients. IMA Journal of Applied Mathematics, 81(2):255–280, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/81/2/
255.

Carr:2016:SEM

[1456] Thomas W. Carr. An SIRS epidemic model with a rapidly decreasing probability for immunity. IMA Journal of Applied Mathematics, 81(2):281– 307, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
281.

Gibbon:2016:HLF

[1457] J. D. Gibbon. High-low frequency slaving and regularity issues in the 3D Navier–Stokes equations. IMA Journal of Applied Mathematics, 81(2):308– 320, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
308.

Wang:2016:DSE

[1458] Jinliang Wang, Ran Zhang, and Toshikazu Kuniya. The dynamics of an SVIR epidemiological model with infection age. IMA Journal of Applied Mathematics, 81(2):321– 343, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
321.

Almeida:2016:RDM

[1459] Rui M. P. Almeida, Stanislav N. Antontsev, José C. M. Duque, and Jorge Ferreira. A reaction-diffusion model for the non-local coupled system: existence, uniqueness, longtime behaviour and localization properties of solutions. IMA Journal of Applied Mathematics, 81(2):344– 364, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
344.

Rogosin:2016:CMF

[1460] Sergei Rogosin and Gennady Mishuris.

Constructive methods for factorization of matrix-functions. IMA Journal of Applied Mathematics, 81(2):365– 391, April 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/2/
365.

Champneys:2016:SHY

[1461] Alan R. Champneys, Darren Crowdy, and Demetrius Papageorgiou. Some highlights from 50 years of the IMA Journal of Applied Mathematics. IMA Journal of Applied Mathematics, 81(3):393–408, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/81/3/
393.

Lai:2016:TTS

[1462] Eric L. Lai, Daniel Moyer, Baichuan Yuan, Eric Fox, Blake Hunter, Andrea L. Bertozzi, and P. Jeffrey Brantingham. Topic time series analysis of microblogs. IMA Journal of Applied Mathematics, 81(3):409– 431, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
409.

Feng:2016:WBD

[1463] S. Feng and P. Holmes. Will big data yield new mathematics? An evolving synergy with neuroscience. IMA Journal of Applied Mathematics, 81(3):432– 456, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
432.

Knobloch:2016:LSF

[1464] E. Knobloch. Localized structures and front propagation in systems with a conservation law. IMA Journal of Applied Mathematics, 81(3):457– 487, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
457.

Pedley:2016:SSM

[1465] T. J. Pedley. Spherical squirmers: models for swimming microorganisms. IMA Journal of Applied Mathematics, 81(3):488–521, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
488.

Renardy:2016:TYS

[1466] Michael Renardy and Yuriko Renardy. Thixotropy in yield stress fluids as a limit of viscoelasticity. IMA Journal of Applied Mathematics, 81(3):522– 537, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
522.

Champneys:2016:PPC

[1467] Alan R. Champneys and Péter L. Várkonyi. The Painlevé paradox in contact mechanics. IMA Journal of Applied Mathematics, 81(3):538– 588, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
538.

Crowdy:2016:SKP

[1468] D. G. Crowdy, E. H. Kropf, C. C.

Green, and M. M. S. Nasser. The Schottky–Klein prime function: a theoretical and computational tool for applications. IMA Journal of Applied Mathematics, 81(3):589–628, June 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/3/
589.

Champneys:2016:AAN

[1469] Alan Champneys, Mike Jeffrey, Mario di Bernardo, and Yulia Kyrychko. ‘Advances in applied nonlinear mathematics’: a special issue in honour of John Hogan’s 60th birthday. IMA Journal of Applied Mathematics, 81(4):629– 630, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/4/
629.

Stuhlmeier:2016:AHW

[1470] Raphael Stuhlmeier and Michael Stiassnie. Adapting Havelock’s wavemaker theorem to acoustic-gravity waves. IMA Journal of Applied Mathematics, 81(4):631–646, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.

                                        oxfordjournals.org/content/81/4/
                                        631.

Chillingworth:2016:PHC

[1471] D. R. G. Chillingworth. Perturbed hedgehogs: continuous deformation of point defects in biaxial nematic liquid crystals. IMA Journal of Applied Mathematics, 81(4):647–661, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/4/

647. See [1517].

Jiang:2016:GIN

[1472] Haibo Jiang and Marian Wiercigroch. Geometrical insight into nonsmooth bifurcations of a soft impact oscillator. IMA Journal of Applied Mathematics, 81(4):662–678, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/4/
                                        662.

Ockendon:2016:MTC

[1473] J. R. Ockendon and J. R. Barber. A model for thermoelastic contact oscillations. IMA Journal of Applied Mathematics, 81(4):679–687, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/4/
                                        679.

Carrero:2016:ADC

[1474] Niliana Carrero, Carles Batlle, and Enric Fossas. Averaged dynamics of a coupled-inductor boost converter under sliding mode control using a piecewise linear complementarity model. IMA Journal of Applied Mathematics, 81(4):688–698, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/4/
688.

Glendinning:2016:BSF

[1475] Paul Glendinning. Bifurcation from stable fixed point to 2D attractor in the border collision normal form. IMA Journal of Applied Mathematics, 81(4):699–710, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/4/
699.

Glendinning:2016:MAG

[1476] Paul Glendinning, Piotr Kowalczyk, and Arne B. Nordmark. Multiple attractors in grazing-sliding bifurcations in Filippov-type flows. IMA Journal of Applied Mathematics, 81(4):711– 722, August 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/4/
711.

Marshall:2016:ESD

[1477] J. S. Marshall. Exact solutions describing the injection-driven growth of a doubly-periodic fluid region in a Hele–Shaw cell. IMA Journal of Applied Mathematics, 81(5):723–749, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/5/
723.

Alharbi:2016:WSA

[1478] Amnah M. Alharbi and Nigel H. Scott. Wave stability in anisotropic generalized temperature-rate-dependent thermoelasticity. IMA Journal of Applied Mathematics, 81(5):750–778, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/5/
                                        750.

Haddar:2016:AKM

[1479] Houssem Haddar, Jing Rebecca Li, and Simona Schiavi. Adapting the Kärger model to account for finite diffusion-encoding pulses in diffusion MRI. IMA Journal of Applied Mathematics, 81(5):779–794, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/5/
                                        779.

Zhao:2016:TWF

[1480] Lin Zhao and Zhi-Cheng Wang. Traveling wave fronts in a diffusive epidemic model with multiple parallel infectious stages. IMA Journal of Applied Mathematics, 81(5):795–823, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/5/
                                        795.

Gupta:2016:DWS

[1481] Neelam Gupta and V. D. Sharma. Diffraction of a weak shock by a wedge in a van der Waals gas. IMA Journal of Applied Mathematics, 81(5):824– 841, October 2016. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/5/
824.

Elliott:2016:POM

[1482] Robert J. Elliott, Leunglung Chan, and Tak Kuen Siu. Pricing options in a Markov regime switching model with a random acceleration for the volatility. IMA Journal of Applied Mathematics, 81(5):842–859, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/5/
842.

Black:2016:WCC

[1483] Tobias Black, Johannes Lankeit, and Masaaki Mizukami. On the weakly competitive case in a two-species chemotaxis model. IMA Journal of Applied Mathematics, 81(5):860–876, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/5/
860.

Kumar:2016:HPS

[1484] K. Kumar, M. Neuss-Radu, and I. S.

Pop. Homogenization of a pore scale model for precipitation and dissolution in porous media. IMA Journal of Applied Mathematics, 81(5):877– 897, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/5/
877.

Redeker:2016:UTP

[1485] Magnus Redeker, Christian Rohde, and Iuliu Sorin Pop. Upscaling of a triphase phase-field model for precipitation in porous media. IMA Journal of Applied Mathematics, 81(5):898– 939, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/5/
                                        898.

Towers:2016:VMW

[1486] P. D. Towers, Z. Hussain, P. T. Griffiths, and S. J. Garrett. Viscous modes within the compressible boundary-layer flow due to a broad rotating cone. IMA Journal of Applied Mathematics, 81(5):940–960, October 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/5/
                                        940.

Claeys:2016:ESL

[1487] Xavier Claeys. Essential spectrum of local multi-trace boundary integral operators. IMA Journal of Applied Mathematics, 81(6):961–983, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

                                        oxfordjournals.org/content/81/6/
                                        961.

Hocking:2016:MFS

[1488] G. C. Hocking and S. L. Mitchell. A model of flow in a sugar diffuser. IMA Journal of Applied Mathematics, 81(6):984–998, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
984.

Hattam:2016:SPS

[1489] L. L. Hattam and S. R. Clarke.

Steady periodic solutions of a variable coefficient KS-KdV equation. IMA Journal of Applied Mathematics, 81(6):999–1019, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/81/6/
999.

Al-Musallam:2016:EME

[1490] Fadhel Al-Musallam, Durga Prasad Challa, and Mourad Sini. The equivalent medium for the elastic scattering by many small rigid bodies and applications. IMA Journal of Applied Mathematics, 81(6):1020–1050, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1020.

Maddalena:2016:LNL

[1491] Francesco Maddalena, Danilo Percivale, and Franco Tomarelli. Local and non-local energies in adhesive interaction. IMA Journal of Applied Mathematics, 81(6):1051–1075, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1051.

Soldatos:2016:RDF

[1492] K. P. Soldatos and A. F. Farhat.

On Reissner’s displacement field in modelling thin elastic plates with embedded fibres resistant in bending. IMA Journal of Applied Mathematics, 81(6):1076–1095, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1076.

Luo:2016:GEW

[1493] Wei Luo and Zhaoyang Yin. Global existence of weak solutions for a three-component Camassa–Holm system with N-peakon solutions. IMA Journal of Applied Mathematics, 81(6):1096–1111, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://imamat.
oxfordjournals.org/content/81/6/
1096.

Zemlyanova:2016:CMI

[1494] Anna Y. Zemlyanova. Curvilinear mode-I/mode-II interface fracture with a curvature-dependent surface tension on the boundary. IMA Journal of Applied Mathematics, 81(6):1112–1136, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1112.

Gou:2016:IPA

[1495] Jia Gou, Wayne Nagata, and Yue-Xian Li. Interaction of in-phase and anti-phase synchronies in a coupled compartment-bulk diffusion model at a double Hopf bifurcation. IMA Journal of Applied Mathematics, 81(6):1137–1162, December 2016. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1137.

Navickas:2016:ESO

[1496] Z. Navickas, R. Marcinkevicius, T. Telksnys, and M. Ragulskis. Existence of second order solitary solutions to Riccati differential equations coupled with a multiplicative term. IMA Journal of Applied Mathematics, 81(6):1163–1190, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1163.

Maling:2016:AAB

[1497] B. Maling and R. V. Craster. Asymptotic analysis of a Bragg fibre: a multiple scales approach. IMA Journal of Applied Mathematics, 81(6):1191–1208, December 2016. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://imamat.

oxfordjournals.org/content/81/6/
1191.

Barlow:2017:AST

[1498] Nathaniel S. Barlow, Brian T. Helenbrook, and Steven J. Weinstein. Algorithm for spatio-temporal analysis of the signalling problem. IMA Journal of Applied Mathematics, 82 (1):1–32, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/1/1/2884399/Algorithm-for-spatio-
temporal-analysis-of-the.

Zhen:2017:STD

[1499] Qiang Zhen, Johan S. H. van Leeuwaarden, and Charles Knessl. On the sojourn time distribution in a finite population Markovian processor sharing queue. IMA Journal of Applied Mathematics, 82(1):33–59, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/82/1/33/2884400/On-the-
                                        sojourn-time-distribution-in-a-
                                        finite.

Dunster:2017:MMT

[1500] J. L. Dunster and J. R. King. Mathematical modelling of thrombin generation: asymptotic analysis and pathway characterization. IMA Journal of Applied Mathematics, 82(1):60–96, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/82/1/60/
                                        2884401/Mathematical-modelling-
                                        of-thrombin-generation.

Geng:2017:AGC

[1501] Xianguo Geng and Hui Wang. Algebrogeometric constructions of quasiperiodic flows of the Newell hierarchy and applications. IMA Journal of Applied Mathematics, 82(1):97–130, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/1/97/2884402/
                                        Algebro-geometric-constructions-
                                        of-quasi-periodic.

Pichler:2017:FNL

[1502] Monika Pichler and Gino Biondini.

On the focusing non-linear Schrödinger equation with non-zero boundary conditions and double poles. IMA Journal of Applied Mathematics, 82(1): 131–151, February 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/1/131/2884403/On-the-
focusing-non-linear-Schrodinger-
equation.

Armbruster:2017:EPC

[1503] Benjamin Armbruster and Ekkehard Beck. An elementary proof of convergence to the mean-field equations for an epidemic model. IMA Journal of Applied Mathematics, 82(1):152–157, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/82/1/152/
2884404/An-elementary-proof-of-
convergence-to-the-mean.

Gomes:2017:SNT

[1504] Susana N. Gomes, Demetrios T. Papageorgiou, and Grigorios A. Pavliotis. Stabilizing non-trivial solutions of the generalized Kuramoto–Sivashinsky equation using feedback and optimal control: Lighthill–Thwaites Prize. IMA Journal of Applied Mathematics, 82(1):158–194, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/82/1/158/
2884405/Stabilizing-non-trivial-
solutions-of-the.

Rachah:2017:MMS

[1505] A. Rachah, D. Noll, F. Espitalier, and F. Baillon. Mathematical modelling for solvated crystallization. IMA Journal of Applied Mathematics, 82(1):195–223, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/82/1/195/
                                        2884406/Mathematical-modelling-
                                        for-solvated.

Harris:2017:WSA

[1506] Pietro Artale Harris, Emilio N. M. Cirillo, and Adrian Muntean. Weak solutions to Allen–Cahn-like equations modelling consolidation of porous media. IMA Journal of Applied Mathematics, 82(1):224–250, February 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/82/1/224/
                                        2884407/Weak-solutions-to-Allen-
                                        -Cahn-like-equations.

Black:2017:QME

[1507] Jonathan P. Black, Christopher J. W. Breward, and Peter D. Howell. Quantum mechanical effects in continuum charge flow models. IMA Journal of Applied Mathematics, 82(2):251–279, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/2/251/2669854/
                                        Quantum-mechanical-effects-in-continuum-
                                        charge.

Ayton:2017:ASG

[1508] Lorna J. Ayton. An analytic solution for gust–aerofoil interaction noise including effects of geometry. IMA Journal of Applied Mathematics, 82(2):280–304, April 1, 2017.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL https://academic.oup.com/
imamat/article/82/2/280/2669860/
An-analytic-solution-for-gust-aerofoil-
interaction.

Craske:2017:PIM

[1509] John Craske. The properties of integral models for planar and axisymmetric unsteady jets. IMA Journal of Applied Mathematics, 82(2): 305–333, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/2/305/2419535/The-properties-
of-integral-models-for-planar-and.

Matei:2017:MVF

[1510] Andaluzia Matei and Mircea Sofonea.

A mixed variational formulation for a piezoelectric frictional contact problem. IMA Journal of Applied Mathematics, 82(2):334–354, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL https://academic.oup.com/
imamat/article/82/2/334/2739334/
A-mixed-variational-formulation-
for-a.

Cox:2017:AHC

[1511] Stephen M. Cox, Jun Yu, Wang Ling Goh, and Meng Tong Tan. Analysis of a hysteresis-controlled selfoscillating class-D amplifier. IMA Journal of Applied Mathematics, 82 (2):355–370, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/2/355/2739337/Analysis-of-a-
hysteresis-controlled-self.

Borcea:2017:DD

[1512] Ciprian S. Borcea and Ileana Streinu. Deforming diamond. IMA Journal of Applied Mathematics, 82(2):371–383, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/82/2/371/
                                        2739341/Deforming-diamond.

Dai:2017:CFS

[1513] Ming Dai, Cun-Fa Gao, and Peter Schiavone. Closed-form solution for a circular nano-inhomogeneity with interface effects in an elastic plane under uniform remote heat flux. IMA Journal of Applied Mathematics, 82 (2):384–395, April 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/82/2/384/2739343/Closed-
                                        form-solution-for-a-circular-nano.

Liu:2017:SDT

[1514] Meng Liu and Meng Fan. Stability in distribution of a three-species stochastic cascade predator-prey system with time delays. IMA Journal of Applied Mathematics, 82(2):396–423, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/2/396/3073985/
                                        Stability-in-distribution-of-a-
                                        three-species.

Wu:2017:HSI

[1515] Bin Wu and Jun Yu. Hölder stability of an inverse problem for a strongly coupled reaction–diffusion system. IMA Journal of Applied Mathematics, 82(2):424–444, April 1, 2017.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/2/424/2938062/Holder-
stability-of-an-inverse-problem-
for-a.

Chen:2017:RDA

[1516] Qiaoling Chen, Fengquan Li, and Feng Wang. A reaction–diffusion– advection competition model with two free boundaries in heterogeneous time-periodic environment. IMA Journal of Applied Mathematics, 82(2):445–470, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL https://academic.oup.com/
imamat/article/82/2/445/2894348/
A-reaction-diffusion-advection-
competition-model.

Chillingworth:2017:EPH

[1517] D. R. J. Chillingworth. Erratum to Perturbed hedgehogs: continuous deformation of point defects in biaxial nematic liquid crystals. IMA Journal of Applied Mathematics, 82(2):471, April 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL https://academic.oup.com/
imamat/article/82/2/471/2884410/
Erratum-to-Perturbed-hedgehogs-

continuous. See [1471].

Murphy:2017:MMC

[1518] Ellen A. Murphy and William T. Lee.

Mathematical modelling of contact lens moulding. IMA Journal of Applied Mathematics, 82(3):473–495, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/82/3/473/
                                        3092317/Mathematical-modelling-
                                        of-contact-lens-moulding.

Chugainova:2017:PAD

[1519] Anna P. Chugainova, Andrey T. Il’ichev, Andrey G. Kulikovskii, and Vladimir A. Shargatov. Problem of arbitrary discontinuity disintegration for the generalized Hopf equation: selection conditions for a unique solution. IMA Journal of Applied Mathematics, 82(3):496–525, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/3/496/3098301/
                                        Problem-of-arbitrary-discontinuity-
                                        disintegration.

Florio:2017:ICM

[1520] Brendan J. Florio, Andrew P. Bassom, Konstantinos Sakellariou, and Thomas Stemler. Interacting convection modes in a saturated porous medium of nearly square planform: four modes. IMA Journal of Applied Mathematics, 82(3):526–547, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/82/3/526/
                                        3056509/Interacting-convection-
                                        modes-in-a-saturated-porous.

Wihler:2017:APS

[1521] Thomas P. Wihler. Animal population social structure models. IMA Journal of Applied Mathematics, 82 (3):548–560, June 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/3/548/3744561/Animal-
population-social-structure-models.

Carvalho:2017:BLC

[1522] Tiago Carvalho, Rodrigo D. Euzébio, Marco Antonto Teixeira, and Durval José Tonon. Birth of limit cycles from a 3D triangular center of a piecewise smooth vector field. IMA Journal of Applied Mathematics, 82(3):561–578, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL https://academic.oup.com/
imamat/article/82/3/561/3056510/
Birth-of-limit-cycles-from-a-3D-
triangular-center.

Zhang:2017:RPT

[1523] Zhidong Zhang and Zhi Zhou. Recovering the potential term in a fractional diffusion equation. IMA Journal of Applied Mathematics, 82(3): 579–600, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/3/579/3744562/Recovering-the-
potential-term-in-a-fractional.

Zhang:2017:DST

[1524] Yalin Zhang and Guoliang Shi.

The distributions of some transmission eigenvalues in one dimension. IMA Journal of Applied Mathematics, 82(3):601–628, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL https://academic.oup.com/
imamat/article/82/3/601/3782818/
The-distributions-of-some-transmission-
eigenvalues.

Yang:2017:TDA

[1525] Junyuan Yang, Yuming Chen, and Toshikazu Kuniya. Threshold dynamics of an age-structured epidemic model with relapse and nonlinear incidence. IMA Journal of Applied Mathematics, 82(3):629–655, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/3/629/3749197/
                                        Threshold-dynamics-of-an-age-structured-
                                        epidemic.

Shariff:2017:SAR

[1526] M. H. B. M. Shariff, R. Bustamante, and J. Merodio. On the spectral analysis of residual stress in finite elasticity. IMA Journal of Applied Mathematics, 82(3):656–680, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/82/3/656/3792106/On-the-
                                        spectral-analysis-of-residual-stress-
                                        in.

Wattis:2017:STL

[1527] Jonathan A. D. Wattis. Shape of transition layers in a differential-delay equation. IMA Journal of Applied Mathematics, 82(3):681–696, June 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/3/681/3797159/
                                        Shape-of-transition-layers-in-a-
                                        differential-delay.

Boldrighini:2017:BSC

[1528] C. Boldrighini, S. Frigio, and P. Maponi. On the blow-up of some complex solutions of the 3D Navier–Stokes equations: theoretical predictions and computer simulations. IMA Journal of Applied Mathematics, 82(4):697–716, August 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/82/4/697/
3824630/On-the-blow-up-of-some-
complex-solutions-of-the-3D.

Leach:2017:NLT

[1529] J. A. Leach. A note on the large-time development of the solution to an initial-value problem for a variable coefficient Korteweg– de Vries equation. IMA Journal of Applied Mathematics, 82(4):717– 725, August 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/4/717/3835409/A-note-on-the-
large-time-development-of-the.

Konrad:2017:TSS

[1530] Isabell Konrad, Malte A. Peter, and John M. Stockie. A two-scale Stefan problem arising in a model for tree sap exudation. IMA Journal of Applied Mathematics, 82(4): 726–762, August 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/4/726/3782821/A-two-
scale-Stefan-problem-arising-in-
a-model-for.

Homberg:2017:RJM

[1531] Dietmar Hömberg, Francesco Saverio Patacchini, Kenichi Sakamoto, and Johannes Zimmer. A revisited Johnson– Mehl–Avrami–Kolmogorov model and the evolution of grain-size distributions in steel. IMA Journal of Applied Mathematics, 82(4):763–780, August 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/4/763/3832242/
                                        A-revisited-Johnson-Mehl-Avrami-
                                        Kolmogorov-model.

Cowley:2017:NEA

[1532] James Cowley, Anthony J. Mulholland, and Anthony Gachagan. A nonlinear elasticity approach to modelling the collapse of a shelled microbubble. IMA Journal of Applied Mathematics, 82(4):781–801, August 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/82/4/781/
                                        3835410/A-nonlinear-elasticity-
                                        approach-to-modelling-the.

Grimshaw:2017:RSW

[1533] R. H. J. Grimshaw, K. R. Khusnutdinova, and K. R. Moore. Radiating solitary waves in coupled Boussinesq equations. IMA Journal of Applied Mathematics, 82(4):802–820, August 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

                                        URL https://academic.oup.com/
                                        imamat/article/82/4/802/3852583/
                                        Radiating-solitary-waves-in-coupled-
                                        Boussinesq.

Doak:2017:SGW

[1534] Alex Doak and Jean-Marc Vanden-Broeck. Solitary gravity waves and free surface flows past a point vortex. IMA Journal of Applied Mathematics, 82(4):821–835, August 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/82/4/821/
3859007/Solitary-gravity-waves-
and-free-surface-flows-past.

Zhao:2017:FDM

[1535] Hui Zhao and Hongjiong Tian. Finite difference methods of the spatial fractional Black–Scholes equation for a European call option. IMA Journal of Applied Mathematics, 82 (4):836–848, August 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/4/836/3897366/Finite-
difference-methods-of-the-spatial.

Braga:2017:LCB

[1536] Denis de Carvalho Braga, Tiago Carvalho, and Luis Fernando Mello. Limit cycles bifurcating from discontinuous centres. IMA Journal of Applied Mathematics, 82(4): 849–863, August 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/4/849/3920515/Limit-
cycles-bifurcating-from-discontinuous.

Guo:2017:STP

[1537] Shangjiang Guo. Spatio-temporal patterns in a diffusive model with non-local delay effect. IMA Journal of Applied Mathematics, 82(4): 864–908, August 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/4/864/3869472/Spatio-

                                        temporal-patterns-in-a-diffusive-
                                        model-with.

Ding:2017:NSO

[1538] Hengfei Ding, Changpin Li, and Qian Yi. A new second-order midpoint approximation formula for Riemann–Liouville derivative: algorithm and its application. IMA Journal of Applied Mathematics, 82(5): 909–944, October 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/82/5/909/3869935/A-new-
                                        second-order-midpoint-approximation-
                                        formula.

Wang:2017:SEM

[1539] Jinliang Wang, Min Guo, and Shengqiang Liu. SVIR epidemic model with age structure in susceptibility, vaccination effects and relapse. IMA Journal of Applied Mathematics, 82(5): 945–970, October 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        82/5/945/3883927/SVIR-epidemic-
                                        model-with-age-structure-in.

Roberts:2017:SVM

[1540] A. J. Roberts and J. E. Bunder. Slowly varying, macroscale models emerge from microscale dynamics over multiscale domains. IMA Journal of Applied Mathematics, 82(5):971– 1012, October 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        82/5/971/3924902/Slowly-varying-
                                        macroscale-models-emerge-from.

Li:2017:EIT

[1541] Xiaofei Li, Jingzhi Li, Hongyu Liu, and Yuliang Wang. Electromagnetic interior transmission eigenvalue problem for inhomogeneous media containing obstacles and its applications to near cloaking. IMA Journal of Applied Mathematics, 82(5): 1013–1042, October 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/5/1013/3924909/Electromagnetic-
interior-transmission-eigenvalue.

Liu:2017:MKS

[1542] Bin Liu, Hai Huyen Heidi Dam, Kok Lay Teo, and David John Hill. KL_{∗}-stability for a class of hybrid dynamical systems. IMA Journal of Applied Mathematics, 82(5):1043– 1060, October 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/5/1043/4056082/mathcal-KL-stability-
for-a-class-of-hybrid.

Muriel:2017:LSI

[1543] C. Muriel, J. L. Romero, and A. Ruiz. λ-Symmetries and integrability by quadratures. IMA Journal of Applied Mathematics, 82(5): 1061–1087, October 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/5/1061/4083276/lambda-Symmetries-
and-integrability-by-quadratures.

Wang:2017:SII

[1544] Xu Wang and Peter Schiavone. Surface instability of imperfectly bonded multi-layered curved films under van der Waals forces. IMA Journal of Applied Mathematics, 82(5):1088– 1103, October 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/5/1088/4100600/Surface-instability-
of-imperfectly-bonded-multi.

Pedneault:2017:SCD

[1545] Michael Pedneault, Catalin Turc, and Yassine Boubendir. Schur complement domain decomposition methods for the solution of multiple scattering problems. IMA Journal of Applied Mathematics, 82(5): 1104–1134, October 1, 2017. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/5/1104/4102357/Schur-
complement-domain-decomposition-
methods-for.

Walker:2017:POI

[1546] Alan J. Walker and Anthony J. Mulholland. A pipe organ-inspired ultrasonic transducer. IMA Journal of Applied Mathematics, 82(6):1135– 1150, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/6/1135/4103631.

Ratliff:2017:RCN

[1547] Daniel J. Ratliff. On the reduction of coupled NLS equations to non-linear phase equations via modulation of a two-phase wavetrain. IMA Journal of Applied Mathematics, 82(6):1151– 1170, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/6/1151/4237505.

Wang:2017:FBM

[1548] Yanjiao Wang and Feng Ding. A filtering based multi-innovation gradient estimation algorithm and performance analysis for nonlinear dynamical systems. IMA Journal of Applied Mathematics, 82(6):1171– 1191, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/6/1171/4599949.

Gomes:2017:NSS

[1549] Susana N. Gomes and Stephen James Tate. On the numerical solution of a T -Sylvester type matrix equation arising in the control of stochastic partial differential equations. IMA Journal of Applied Mathematics, 82(6):1192– 1208, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
82/6/1192/4210980.

Upadhyay:2017:IEF

[1550] Ranjit Kumar Upadhyay, Swati Mishra, Rana D. Parshad, Jingjing Lyu, and Aladeen Al Basheer. Investigation of an explosive food chain model with interference and inhibitory effects. IMA Journal of Applied Mathematics, 82(6):1209–1237, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/82/6/1209/4641688.

Sowa:2017:RZF

[1551] Artur Sowa. Riemann’s zeta function and the broadband structure of pure harmonics. IMA Journal of Applied Mathematics, 82(6):1238– 1252, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        82/6/1238/4643393.

Roberts:2017:MMM

[1552] Lewis Roberts, Thom Griffith, Alan Champneys, Martina Piano, Janice Kiely, and Richard Luxton. Mathematical modelling of a magnetic immunoassay. IMA Journal of Applied Mathematics, 82(6):1253– 1282, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        82/6/1253/4638324.

Destuynder:2017:CSA

[1553] Philippe Destuynder and Caroline Fabre. On control strategies for avoiding loss of information in nondestructive testing. IMA Journal of Applied Mathematics, 82(6):1283– 1302, December 1, 2017. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        82/6/1283/4695478.

Khapalov:2018:DDM

[1554] Alexander Khapalov and Sergey Lapin. Dynamic discrete models for the granular matter formation process. IMA Journal of Applied Mathematics, 83 (1):1–23, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/1/1/2884408.

Wang:2018:UEF

[1555] Qiming Wang and Demetrios T. Papageorgiou. Using electric fields to induce patterning in leaky dielectric fluids in a rod-annular geometry. IMA Journal of Applied Mathematics, 83 (1):24–52, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/1/24/2884409.

Lyalinov:2018:ASS

[1556] Mikhail A. Lyalinov. Acoustic scattering by a semi-infinite angular sector with impedance boundary conditions. IMA Journal of Applied Mathematics, 83(1):53–91, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/1/53/4796912.

Bastos:2018:FCM

[1557] M. A. Bastos, P. A. Lopes, and A. F.

dos Santos. Factorization of a class of 2 × 2 matrix symbols by reduction to a scalar factorization. IMA Journal of Applied Mathematics, 83(1): 92–105, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/1/92/4772213.

Planella:2018:ISS

[1558] Ferran Brosa Planella, Colin P. Please, and Robert A. Van Gorder. Instability in the self-similar motion of a planar solidification front. IMA Journal of Applied Mathematics, 83(1): 106–130, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/1/106/4772214.

Rabieifar:2018:TAT

[1559] Arman Rabieifar, Mehdi Pourseifi, and Hesameldien Derili. Transient analysis for torsional impact of multiple axisymmetric cracks in the functionally graded orthotropic medium. IMA Journal of Applied Mathematics, 83(1): 131–147, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/1/131/4780156.

Liu:2018:VDN

[1560] Huan Liu and Xianguo Geng. The vector derivative nonlinear Schrödinger equation on the half-line. IMA Journal of Applied Mathematics, 83(1): 148–173, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/1/148/4772215.

Alsharif:2018:DFB

[1561] Abdullah Madhi Alsharif. Drop formation and break-up of rotating viscoelastic liquid jets in the Giesekus model. IMA Journal of Applied Mathematics, 83(1):174–187, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/1/174/4797574.

Page:2018:FST

[1562] Michael A. Page and Stephen J. Cowley. On the formation of small-time curvature singularities in vortex sheets. IMA Journal of Applied Mathematics, 83(1):188–203, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/1/188/4772216.

Fernandez:2018:UAS

[1563] A. Fernandez, E. A. Spence, and A. S.

Fokas. Uniform asymptotics as a stationary point approaches an endpoint. IMA Journal of Applied Mathematics, 83(1):204–242, January 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/1/204/4788352.

Doschoris:2018:QED

[1564] Michael Doschoris and Foteini Kariotou. Quantifying errors during the source localization process in electroencephalography for confocal systems. IMA Journal of Applied Mathematics, 83(2):243–260, March 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/2/243/4817458.

Chugreeva:2018:SSR

[1565] Olga Chugreeva and Christof Melcher.

Strong solvability of regularized stochastic Landau–Lifshitz–Gilbert equation. IMA Journal of Applied Mathematics, 83(2):261–282, March 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.

                                        oup.com/imamat/article/83/2/261/
                                        4817459.

Wang:2018:FMI

[1566] Yan Wang, Jun Guo, and Guozheng Yan. The factorization method for the inverse obstacle scattering problem with oblique derivative boundary condition. IMA Journal of Applied Mathematics, 83(2):283–301, March 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/2/283/4859458.

Denk:2018:ESC

[1567] Robert Denk and Felix Kammerlander. Exponential stability for a coupled system of damped–undamped plate equations. IMA Journal of Applied Mathematics, 83(2):302–322, March 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/2/302/4877164.

Bykov:2018:SPP

[1568] V. Bykov, Y. Cherkinsky, V. Gol’dshtein, N. Krapivnik, and U. Maas. Singularly perturbed profiles. IMA Journal of Applied Mathematics, 83(2): 323–346, March 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/2/323/4925456.

Chen:2018:BCM

[1569] Chen Chen, A. J. Roberts, and J. E. Bunder. Boundary conditions for macroscale waves in an elastic system with microscale heterogeneity. IMA Journal of Applied Mathematics, 83 (3):347–379, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/3/347/4944047.

Sakajo:2018:TRS

[1570] Takashi Sakajo and Tomoo Yokoyama.

Tree representations of streamline topologies of structurally stable 2D incompressible flows. IMA Journal of Applied Mathematics, 83(3):380–411, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/83/3/380/
4953824.

Baranska:2018:GES

[1571] Joanna Barańska and Yuri Kozitsky.

The global evolution of states of a continuum Kawasaki model with repulsion. IMA Journal of Applied Mathematics, 83(3):412–435, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/3/412/4944048.

Gao:2018:NCT

[1572] Tao Gao, Jean-Marc Vanden-Broeck, and Zhan Wang. Numerical computations of two-dimensional flexuralgravity solitary waves on water of arbitrary depth. IMA Journal of Applied Mathematics, 83(3):436–450, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/83/3/436/
4951679.

Meur:2018:DVK

[1573] H. V. J. Le Meur. Derivation of a viscous KP equation including surface tension, and related equations. IMA Journal of Applied Mathematics, 83 (3):451–470, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/3/451/4952145.

Deliktas:2018:SSH

[1574] Ekin Deliktas and Mevlut Teymur. Surface shear horizontal waves in a double-layered nonlinear elastic half space. IMA Journal of Applied Mathematics, 83(3):471–495, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/3/471/4967637.

Soffer:2018:MEA

[1575] Avy Soffer and Xiaofei Zhao. Modulation equations approach for solving vortex and radiation in nonlinear Schrödinger equation. IMA Journal of Applied Mathematics, 83(3): 496–513, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/3/496/4993776.

Lazowski:2018:EMW

[1576] Andrew Lazowski and Stephen Shea. The effects on monotonicity when combining alternative vote with plurality. IMA Journal of Applied Mathematics, 83(3):514–525, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/3/514/4993777.

Burger:2018:FIS

[1577] Raimund Bürger, Julio Careaga, and Stefan Diehl. Flux identification of scalar conservation laws from sedimentation in a cone. IMA Journal of Applied Mathematics, 83(3):526–552, May 23, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/83/3/526/
4992915.

Smith:2018:SIH

[1578] David J. Smith, Timothy J. Pedley, and Darren G. Crowdy. Special issue in honour of Professor John Blake FIMA CMath. IMA Journal of Applied Mathematics, 83(4):553–555, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/4/553/5054622.

Lauterborn:2018:BMR

[1579] Werner Lauterborn, Christiane Lechner, Max Koch, and Robert Mettin. Bubble models and real bubbles: Rayleigh and energy-deposit cases in a Tait-compressible liquid. IMA Journal of Applied Mathematics, 83 (4):556–589, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/4/556/4995067.

Dauparas:2018:LOS

[1580] Justas Dauparas and Eric Lauga.

Leading-order Stokes flows near a corner. IMA Journal of Applied Mathematics, 83(4):590–633, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/4/590/4995066.

Gadelha:2018:FBE

[1581] Hermes Gadêlha. The filament-bundle elastica. IMA Journal of Applied Mathematics, 83(4):634–654, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/4/634/5054628.

Ishimoto:2018:ESS

[1582] Kenta Ishimoto and Eamonn A. Gaffney. An elastohydrodynamical simulation study of filament and spermatozoan swimming driven by internal couples. IMA Journal of Applied Mathematics, 83(4):655–679, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/4/655/5054625.

Delmotte:2018:SBT

[1583] Blaise Delmotte, Eric E. Keaveny, Eric Climent, and Franck Plouraboué. Simulations of Brownian tracer transport in squirmer suspensions. IMA Journal of Applied Mathematics, 83(4): 680–699, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/4/680/5054626.

Kirkegaard:2018:RTF

[1584] Julius B. Kirkegaard and Raymond E. Goldstein. The role of tumbling frequency and persistence in optimal runand-tumble chemotaxis. IMA Journal of Applied Mathematics, 83(4): 700–719, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/4/700/4990631.

Gravatt:2018:IIC

[1585] Michael Gravatt, Vinod Suresh, Alys Clark, and Richard Clarke. Importance of irrotational components of swimming flows on the stability of a suspension of weakly-squirming microorganisms. IMA Journal of Applied Mathematics, 83(4):720–742, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/4/720/5054623.

Clifton:2018:ESE

[1586] W. Clifton, R. N. Bearon, and M. A.

Bees. Enhanced sedimentation of elongated plankton in simple flows. IMA Journal of Applied Mathematics, 83 (4):743–766, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/4/743/5032334.

Matsunaga:2018:FFT

[1587] Daiki Matsunaga, Andreas Zöttl, Fanlong Meng, Ramin Golestanian, and Julia M. Yeomans. Far-field theory for trajectories of magnetic ellipsoids in rectangular and circular channels. IMA Journal of Applied Mathematics, 83(4):767–782, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/4/767/5054627.

Brennen:2018:NAP

[1588] C. E. Brennen, G. Keady, and J. Imberger. A note on algal population dynamics. IMA Journal of Applied Mathematics, 83(4):783–796, July 25, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/4/783/5055018.

Champneys:2018:I

[1589] Alan R. Champneys. Introduction. IMA Journal of Applied Mathematics, 83(5):797–798, September 24, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/83/5/797/5095480.

OKiely:2018:GSR

[1590] Doireann O’Kiely, Chris J. W. Breward, Ian M. Griffiths, Peter D. Howell, and Ulrich Lange. Glass sheet redraw through a long heater zone. IMA Journal of Applied Mathematics, 83(5): 799–820, September 24, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/5/799/5021699.

Fadai:2018:IDC

[1591] Nabil T. Fadai, Zahra Akram, Fabien Guilmineau, John Melrose, Colin P. Please, and Robert A. Van Gorder. The influence of distributed chemical reaction groups in a multiphase coffee bean roasting model. IMA Journal of Applied Mathematics, 83(5):821– 848, September 24, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/5/821/5021697.

Allanson:2018:IPC

[1592] Oliver Allanson, Sascha Troscheit, and Thomas Neukirch. On the inverse problem for Channell collisionless plasma equilibria. IMA Journal of Applied Mathematics, 83(5):849–873, September 24, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/83/5/849/
5033827.

Wilmott:2018:SFT

[1593] Z. M. Wilmott, C. J. W. Breward, and S. J. Chapman. Slip flow through channels with varying elliptic cross section. IMA Journal of Applied Mathematics, 83(5):874–893, September 24, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/5/874/5032333.

Majic:2018:LEP

[1594] Matt R. A. Majić, Baptiste Auguié, and Eric C. Le Ru. Laplace’s equation for a point source near a sphere: improved internal solution using spheroidal harmonics. IMA Journal of Applied Mathematics, 83(6):895– 907, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/6/895/5038895.

Moya-Cessa:2018:DFF

[1595] Héctor M. Moya-Cessa and Francisco Soto-Eguibar. Discrete fractional Fourier transform: Vandermonde approach. IMA Journal of Applied Mathematics, 83(6):908–916, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/83/6/908/
5041981.

Luca:2018:LSF

[1596] Elena Luca, Jonathan Marshall, and Georgios Karamanis. Longitudinal shear flow over a bubble mattress with curved menisci: arbitrary protrusion angle and solid fraction. IMA Journal of Applied Mathematics, 83(6):917– 941, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/6/917/5046635.

Luca:2018:TMB

[1597] Elena Luca and Darren G. Crowdy. A transform method for the biharmonic equation in multiply connected circular domains. IMA Journal of Applied Mathematics, 83(6):942–976, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/83/6/942/
5048781.

Berlyand:2018:MGL

[1598] Leonid Berlyand, Vladimir Mityushev, and Shawn D. Ryan. Multiple Ginzburg–Landau vortices pinned by randomly distributed small holes. IMA Journal of Applied Mathematics, 83(6): 977–1006, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/6/977/5066467.

Falocchi:2018:TIN

[1599] Alessio Falocchi. Torsional instability in a nonlinear isolated model for suspension bridges with fixed cables and extensible hangers. IMA Journal of Applied Mathematics, 83(6):1007– 1036, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/6/1007/5068239.

Yan:2018:AAS

[1600] Dongxue Yan and Xianlong Fu. Analysis of an age-structured HIV infection model with logistic target-cell growth and antiretroviral therapy. IMA Journal of Applied Mathematics, 83 (6):1037–1065, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/83/6/1037/5068238.

Feng:2018:CSD

[1601] Liuyang Feng, Nan Qi, Hao Gao, Wei Sun, Mariano Vazquez, Boyce E. Griffith, and Xiaoyu Luo. On the chordae structure and dynamic behaviour of the mitral valve. IMA Journal of Applied Mathematics, 83(6):1066– 1091, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
83/6/1066/5079152.

Rickaby:2018:CSI

[1602] Stephen R. Rickaby and Nigel H. Scott.

On the complex singularities of the inverse Langevin function. IMA Journal of Applied Mathematics, 83(6):1092– 1116, November 27, 2018. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        83/6/1092/5098281.

vanRensburg:2019:EUS

[1603] N. F. J. van Rensburg and B. Stapelberg. Existence and uniqueness of solutions of a general linear secondorder hyperbolic problem. IMA Journal of Applied Mathematics, 84(1): 1–22, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/1/1/5098280.

Bhattacharyya:2019:GDM

[1604] Samit Bhattacharyya and Timothy Reluga. Game dynamic model of social distancing while cost of infection varies with epidemic burden. IMA Journal of Applied Mathematics, 84 (1):23–43, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/1/23/5101360.

Tong:2019:UGP

[1605] William Tong and Holger R. Dullin. Using the geometric phase to optimize planar somersaults. IMA Journal of Applied Mathematics, 84(1): 44–62, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/1/44/5115487.

Anguiano:2019:NRE

[1606] María Anguiano and Francisco Javier Suárez-Grau. Nonlinear Reynolds equations for non-Newtonian thin-film fluid flows over a rough boundary. IMA Journal of Applied Mathematics, 84(1): 63–95, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/1/63/5115485.

Alsharif:2019:TSC

[1607] Abdullah Madhi Alsharif, Stephen P.

Decent, Emilian I. Parau, Mark J. H. Simmons, and Jamal Uddin. The trajectory of slender curved liquid jets for small Rossby number. IMA Journal of Applied Mathematics, 84(1): 96–117, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/1/96/5115486.

Sugimoto:2019:MCO

[1608] Nobumasa Sugimoto. Marginal conditions for the onset of thermoacoustic oscillations due to instability of heat flow. IMA Journal of Applied Mathematics, 84(1):118–144, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/84/1/118/
5123590.

Liu:2019:MAW

[1609] Qihuai Liu, Pedro J. Torres, and Mingyan Xing. Modulated amplitude waves with non-trivial phase of multi-component Bose–Einstein condensates in optical lattices. IMA Journal of Applied Mathematics, 84(1): 145–170, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/1/145/5123754.

Wang:2019:UST

[1610] Yang Wang, Lingling Shi, Guirong Liu, and Zhaohai Ma. The uniqueness and stability of travelling fronts of competitive systems with different diffusive coefficients. IMA Journal of Applied Mathematics, 84(1): 171–196, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/1/171/5123591.

Ren:2019:RAD

[1611] Jingli Ren and Dandan Zhu. On a reaction–advection–diffusion equation with double free boundaries and m th-order Fisher non-linearity. IMA Journal of Applied Mathematics, 84(1): 197–227, January 25, 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/1/197/5132953.

Li:2019:IEW

[1612] Jingzhi Li, Hongyu Liu, and Hongpeng Sun. On an inverse elastic wave imaging scheme for nearly incompressible materials. IMA Journal of Applied Mathematics, 84(2):229– 257, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/84/2/229/
                                        5123589.

Tuan:2019:RSS

[1613] Nguyen Huy Tuan, Daniel Lesnic, Tran Quoc Viet, and Vo Van Au.

Regularization of the semilinear sideways heat equation. IMA Journal of Applied Mathematics, 84(2):258–291, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/84/2/258/
5136502.

Mindrinos:2019:ESP

[1614] Leonidas Mindrinos. The electromagnetic scattering problem by a cylindrical doubly connected domain at oblique incidence: the direct problem. IMA Journal of Applied Mathematics, 84(2):292–311, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/2/292/5161356.

Sargent:2019:TMH

[1615] Cristina V. Sargent and A. J. Mestel. Trapped modes of the Helmholtz equation in infinite waveguides with wall indentations and circular obstacles. IMA Journal of Applied Mathematics, 84(2):312–344, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/2/312/5199420.

Zelentsov:2019:DWT

[1616] V. B. Zelentsov, B. I. Mitrin, I. A.

Lubyagin, and I. I. Kudish. Diagnostics of wear thermoelastic instability based on sliding contact parameter monitoring. IMA Journal of Applied Mathematics, 84(2):345–365, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/2/345/5233906.

Giddings:2019:EID

[1617] Josef Adam Giddings and John Billingham. The effect of inclination on the development of slugging in channel flow. IMA Journal of Applied Mathematics, 84(2):366–384, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/84/2/366/5233396.

Whittaker:2019:STP

[1618] Robert J. Whittaker and Simon Cox. Stability of a twisted plateau border with line tension and bending stiffness. IMA Journal of Applied Mathematics, 84(2):385–415, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/2/385/5288764.

Ambrosio:2019:LTB

[1619] B. Ambrosio, M. A. Aziz-Alaoui, and V. L. E. Phan. Large time behaviour and synchronization of complex networks of reaction-diffusion systems of FitzHugh–Nagumo type. IMA Journal of Applied Mathematics, 84(2):416– 443, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/84/2/416/
                                        5289480.

Hagwood:2019:LDR

[1620] Charles Hagwood and George Mulholland. Limiting dynamics for a rotating system. IMA Journal of Applied Mathematics, 84(2):444–453, April 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/2/444/5289471.

Marshall:2019:MLB

[1621] Nicholas F. Marshall and Ronald R.

Coifman. Manifold learning with bistochastic kernels. IMA Journal of Applied Mathematics, 84(3):455–482, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/84/3/455/
5289465.

Montecinos:2019:NPC

[1622] Gino I. Montecinos, Juan C. López-Ríos, Jaime H. Ortega, and Rodrigo Lecaros. A numerical procedure and coupled system formulation for the adjoint approach in hyperbolic PDEconstrained optimization problems. IMA Journal of Applied Mathematics, 84(3):483–516, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/3/483/5289479.

Kay:2019:TBL

[1623] Anthony Kay. The thermal boundary layer due to viscous dissipation in impulsively started Poiseuille flow. IMA Journal of Applied Mathematics, 84(3):517–532, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/3/517/5301756.

Vynnycky:2019:AAD

[1624] Michael Vynnycky, Sean McKee, Martin Meere, Christopher McCormick, and Sean McGinty. Asymptotic analysis of drug dissolution in two layers having widely differing diffusivities. IMA Journal of Applied Mathematics, 84(3):533–554, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/3/533/5304593.

Wang:2019:EHI

[1625] Xu Wang, Liang Chen, and Peter Schiavone. An elastic harmonic inclusion near a rigid harmonic inclusion loaded by a couple. IMA Journal of Applied Mathematics, 84(3):555– 566, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/84/3/555/
5308136.

Shin:2019:ICA

[1626] Ho-Chul Shin. Image conditions and addition theorems for prolate and oblate spheroidal-coordinate separation-of-variables acoustic multiple scattering models with perfectlyreflecting flat surfaces. IMA Journal of Applied Mathematics, 84(3):567– 604, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/84/3/567/
5310421.

Algehyne:2019:AFU

[1627] Ebrahem A. Algehyne and Anthony J. Mulholland. Analysis of a fractal ultrasonic transducer with a range of piezoelectric length scales. IMA Journal of Applied Mathematics, 84(3):605– 631, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.

oup.com/imamat/article/84/3/605/
5381041.

Bailey:2019:ERF

[1628] Nicola Y. Bailey, Stephen Hibberd, Michael V. Tretyakov, and Henry Power. Effect of random forcing on fluid-lubricated bearing. IMA Journal of Applied Mathematics, 84(3):632– 649, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/84/3/632/
5482273.

Gautam:2019:TCC

[1629] K. Gautam, P. A. L. Narayana, and A. A. Hill. Thermo-convective carbon sequestration in horizontal porous layers. IMA Journal of Applied Mathematics, 84(3):650–668, June 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/3/650/5487608.

ONaraigh:2019:MFD

[1630] Lennon Ó Náraigh and Khang Ee Pang.

A mathematical framework for determining the stability of steady states of reaction–diffusion equations with periodic source terms. IMA Journal of Applied Mathematics, 84(4):669–678, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/84/4/669/
5522254.

Tan:2019:EIT

[1631] Linda S. L. Tan. Explicit inverse of tridiagonal matrix with applications in autoregressive modelling. IMA Journal of Applied Mathematics, 84 (4):679–695, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/4/679/5525701.

Wang:2019:MDV

[1632] Qianxi Wang, WenKe Liu, David M. Leppinen, and A. D. Walmsley. Microbubble dynamics in a viscous compressible liquid near a rigid boundary. IMA Journal of Applied Mathematics, 84(4):696–711, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/4/696/5528537.

Wang:2019:LRS

[1633] JinRong Wang, Mengmeng Li, and Donal O’Regan. Lyapunov regularity and stability of linear noninstantaneous impulsive differential systems. IMA Journal of Applied Mathematics, 84(4):712–747, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/84/4/712/
                                        5537171.

Greco:2019:UDC

[1634] Carlo Greco. Uniqueness of director configuration states for liquid crystals in the case of weak anchoring. IMA Journal of Applied Mathematics, 84 (4):748–762, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/4/748/5536910.

Junior:2019:HEW

[1635] D. S. Almeida Júnior, I. Elishakoff, A. J. A. Ramos, and L. G. Rosário Miranda. The hypothesis of equal wave speeds for stabilization of Timoshenko beam is not necessary anymore: the time delay cases. IMA Journal of Applied Mathematics, 84 (4):763–796, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/4/763/5536911.

Behi-Gornostaeva:2019:TWS

[1636] E. El Behi-Gornostaeva, K. Mitra, and B. Schweizer. Traveling wave solutions for the Richards equation with hysteresis. IMA Journal of Applied Mathematics, 84(4):797–812, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/4/797/5536912.

Schnitzer:2019:GQL

[1637] Ory Schnitzer. Geometric quantization of localized surface plasmons. IMA Journal of Applied Mathematics, 84 (4):813–832, August 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/4/813/5536913.

Colombo:2019:MCM

[1638] Rinaldo M. Colombo and Elena Rossi.

Modelling crowd movements in domains with boundaries. IMA Journal of Applied Mathematics, 84(5):833– 853, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/84/5/833/
5560063.

Ozcakir:2019:CMP

[1639] Ozge Ozcakir, Philip Hall, and Saleh Tanveer. Centre modes in pipe flow. IMA Journal of Applied Mathematics, 84(5):854–872, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/5/854/5574026.

Shubov:2019:AES

[1640] Marianna A. Shubov and Laszlo P. Kindrat. Asymptotics of the eigenmodes and stability of an elastic structure with general feedback matrix. IMA Journal of Applied Mathematics, 84(5):873–911, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/5/873/5582261.

Martin:2019:TDB

[1641] P. A. Martin. Two-dimensional Brinkman flows and their relation to analogous Stokes flows. IMA Journal of Applied Mathematics, 84(5):912– 929, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/84/5/912/
5574707.

Burger:2019:SCL

[1642] Raimund Bürger, Stefan Diehl, and María del Carmen Martí. A system of conservation laws with discontinuous flux modelling flotation with sedimentation. IMA Journal of Applied Mathematics, 84(5):930–973, October 2019.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/5/930/5582259.

Cantin:2019:LTD

[1643] Guillaume Cantin, M. A. Aziz-Alaoui, and Nathalie Verdière. Large-time dynamics in complex networks of reaction-diffusion systems applied to a panic model. IMA Journal of Applied Mathematics, 84(5):974–1000, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/84/5/974/
5574708.

Sayli:2019:SNF

[1644] Mustafa Sayli, Yi Ming Lai, Rüdiger Thul, and Stephen Coombes. Synchrony in networks of Franklin bells. IMA Journal of Applied Mathematics, 84(5):1001–1021, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/5/1001/5574709.

Edwards:2019:MPE

[1645] D. A. Edwards, M. E. Mackay, Z. R.

Swain, C. R. Banbury, and D. D. Phan. Maximal 3D printing extrusion rates. IMA Journal of Applied Mathematics, 84(5):1022–1043, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/84/5/1022/5582258.

Fokas:2019:EET

[1646] Athanasios S. Fokas, Beatrice Pelloni, and Baoqiang Xia. Evolution equations on time-dependent intervals. IMA Journal of Applied Mathematics, 84(5): 1044–1060, October 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/5/1044/5582260.

DiCrescenzo:2019:ARW

[1647] Antonio Di Crescenzo, Claudio Macci, Barbara Martinucci, and Serena Spina. Analysis of random walks on a hexagonal lattice. IMA Journal of Applied Mathematics, 84(6): 1061–1081, December 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        84/6/1061/5628131.

Baravdish:2019:DSO

[1648] G. Baravdish, O. Svensson, M. Gulliksson, and Y. Zhang. Damped second order flow applied to image denoising. IMA Journal of Applied Mathematics, 84(6):1082–1111, December 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/84/6/1082/5641103.

Tomlin:2019:PAF

[1649] Ruben J. Tomlin and Susana N. Gomes. Point-actuated feedback control of multidimensional interfaces. IMA Journal of Applied Mathematics, 84(6):1112–1142, December 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/84/6/1112/5678719.

Emmrich:2019:EWS

[1650] Etienne Emmrich and Robert Lasarzik. Existence of weak solutions to a dynamic model for smectic-A liquid crystals under undulations. IMA Journal of Applied Mathematics, 84(6): 1143–1176, December 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/6/1143/5680193.

Bouet:2019:GCI

[1651] V. Bouet and A. Y. Klimenko. Graph clustering in industrial networks. IMA Journal of Applied Mathematics, 84(6): 1177–1202, December 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/6/1177/5680196. See erratum

[1659].

Radjen:2019:RTH

[1652] Anthony Radjen, Gabriele Gradoni, and Richard Tew. Reflection and transmission of high-frequency acoustic, electromagnetic and elastic waves at a distinguished class of irregular, curved boundaries. IMA Journal of Applied Mathematics, 84(6): 1203–1219, December 2019. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
84/6/1203/5681665.

BritodaSilva:2020:TPM

[1653] Romulo Brito da Silva, I-Shih Liu, and Mauro Antonio Rincon. A theory of porous media and harmonic wave propagation in poroelastic body. IMA Journal of Applied Mathematics, 85(1):1–26, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/

85/1/1/5721767. See erratum [1667].

James:2020:KKN

[1654] Guillaume James, Kirill Vorotnikov, and Bernard Brogliato. Kuwabara– Kono numerical dissipation: a new method to simulate granular matter. IMA Journal of Applied Mathematics, 85(1):27–66, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        85/1/27/5721768.

Bykov:2020:FSV

[1655] V. Bykov, Y. Cherkinsky, V. Gol’dshtein, N. Krapivnik, and U. Maas. Fast–slow vector fields of reaction–diffusion systems. IMA Journal of Applied Mathematics, 85(1):67–86, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/85/1/67/5736272.

Lyalinov:2020:ASS

[1656] Mikhail A. Lyalinov. Acoustic scattering by a semi-infinite angular sector with impedance boundary conditions, II: the far-field asymptotics. IMA Journal of Applied Mathematics, 85 (1):87–112, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        85/1/87/5760769.

Wootton:2020:SOA

[1657] Peter T. Wootton, Julius Kaplunov, and Danila Prikazchikov. A secondorder asymptotic model for Rayleigh waves on a linearly elastic half plane.

IMA Journal of Applied Mathematics, 85(1):113–131, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/85/1/113/5743496.

Nicolopoulos:2020:DEE

[1658] Anouk Nicolopoulos, Martin Campos Pinto, Bruno Després, and Patrick Ciarlet, Jr. Degenerate elliptic equations for resonant wave problems. IMA Journal of Applied Mathematics, 85 (1):132–159, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
85/1/132/5739942.

Bouet:2020:EGC

[1659] V. Bouet and A. Y. Klimenko. Erratum to: Graph clustering in industrial networks. IMA Journal of Applied Mathematics, 85(1):160, February 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/85/1/160/

5716256. See [1651].

Russell:2020:HAU

[1660] Matthew J. Russell and Oliver E.

Jensen. Homogenization approximations for unidirectional transport past randomly distributed sinks. IMA Journal of Applied Mathematics, 85(2):231– 262, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/85/2/231/
5324989.

Ohl:2020:BBB

[1661] Siew-Wan Ohl, Md Haiqal Haqim Bin Md. Rahim, Evert Klaseboer, and Boo Cheong Khoo. Blake, bubbles and boundary element methods. IMA Journal of Applied Mathematics, 85 (2):190–213, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        85/2/190/5680197.

Masuda:2020:ASI

[1662] Naoki Masuda, Victor M. Preciado, and Masaki Ogura. Analysis of the susceptible-infected-susceptible epidemic dynamics in networks via the non-backtracking matrix. IMA Journal of Applied Mathematics, 85 (2):214–230, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        85/2/214/5809070.

Ainsworth:2020:FPF

[1663] Mark Ainsworth and Zhiping Mao. Fractional phase-field crystal modelling: analysis, approximation and pattern formation. IMA Journal of Applied Mathematics, 85(2):231–262, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://academic.
                                        oup.com/imamat/article/85/2/231/
                                        5809071.

Griffiths:2020:NNC

[1664] P. T. Griffiths. Non-Newtonian channel flow-exact solutions. IMA Journal of Applied Mathematics, 85(2):263– 279, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/85/2/263/
5809065.

Guo:2020:MST

[1665] Hongxia Guo, Changfeng Gui, Ping Lin, and Mingfeng Zhao. Multiple solutions and their asymptotics for laminar flows through a porous channel with different permeabilities. IMA Journal of Applied Mathematics, 85 (2):280–308, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
85/2/280/5809067.

Huber:2020:SWC

[1666] Colin Huber, Meaghan Hoitt, Nathaniel S.

Barlow, Nicole Hill, Kimberlee Keithley, and Steven J. Weinstein. On the stability of waves in classically neutral flows. IMA Journal of Applied Mathematics, 85(2):309–340, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/85/2/309/5814602.

daSilva:2020:ETP

[1667] Romulo Brito da Silva, I-Shih Liu, and Mauro Antonio Rincon. Erratum to: A theory of porous media and harmonic wave propagation in poroelastic body. IMA Journal of Applied Mathematics, 85(2):341, April 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
85/2/341/5808790. See [1653].

Dambrine:2020:RTC

[1668] Julien Dambrine, Evi Noviani, and Morgan Pierre. Rankine-type cylinders having zero wave resistance in infinitely deep flows. IMA Journal of Applied Mathematics, 85(3):343–364, June 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/3/343/
5810958.

Wang:2020:NIN

[1669] Mian Wang, Zhan Wang, and Hennes Hajduk. Nonlinear interactions of nearly non-dispersive equatorial shallow-water waves. IMA Journal of Applied Mathematics, 85(3):365–384, June 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/3/365/
5820063.

Devine:2020:AMF

[1670] K. M. Devine, M. Vynnycky, S. L. Mitchell, and S. B. G. O’Brien. Analysis of a model for the formation of fold-type oscillation marks in the continuous casting of steel. IMA Journal of Applied Mathematics, 85(3):385– 420, June 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/85/3/385/
5822756.

Nethercote:2020:HCA

[1671] M. A. Nethercote, R. C. Assier, and I. D. Abrahams. High-contrast approximation for penetrable wedge diffraction. IMA Journal of Applied Mathematics, 85(3):421–466, June 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/85/3/421/5822762.

Plociniczak:2020:AAI

[1672] Lukasz Płociniczak. Asymptotic analysis of internal relaxation oscillations in a conceptual climate model. IMA Journal of Applied Mathematics, 85 (3):467–494, June 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/3/467/5827996.

Bardall:2020:GID

[1673] Aaron Bardall, Shih-Yuan Chen, Karen E. Daniels, and Michael Shearer. Gradient-induced droplet motion over soft solids. IMA Journal of Applied Mathematics, 85(3):495–512, June 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/3/495/
5828284.

Charette:2020:PFS

[1674] Laurent Charette, Colin B. Macdonald, and Wayne Nagata. Pattern formation in a slowly flattening spherical cap: delayed bifurcation. IMA Journal of Applied Mathematics, 85 (4):513–541, August 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/4/513/5835988.

Yu:2020:THE

[1675] H. Y. Yu and Sanboh Lee. Timeharmonic elastic singularities and oscillating indentation of layered solids. IMA Journal of Applied Mathematics, 85(4):542–563, August 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/4/542/5854483.

Kory:2020:ECB

[1676] Jakub Köry, Armin U. Krupp, Colin P. Please, and Ian M. Griffiths. The effect of compressibility on the behaviour of filter media. IMA Journal of Applied Mathematics, 85(4):564–583, August 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/4/564/
5868232.

Kamath:2020:IDA

[1677] Gopal Krishna Kamath, Krishna Jagannathan, and Gaurav Raina. Impact of delayed acceleration feedback on the classical car-following model. IMA Journal of Applied Mathematics, 85(4):584–604, August 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/4/584/5870647.

Kolokolnikov:2020:SAS

[1678] Theodore Kolokolnikov, Frédéric Paquin-Lefebvre, and Michael J. Ward. Stable asymmetric spike equilibria for the Gierer–Meinhardt model with a precursor field. IMA Journal of Applied Mathematics, 85(4):605–634, August 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/85/4/605/5877658.

Tseng:2020:GSC

[1679] Jui-Pin Tseng. Global synchronization of coupled reaction–diffusion neural networks with general couplings via an iterative approach. IMA Journal of Applied Mathematics, 85(4):635– 669, August 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/85/4/635/
5882128.

Champneys:2020:ISI

[1680] Alan Champneys. Introduction to the special issue for the IMA Lighthill– Thwaites Prize. IMA Journal of Applied Mathematics, 85(5):671–672, October 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/5/671/
5902682.

Butler:2020:DCA

[1681] Matthew D. Butler and Dominic Vella.

Detachment in capillary adhesion: the relative roles of tilting and separation. IMA Journal of Applied Mathematics, 85(5):673–702, October 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/5/673/5890187.

Williams:2020:LPM

[1682] J. G. Williams, L. Rouse, B. W. Turney, S. L. Waters, and D. E. Moulton. A lumped-parameter model for kidney pressure during stone removal. IMA Journal of Applied Mathematics, 85(5):703–723, October 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/5/703/5877807.

Franssen:2020:MMO

[1683] Linnea C. Franssen and Mark A. J. Chaplain. A mathematical multi-organ model for bidirectional epithelialmesenchymal transitions in the metastatic spread of cancer. IMA Journal of Applied Mathematics, 85(5):724–761, October 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/85/5/724/
                                        5901199.

Eggl:2020:SOS

[1684] Maximilian F. Eggl and Peter J. Schmid. Shape optimization of stirring rods for mixing binary fluids. IMA Journal of Applied Mathematics, 85 (5):762–789, October 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        85/5/762/5836940.

Colbrook:2020:FSA

[1685] Matthew J. Colbrook and Matthew J. Priddin. Fast and spectrally accurate numerical methods for perforated screens (with applications to Robin boundary conditions). IMA Journal of Applied Mathematics, 85(5):790–821, October 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/85/5/790/
                                        5875045.

Xu:2020:UIE

[1686] Xiaoxu Xu, Bo Zhang, and Haiwen Zhang. Uniqueness in inverse electromagnetic scattering problem with phaseless far-field data at a fixed frequency. IMA Journal of Applied Mathematics, 85(6):823–839, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/6/823/
5890186.

Jiang:2020:LSI

[1687] Yu Jiang, Gen Nakamura, and Haibing Wang. Locating small inclusions in diffuse optical tomography by a direct imaging method. IMA Journal of Applied Mathematics, 85(6): 840–864, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/6/840/5896976.

Barnard:2020:MMA

[1688] Richard C. Barnard, Kai Huang, and Cory Hauck. A mathematical model of asynchronous data flow in parallel computers. IMA Journal of Applied Mathematics, 85(6):865–891, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/85/6/865/5910101.

Huang:2020:MCM

[1689] Vincent Huang and James Unwin.

Markov chain models of refugee migration data. IMA Journal of Applied Mathematics, 85(6):892–912, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/85/6/892/
5912890.

Liu:2020:RNB

[1690] Chein-Shan Liu and Chih-Wen Chang.

Reducing the near boundary errors of nonhomogeneous heat equations by boundary consistent methods. IMA Journal of Applied Mathematics, 85 (6):913–932, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
85/6/913/5903960.

DiCristo:2020:RIE

[1691] Michele Di Cristo and Giacomo Milan. Reconstruction of inclusions in electrical conductors. IMA Journal of Applied Mathematics, 85(6):933–950, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/85/6/933/
5912815.

Zelati:2020:HHF

[1692] Michele Coti Zelati and Grigorios A. Pavliotis. Homogenization and hypocoercivity for Fokker–Planck equations driven by weakly compressible shear flows. IMA Journal of Applied Mathematics, 85(6):951–979, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/85/6/951/5918282.

Liu:2020:SST

[1693] Weifan Liu and Thomas P. Witelski. Steady states of thin film droplets on chemically heterogeneous substrates. IMA Journal of Applied Mathematics, 85(6):980–1020, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/85/6/980/5919193.

Johnstone:2020:FSC

[1694] Eleanor C. Johnstone and Philip Hall.

Free-stream coherent structures in the unsteady Rayleigh boundary layer. IMA Journal of Applied Mathematics, 85(6):1021–1040, December 2020. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/85/6/1021/5919894.

Xiang:2021:IES

[1695] Jianli Xiang and Guozheng Yan. The inverse elastic scattering by an impenetrable obstacle and a crack. IMA Journal of Applied Mathematics, 86 (1):1–29, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/1/1/5918286.

Jang:2021:NSA

[1696] T. S. Jang. A new solution approach to the Serre equations. IMA Journal of Applied Mathematics, 86 (1):30–57, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/1/30/5920451.

Alhushaybari:2021:CAI

[1697] A. Alhushaybari and J. Uddin. Convective and absolute instability of falling viscoelastic liquid jets surrounded by a gas. IMA Journal of Applied Mathematics, 86(1):58–75, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/1/58/5902224.

Morupisi:2021:APF

[1698] Kgomotso S. Morupisi and Chris Budd. An analysis of the periodically forced PP04 climate model, using the theory of non-smooth dynamical systems. IMA Journal of Applied Mathematics, 86(1):76–120, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/1/76/5922977.

Davies:2021:SFA

[1699] Penny J. Davies and Ingolf Sack. A stacked frequency approach for inhomogeneous time-dependent MRE: an inverse problem for the elastic shear modulus. IMA Journal of Applied Mathematics, 86(1):121–145, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/86/1/121/
5983111.

Lin:2021:SRF

[1700] Junshan Lin, Sang-Hyun Oh, and Hai Zhang. Sensitivity of resonance frequency in the detection of thin layer using nano-slit structures. IMA Journal of Applied Mathematics, 86(1): 146–164, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/1/146/5983112.

Holloway:2021:CEE

[1701] Ian Holloway and Sivaguru S. Sritharan. Compressible Euler equations on a sphere and elliptic-hyperbolic property. IMA Journal of Applied Mathematics, 86(1):165–187, February 2021.

CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/1/165/6029299.

Middleton:2021:HCN

[1702] Amy L. L. Middleton, Antal A. Járai, Jonathan H. P. Dawes, and Keith Briggs. How close is the nearest node in a wireless network? IMA Journal of Applied Mathematics, 86 (1):188–219, February 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/1/188/6055473.

Baldassari:2021:RSA

[1703] Lorenzo Baldassari and Andrea Scapin.

Reconstruction of a small acoustic inclusion via time-dependent polarization tensors. IMA Journal of Applied Mathematics, 86(2):221–246, April 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/86/2/221/
6134257.

Yang:2021:RSS

[1704] Jichen Yang and Jens D. M. Rademacher.

Reaction–subdiffusion systems and memory: spectra, Turing instability and decay estimates. IMA Journal of Applied Mathematics, 86(2):247– 293, April 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/2/247/
6151206.

Gin:2021:SRR

[1705] Craig Gin and Prabir Daripa. Stability results on radial porous media and Hele–Shaw flows with variable viscosity between two moving interfaces. IMA Journal of Applied Mathematics, 86(2):294–319, April 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/2/294/6175327.

Yang:2021:GRW

[1706] Bo Yang and Jianke Yang. General rogue waves in the three-wave resonant interaction systems. IMA Journal of Applied Mathematics, 86(2):378– 425, April 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/2/378/
6218144.

Devine:2021:VAT

[1707] K. M. Devine, M. Vynnycky, S. L. Mitchell, and S. B. G. O’Brien. A validated asymptotic thermomechanical model for air-gap formation in tapered moulds in the continuous casting of steel. IMA Journal of Applied Mathematics, 86(2):349–377, April 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/2/349/6218145.

Kow:2021:RIO

[1708] Pu-Zhao Kow and Jenn-Nan Wang. Reconstruction of an impenetrable obstacle in anisotropic inhomogeneous background. IMA Journal of Applied Mathematics, 86(2):320–348, April 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/86/2/320/

6218146. See corrigendum [1751].

Yan:2021:IBV

[1709] Zhenya Yan. An initial-boundary value problem for the general threecomponent nonlinear Schrödinger equations on a finite interval. IMA Journal of Applied Mathematics, 86 (3):427–489, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/3/427/6218143.

Yariv:2021:LTS

[1710] Ehud Yariv and Toby L. Kirk. Longitudinal thermocapillary slip about a dilute periodic mattress of protruding bubbles. IMA Journal of Applied Mathematics, 86(3):490–501, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/3/490/6243396.

Yang:2021:DBS

[1711] Mimi X. Yang, Fuqian Yang, and Sanboh Lee. Dielectric breakdown sizes of conducting plates. IMA Journal of Applied Mathematics, 86(3):502– 513, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/3/502/
6278613.

Burger:2021:MLF

[1712] Raimund Bürger, Julio Careaga, and Stefan Diehl. A method-of-lines formulation for a model of reactive settling in tanks with varying cross-sectional area. IMA Journal of Applied Mathematics, 86(3):514–546, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/3/514/6278612.

Keady:2021:SSF

[1713] Grant Keady. Steady slip flow of Newtonian fluids through tangential polygonal microchannels. IMA Journal of Applied Mathematics, 86(3):547– 564, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/3/547/
6279504.

Quintanilla:2021:INS

[1714] R. Quintanilla and K. R. Rajagopal. On the instability, nonexistence and spatial behaviour of the one-dimensional response of a new class of elastic bodies. IMA Journal of Applied Mathematics, 86(3):565– 576, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/3/565/
6280577.

Wang:2021:SSP

[1715] Zhi-An Wang and Xin Xu. Steady states and pattern formation of the density-suppressed motility model. IMA Journal of Applied Mathematics, 86(3):577–603, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/3/577/6283599.

Markaki:2021:DSM

[1716] Vanessa Markaki, Drosos Kourounis, and Antonios Charalambopoulos. A dual self-monitored reconstruction scheme on the TV-regularized inverse conductivity problem. IMA Journal of Applied Mathematics, 86(3):604– 630, June 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/3/604/
6284268.

Wang:2021:MLE

[1717] Yijun Wang, Stephen Guzik, Milija Zupanski, and Xinfeng Gao. The maximum likelihood ensemble filter for computational flame and fluid dynamics. IMA Journal of Applied Mathematics, 86(4):631–661, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/4/631/6291452.

Xiang:2021:FMS

[1718] Jianli Xiang and Guozheng Yan. The factorization method for the scattering by a mixed inhomogeneous medium. IMA Journal of Applied Mathematics, 86(4):662–687, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/4/662/6291511.

Fritz:2021:STG

[1719] Marvin Fritz, Christina Kuttler, Mabel L. Rajendran, Barbara Wohlmuth, and Laura Scarabosio. On a subdiffusive tumour growth model with fractional time derivative. IMA Journal of Applied Mathematics, 86(4):688– 729, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/4/688/
6294804.

Horikis:2021:IRS

[1720] Theodoros P. Horikis. Integrable reduction and solitons of the Fokas– Lenells equation. IMA Journal of Applied Mathematics, 86(4):730–738, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/86/4/730/
                                        6309916.

Fazly:2021:CDS

[1721] Mostafa Fazly. Critical domain sizes of a discrete-map hybrid and reactiondiffusion model on hostile exterior domains. IMA Journal of Applied Mathematics, 86(4):739–760, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/86/4/739/6310998.

Prugger:2021:ESF

[1722] Artur Prugger and Jens D. M. Rademacher. Explicit superposed and forced plane wave generalized Beltrami flows. IMA Journal of Applied Mathematics, 86(4):761–784, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/86/4/761/6310999.

deOliveira:2021:OSC

[1723] L. L. A. de Oliveira and M. V. Travaglia. Optimizing the spring constants of forced, damped and circular spring-mass systems-characterization of the discrete and periodic bi-Laplacian operator. IMA Journal of Applied Mathematics, 86(4):785–807, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/4/785/
6314452.

Saadi:2021:SOL

[1724] Fahad Al Saadi, Alan Champneys, Annette Worthy, and Ahmed Msmali. Stationary and oscillatory localized patterns in ratio-dependent predator–prey systems. IMA Journal of Applied Mathematics, 86(4):808– 827, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/4/808/
6311547.

Tiryakioglu:2021:RSW

[1725] Burhan Tiryakioglu and Ayse Tiryakioglu.

Radiation of sound waves from a coaxial duct with perforated screen. IMA Journal of Applied Mathematics, 86 (4):828–844, August 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/4/828/6312957.

Champneys:2021:EHS

[1726] Alan Champneys. Editorial to homoclinic snaking at 21: in memory of Patrick Woods. IMA Journal of Applied Mathematics, 86(5):845–855, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/5/845/
6383556.

Parra-Rivas:2021:OBS

[1727] P. Parra-Rivas, E. Knobloch, L. Gelens, and D. Gomila. Origin, bifurcation structure and stability of localized states in Kerr dispersive optical cavities. IMA Journal of Applied Mathematics, 86(5):856–895, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/5/856/6347202.

Holl:2021:LSP

[1728] Max Philipp Holl, Andrew J. Archer, Svetlana V. Gurevich, Edgar Knobloch, Lukas Ophaus, and Uwe Thiele. Localized states in passive and active phase-field-crystal models. IMA Journal of Applied Mathematics, 86(5):896– 923, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/5/896/
6318163.

Frohoff-Hulsmann:2021:LSC

[1729] Tobias Frohoff-Hülsmann and Uwe Thiele. Localized states in coupled Cahn–Hilliard equations. IMA Journal of Applied Mathematics, 86(5):924– 943, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/86/5/924/
6316625.

Bentley:2021:LPG

[1730] David C. Bentley and Alastair M. Rucklidge. Localized patterns in a generalized Swift–Hohenberg equation with a quartic marginal stability curve. IMA Journal of Applied Mathematics, 86(5):944–983, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/944/6337890.

Alrihieli:2021:SLB

[1731] Haifaa Alrihieli, Alastair M. Rucklidge, and Priya Subramanian. Spatial localization beyond steady states in the neighbourhood of the Takens– Bogdanov bifurcation. IMA Journal of Applied Mathematics, 86(5): 984–1009, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/984/6328308.

Groh:2021:LSA

[1732] Rainer M. J. Groh and Giles W. Hunt.

Localization and snaking in axially compressed and internally pressurized thin cylindrical shells. IMA Journal of Applied Mathematics, 86(5): 1010–1030, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/1010/6322316.

AlSaadi:2021:LPS

[1733] Fahad Al Saadi, Alan Champneys, and Nicolas Verschueren. Localized patterns and semi-strong interaction, a unifying framework for reaction-diffusion systems. IMA Journal of Applied Mathematics, 86(5): 1031–1065, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/1031/6352382.

Knobloch:2021:SPM

[1734] Edgar Knobloch and Arik Yochelis. Stationary peaks in a multivariable reaction–diffusion system: foliated snaking due to subcritical Turing instability. IMA Journal of Applied Mathematics, 86(5): 1066–1093, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/1066/6319941.

Gomila:2021:CER

[1735] Damià Gomila and Edgar Knobloch. Curvature effects and radial homoclinic snaking. IMA Journal of Applied Mathematics, 86(5):1094– 1111, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/1094/6352381.

Tian:2021:SBL

[1736] Moyi Tian, Jason J. Bramburger, and Björn Sandstede. Snaking bifurcations of localized patterns on ring lattices. IMA Journal of Applied Mathematics, 86(5):1112–1140, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/5/1112/6311546.

Hasan:2021:NCS

[1737] Cris R. Hasan, Hinke M. Osinga, Claire M. Postlethwaite, and Alastair M. Rucklidge. Numerical continuation of spiral waves in heteroclinic networks of cyclic dominance. IMA Journal of Applied Mathematics, 86(5): 1141–1163, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/1141/6321368.

Subramanian:2021:SSG

[1738] Priya Subramanian, Andrew J. Archer, Edgar Knobloch, and Alastair M. Rucklidge. Snaking without subcriticality: grain boundaries as nontopological defects. IMA Journal of Applied Mathematics, 86(5): 1164–1180, October 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/5/1164/6316086.

Berntsson:2021:ADR

[1739] Fredrik Berntsson, Jennifer Chepkorir, and Vladimir Kozlov. Accelerated Dirichlet–Robin alternating algorithms for solving the Cauchy problem for the Helmholtz equation. IMA Journal of Applied Mathematics, 86(6): 1181–1203, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/6/1181/6312956.

Keady:2021:TRT

[1740] Grant Keady. Torsional rigidity for tangential polygons. IMA Journal of Applied Mathematics, 86(6): 1204–1211, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/6/1204/6314451.

Balac:2021:AWG

[1741] Stéphane Balac, Monique Dauge, and Zoïs Moitier. Asymptotics for 2D whispering gallery modes in optical microdisks with radially varying index. IMA Journal of Applied Mathematics, 86(6): 1212–1265, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/6/1212/6347564.

Rame:2021:SAL

[1742] Enrique Ramé, Steven J. Weinstein, and Nathaniel S. Barlow. On the shape of air–liquid interfaces with surface tension that bound rigidly rotating liquids in partially filled containers. IMA Journal of Applied Mathematics, 86(6): 1266–1286, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/6/1266/6364550.

Hauge:2021:NAC

[1743] Jordan C. Hauge and Darren Crowdy. A new approach to the complex Helmholtz equation with applications to diffusion wave fields, impedance spectroscopy and unsteady Stokes flow. IMA Journal of Applied Mathematics, 86(6):1287–1326, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/6/1287/6364803.

Anderson:2021:BGS

[1744] Jeffrey R. Anderson and Keng Deng. Blow up and global solvability for an absorptive porous medium equation with memory at the boundary. IMA Journal of Applied Mathematics, 86(6): 1327–1348, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
86/6/1327/6374654.

Millet:2021:BSF

[1745] Annie Millet, Svetlana Roudenko, and Kai Yang. Behaviour of solutions to the 1D focusing stochastic L^{2}-critical and supercritical nonlinear Schrödinger equation with space–time white noise. IMA Journal of Applied Mathematics, 86(6):1349–1396, December 2021. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/86/6/1349/6380453.

Kallyadan:2022:DAR

[1746] Sreethin Sreedharan Kallyadan and Priyanka Shukla. Dynamical aspects of a restricted three-vortex problem. IMA Journal of Applied Mathematics, 87(1):1–19, February 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
87/1/1/6412928.

Eswarathasan:2022:LBS

[1747] Suresh Eswarathasan and Theodore Kolokolnikov. Laplace–Beltrami spectrum of ellipsoids that are close to spheres and analytic perturbation theory. IMA Journal of Applied Mathematics, 87(1):20–49, February 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/87/1/20/6423527.

Cao:2022:DSV

[1748] Daomin Cao, Weicheng Zhan, and Changjun Zou. On desingularization of steady vortex for the lake equations. IMA Journal of Applied Mathematics, 87(1):50–79, February 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        87/1/50/6425815.

Goddard:2022:NBC

[1749] B. D. Goddard, B. Gooding, H. Short, and G. A. Pavliotis. Noisy bounded confidence models for opinion dynamics: the effect of boundary conditions on phase transitions. IMA Journal of Applied Mathematics, 87 (1):80–110, February 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        87/1/80/6427747.

Leach:2022:LTS

[1750] J. A. Leach and Andrew P. Bassom. Long-time solutions of scalar hyperbolic reaction equations incorporating relaxation and the Arrhenius combustion nonlinearity. IMA Journal of Applied Mathematics, 87(1):111–128, February 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/87/1/111/
                                        6511199.

Kow:2022:CRI

[1751] Pu-Zhao Kow and Jenn-Nan Wang. Corrigendum to: Reconstruction of an impenetrable obstacle in anisotropic inhomogeneous background. IMA Journal of Applied Mathematics, 87(1): 129–130, February 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        87/1/129/6431881. See [1708].

Goh:2022:DHB

[1752] Ryan Goh, Tasso J. Kaper, and Theodore Vo. Delayed Hopf bifurcation and space-time buffer curves in the complex Ginzburg–Landau equation. IMA Journal of Applied Mathematics, 87(2):131–186, April 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
87/2/131/6540628.

Volkov:2022:PSA

[1753] Darko Volkov. A parallel sampling algorithm for some nonlinear inverse problems. IMA Journal of Applied Mathematics, 87(2):187–206, April 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/87/2/187/
6544229.

Gao:2022:SAS

[1754] Chao Gao and Xing-Min Chen. Stability analysis of a standby system with an unreliable server and switching failure. IMA Journal of Applied Mathematics, 87(2):207–230, April 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
87/2/207/6546852.

Malham:2022:ILN

[1755] Simon J. A. Malham. Integrability of local and non-local noncommutative fourth-order quintic nonlinear Schrödinger equations. IMA Journal of Applied Mathematics, 87 (2):231–259, April 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
87/2/231/6550296.

Ye:2022:FMI

[1756] Jianguo Ye and Guozheng Yan. The factorization method for inverse scattering by a two-layered cavity with conductive boundary condition. IMA Journal of Applied Mathematics, 87 (2):260–290, April 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
87/2/260/6562946.

Edwards:2022:QMW

[1757] D. A. Edwards. Quasisteady models for weld temperatures in fused filament fabrication. IMA Journal of Applied Mathematics, 87(2):291–313, April 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/87/2/291/
6562947.

Hill:2022:ELR

[1758] Dan J. Hill. Existence of localized radial patterns in a model for dryland vegetation. IMA Journal of Applied Mathematics, 87(3):315– 353, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/87/3/315/
6580514.

Fowler:2022:NEG

[1759] A. C. Fowler. Non-equilibrium growth of a binary alloy. IMA Journal of Applied Mathematics, 87(3):354– 379, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/87/3/354/
6583409.

Alsharif:2022:TIC

[1760] Abdullah M. Alsharif and Emilian I.

Parau. Temporal instability of curved viscous fibers with a radial electric field. IMA Journal of Applied Mathematics, 87(3):380–406, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/87/3/380/6583410.

Xiang:2022:FMM

[1761] Jianli Xiang and Guozheng Yan. The factorization method for a mixed inverse elastic scattering problem. IMA Journal of Applied Mathematics, 87 (3):407–437, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/3/407/6603923.

Kumar:2022:LWE

[1762] Suraj Kumar and S. K. Tomar. Lamb waves in elastic-plastic plate containing voids. IMA Journal of Applied Mathematics, 87(3):438–461, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/87/3/438/6619276.

Budd:2022:GBT

[1763] Chris Budd and Kgomotso S. Morupisi.

Grazing bifurcations and transitions between periodic states of the PP04 model for the glacial cycle. IMA Journal of Applied Mathematics, 87(3):462– 491, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/87/3/462/
6638812.

Luckins:2022:MAC

[1764] Ellen K. Luckins, James M. Oliver, Colin P. Please, Benjamin M. Sloman, Aasgeir M. Valderhaug, and Robert A. Van Gorder. Modelling alternating current effects in a submerged arc furnace. IMA Journal of Applied Mathematics, 87(3):492–520, June 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/87/3/492/6644717.

Kozyreff:2022:ASS

[1765] Gregory Kozyreff. Asymptotic solutions of the SIR and SEIR models well above the epidemic threshold. IMA Journal of Applied Mathematics, 87 (4):521–536, August 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/4/521/6649784.

Oliveira:2022:GMR

[1766] H. L. Oliveira, S. McKee, G. C. Buscaglia, J. A. Cuminato, I. W. Stewart, and D. J. Wheatley. A generalized mathematical representation of the shape of the Wheatley heart valve and the associated static stress fields upon opening and closing. IMA Journal of Applied Mathematics, 87 (4):537–567, August 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/4/537/6652561.

Moore:2022:PDL

[1767] Joshua W. Moore, Trevor C. Dale, and Thomas E. Woolley. Polarity-driven laminar pattern formation by lateralinhibition in 2D and 3D bilayer geometries. IMA Journal of Applied Mathematics, 87(4):568–606, August 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/87/4/568/6661184.

AlJebawy:2022:IPT

[1768] Hanin Al Jebawy and Abdellatif El Badia. On an inverse photoacoustic tomography problem of small absorbers with inhomogeneous sound speed. IMA Journal of Applied Mathematics, 87 (4):607–646, August 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/4/607/6657673.

Dominguez:2022:BIE

[1769] Víctor Domínguez and Catalin Turc.

Boundary integral equation methods for the solution of scattering and transmission 2D elastodynamic problems. IMA Journal of Applied Mathematics, 87(4):647–706, August 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/4/647/6674490.

Erbas:2022:AAD

[1770] Baris Erbas, Julius Kaplunov, and Gökçe Kiliç. Asymptotic analysis of 3D dynamic equations in linear elasticity for a thin layer resting on a Winkler foundation. IMA Journal of Applied Mathematics, 87(5):707–721, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/87/5/707/
                                        6712386.

Kudish:2022:LCL

[1771] Ilya I. Kudish, Eugene Pashkovski, and Reid Patterson. Line contact lubricated by a fluid described by non-Newtonian Giesekus model. IMA Journal of Applied Mathematics, 87 (5):722–756, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        87/5/722/6713907.

Sheng:2022:RPT

[1772] Wancheng Sheng and Qinglong Zhang. The Riemann problem for a traffic flow model on a road with variable widths. IMA Journal of Applied Mathematics, 87(5):757–785, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        87/5/757/6713908.

Zhou:2022:NDR

[1773] Liangqiang Zhou and Fangqi Chen. Nonlinear dynamic responses of an inclined beam to harmonic excitation in temperature field. IMA Journal of Applied Mathematics, 87(5):786–803, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/87/5/786/
                                        6725992.

Belinskiy:2022:SFO

[1774] Boris P. Belinskiy, Don B. Hinton, and Roger A. Nichols. Singular fourthorder Sturm–Liouville operators and acoustic black holes. IMA Journal of Applied Mathematics, 87(5):804–851, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/87/5/804/
6754048.

Geng:2022:HSS

[1775] Xianguo Geng, Kedong Wang, and Mingming Chen. The Hermitian symmetric space Fokas-Lenells equation: spectral analysis and long-time asymptotics. IMA Journal of Applied Mathematics, 87(5):852–905, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/87/5/852/6695022.

Giddings:2022:SHC

[1776] Josef Adam Giddings, John Billingham, and Stephen Michael Cox. Slugging in horizontal channel flow. IMA Journal of Applied Mathematics, 87 (5):906–933, October 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/5/906/6761024.

Hurst:2022:MLE

[1777] Christopher Hurst, Milija Zupanski, and Xinfeng Gao. The maximum likelihood ensemble smoother for the Kuramoto–Sivashinsky equation. IMA Journal of Applied Mathematics, 87 (6):935–963, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/6/935/6722604.

Papageorgiou:2022:OAE

[1778] Demetrios T. Papageorgiou, Yiorgos-Sokratis Smyrlis, and Ruben J. Tomlin. Optimal analyticity estimates for nonlinear active-dissipative evolution equations. IMA Journal of Applied Mathematics, 87(6):964–984, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/87/6/964/6760949.

Berthier:2022:AGA

[1779] Raphaël Berthier and Mufan (Bill) Li. Acceleration of gossip algorithms through the Euler–Poisson–Darboux equation. IMA Journal of Applied Mathematics, 87(6):985–1009, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://academic.
                                        oup.com/imamat/article/87/6/985/
                                        6775269.

Fokas:2022:EDF

[1780] A. S. Fokas and K. Kalimeris. Extensions of the d’Alembert formulae to the half line and the finite interval obtained via the unified transform. IMA Journal of Applied Mathematics, 87(6): 1010–1042, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        87/6/1010/6779978.

Cassidy:2022:NMH

[1781] Tyler Cassidy, Peter Gillich, Antony R. Humphries, and Christiaan H. van Dorp. Numerical methods and hypoexponential approximations for gamma distributed delay differential equations.

IMA Journal of Applied Mathematics, 87(6):1043–1089, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/87/6/1043/6794093.

Bonner:2022:REP

[1782] Gage Bonner, Jean-Luc Thiffeault, and Benedek Valkó. On a random entanglement problem. IMA Journal of Applied Mathematics, 87(6): 1090–1120, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/6/1090/6798415.

Ruhomally:2022:SMM

[1783] Yusra Bibi Ruhomally, Muhammad Zaid Dauhoo, and Laurent Dumas. Stochastic modelling of marijuana use in Washington: pre- and post-Initiative-502 (I-502). IMA Journal of Applied Mathematics, 87(6): 1121–1150, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/6/1121/6845440.

Burger:2022:DCD

[1784] Raimund Bürger, Stefan Diehl, M. Carmen Martí, and Yolanda Vásquez. A degenerating convection–diffusion system modelling froth flotation with drainage. IMA Journal of Applied Mathematics, 87(6):1151– 1190, December 2022. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
87/6/1151/6947314. See correction

[1790].

Colbrook:2023:CSP

[1785] Matthew J. Colbrook, Andrew Horning, Kyle Thicke, and Alexander B. Watson. Computing spectral properties of topological insulators without artificial truncation or supercell approximation. IMA Journal of Applied Mathematics, 88(1): 1–42, February 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        88/1/1/7017667.

Naghshineh:2023:UAM

[1786] Nastaran Naghshineh, W. Cade Reinberger, Nathaniel S. Barlow, Mohamed A. Samaha, and Steven J. Weinstein. On the use of asymptotically motivated gauge functions to obtain convergent series solutions to nonlinear ODEs. IMA Journal of Applied Mathematics, 88(1):43–66, February 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/88/1/43/7022108. See cor-

rection [1807].

Zanoni:2023:HRG

[1787] Andrea Zanoni. Homogenization results for the generator of multiscale Langevin dynamics in weighted Sobolev spaces. IMA Journal of Applied Mathematics, 88(1):67–101, February 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/88/1/67/
                                        7022107.

Bourgeois:2023:SPO

[1788] Laurent Bourgeois, Sonia Fliss, Jean-François Fritsch, Christophe Hazard, and Arnaud Recoquillay. Scattering in a partially open waveguide: the forward problem. IMA Journal of Applied Mathematics, 88(1):102–151, February 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/88/1/102/
7024678.

Farkas:2023:AES

[1789] Matthew Farkas, Jorge Cisneros, and Bernard Deconinck. The analytic extension of solutions to initial-boundary value problems outside their domain of definition. IMA Journal of Applied Mathematics, 88(1):152–212, February 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/88/1/152/
7035423.

Anonymous:2023:CDC

[1790] Anonymous. Correction to: A degenerating convection–diffusion system modelling froth flotation with drainage. IMA Journal of Applied Mathematics, 88(1):213, February 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/1/213/7034470. See [1784].

Carvalho:2023:SRU

[1791] Camille Carvalho and Zoïs Moitier.

Scattering resonances in unbounded transmission problems with signchanging coefficient. IMA Journal of Applied Mathematics, 88(2):215– 257, April 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/88/2/215/
                                        7048617.

Brander:2023:IPM

[1792] Tommi Brander, Daniel Lesnic, and Kai Cao. Inverse problems for a model of biofilm growth. IMA Journal of Applied Mathematics, 88(2):258– 281, April 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/88/2/258/
                                        7097857.

Nika:2023:HEM

[1793] Grigor Nika. On a hierarchy of effective models for the biomechanics of human compact bone tissue. IMA Journal of Applied Mathematics, 88 (2):282–307, April 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        88/2/282/7103513.

Yuan:2023:DAS

[1794] Yuan Yuan and Xianlong Fu. Dynamics of an age-structured HIV model with general nonlinear infection rate. IMA Journal of Applied Mathematics, 88(2):308–353, April 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        88/2/308/7103506.

Wang:2023:GTD

[1795] Kai Wang, Hao Wang, and Hongyong Zhao. Global threshold dynamics of a spatial chemotactic mosquitoborne disease model. IMA Journal of Applied Mathematics, 88(2):354– 377, April 2023. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/88/2/354/
7103511.

Baravdish:2023:IRP

[1796] G. Baravdish, B. T. Johansson, O. Svensson, and W. Ssebunjo. Identifying a response parameter in a model of brain tumour evolution under therapy. IMA Journal of Applied Mathematics, 88(2):378–404, April 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/88/2/378/7103503.

Breschi:2023:GTS

[1797] G. Breschi and M. A. Fontelos. On global in time self-similar solutions of Smoluchowski equation with multiplicative kernel. IMA Journal of Applied Mathematics, 88(2):405–428, April 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/88/2/405/
7103505.

Wroblewska:2023:SFP

[1798] Zofia Wróblewska, Piotr Kowalczyk, and Łukasz Płociniczak. Stability of fixed points in an approximate solution of the spring–mass running model. IMA Journal of Applied Mathematics, 88(3):429–454, June 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/3/429/7116029.

Hu:2023:TES

[1799] Pipi Hu, Peng Xie, and Yi Zhu. Traveling edge states in massive Dirac equations along slowly varying edges. IMA Journal of Applied Mathematics, 88(3):455–471, June 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/3/455/7135734.

Miyoshi:2023:ECC

[1800] Hiroyuki Miyoshi and Darren G. Crowdy. Estimating conformal capacity using asymptotic matching. IMA Journal of Applied Mathematics, 88 (3):472–497, June 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/3/472/7206011.

Biber:2023:APC

[1801] Stanisław W. Biber, Alan R. Champneys, and Robert Szalai. Analysis of point-contact models of the bounce of a hard spinning ball on a compliant frictional surface. IMA Journal of Applied Mathematics, 88(3):498– 523, June 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/88/3/498/

7235664. See correction [1850].

Chaikovskii:2023:SFI

[1802] Dmitrii Chaikovskii and Ye Zhang. Solving forward and inverse problems involving a nonlinear three-dimensional partial differential equation via asymptotic expansions. IMA Journal of Applied Mathematics, 88(4):525–557, August 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/88/4/525/
7252977.

Xiang:2023:URI

[1803] Jianli Xiang and Guozheng Yan.

Uniqueness of refractive indices and transmission coefficients by an inhomogeneous medium in acoustic scattering. IMA Journal of Applied Mathematics, 88(4):558–575, August 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/4/558/7258924.

Hu:2023:RAS

[1804] Wenjie Hu, Quanxin Zhu, and Tomás Caraballo. Random attractors for a stochastic nonlocal delayed reaction– diffusion equation on a semi-infinite interval. IMA Journal of Applied Mathematics, 88(4):576–601, August 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/88/4/576/7276455.

Brudvik-Lindner:2023:ORS

[1805] Larkspur Brudvik-Lindner, Dimitrios Mitsotakis, and Athanasios E. Tzavaras. Oscillatory and regularized shock waves for a dissipative Peregrine–Boussinesq system. IMA Journal of Applied Mathematics, 88 (4):602–631, August 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/4/602/7323593.

Pena:2023:IBV

[1806] Y. Pérez Peña, J. Ortíz Sánchez, F. J. Ariza Hernández, and M. P. Árciga Alejandre. Initial-boundary value problem for a fractional heat equation on an interval. IMA Journal of Applied Mathematics, 88(4):632–643, August 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/88/4/632/7317740. See cor-

rection [1815].

Anonymous:2023:CUA

[1807] Anonymous. Correction to: On the use of asymptotically motivated gauge functions to obtain convergent series solutions to nonlinear ODEs. IMA Journal of Applied Mathematics, 88(4): 644, August 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/88/4/644/

7279225. See [1786].

Han:2023:RLG

[1808] Yucen Han, Baoming Shi, Lei Zhang, and Apala Majumdar. A reduced Landau–de Gennes study for nematic equilibria in three-dimensional prisms. IMA Journal of Applied Mathematics, 88(5):645–676, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/5/645/7382157.

Fowler:2023:IGM

[1809] A. C. Fowler and Marian B. Holness. Interfacial growth morphologies in dense eutectic crystal mushes. IMA Journal of Applied Mathematics, 88 (5):677–701, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/5/677/7457444.

Babasola:2023:MAT

[1810] Oluwatosin Babasola and Chris Budd.

Mathematical analysis of a timedelayed model for cocoa yield. IMA Journal of Applied Mathematics, 88 (5):702–734, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/5/702/7459769.

Naghshineh:2023:SAM

[1811] Nastaran Naghshineh, W. Cade Reinberger, Nathaniel S. Barlow, Mohamed A. Samaha, and Steven J. Weinstein. The shape of an axisymmetric meniscus in a static liquid pool: effective implementation of the Euler transformation. IMA Journal of Applied Mathematics, 88(5):735–764, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/88/5/735/
7529154.

Murphy:2023:RDE

[1812] S. Murphy, M. Vynnycky, S. L.

Mitchell, and D. O’Kiely. Reaction dynamics and early-time behaviour of chemical decontamination. IMA Journal of Applied Mathematics, 88 (5):765–804, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/5/765/7589765.

Song:2023:BMG

[1813] Zilong Song, Robert Eisenberg, Shixin Xu, and Huaxiong Huang. A bubble model for the gating of K_{v} channels. IMA Journal of Applied Mathematics, 88(5):805–836, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        88/5/805/7589766.

Pei:2023:SMK

[1814] Hansen Pei, Pak-Wing Fok, and Louis Rossi. Stochastic modelling of Karlotoxin influence on prey. IMA Journal of Applied Mathematics, 88(5):837– 866, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/88/5/837/
                                        7596266.

Anonymous:2023:CIB

[1815] Anonymous. Correction to: Initialboundary value problem for a fractional heat equation on an interval. IMA Journal of Applied Mathematics, 88(5):867, October 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        88/5/867/7457913. See [1806].

Koens:2023:SVF

[1816] Lyndon Koens, Rohan Vernekar, Timm Krüger, Maciej Lisicki, and David W. Inglis. The slow viscous flow around a general rectangular doubly-periodic arrays of infinite slender cylinders. IMA Journal of Applied Mathematics, 88 (6):869–887, December 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        88/6/869/7595624.

Rautio:2023:LMP

[1817] Siiri Rautio, Rashmi Murthy, Tatiana A. Bubba, Matti Lassas, and Samuli Siltanen. Learning a microlocal prior for limited-angle tomography. IMA Journal of Applied Mathematics, 88(6):888–916, December 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/88/6/888/7603164.

Nitschke:2023:TTF

[1818] Ingo Nitschke, Souhayl Sadik, and Axel Voigt. Tangential tensor fields on deformable surfaces-how to derive consistent L^{2}-gradient flows. IMA Journal of Applied Mathematics, 88(6): 917–958, December 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/6/917/7610468.

Deliktas-Ozdemir:2023:CBE

[1819] Ekin Deliktas-Ozdemir. The comparison between effects of heterogeneous and homogeneous half-spaces underlying homogeneous layer on solitary love waves. IMA Journal of Applied Mathematics, 88(6):959–977, December 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/88/6/959/7612070.

Liu:2023:GEB

[1820] Chao Liu and Bin Liu. Global existence and boundedness in a two-species chemotaxis-fluid system with indirect pursuit-evasion interaction. IMA Journal of Applied Mathematics, 88(6): 978–1008, December 2023. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
88/6/978/7618447.

Lesnic:2024:EFS

[1821] D. Lesnic and M. J. Ehrhardt. Editorial forward to the special issue on “The Applied Mathematics of Machine Learning”. IMA Journal of Applied Mathematics, 89(1):1, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/89/1/1/7676631.

Cui:2024:MDR

[1822] Zhaofang Cui and Peter Grindrod. Mappings, dimensionality and reversing out of deep neural networks. IMA Journal of Applied Mathematics, 89 (1):2–11, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        89/1/2/7206013.

Benning:2024:TYS

[1823] Martin Benning, Tatiana A. Bubba, Luca Ratti, and Danilo Riccio. Trust your source: quantifying source condition elements for variational regularisation methods. IMA Journal of Applied Mathematics, 89(1):12–43, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/89/1/12/
                                        7626941.

Tyukin:2024:FIS

[1824] Ivan Y. Tyukin, Desmond J. Higham, Alexander Bastounis, Eliyas Woldegeorgis, and Alexander N. Gorban. The feasibility and inevitability of stealth attacks. IMA Journal of Applied Mathematics, 89(1):44–84, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/89/1/44/7323594.

losReyes:2024:IML

[1825] Juan Carlos De los Reyes and David Villacís. Interpretable model learning in variational imaging: a bilevel optimization approach. IMA Journal of Applied Mathematics, 89(1):85– 122, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/89/1/85/
7258921.

Hao:2024:LRW

[1826] Dinh Nho Hào, Duong Xuan Hiep, and Pham Quy Muoi. Learning river water quality models by l^{1}-weighted regularization. IMA Journal of Applied Mathematics, 89(1):123–142, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/89/1/123/7258925.

Grossmann:2024:CPI

[1827] Tamara G. Grossmann, Urszula Julia Komorowska, Jonas Latz, and Carola-Bibiane Schönlieb. Can physicsinformed neural networks beat the finite element method? IMA Journal of Applied Mathematics, 89(1):143– 174, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/89/1/143/
7680268.

Beerens:2024:AIC

[1828] Lucas Beerens and Desmond J.

Higham. Adversarial ink: componentwise backward error attacks on deep learning. IMA Journal of Applied Mathematics, 89(1):175–196, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic).  URL http://academic.
                                        oup.com/imamat/article/89/1/175/
                                        7209578.

Bui-Thanh:2024:UCF

[1829] Tan Bui-Thanh. A unified and constructive framework for the universality of neural networks. IMA Journal of Applied Mathematics, 89(1):197– 230, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/89/1/197/
                                        7413194.

Park:2024:NIP

[1830] Hyoung Suk Park, Chang Min Hyun, and Jin Keun Seo. Nonlinear illposed problem in low-dose dental conebeam computed tomography. IMA Journal of Applied Mathematics, 89 (1):231–253, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        89/1/231/7206012.

Ehrhardt:2024:AIH

[1831] Matthias J. Ehrhardt and Lindon Roberts. Analyzing inexact hypergradients for bilevel learning. IMA Journal of Applied Mathematics, 89 (1):254–278, January 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        89/1/254/7456287.

Katsiavria:2024:SAV

[1832] Anna Katsiavria and Demetrios T.

Papageorgiou. Stability analysis of viscous multi-layer shear flows with interfacial slip. IMA Journal of Applied Mathematics, 89(2):279–317, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic).  URL http://academic.
oup.com/imamat/article/89/2/279/
7692945.

Shelton:2024:MOE

[1833] Josh Shelton and Philippe H. Trinh. A model ODE for the exponential asymptotics of nonlinear parasitic capillary ripples. IMA Journal of Applied Mathematics, 89(2):318–342, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://academic.oup.com/imamat/
article/89/2/318/7689119.

Rees-Zimmerman:2024:SDF

[1834] Clare R. Rees-Zimmerman and Alexander F. Routh. Stratification in drying films: diffusiophoresis driven by nanoparticles and their counterions. IMA Journal of Applied Mathematics, 89(2):343–373, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
89/2/343/7687988.

McDonald:2024:ENS

[1835] N. R. McDonald and Samuel J. Harris. Exact and numerical solutions of a free boundary problem with a reciprocal growth law. IMA Journal of Applied Mathematics, 89(2):374– 386, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/89/2/374/
                                        7687482.

Song:2024:GES

[1836] Cui Song, Zhi-Cheng Wang, and Zhaosheng Feng. Global existence and steady states of the density-suppressed motility model with strong Allee effect. IMA Journal of Applied Mathematics, 89(2):387–425, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        89/2/387/7687987.

Bustamante:2024:IMS

[1837] R. Bustamante, P. Arrue, O. Orellana, and R. Meneses. An iteration method to study nonlinear wave propagation for a non-Green elastic 1D bar. IMA Journal of Applied Mathematics, 89(2):426–440, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

                                        academic.oup.com/imamat/article/
                                        89/2/426/7691191.

Elling:2024:CSP

[1838] Volker W. Elling. Convexity of shock polars. IMA Journal of Applied Mathematics, 89(2):441–462, April 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        http://academic.oup.com/imamat/
                                        article/89/2/441/7698155.

Kunz:2024:DRA

[1839] Valentin D. Kunz and Raphael C. Assier. Diffraction by a right-angled no-contrast penetrable wedge: recovery of far-field asymptotics. IMA Journal of Applied Mathematics, 89(3):463– 497, June 2024. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/89/3/463/

7758664. See correction [1855].

Finney:2024:ICD

[1840] Simon M. Finney, Matthew G. Hennessy, Andreas Münch, and Sarah L. Waters. The impact of confinement on the deformation of an elastic particle under axisymmetric tube flow. IMA Journal of Applied Mathematics, 89(3):498–532, June 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL http://

academic.oup.com/imamat/article/
89/3/498/7759751.

Neto:2024:EMS

[1841] Antônio Francisco Neto. An explicit Maclaurin series solution to nonautonomous and non-homogeneous evolution equation, Omega Calculus and associated applications. IMA Journal of Applied Mathematics, 89(3):533– 573, June 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/89/3/533/
7710816.

Hulse:2024:TBT

[1842] Jesse J. Hulse, Loredana Lanzani, Stefan G. Llewellyn Smith, and Elena Luca. A transform-based technique for solving boundary value problems on convex planar domains. IMA Journal of Applied Mathematics, 89(3):574– 597, June 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://academic.
oup.com/imamat/article/89/3/574/
7695384.

Kamizawa:2024:IQP

[1843] Takeo Kamizawa, Andrzej Jamiołkowski, Takashi Matsuoka, and Noboru Watanabe. On the p-irreducibility of quintic positive polynomials. IMA Journal of Applied Mathematics, 89(3):598– 621, June 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL http://academic.
                                        oup.com/imamat/article/89/3/598/
                                        7735383.

Rodriguez-Broadbent:2024:AEF

[1844] Henry Rodriguez-Broadbent, Hiroyuki Miyoshi, and Darren G Crowdy. Asymptotically exact formulas for channel flows over liquid-infused surfaces. IMA Journal of Applied Mathematics, 89(4):623–660, August 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/89/4/623/7817827.

Zhang:2024:LTA

[1845] Xiao-Fan Zhang and Shou-Fu Tian. On long-time asymptotic behavior and Painlevé asymptotic to the matrix Hirota equation. IMA Journal of Applied Mathematics, 89(4):661–704, August 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/89/4/661/
                                        7818314.

Toumi:2024:CSM

[1846] Samir Toumi and Rhouma Mlayeh. Control strategies for mitigating torsional and axial vibrations in rotary oilwell drilling systems. IMA Journal of Applied Mathematics, 89(4):705– 724, August 2024. CODEN IJAMDM.

ISSN 0272-4960 (print), 1464-3634

(electronic). URL https://academic.
oup.com/imamat/article/89/4/705/
7821057.

Grinfeld:2024:PSR

[1847] Michael Grinfeld and David Pritchard.

Pendent steady rivulets and droplets: from lubrication to bifurcation. IMA Journal of Applied Mathematics, 89 (4):725–744, August 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
89/4/725/7826680.

Tasse:2024:MMI

[1848] Arsène Jaurès Ouemba Tassé, Berge Tsanou, Cletus Kwa Kum, and Jean Lubuma. A mathematical model on the impact of awareness and traditional medicine in the control of Ebola: case study of the 2014–2016 outbreaks in Sierra Leone and Liberia. IMA Journal of Applied Mathematics, 89 (4):745–775, August 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
89/4/745/7817826.

Filelis-Papadopoulos:2024:PRF

[1849] Christos Filelis-Papadopoulos, John P Morrison, and Philip O’Reilly. A parallel recursive framework for modelling time series. IMA Journal of Applied Mathematics, 89(4):776–805, August 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/89/4/776/
7818310.

Anonymous:2024:CAP

[1850] Anonymous. Correction to: Analysis of point-contact models of the bounce of a hard spinning ball on a compliant frictional surface. IMA Journal of Applied Mathematics, 89(4):806, August 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/89/4/806/

7818385. See [1801].

Woods:2024:MEI

[1851] Tilly Woods and Ian J. Hewitt. Modelling the evolution of an ice sheet’s weathering crust. IMA Journal of Applied Mathematics, 89(5):807–832, October 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/89/5/807/

7903156. See correction [1856].

Bertoli:2024:SSS

[1852] Benedetta Bertoli, Benjamin D. Goddard, and Grigorios A. Pavliotis. Stability of stationary states for mean field models with multichromatic interaction potentials. IMA Journal of Applied Mathematics, 89(5):833–859, October 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/89/5/833/

7953227. See correction [1869].

Kahlaoui:2024:ATS

[1853] Hamza Kahlaoui, Sabeur Mansouri, and Mohamed BenSalah. Application of the topological sensitivity method to the detection of breast cancer. IMA Journal of Applied Mathematics, 89 (5):860–890, October 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
89/5/860/7328856.

Zhang:2024:NMC

[1854] Yupeng Zhang, Yueyang Shen, Rongqian Zhang, Yuyao Liu, Yunjie Guo, Daxuan Deng, and Ivo D. Dinov. Numerical methods for computing the discrete and continuous Laplace transforms. IMA Journal of Applied Mathematics, 89(5):891–950, October 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/89/5/891/7915386.

Anonymous:2024:CDR

[1855] Anonymous. Correction to: Diffraction by a right-angled no-contrast penetrable wedge: recovery of far-field asymptotics. IMA Journal of Applied Mathematics, 89(5):951, October 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/89/5/951/

7913648. See [1839].

Anonymous:2024:CME

[1856] Anonymous. Correction to: Modelling the evolution of an ice sheet’s weathering crust. IMA Journal of Applied Mathematics, 89(5):952, October 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/89/5/952/

7950738. See [1851].

Wu:2024:GRW

[1857] Chengfa Wu, Guangxiong Zhang, Changyan Shi, and Bao-Feng Feng. General rogue wave solutions to the Sasa–Satsuma equation. IMA Journal of Applied Mathematics, 89(6): 953–975, December 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        89/6/953/7994229.

Davies:2024:DAN

[1858] Penny J. Davies. Data approximation by neural nets for the MRE inverse problem in the frequency and time domains. IMA Journal of Applied Mathematics, 89(6):976–988, December 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/89/6/976/8020502.

Kevrekidis:2024:SSS

[1859] P. G. Kevrekidis, D. E. Pelinovsky, and R. M. Ross. Stability of smooth solitary waves under intensity-dependent dispersion. IMA Journal of Applied Mathematics, 89(6):989–1005, December 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/89/6/989/
                                        8089912.

Strickland:2024:RHO

[1860] Stephen L. Strickland, Karen E. Daniels, and Michael Shearer. The role of higher-order viscous and interfacial effects on the onset of surfactantcovered Faraday waves. IMA Journal of Applied Mathematics, 89(6): 1006–1040, December 2024. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
89/6/1006/8110040.

Bhat:2025:NAM

[1861] Manasa Bhat and Santanu Manna. A novel approach to modelling non-local surface waves: refinements in boundary and interface conditions. IMA Journal of Applied Mathematics, 90 (1):1–28, January 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
90/1/1/8100462.

Zhou:2025:SGK

[1862] Baoquan Zhou, Hao Wang, Tianxu Wang, and Daqing Jiang. Stochastic generalized Kolmogorov systems with small diffusion. II: Explicit approximations for periodic solutions in distribution. IMA Journal of Applied Mathematics, 90(1):29–74, January 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/90/1/29/8114607.

Abbass:2025:HAD

[1863] Ghulam Abbass, Haibo Chen, Muhammad Abdullahi, Abubakar Sani Halilu, and Israr Ahmed Memon. Hybrid accelerated double step length approach for convex constrained nonlinear monotone equations with applications. IMA Journal of Applied Mathematics, 90 (1):75–98, January 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
90/1/75/8119976.

Bustamante:2025:PLC

[1864] R. Bustamante, K. R. Rajagopal, and C. Rodriguez. Piecewise linear constitutive relations for stretch-limited elastic strings. IMA Journal of Applied Mathematics, 90(1):99–114, January 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/90/1/99/
                                        8113206.

Mandal:2025:DNN

[1865] Bankim C. Mandal and Deeksha Tomer. Dirichlet–Neumann and Neumann–Neumann waveform relaxation methods for PDEs with time delay. IMA Journal of Applied Mathematics, 90(2):115–139, April 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/90/2/115/8149107.

He:2025:PDN

[1866] Yuan He, Rong Wang, and Fei-Ying Yang. Propagation dynamics in a non-local dispersal system with shifting habitats. IMA Journal of Applied Mathematics, 90(2):140–173, April 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

                                        tronic). URL https://academic.
                                        oup.com/imamat/article/90/2/140/
                                        8152716.

Nethercote:2025:AWD

[1867] M. A. Nethercote, A. V. Kisil, and R. C. Assier. Acoustic wave diffraction by a quadrant of sound-soft scatterers. IMA Journal of Applied Mathematics, 90(2):174–203, April 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
90/2/174/8210302.

Zhu:2025:CMC

[1868] Neng Zhu, Wanwan Wang, and Kun Zhao. Critical mass and competing effect in chemotaxis: insights from an attraction–repulsion model with rotational fluxes. IMA Journal of Applied Mathematics, 90(2):204–230, April 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (elec-

tronic). URL https://academic.
oup.com/imamat/article/90/2/204/
8213837.

Anonymous:2025:CSS

[1869] Anonymous. Correction to: Stability of stationary states for mean field models with multichromatic interaction potentials. IMA Journal of Applied Mathematics, 90(2):231, April 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

academic.oup.com/imamat/article/
90/2/231/8157363. See [1852].

Walters:2025:BBL

[1870] Martin C. Walters and Robert J. Whittaker. Bending boundary layers near the clamped end of a thin-walled elastic tube. IMA Journal of Applied Mathematics, 90(3):233–276, June 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

https://academic.oup.com/imamat/
article/90/3/233/8157712.

Haji-Ali:2025:WCA

[1871] Abdul-Lateef Haji-Ali, Håkon Hoel, and Raúl Tempone. Weak convergence analysis in the particle limit of the McKean–Vlasov equation using stochastic flows of particle systems. IMA Journal of Applied Mathematics, 90(3):277–304, June 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL https://

                                        academic.oup.com/imamat/article/
                                        90/3/277/8210981.

Ren:2025:DAT

[1872] Yu Ren and Liangqiang Zhou. Dynamic analysis of a three-degree-of-freedom collision vibration system with clearances. IMA Journal of Applied Mathematics, 90(3):305–327, June 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

                                        https://academic.oup.com/imamat/
                                        article/90/3/305/8220720.

Zimmermann:2025:ESL

[1873] Sebastian Zimmermann, Clarissa Schönecker, and Darren G Crowdy. Exact solutions for laminar pressuredriven flow through a pipe with a concentric internal grating. IMA Journal of Applied Mathematics, 90(3):328– 346, June 2025. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

                                        (electronic). URL https://academic.
                                        oup.com/imamat/article/90/3/328/
                                        8241288.